### v1

# Post-detection inference for sequential changepoint localization

Abstract
This paper addresses a fundamental but largely unexplored challenge in sequential changepoint analysis: conducting inference following a detected change. We study the problem of localizing the changepoint using only the data observed up to a data-dependent stopping time at which a sequential detection algorithm $\mathcal{A}$ declares a change. We first construct confidence sets for the unknown changepoint when pre- and post-change distributions are assumed to be known. We then extend our framework to composite pre- and post-change scenarios. We impose no conditions on the observation space or on $\mathcal{A}$ — we only need to be able to run $\mathcal{A}$ on simulated data sequences. In summary, this work offers both theoretically sound and practically effective tools for sequential changepoint localization.
1 Introduction
We consider the following problem of sequential change analysis in a general space $\mathcal{X}$: a sequence of $\mathcal{X}$-valued observations $X_{1},X_{2},\cdots$ arrives sequentially, such that for some unknown time $T\in\mathbb{N}\cup\{\infty\}$ (the changepoint),
$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},$ (1.1)
for some pre-change distribution $F_{0}\in\mathcal{P}$ and some post-change distribution $F_{1}\in\mathcal{P}$, where $\mathcal{P}$ denotes a class of probability distributions on $\mathcal{X}.$ We start by assuming $F_{0}$ and $F_{1}$ are known, and generalize our methods to encompass settings where $F_{0}$ and $F_{1}$ are only known to lie in some classes, but are not known exactly. A sequential changepoint detection algorithm $\mathcal{A}$ is simply a data-dependent stopping rule $\tau$ to raise an alarm that a change has previously occurred. Our aim in this paper is to construct confidence sets for the unknown changepoint $T$ after stopping at $\tau$, assuming only black-box access to $\mathcal{A}$ (and ideally no further assumptions or restrictions). In other words, we hope to obtain a set $\mathcal{C}\subseteq\{1,\dots,\tau\}$ such that
$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha,$
where $\mathbb{P}_{T}$ is the data distribution in (1.1), i.e. having the true changepoint at $T$. We will show for algorithms $\mathcal{A}$ with a bounded average run length (ARL), no set $\mathcal{C}$ can satisfy such a guarantee, but if $\mathcal{A}$ has probability of false alarm (PFA) bounded above by $\alpha$, such a guarantee is achievable. In either case, another natural criterion for $\mathcal{C}$ is for it to satisfy
$\mathbb{P}_{T}(T\in\mathcal{C}|\tau\geq T)\geq 1-\alpha,$
meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$. We show how to construct such sets $\mathcal{C}$ for both types of algorithms (those with bounded ARL or PFA).
Sequential (also known as the online or quickest) changepoint detection is a well-studied field, where the goal is to detect that a change has occurred as soon as possible, is crucial in applications requiring timely responses to distributional shifts, such as quality control, cybersecurity, medical diagnostics, and so on. The field traces back to early work on statistical process control. Shewhart, (1925) introduced the concept of a control chart, which computes a test statistic for each sample, with values within a threshold indicating the process remains in control. Page, (1954) proposed the CUSUM (cumulative sum) procedure, which computes the cumulative log-likelihood ratio under the assumption of two exactly known distributions. Another influential approach is the Shiryaev-Roberts (SR) procedure (Shiryaev,, 1963; Roberts,, 1966). These early methods, however, assume the pre- and post-change distributions to be fully specified. To relax the assumption of known post-change distributions, the Generalized Likelihood Ratio (GLR) procedure was introduced (Lorden,, 1971; Siegmund and Venkatraman,, 1995). Tartakovsky et al., (2014) provide comprehensive overviews of these classical methods, including CUSUM, SR, and GLR, and their numerous variants. Several nonparametric change-point detection algorithms have been developed, such as kernel-based methods (Harchaoui et al.,, 2008; Li et al.,, 2015) and e-detectors (Shin et al.,, 2023).
Localizing the changepoint is another fundamental and crucial problem. As far as we are aware, the existing approaches to changepoint localization focus on offline or batch settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times. Hinkley, (1970) shows some asymptotic results about the maximum likelihood estimator (MLE) of the changepoint and also provides a confidence set based on likelihood ratio statistics. Yao, (1987) provides an approximation of the distribution of the MLE of the changepoint. Darkhovskh, (1976) presents a nonparametric changepoint estimator based on Mann-Whitney statistics. Carlstein, (1988) proposes a class of strongly consistent estimators for the changepoint in the nonparametric setting. Worsley, (1986) develops likelihood ratio-based confidence regions and tests for a changepoint for exponential family random variables. Dumbgen, (1991) constructs nonparametric confidence regions for changepoint by testing for each candidate time-point $t$ at level $\alpha$ using bootstrapping, achieving asymptotic coverage of $1-\alpha$, as the sample size $n\to\infty$. A more recent method by Jang and Walther, (2024) constructs confidence sets by evaluating a local test statistic on a triplet of time points and combining the results of the local tests with a weighted Bonferroni correction. To re-emphasize, all these methods are for the offline setting, while our main focus is on the sequential problem of constructing confidence sets after stopping, which has been largely overlooked in the literature. This gap is significant because in most real-world applications, data arrives sequentially, and making timely inferences about the changepoint immediately upon detecting a change is essential. Thus, developing methodologies for post-detection inference in a sequential framework is a significant step forward in changepoint analysis.
The key idea behind constructing the confidence interval is simple: for each candidate $t\in\{1,\cdots,\tau\}$, we will test whether $t$ is the true changepoint and form the confidence set by inverting these tests. Our approach is conceptually similar to Dumbgen, (1991), it adapts those ideas to the sequential setting in a nontrivial manner. For example, we show that for change detection algorithms that have finite average run length (ARL), it is impossible to obtain finite length confidence intervals with marginal coverage guarantees. Further, for $1-\alpha$ conditional coverage, we show that we need to test the $t$-th hypothesis at level $\alpha\times\mathbb{P}_{\infty}(\tau\geq t)$ (a quantity we do not know apriori), whereas testing for each $t$ at level $\alpha$ works in the offline setting. We develop simulation based approaches to get around these hurdles.
There are two settings for the changepoint problem: non-partitioned and partitioned. The non-partitioned changepoint problem assumes that the data-generating distribution changes from one unknown distribution $F_{0}\in\mathcal{P}$ to another unknown distribution $F_{1}\in\mathcal{P}$, without requiring any pre-specified partitioning of $\mathcal{P}$ into pre- and post-change distribution classes (eg: detecting a change in a Gaussian mean from any value to any other value). In contrast, the partitioned setting assumes that we know some partition of $\mathcal{P}$ into $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ such that $F_{0}$ and $F_{1}$ are known to lie in $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively. This formulation is useful when changes within $\mathcal{P}_{0}$ are not of interest, and we only seek to detect shifts from $\mathcal{P}_{0}$ to $\mathcal{P}_{1}$ (eg: detecting a change in Gaussian mean from negative to positive). Our method is designed for the partitioned setting. While extending it to the non-partitioned case is an interesting direction, it is left for future work.
Our contributions.
Here we summarize the key contributions of the paper.
  • For detection algorithms that control PFA, we propose Algorithm 1 for constructing a confidence set for the changepoint and we establish its unconditional and conditional coverage guarantees in Theorem 2.2 and Corollary 2.6, assuming both pre- and post-change distributions to be known.
  • For detection algorithms with bounded ARL, we show in Proposition 2.4 that we cannot have an unconditional coverage guarantee and propose Algorithm 2, which is a modification of Algorithm 1. We provide its conditional coverage guarantee in Theorem 2.7.
  • We extend our framework to composite pre- and post-change classes in Algorithm 3 and Algorithm 4 and provide the coverage guarantees in Theorem 3.1 and Theorem 4.2.
  • We demonstrate that the approach is not overly conservative (and thus practical) by examining the achieved coverage and size properties of our constructed confidence sets in extensive simulations.
We point out again that for obtaining the aforementioned guarantees, we just need to have black-box access to the detection algorithm $\mathcal{A}$, meaning that we do not need to know its inner workings (e.g., it can be a heuristic algorithm), and mainly need to be able to run that algorithm on any stream of data.
Paper outline.
The rest of the paper is organized as follows. In Section 2, we construct our confidence sets for the changepoint for known pre- and post-change and study its unconditional and conditional coverage probabilities. Section 3 extend this approach to handle composite post-change, assuming the pre-change distribution to be known. In Section 4, we finally extend this approach to handle composite pre- and post-change setting. Section 5.1 and Section 5.2 present a comprehensive set of simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach. This article is concluded in Section 6. Omitted proofs can be found in Appendix A. Some additional experimental results are provided in Appendix B and some heuristic algorithms are presented in Appendices C and D, which are computationally more efficient. A summary of comparisons of the heuristic and non-heuristic methods is provided in Appendix E, pointing at open avenues for further theoretical work.
2 Known pre- and post-change distributions
We now address the question of constructing confidence sets for the changepoint, assuming that our change detection procedure stopped after a changepoint. We start from the simplest case when both pre- and post-change distributions are known, addressing more general cases in later sections.
Notation.
For $t\in\mathbb{N}$, let $\mathbb{P}_{t}$ denote the joint distribution of the sequence of independent random variables $X_{1},X_{2},\dots$ when there is a changepoint at time t, meaning that $X_{1},\dots,X_{t-1}$ are from $F_{0}$ (with density $f_{0}$) and $X_{t},X_{t+1},\dots$ are from $F_{1}$ (with density $f_{1}$). As corner cases, let $\mathbb{P}_{1}$ denotes the joint distribution where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}$ and $\mathbb{P}_{\infty}$ denotes the case where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}$. Let $\mathbb{E}_{t}$ denote the expectation under $\mathbb{P}_{t}$, for $t\in\mathbb{N}\cup\{\infty\}$. Suppose we have a data sequence $\{X_{n}\}\sim\mathbb{P}_{T}$ for some unknown $T\in\mathbb{N}\cup\{\infty\}$ and some sequential change detection algorithm $\mathcal{A}$ that raises an alarm at the stopping time $\tau$, which is finite for the observed sequence.
In sequential change detection, there are typically two types of Type I error metrics. One is the average run length (ARL), which is defined as $\mathbb{E}_{\infty}[\tau]$, the expected time until a false alarm occurs under $T=\infty$ (i.e., no change). The other one is the probability of false alarm (PFA), that is, $\mathbb{P}_{\infty}(\tau<\infty)$. We will provide solutions for algorithms $\mathcal{A}$ that control either metric.
2.1 A first confidence set
Our main idea is to test whether there is a changepoint at $t$, for each $t\in\{1,\cdots,\tau\}$ and then construct the confidence set consisting of all time points $t$ for which we fail to reject the tests. Formally, we test the following null hypothesis at each candidate time point $t$:
$H_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\text{and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}.$
We consider some test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ such that for each $t\in\mathbb{N}$,
$M_{t}:=\begin{cases}M_{t}^{(\tau)}(X_{1},\cdots,X_{\tau}),&\text{ if }t\leq\tau<\infty,\\ -\infty&\text{ if }t>\tau\text{ or }\tau=\infty,\end{cases}$ (2.1)
where $M_{t}^{(n)}$ is some test statistic based on $n$ many samples, for each $t,n\in\mathbb{N},t\leq n$. While our theoretical guarantees will not depend on any particular choice of $M_{t}$, we provide a prototypical choice in a later subsection that we use for our experiments.
To test $H_{0,t}$, we draw $B$ many independent streams of data under $H_{0,t}$. For the $j$th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at the stopping time $\tau_{j}^{\prime}$ using $\mathcal{A}$ or until time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$ (which we discuss in detail in the next subsection), based on $\tau_{j}^{\prime},L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$, which is denoted as $M^{j}_{t,L}$. We reject $H_{0,t}$ if $M_{t}>\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, and the confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$, i.e,
$\mathcal{C}=\Big{\{}t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\Big{\}}.$ (2.2)
Algorithm 1 provides an overview of the method.
Input : $\alpha\in(0,1)$, $B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, algorithm $\mathcal{A}$ and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
output : Confidence set $\mathcal{C}$ for the changepoint $T$
1 Detect change at $\tau$ using $\mathcal{A}$
2 $\mathcal{C}=\varnothing$
3 for $t=1,\cdots,\tau$ do
4       Compute some test statistic $M_{t}$ (e.g., as defined in (2.9)) based on $X_{1},\cdots,X_{\tau}$
5       for $j=1,\cdots,B$ do
6             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$, or until time $L$, whichever happens first.
7             Subroutine $\mathcal{L}$: Compute $M_{t,L}^{j}$ (defined in (2.3),(2.4)) based on $\tau_{j}^{\prime}$, $L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$
8
9       end for
10      if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$ then
11            $\mathcal{C}=\mathcal{C}\cup\{t\}$
12       end if
13
14 end for
Algorithm 1 CI for known pre-change $F_{0}$ and post-change $F_{1}$
We now provide more details about Subroutine $\mathcal{L}$.
2.2 Subtleties regarding Subroutine $\mathcal{L}$
The problem with $L=\infty$ is that the stopping time may never be reached for the $j$-th sampled sequence, i.e., we may have $\tau^{\prime}_{j}=\infty$. One can safely set $L=\infty$ if the $\mathcal{A}$ (i.e., its stopping time) satisfies the condition $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$, meaning that if there is a change, $\mathcal{A}$ stops almost surely. This is a very mild assumption because any reasonable change detector should always eventually raise an alarm if there is a change. Indeed, all algorithms with finite ARL satisfy the condition, and indeed typically so do algorithms that control PFA. To avoid making even this mild assumption, one can simply pick $L<\infty$.
If $L<\infty$ and the $j$th sequence stopped because the algorithm $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual, but if $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it). Therefore, for $L<\infty$,
$M^{j}_{t,L}:=\begin{cases}-\infty,&\text{ if }t>\tau_{j}^{\prime},\\ M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),&\text{ if }t\leq\tau_{j}^{\prime}\leq L,\\ \infty,&\text{ if }t\leq\tau_{j}^{\prime}\text{ and }\tau_{j}^{\prime}>L.\\ \end{cases}$ (2.3)
For the proof of the upcoming theorem, we still need a definition of $M^{j}_{t,\infty}$ in the case that $L=\infty$ and $\tau_{j}=\infty$. We define
$M^{j}_{t,\infty}:=\begin{cases}M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),&\text{ if }t\leq\tau_{j}^{\prime}<\infty,\\ -\infty,&\text{ if }t>\tau_{j}^{\prime}\text{ or }\tau_{j}^{\prime}=\infty.\\ \end{cases}$ (2.4)
The rationale behind setting $M^{j}_{t,\infty}=-\infty$ when $\tau^{\prime}_{j}=\infty$ is the following. If for all $j=1,\cdots,B$, we have $\tau_{j}=\infty$, then we want to reject $H_{0,t}$, which is achieved by setting $M^{j}_{t,\infty}=-\infty$.
Remark 2.1.
The value of $L$ does not need to be a fixed constant; it can be set adaptively depending on the observed data sequence, for instance, one can choose $L$ to be $2\tau$. The proof of coverage works by arguing validity when $L=\infty$, and then pointing out that any smaller choice of $L$ only leads to more conservatism.
2.3 Prototypical choice of test statistic $M_{t}$
In the offline setting, where data is observed up to some fixed time $n$, the likelihood of data is
$L_{T}^{n}=\prod_{i=1}^{T-1}f_{0}(X_{i})\prod_{i=T}^{n}f_{1}(X_{i}),$ (2.5)
and the maximum likelihood estimator of $T$ (Hinkley,, 1970; Yao,, 1987) is given by
$\hat{T}_{n}=\operatorname*{arg\,max}_{1\leq j\leq n}L_{j}^{n}=\operatorname*{arg\,max}_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ (2.6)
In our sequential setting, where the data is observed up to a data-dependent stopping time $\tau$, we adapt this estimator by replacing $n$ with $\tau$ and define the estimator of the changepoint as:
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ (2.7)
The prototypical test statistic that we employ in our simulations is the likelihood-ratio-based test statistic $M_{t}^{(n)}$, where for all $n\in\mathbb{N},t\leq n$,
$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}^{n}}{L_{t}^{n}}.$ (2.8)
Note that for all $t\in\mathbb{N}$, the test statistic $M_{t}$ for testing $H_{0,t}$ simplifies to
$M_{t}=\begin{cases}L_{\hat{T}}^{\tau}/L_{t}^{\tau},&\text{ if }t\leq\tau<\infty\\ -\infty&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (2.9)
2.4 Unconditional coverage when $\mathcal{A}$ controls PFA
We now investigate the coverage properties and provide a bound on the coverage of the confidence set produced by Algorithm 1.
Theorem 2.2.
Let $\mathcal{C}$ be the confidence set for changepoint $T\in\mathbb{N}$ produced by Algorithm 1, for some $\alpha\in(0,1)$, $B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$ and change detection algorithm $\mathcal{A}$. Then,
$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<T).$
Therefore, if $\mathcal{A}$ controls PFA at level $\delta$,
$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\delta.$
Remark 2.3.
When $T\ll$ ARL, we typically have $\mathbb{\mathbb{P}}_{\infty}(\tau<T)\approx 0$ and so $\mathbb{P}_{T}(T\in\mathcal{C})\gtrapprox 1-\alpha$. The confidence set constructed using our method will approximately achieve the desired unconditional coverage level $1-\alpha,$ in situations where the probability of false detections before the true changepoint is negligible. However, $T$ is not known to us, and hence, we are unable to produce a suitable bound on the unconditional coverage for the detection algorithms that almost surely stop at a finite time, even when there is no change.
2.5 Unconditional coverage when $\mathcal{A}$ has finite ARL
If the detection algorithm has finite ARL (i.e., $\tau$ is finite almost surely), then it is impossible for any confidence set $\mathcal{C}^{*}$, which is a subset of $\{1,\cdots,\tau\}$, to produce an unconditional coverage guarantee at a predetermined level $1-\alpha$. The intuitive reason is that if $\tau$ is random and finite almost surely, it is possibly a false detection, i.e., $\tau<T$, in which case the set $\{1,\cdots,\tau\}$ itself does not cover $T$. The next proposition shows that.
Proposition 2.4.
If a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies an unconditional coverage guarantee $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$, for some $\alpha\in(0,1)$, then $\mathbb{P}_{\infty}(\tau=\infty)\geq 1-\alpha$ (thus, the PFA of the algorithm is at most $\alpha$ and thus the ARL of the algorithm is infinite). Said differently, for any algorithm with finite ARL, it is impossible to construct a valid confidence set that is only a subset of $\{1,\dots,\tau\}$.
The proposition says that if we can have a confidence set with high unconditional coverage, then the detection algorithm is such that it never stops with high probability if there is no change.
Therefore, in order to achieve an unconditional coverage guarantee with $\mathcal{A}$ having finite ARL, we must extend our confidence set beyond the stopping time $\tau$. So, we consider the union of the confidence set constructed using Algorithm 1 and all integers greater than $\tau$. The following proposition establishes that this augmented confidence set achieves the desired coverage guarantee of $1-\alpha$.
Proposition 2.5.
Let $\mathcal{C}$ be the confidence set for changepoint $T\in\mathbb{N}$ produced by Algorithm 1, for some $\alpha\in(0,1)$, $B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$ and change detection algorithm $\mathcal{A}$. Then, $\mathcal{C}^{*}:=\mathcal{C}\cup\{\tau+1,\tau+2,\cdots\}$ satisfies
$\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha.$ (2.10)
So, the confidence set constructed using Algorithm 1 can be interpreted as the pre-detection part of the augmented confidence set, which has the desired coverage guarantee of $1-\alpha$.
2.6 Conditional coverage when $\mathcal{A}$ has finite ARL
In the previous subsection, we established that the conventional notion of coverage is unsuitable for detection algorithms with finite ARL, where the confidence set is a subset of the random set $\{1,\cdots,\tau\}$. Intuitively, it is more appropriate to consider the conditional coverage, given that no false detection occurs, i.e., $T\leq\tau$. This adjustment is crucial because discussing coverage becomes meaningless when $\tau<T$, meaning that the detection occurs before the true changepoint — the target point we wish to cover by our confidence set.
The following is an immediate corollary of Theorem 2.2, where we show a bound on the conditional coverage of the confidence set produced by Algorithm 1.
Corollary 2.6.
For any $\alpha\in(0,1)$, $B\in\mathbb{N}$ and $T\in\mathbb{N}$, and change detection algorithm $\mathcal{A}$ satisfying $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ produced by Algorithm 1 satisfies
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)}.$ (2.11)
Therefore, if $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$,
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{1-\delta}.$ (2.12)
The restriction that $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$ is very weak and clearly necessary for conditional coverage. Since $T$ is arbitrary, the condition means that there is no $S<\infty$ satisfying $\mathbb{P}_{\infty}(\tau<S)=1$. If the algorithm always stops under $\mathbb{P}_{\infty}$ before $S$, then it can never detect any change that occurs after $S$. This is because the information that distinguishes $\mathbb{P}_{\infty}$ from $\mathbb{P}_{T}$ for $T>S$ occurs after $S$, but the algorithm has halted before $S$.
When $\tau$ is finite almost surely (i.e., PFA is $1$), we cannot have a positive lower bound on $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)$ uniformly over $T\in\mathbb{N}$. Therefore, Algorithm 1 does not guarantee conditional coverage at some prespecified level for detection algorithms with finite ARL (which implies that PFA is 1, although the converse is not true). Furthermore, if $\mathcal{A}$ is completely black-box — meaning its PFA is unknown — it becomes infeasible to construct a confidence set with the desired conditional or unconditional coverage using this approach. We propose a simple modification to address the issue in the following subsection, which applies to any black-box detection algorithm.
2.7 Modified confidence set when $\mathcal{A}$ has finite ARL (or $\mathcal{A}$ is a blackbox)
Equation 2.11 indicates that to ensure a conditional coverage guarantee of $1-\alpha$, we observe that we need to test $H_{0,T}$ at level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq T),$ rather than simply $\alpha$. To achieve that, for each $t\in\{1,\cdots,\tau\}$, we test $H_{0,t}$ at level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq t).$ For each $i\in\{1,\cdots,N\}$, we generate a sequence $\{Z_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$ until a change is detected, recording the detection times as $\tau_{i}$, which help to estimate the probability $\mathbb{P}_{\infty}(\tau\geq t)$, for $t=1,\cdots,\tau$. Therefore, it suffices to draw data points only up to $\min\{\tau,\tau_{j}\}$, for $j=1,\cdots,N.$
Then, we draw $B$ many independent streams of data under $H_{0,t}$. For the $j$th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at the stopping time $\tau_{j}^{\prime}$, using the same change detection method $\mathcal{A}$. Then, we compute the test statistic $M_{t,L}^{j}$ as described in Section 2.2, based on $\tau_{j}^{\prime}\wedge L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$. Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$, i.e.,
$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (2.13)
where $\hat{\alpha}_{t}=\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$. Algorithm 2 contains a structured overview of the method.
Input :  $\alpha\in(0,1)$, $N\in\mathbb{N},B\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
output : A confidence set $\mathcal{C}$ for the changepoint $T$
1 $\mathcal{C}=\varnothing$
2 for $i=1,\cdots,N$ do
3       Simulate $\{Z_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until $\mathcal{A}$ detects a change at $\tau_{j}$, or until $\tau$, whichever happens first.
4
5 end for
6for $t=1,\cdots,\tau$ do
7       Compute some test statistic $M_{t}$ (e.g., as defined in (2.9)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
8       for $j=1,\cdots,B$ do
9             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{1}$, until $\mathcal{A}$ detects a change at $\tau_{j}^{\prime}$, or time $L$, whichever happens first.
10             Subroutine $\mathcal{L}$: Compute $M_{t,L}^{j}$ (defined in (2.3),(2.4)) based on $\tau_{j}^{\prime}$, $L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$
11
12       end for
13      Compute $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$
14       if $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$ then
15            $\mathcal{C}=\mathcal{C}\cup\{t\}$
16       end if
17
18 end for
Algorithm 2 CI for known pre-change $F_{0}$ and post-change $F_{1}$
By suitably adjusting the test levels for each $t$, this approach ensures that the desired conditional coverage guarantee is achieved. Notably, this guarantee does not depend on any specific property of $\mathcal{A}$ (e.g., its PFA or ARL). We prove it in the following theorem.
Theorem 2.7.
For any fixed $\alpha\in(0,1)$, $B,N\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, $T\in\mathbb{N}$ and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (2.13) satisfies
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-{\alpha}.$ (2.14)
Proof.
Step 1 [Establish the relationship between the level of testing ${H}_{0,T}$ and the conditional coverage]: For each $t$, define $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$. Then, the level of our test for ${H}_{0,T}$ is $P_{T}(M_{T}>C_{t,L}).$ Now,
$\mathbb{P}_{T}(M_{T}\leq C_{t,L})$ $=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\})$
$=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau\geq T)+\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau<T)$
$\leq\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\mid\tau\geq T)\mathbb{P}_{T}(\tau\geq T)+\mathbb{\mathbb{P}}_{\infty}(\tau<T)$
$=\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{\infty}(\tau\geq T).$
The inequality in the third line holds because
$\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t}^{N}\},\tau<T)\leq\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$
where the last equality holds since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$ and by the same argument, $\mathbb{P}_{t}(\tau\geq t)=\mathbb{\mathbb{P}}_{\infty}(\tau\geq t)$ for all $t$. So, we obtain
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{t,L})}{P_{\infty}(\tau\geq T)}.$ (2.15)
Step 2 [Establish that our test for each ${H}_{0,t}$ has level $\alpha\times\mathbb{P}_{\infty}(\tau\geq t)$]: First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d. under $H_{0,t}$. Hence, $M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})\iff M_{t}$ is one among the $\lceil(1-c)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$. So, for any fixed $c\in(0,1)$,
$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c\text{ for all }t.$ (2.16)
Then, it follows that
$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t})\geq\frac{\lceil(1-\hat{\alpha}_{t})(B+1)\rceil}{B+1}\geq 1-\hat{\alpha}_{t},\text{ almost surelyfor all }t.$ (2.17)
Therefore, we have
$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})=\mathbb{E}_{t}(\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t}))\geq\mathbb{E}_{t}(1-\hat{\alpha}_{t})=1-\alpha\mathbb{P}_{t}(\tau\geq t)=1-\alpha\mathbb{P}_{\infty}(\tau\geq t).$ (2.18)
Also, for all $L\in\mathbb{N}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$ and hence, $C_{t,L}\geq C_{t,\infty}$.
Thus, we obtain $\mathbb{P}_{t}(M_{t}>C_{t,L})\leq\mathbb{P}_{t}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{\infty}(\tau\geq t)$, for all $t$.
Therefore, from (2.15), we have $\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
Remark 2.8.
As mentioned before, the assumption $\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ is essential; otherwise, conditioning the coverage probability on $\{\tau\geq T\}$ becomes undefined. If this assumption does not hold, we have $\mathbb{P}_{\theta_{0},\infty}(\tau<T)=\mathbb{P}_{T}(\tau<T)=1$, implying that the detection is almost surely false. In such a scenario, constructing a confidence interval would not be meaningful.
Remark 2.9.
We do not require any specific property of the sequential change detection algorithm; it can function as an entirely black-box or heuristic method. The only necessity is a stopping rule that can be applied to generated data sequences.
Remark 2.10.
For a fixed $t$, the sampled sequences $\{Z^{j}_{n}\}_{n}$ must be independent across all $j$ to ensure the validity of the above theorem. However, for testing the $t$-th hypothesis ($t\geq 3$), one can reuse the first $t-2$ data points from the sequences generated for testing the $(t-1)$-th hypothesis, thereby reducing computational redundancy. Nevertheless, in our experiments, we sampled the sequences independently across $t$.
3 Known pre-change but unknown post-change distribution
In this section, we assume that the post-change distribution follows some parametric composite model $\mathcal{P}_{1}=\{F_{\theta}:{\theta}\in\Theta_{1}\}$ and the pre-change distribution, $F_{\theta_{0}}(\theta_{0}\notin\Theta_{1})$ is known.
Notation.
For $t\in\mathbb{N}$, let $\mathbb{P}_{\theta,t,\theta^{\prime}}$ denote the joint distribution of the sequence of independent random variables $\{X_{n}\}_{n\in\mathbb{N}}$, when $X_{1},\dots,X_{t-1}$ are from $F_{\theta}$ (with density $f_{\theta}$) and $X_{t},X_{t+1},\dots$ are from $F_{\theta^{\prime}}$ (with density $f_{\theta^{\prime}}$). As corner cases, let $\mathbb{P}_{1,\theta^{\prime}}$ denotes the joint distribution where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta^{\prime}}$ and $\mathbb{P}_{\theta,\infty}$ denotes the case where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta}$. Let $\mathbb{E}_{\theta,t,\theta^{\prime}}$ denote the expectation under $\mathbb{P}_{\theta,t,\theta^{\prime}}$, for $t\in\mathbb{N}\cup\{\infty\}$, for $t\in\mathbb{N}\cup\{\infty\}$ and $\mathbb{E}_{\theta,\infty}$ denote the expectation under $\mathbb{P}_{\theta,\infty}$.
Suppose we have the data sequence $\{X_{n}\}_{n}\sim\mathbb{P}_{\theta_{0},T,\theta_{1}}$, for some unknown $T\in\mathbb{N}\cup\{\infty\},\theta_{1}\in\Theta_{1}$ and some sequential change detection algorithm $\mathcal{A}$ that raises an alarm at the stopping time $\tau$ in the sequence. We define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{{\theta_{0}}},\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta_{1}},\text{ for some }{\theta_{1}}\in\Theta_{1}.$
For testing the null hypothesis $H_{0,t}$, we consider some test statistic $M_{t}:=M_{t}^{(\tau)}(X_{1},\cdots,X_{\tau})$ based on data up to time $\tau,$ if $t\leq\tau<\infty$ and $M_{t}:=-\infty$, otherwise, for each $t\in\mathbb{N}$, where $M_{t}^{(n)}$ is some test statistic based on $n$ many data points, for $n,t\in\mathbb{N},t\leq n$. An example is provided in Section 3.1.
Suppose the change detection algorithm, $\mathcal{A}$ has finite ARL or $\mathcal{A}$ is possibly a blackbox (i.e., its properties or inner workings are unknown to us). In that case, we generate a sequence $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$ until a change is detected, recording the detection times as $\tau_{i}$. These will help empirically estimate the probability $\mathbb{P}_{\infty}(\tau\geq t)$, for $t=1,\cdots,\tau$. Therefore, it suffices to draw data points only up to $\min\{\tau,\tau_{j}\}$, for $j=1,\cdots,N$. However, if $\mathcal{A}$ controls PFA at a known level $\delta\in(0,1)$, we can avoid this step.
Since the post-change parameter is unknown, for each $t$, we estimate it based on $X_{t},\cdots,X_{\tau}$. However, note that the sample size used for estimate might be too small to have a reliable estimate. To address this issue, we would use interval estimates instead of point estimates so that we can hope to achieve a predetermined control over the error. Notably, conventional confidence intervals are invalid in this context, since $\tau$ depends on the data. Therefore, we rely on confidence sequences (Darling and Robbins,, 1967; Jennison and Turnbull,, 1989; Howard et al.,, 2021). A confidence sequence having coverage $1-c$ is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform coverage guarantee: $\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage. Otherwise, we construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage. Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS has coverage probability $\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; we formalize this more explicitly in the proof of the Theorem 3.1.
For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ having joint distribution as $\mathbb{P}_{\theta_{0},t,\theta}$. Note that for a fixed $t$ and $\theta$, the sequences must be independent across $j$. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta_{0},\theta)\}_{n}$ are allowed to be dependent across different values of $t$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section 3.2. We draw the datapoints until a change is detected at the stopping time $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau_{j,t}^{\theta}$ and $X^{j}_{1}(\theta_{0},t,\theta),\cdots,X^{j}_{\tau_{j,t}^{\theta}\wedge L}(\theta_{0},t,\theta)$, which is denoted as $M^{j}_{t,L}(\theta)$. For $L<\infty$,
$M^{j}_{t,L}(\theta):=\begin{cases}-\infty,&\text{ if }t>\tau_{j,t}^{\theta},\\ M_{t}^{(\tau_{j,t}^{\theta})}(X_{1}^{j}(\theta_{0},t,\theta),\cdots,X_{\tau_{j,t}^{\theta}}^{j}(\theta_{0},t,\theta)),&\text{ if }t\leq\tau_{j,t}^{\theta}\leq L\\ \infty,&\text{ if }t\leq\tau_{j,t}^{\theta}\text{ and }\tau_{j,t}^{\theta}>L.\end{cases}$ (3.1)
[…]
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\},$ (3.2)
where $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$. If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$, we can have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
$\mathcal{C}_{2}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha;M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\},$ (3.3)
Algorithm 3 contains an overview of the method.
Note that, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set, it can be computationally challenging to approximate the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$. However, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient. We provide details in Section 3.2 and a concrete Gaussian mean shift example in Section 3.3.
Input :  $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N},L\in\mathbb{N}\cup\{\infty\},$ a confidence sequence procedure $\operatorname{CS(data;coverage)}$, a change detection algorithm $\mathcal{A}$, and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
output : Confidence sets $\mathcal{C}$ for changepoint $T$
1 $r_{1}=r_{2}=\cdots=r_{\tau}=1$
2 if $\mathcal{A}$ has finite ARL then
3       for $i=1,\cdots,N$ do
4             Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first.
5
6       end for
7      $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, for $t=1,\cdots,\tau$
8 end if
[…]
10 for $t=1,\cdots,\tau$ do
11       Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$ (e.g., as defined in (3.7))
12       Find a confidence sequence for $\theta_{1}$: $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$
13       for $j=1,\cdots,B$ do
14             Subroutine 2: for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}(X_{t},\cdots,X_{\tau})$ do
15                   Get the sequence $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ following $\mathbb{P}_{\theta_{0},t,\theta}$ (independently for all $j$), until its detection time $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until time $L$, whichever happens first.
16                   Compute $M_{t,L}^{j}(\theta)$ based on $L,\tau_{j,t}^{\theta}$ and $X^{j}_{1}(\theta_{0},t,\theta),\cdots,X^{j}_{\tau_{j,t}^{\theta}\wedge L}(\theta_{0},t,\theta)$
17
18             end for
19
20       end for
21      Subroutine 3: if there exists $\text{there exists }\theta\in\mathcal{S}^{\prime}_{\tau-t+1}:M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ then
22            $\mathcal{C}=\mathcal{C}\cup\{t\}$
23       end if
24
25 end for
26
Algorithm 3 CI for known pre-change ($F_{\theta_{0}}$) and composite post-change
Theorem 3.1.
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be the confidence sets for changepoint $T\in\mathbb{N}$ as defined in (3.2) and (3.3) respectively, for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$. Then for any given change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$,
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$
If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, then for any $\theta_{1}\in\Theta_{1}$,
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\delta\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{1-\delta},\text{ when }\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0.$
3.1 Prototypical choice of test statistic $M_{t}$
Let $L_{t}(\theta_{1};X_{1},\cdots,X_{n})$ be the likelihood of data $X_{1},\cdots,X_{n}$ under $\mathcal{H}_{0,t}$. Then, for $n\geq t$,
$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (3.4)
We define $\hat{\theta}_{1,j:n}$ to be a maximum likelihood estimator (MLE) of the post-change parameter $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$. The point estimate of $T$ can be naturally extended to the following form:
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (3.5)
The prototypical test statistic that we employ in our simulations is the likelihood-ratio-based test statistic $M_{t}^{(n)}$, where for $t,n\in\mathbb{N},t\leq n$,
$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (3.6)
Note that the test statistic $M_{t}$ for testing $\mathcal{H}_{0,t}$ reduces to
$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{ if }t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (3.7)
3.2 Computationally and statistically efficient subroutines
For computational and statistical efficiency, we generate sequences from all $t\in\{1,\cdots,\tau\}$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ simultaneously using a single shared source of randomness. To achieve this, we assume that we have a fixed element $\theta^{\prime\prime}\in\Theta$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta$. Therefore, for each $j=1,\cdots,B$, we generate only one data sequence $\{X_{n}^{j}\}_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{\theta^{\prime\prime}}$, and then for each $t\in\{1,\cdots,\tau\}$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ (having joint distribution $\mathbb{P}_{\theta_{0},t,\theta}$) by applying the function $G(.,\theta_{0})$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta)$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
$Y_{n}^{t}(x,\theta):=\begin{cases}G(x,\theta_{0}),&\text{ if }n<t\\ G(x,\theta),&\text{ otherwise. }\end{cases}$ (3.8)
Let us provide a few examples of the function $G$. If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}}(x))$ works. In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$.
Let us now fix $t\in\{1,\cdots,\tau\}$. Computing the detection time $\tau_{j,t}^{\theta}$ for each $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether there exists $\text{there exists }\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set. However, it is often straightforward to compute $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$. Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient. For example, we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$. Again, we bound $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)$ by $V_{t,L}^{j}$, where for $L<\infty$,
$V_{t,L}^{j}:=\begin{cases}-\infty&\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{ if }t\leq t_{2}^{j}\leq L\\ \infty&\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}$ (3.9)
And for $L=\infty$,
$V_{t,\infty}^{j}:=\begin{cases}\displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{ if }t\leq t_{2}^{j}\text{ and }t_{1}^{j}<\infty\\ -\infty&\text{ if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \end{cases}$ (3.10)
which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, as we explain in Section 3.3). We define the confidence interval as
$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (3.11)
$\mathcal{C}^{\prime}_{2}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}.$ (3.12)
We observe that $\mathcal{C}^{\prime}\supseteq\mathcal{C}$, where $\mathcal{C}$ is the set produced by Algorithm 3, since
$\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t);M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$ $\geq\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t);M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
$\geq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t);M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}).$
Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 3.1.
Corollary 3.2.
$\mathcal{C}^{\prime}_{1}$ and $\mathcal{C}^{\prime}_{2}$ defined in (3.11) and (3.12) respectively, satisfy the same coverage guarantees as $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ in Theorem 3.1.
3.3 A concrete example: Gaussian mean shift problem
We consider the most common changepoint model, where the pre-change distribution is $F_{0}$ is $N(\theta_{0},1)$ and the post-change model is $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
In this setting, we have the following example of a $1-\beta$ confidence sequence (Howard et al.,, 2021) for $\theta_{1}$
$\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (3.13)
where $\bar{X}_{t:\tau}=\frac{1}{\tau-t+1}\sum_{i=t}^{\tau}X_{i}$ and $s_{n}(\beta)=\sqrt{\log\log(2n)+0.72\log(10.4/\beta)}$.
We fix $t\in\{1,\cdots,\tau\}$. Note that we can consider $\theta^{\prime\prime}=\theta_{0}$ and $G(x,\theta)=x+\theta-\theta_{0}$. Therefore, only need to generate the sequence $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(\theta_{0},1)$ for $j=1,\cdots,B$. For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (3.14)
where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (2.8), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let, $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$, where $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark 3.3) and they are finite. So, we can have $L=\infty$.
If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,V_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
$V_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}{L_{t}(\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j}+\theta;Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}$
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta_{t^{\prime}}^{i});Y_{1}^{t}(X_{1}^{j},\theta_{t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta_{t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$ (3.15)
where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta)=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta),$ and $\theta_{t^{\prime}}^{i}$ is the unique element that maximizes $L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)$, or equivalently minimizes
$-2\log L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)=$
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\text{constant (wrt }\theta),\text{ if }i<t,\\ \text{constant (wrt }\theta),\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{constant (wrt }\theta),\text{ if }i>t,\end{cases}$
over $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. Note that $\theta_{t^{\prime}}^{i}$ has a simple closed-form expression since the above equation is a quadratic function of $\theta$. And hence, $V_{t}^{j}$ (3.3) can be computed easily.
Remark 3.3.
[…]
$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (3.16)
[…]
For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{\tau}(2X_{i}-\theta_{0}-\theta)\}$ increses as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution. One can easily check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section 5.2.
The Gaussian mean shift example we provided serves as a concrete and illustrative case to clarify the methodology. It is worth noting that $C^{\prime}$, as defined in (3.11), is practically implementable as long as the following relatively weak conditions are met:
  1. 1.
    The coupling function $G$ is available (which is always the case when distributions are continuous).
  2. 2.
    We can construct a confidence sequence for the parameter of interest (which is always achievable as long as one can construct a confidence interval for the parameter).
  3. 3.
    The quantities $t_{j}^{1}$ and $t_{j}^{2}$ are computable (either exactly or numerically).
  4. 4.
    If we employ the likelihood-rato-based test statistic (2.8), then we have to be able to maximize the function $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
For instance, one can easily verify that this approach applies to Gaussian, Laplace, or exponential scale change problems.
4 Composite pre- and post-change
We now assume that both pre- and post-change distributions follow some parametric composite models $\mathcal{P}_{0}=\{F_{\theta}:\theta\in\Theta_{0}\}$ and $\mathcal{P}_{1}=\{F_{\theta}:\theta\in\Theta_{1}\}$ respectively such that $\Theta_{0}\cap\Theta_{1}=\emptyset$.
In this setting, the average run length (ARL) of the detection algorithm is defined as $\inf_{\theta\in\Theta_{0}}\mathbb{E}_{\theta,\infty}[\tau]$, which represents the least expected time until a false alarm occurs under $T=\infty$ (i.e., no change). On the other hand, the probability of false alarm (PFA) is defined as the worst-case probability of triggering a false alarm, that is $\sup_{\theta\in\Theta_{0}}\mathbb{P}_{\theta,\infty}(\tau<\infty)$.
Suppose we have the data sequence $\{X_{n}\}_{n}\sim\mathbb{P}_{\theta_{0},T,\theta_{1}}$, for some unknown $T\in\mathbb{N}\cup\{\infty\},\theta_{i}\in\Theta_{i}(i=0,1)$ and some sequential change detection algorithm $\mathcal{A}$ that raises an alarm at the stopping time $\tau$ in the sequence. Following the approach from the previous sections, we define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta_{0}},\text{ for some }\theta_{0}\in\Theta_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta_{1}},\text{ for some }{\theta_{1}}\in\Theta_{1}.$
We consider some test statistic $M_{t}:=M_{t}(X_{1},\cdots,X_{\tau};\tau)$ based on data up to time $\tau,$ for testing the null hypothesis $\mathcal{H}_{0,t}$. An example is provided in Section 4.1.
Suppose the change detection algorithm, $\mathcal{A}$ has finite ARL or it is a blackbox algorithm so that its PFA is unknown. In that case, we have a problem in generating i.i.d. copies of the detection times under $T=\infty$ (i.e., no change), since the pre-change parameter $\theta_{0}$ is unknown. Therefore, we need to assume the following:
Assumption 4.1.
There exists $\theta_{0}^{*}\in\Theta$ such that $\inf_{\theta\in\Theta_{0}}\mathbb{P}_{\theta,\infty}(\tau\geq t)\geq\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$, for all $t\in\mathbb{N}$.
This assumption holds for a variety of change detection problems. For instance, we illustrate this in Remark 4.6 for some Gaussian mean change problems with CUSUM and SR-type detectors. One can easily check that it holds for analogous Gaussian, Laplace, and exponential scale change problems as well.
We now generate $N$ many i.i.d. sequences under $F_{\theta_{0}^{*}}$ until a change is detected using $\mathcal{A}$, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$, that will help us estimate $\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$, which is a lower bound on $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$. If $\mathcal{A}$ controls PFA at a known level $\delta\in(0,1)$, we can avoid this step and obtain a conditional/unconditional coverage guarantee without any such assumption.
Since both pre- and post-change parameters are unknown, for each $t$, we estimate them based on $X_{1},\cdots,X_{t-1}$ and $X_{t},\cdots,X_{\tau}$ respectively. However, as noted earlier, the sample size used for the estimates might be too small to yield reliable estimates. Therefore, instead of point estimates, we rely on interval estimates in order to achieve a predetermined control over the error.
Suppose that we have some confidence interval procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$, which constructs $1-c$ confidence interval using $X_{1},\cdots,X_{n}$ for $\theta_{0}$ and a confidence sequence procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, which constructs $1-c$ confidence sequence using $\{X_{n}\}_{n}$ for $\theta_{1}$. If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage and $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma)$ for $\theta_{0}$ having $1-\gamma$ coverage. Otherwise, we construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$. Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS and CI have coverage probabilities $1-\beta\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ and $1-\gamma\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ respectively; we formalize this more explicitly in the proof of Theorem 4.2.
Note that for $t=1$, the only plausible confidence interval of $\theta_{0}$ is $\mathcal{S}_{0}=\Theta_{0}$. Therefore, if $\Theta_{0}$ is unbounded, then it becomes infeasible. To avoid this issue, one can assume that $T\in\mathbb{N}\setminus\{1\}$ and test only for $t=2,\cdots,\tau.$
For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ having joint distribution as $\mathbb{P}_{\theta,t,\theta^{\prime}}$. Note that the sequences must be independent as $j$ varies. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ are allowed to be dependent across different values of $t$ and $\theta\in\mathcal{S}_{t-1},\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section 4.2. We draw the datapoints until a change is detected at the stopping time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$, which is denoted as $M^{j}_{t,L}(\theta,\theta^{\prime})$. For $L<\infty$,
$M^{j}_{t,L}(\theta,\theta^{\prime}):=\begin{cases}-\infty,&\text{ if }t>\tau^{j,t}_{\theta,\theta^{\prime}},\\ M_{t}(X_{1}^{j}(\theta,t,\theta^{\prime}),\cdots,X_{\tau^{j,t}_{\theta,\theta^{\prime}}}^{j}(\theta,t,\theta^{\prime})),&\text{ if }t\leq\tau^{j,t}_{\theta,\theta^{\prime}}\leq L,\\ \infty,&\text{ if }t\leq\tau^{j,t}_{\theta,\theta^{\prime}}\text{ and }\tau^{j,t}_{\theta,\theta^{\prime}}>L.\\ \end{cases}$ (4.1)
And for $L=\infty$,
$M^{j}_{t,\infty}(\theta,\theta^{\prime}):=\begin{cases}M_{t}(X_{1}^{j}(\theta,t,\theta^{\prime}),\cdots,X_{\tau^{j,t}_{\theta,\theta^{\prime}}}^{j}(\theta,t,\theta^{\prime})),&\text{ if }t\leq\tau^{j,t}_{\theta,\theta^{\prime}}<\infty,\\ -\infty,&\text{ if }\tau^{j,t}_{\theta,\theta^{\prime}}=\infty\text{ or }t>\tau^{j,t}_{\theta,\theta^{\prime}}.\\ \end{cases}$ (4.2)
Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$ at level $\alpha\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$,
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\},$ (4.3)
where $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$. If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$ , we have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
$\mathcal{C}_{2}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}.$ (4.4)
Algorithm 4 contains an overview of the method.
Note that, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are finite sets, it can be computationally challenging to approximate the threshold $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. However, in many cases, we can derive a suitable upper bound of this threshold, which is computationally and statistically efficient. We provide the details in Section 4.2. We offer a concrete example of the Gaussian mean shift problem in Section 4.3.
Input :  $\theta_{0}^{*},\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N},L\in\mathbb{N}\cup\{\infty\}$, change detection algorithm $\mathcal{A}$, a confidence interval procedure $\operatorname{CI(data;coverage)},$ a confidence sequence procedure $\operatorname{CS(data;coverage)}$, and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
output :  A confidence set $\mathcal{C}$ for changepoint $T$
1 $r_{1}=r_{2}=\cdots=r_{\tau}=1$
2 if $\mathcal{A}$ has finite ARL then
3       for $i=1,\cdots,N$ do
4             Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}^{*}}$, define $\tau_{i}$ as the time when $\mathcal{A}$ detects a change, or $\tau$, whichever occurs first
5
6       end for
7      Define $r_{t}=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$ for $t=1,\dots,\tau$;
8 end if
[…]
10
11for $t=1,\cdots,\tau$ do
12       Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$ (e.g., as defined in (4.8))
13       Find a confidence interval for $\theta_{0}$: $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})$
14       Find a confidence sequence or $\theta_{1}$: $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$
15       for $j=1,\cdots,B$ do
16             Subroutine 1: for $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ do
17                   Simulate a sequence $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ following $\mathbb{P}_{\theta,t,\theta^{\prime}}$ (independently for all $j$) until its detection time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever happens first.
18                   Compute $M_{t,L}^{j}(\theta,\theta^{\prime})$ based on $L,\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$
19
20             end for
21
22       end for
23      Subroutine 2: if $M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ then
24            $\mathcal{C}=\mathcal{C}\cup\{t\}$
25       end if
26
27 end for
Algorithm 4 CI for composite pre-change and post-change
Theorem 4.2.
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.3) and (4.4), for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$. For any $\theta_{i}\in\Theta_{i}(i=0,1$), and any change detection algorithm $\mathcal{A}$ satisfying Assumption 4.1 and having $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$, the following conditional coverage guarantee holds:
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, $\mathcal{C}_{2}$ satisfies the following guarantee for any $\theta_{i}\in\Theta_{i}(i=0,1)$:
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\gamma-\delta,\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta+\gamma}{1-\delta},\text{ when }\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0.$
Remark 4.3.
As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets for $i=0,1$, we can have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$. So, in this case, Algorithms 4 and 3 reduces to our algorithm designed for the known pre- and post-change setting and similarly, Theorems 3.1 and 4.2 reduces to Theorem 2.7.
4.1 Prototypical choice of test statistic $M_{t}$
Let $L_{t}^{n}(\theta_{0},\theta_{1})$ be the likelihood of data up to time $n\geq t$ under $\mathcal{H}_{0,t}$. Then,
$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (4.5)
We define $\hat{\theta}_{0,1:j}$ to be the maximum likelihood estimator (MLE) of the null $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$. Similarly, we define $\hat{\theta}_{1,j:n}$ to be the MLE of the alternative $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$. The point estimate of $T$ can be naturally extended to the following form:
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (4.6)
As a prototypical example, we consider the likelihood-ratio-based test statistic $M_{t}^{(n)}$, which is defined for $t,n\in\mathbb{N},t\leq n$ as below.
$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (4.7)
Note that the test statistic $M_{t}$ for testing the null hypothesis $\mathcal{H}_{0,t}$ reduces to
$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{0,1:\hat{T}-1},\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{ if }t\leq\tau<\infty,\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\\ \end{cases}$ (4.8)
4.2 Computationally and statistically efficient subroutines
As in Section 3.2, we assume that we have a fixed element $\theta^{\prime\prime}$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta_{0}\cup\Theta_{1}$. Therefore, for each $j=1,\cdots,B$, we generate data sequence $X_{1}^{j},X_{2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta^{\prime\prime}}$ only once. Then, for each $t\in\{1,\cdots,\tau\}$ and $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta,\theta^{\prime})\}_{n}$ having joint distribution $\mathbb{P}_{\theta,t,\theta^{\prime}},$ by applying the function $G(.,\theta)$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta^{\prime})$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
$Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{ if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (4.9)
Let us now fix $t\in\{1,\cdots,\tau\}$. Computing the detection time $\tau^{j}_{\theta,\theta^{\prime}}$ for each $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether there exists ($\text{there exists }(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are a finite sets.
However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$ and $t_{2}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$. Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient. E.g., we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where for $L<\infty$,
$U_{t,L}^{j}:=\begin{cases}-\infty&\text{if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),&\text{if}t\leq t_{2}^{j}\leq L\\ \infty&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (4.10)
And for $L=\infty$,
$U_{t,\infty}^{j}:=\begin{cases}\displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),&\text{if }t\leq t_{2}^{j}\text{ and }t_{1}^{j}<\infty\\ -\infty&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty.\end{cases}$ (4.11)
We consider the following confidence sets:
$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (4.12)
$\mathcal{C}^{\prime\prime}_{2}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}.$ (4.13)
Observe that $\mathcal{C}^{\prime\prime}_{i}\supseteq\mathcal{C}_{i}$, for $i=1,2$, where $\mathcal{C}_{i}$’s are the sets produced by Algorithm 4, since
$\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{U_{t,L}^{j}\}_{j=1}^{B})\}$ $\geq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$
$\geq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B}).$
Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.2.
Corollary 4.4.
$\mathcal{C}^{\prime\prime}_{1}$ and $\mathcal{C}^{\prime\prime}_{2}$ defined in (4.12) and (4.13) respectively satisfy the same coverage guarantees as $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ in Theorem 4.2.
4.3 A concrete example: Gaussian mean shift problem
We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$.
In this setting, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$
$\mathcal{S}_{t-1}=\left[\bar{X}_{1:t-1}-\frac{q_{\gamma/2}}{\sqrt{t-1}},\bar{X}_{1:t-1}+\frac{q_{\gamma/2}}{\sqrt{{t-1}}}\right]\cap\Theta_{0},$ (4.14)
and the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (3.13).
We fix $t\in\{2,\cdots,\tau\}$. Note that we can consider $\theta^{\prime\prime}=0$ and $G(x,\theta)=x+\theta$. Therefore, only need to generate the sequence $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$. For $\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}$, we define
$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (4.15)
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (4.8), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms as we illustrate in Remark 4.5.
For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
$U_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j}+\theta_{0},\bar{\epsilon}_{t:t^{\prime}}^{j}+\theta_{1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}$
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\bar{Y}_{i:t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i});Y_{1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime})=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime}),$ and $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ is the one such pair that maximizes $L_{i}\left(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)$, or equivalently minimizes
$-2\log L_{i}\left(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)=$
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{ constant },\text{ if }i<t,\\ \text{ constant},\text{ if }i=t,\\ \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{ constant },\text{ if }i>t,\end{cases}$ (4.16)
over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (4.3) is a quadratic function of $\theta_{1}-\theta_{0}$. Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
Remark 4.5.
[…]
$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\}$ (4.17)
[…]
where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For example, consider $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\geq b\}$, for some $b\geq a$. Observe that for all $\theta\in\Theta_{1},f_{\theta}^{*}=f_{a}$, and $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}=\exp\{\frac{1}{2}(\theta-a)\times\sum_{i=j}^{\tau}(2X_{i}-a-\theta)\}$ increses as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore,
$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
Similarly, it is easy to check that for $\mathcal{P}_{0}=\{N(\theta,1):\theta\geq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\leq b\}$, for some $b\leq a$,
$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
Remark 4.6.
For both types of pre- and post-change classes discussed above, and for CUSUM and SR-type detectors defined in (4.17) and (4.18), as well as the PFA-controlling change detectors used in Section 5.2, assumption 4.1 holds with $\theta_{0}^{*}=a$. We provide the explanation for the CUSUM detector, and the argument for the SR detector is analogous. For example, consider $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\geq b\}$, for some $b\geq a$. Then,
$\mathbb{P}_{\theta_{0},\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}_{\theta_{0},\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)<A\right]$
$=\mathbb{P}_{\theta_{0},\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta}(X_{i})}{f_{a}(X_{i})}dW(\theta)<A\right]$
[…]
Then, $\exp\{\frac{1}{2}(\theta-a)\times\sum_{i=j}^{\tau}(2Z_{i}+2\theta_{0}-a-\theta)\}\leq\exp\{\frac{1}{2}(\theta-a)\times\sum_{i=j}^{\tau}(2Z_{i}+a-\theta)\},\forall\theta\in\Theta_{1},\forall\theta_{0}\in\Theta_{0}.$ Therefore, $\mathbb{P}_{\theta_{0},\infty}[\tau_{\text{wcs-ripr}}\geq t]\geq\mathbb{P}_{a,\infty}[\tau_{\text{wcs-ripr}}\geq t],\forall\theta_{0}\in\Theta_{0}.$
The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology. It is worth noting that $\mathcal{C}^{\prime\prime}$ in (4.12) is computable for parametric families if the following conditions are met:
  1. 1.
    The coupling function $G$ is available (which is always the case when distributions are continuous).
  2. 2.
    We can construct a confidence interval (and hence confidence sequence) for the parameter of interest.
  3. 3.
    The quantities $t_{j}^{1}$ and $t_{j}^{2}$ are computable (either exactly or numerically).
  4. 4.
    If we employ the likelihood-ratio-based test statistic (4.8), then we have to be able to maximize the function $L_{i}(\hat{\theta}_{1:i-1},\hat{\theta}_{i:t^{\prime}};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))$ over $(\theta,\theta^{\prime})$ on a set (either exactly or numerically).
  5. 5.
    If the detection algorithm $\mathcal{A}$ controls ARL (or is a black box), we need to satisfy Assumption 4.1. If $\mathcal{A}$ controls PFA, we do not need this assumption.
For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions.
5 Experiments
We now experimentally validate our algorithms for constructing confidence sets in order to evaluate their performance across various settings and demonstrate their practical utility.
5.1 Experiments with detection algorithms that control ARL
5.1.1 Known pre- and post-change (Setting I)
We consider both Normal and Poisson mean-change scenarios having the true changepoint at $T=100,500$, using the CUSUM change detection method defined in (B.1) with thresholds at $A=1000,10000$ (which controls ARL at level $A$ (Tartakovsky et al.,, 2014)). After detection, we employ Algorithm 2 to construct confidence sets for $T$. Figures 5.1 and 5.2 provide visualizations of the confidence sets and point estimates (2.7) of the changepoints across $5$ runs.
Figure 5.1: Setting I (known pre- and post-change Gaussian distributions): The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$ for $T=100$ (top row) and $T=500$ (bottom row). Change is detected using the CUSUM detector (B.1) with $A=10000$ (top row) or $A=1000$ (bottom row). The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
Figure 5.2: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois$(1)$ and the rest from Pois$(2)$, with $T=100$ (top row) or $T=500$ (bottom row). Change is detected using CUSUM detector (B.1) with $A=10000$ (top row) or $A=1000$ (bottom row). The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.3.
Table 5.3: Setting I: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (B.1) with threshold $A=1000$, while varying $T$.
$T$ $F_{0}$ $F_{1}$
Conditional
coverage
Unconditional
coverage
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 $N(0,1)$ $N(1,1)$ 0.96 0.94 12.34 2.85 13.97
100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.95 17.45 3.98 15.72
500 $N(0,1)$ $N(1,1)$ 0.95 0.86 12.57 2.62 13.22
500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.83 18.72 3.96 16.14
For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the empirical conditional coverage rates meet the desired level, supporting our theory. We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
5.1.2 Known pre-change and composite post-change (Setting II)
We perform experiments for Gaussian mean-change scenarios having pre-change distribution as $N(0,1)$, and the true (unknown) post-change data distribution as $N(1,1)$, with the changepoint at $T=100,500.$ For detection, we employ a weighted CUSUM detector (3.16) with thresholds at $A=1000,10000$ (which controls ARL at $A$). For weighted CUSUM, we employ a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.75\}$ and $\theta=0.9$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we construct confidence sets (as defined in (3.11)) for $T$. Fig. 5.4 provides visualizations of the confidence sets (3.11) and point estimates (3.5) of the changepoint across $5$ random runs.
Figure 5.4: Setting II: The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$ with $T=100$ (top row) or $T=500$ (bottom row). Change is detected using a weighted CUSUM detector (3.16) with $A=10000$ (top row) or $A=1000$ (bottom row). The point estimate (3.5) is shown in a vertical red dashed line and the confidence set (3.11) is shown in red points. $F_{0}=N(0,1)$, $\mathcal{P}_{1}=\{N(\mu,1):\mu\geq 0.9\}$, $N=B=100$, $\alpha=0.05,\beta=0.025$ and $L=\infty$. Results of $5$ independent simulations are shown.
We report the average size and the conditional and unconditional coverage across $100$ independent runs in (5.5). For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.3). We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table. We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
Table 5.5: Setting II: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.11) with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ (3.5) and average detection delay (given no false detection) of the change detection method used (3.16) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and composite post-change model.
T
Composite
post-change
Conditional
coverage
Unconditional
coverage
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 23.25 3.95 16.87
100 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 20.11 3.67 16.21
500 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.87 30.83 3.45 17.63
500 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.86 28.56 3.48 15.81
5.1.3 Composite pre- and post-change (Setting III)
Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, with the changepoint at $T=100$ and $T=500$. For change detection, we employ a weighted CUSUM-type detector (defined in (4.17)), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in setting II is considered. We set the thresholds at $A=1000,10000$. This is a special case of “e-detector” and is known to control ARL at level $A$ (Shin et al.,, 2023).
After detection, we construct confidence sets (as defined in (4.12)) for $T$ (assuming $T\neq 1$). We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs in Table 5.6.
Table 5.6: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ (4.6) and average detection delay (given no false detection) of the change detection method used (4.17) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and pre- and post-change model.
T
Composite
pre-change
Composite
post-change
Conditional
coverage
Unconditional
coverage
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 25.51 4.36 25.81
100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 22.97 4.19 23.15
500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.95 0.95 30.74 4.03 24.48
500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 29.16 4.07 23.07
However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.3). We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table 5.3), exhibiting both higher coverage and larger size.
Fig. 5.7 visualizes the confidence sets (4.12) and point estimates (4.6) of the changepoint across $5$ random runs.
Figure 5.7: Setting III: The first $T-1$ observations are drawn from $N(0,1)$ and the remaining observations from $N(1,1)$ with $T=100$ (top row) or $T=500$ (bottom row). Change is detected using the CUSUM detector defined in (4.17) with $A=1000$. The point estimate (4.6) is shown in vertical red dashed line and the confidence set (3.11) is shown in red points. $\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.1\}$, $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, $N=B=100$, $\alpha=0.05,\beta=\gamma=0.025$ and $L=\infty$. Results of $5$ independent simulations are shown.
5.2 Experiments with detection algorithms that control PFA
5.2.1 Known pre- and post-change (Setting A)
We perform experiments for both Normal and Poisson mean-change scenarios having the true changepoint at $T=100,500$, using a likelihood ratio detector defined as
$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (5.1)
with thresholds at $A=1000$. Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality (Ville,, 1939), the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we employ Algorithm 1 to construct confidence sets for $T$. We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.8. We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
Table 5.8: Setting A: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 1 with $B=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (defined in (5.1)) with $A=1000$, while varying the changepoint $T$.
$T$ $F_{0}$ $F_{1}$
Conditional
coverage
Unconditional
coverage
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 $N(0,1)$ $N(1,1)$ 0.95 0.94 12.60 2.74 110.91
100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 17.95 4.98 94.33
500 $N(0,1)$ $N(1,1)$ 0.97 0.97 12.65 2.24 514.12
500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 16.82 3.18 417.44
5.2.2 Known pre- and unknown post-change (Setting B)
We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector defined as
$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (5.2)
with thresholds at $A=1000$. The weights are chosen just as described in Section 5.1.2. Similarly, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we construct confidence sets for $T$ as defined in (3.12). We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.9. We also report the average absolute deviation of $\hat{T}$ (defined in (3.5)) and delay (given no false detection) in the table.
Table 5.9: Setting B: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.12) with $B=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation of $\hat{T}$ (defined in (3.5)) and average detection delay (given no false detection) of the change detector used (as defined in (5.2)) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and composite post-change model.
T
Composite
post-change
Conditional
coverage
Unconditional
coverage
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 $\{N(\mu,1):\mu>0.75\}$ 0.95 0.95 26.03 3.68 118.76
100 $\{N(\mu,1):\mu>0.9\}$ 0.94 0.94 24.84 3.36 115.64
500 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 30.95 3.18 530.83
500 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 30.09 3.23 528.92
5.2.3 Composite pre- and post-change (Setting C)
We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) Larsson et al., (2025); Grünwald et al., (2024); Lardy et al., (2024) being the representative of the pre-change class) defined as
$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (5.3)
with $f^{*}_{\theta}$ being the RIPr of $f_{\theta}$ on the pre-change class $\mathcal{P}_{0}$ and with thresholds at $A=1000$. This is a stop time of an ”e-process” (Ramdas et al.,, 2023) and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. The weights are chosen just as described in Section 5.1.3. After detection, we construct confidence sets for $T$ as defined in (4.13) (assuming $T\neq 1$). We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.10. We also report the average absolute deviation of $\hat{T}$ (defined in (4.6)) and delay (given no false detection) in the table.
Table 5.10: Setting C: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $B=100$ and $\alpha=0.05,\beta=\gamma=0.025,L=\infty$, the average absolute deviation of $\hat{T}$ (defined in (4.6)) and average detection delay (given no false detection) of the change detector used (as defined in (5.3)) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and pre- and post-change models.
T
Composite
pre-change
Composite
post-change
Conditional
coverage
Unconditional
coverage
Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.95 0.95 26.72 3.97 125.64
100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 25.48 3.88 121.28
500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 32.18 3.82 544.91
500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 31.45 3.75 536.76
5.3 Comparison of results with ARL and PFA controlling detection algorithms
As expected, algorithms that control PFA exhibit a significantly higher average detection delay compared to those controlling ARL. This delay arises because stricter false alarm constraints force the detection algorithm to accumulate stronger evidence before signaling a change.
In terms of inference, we observe that PFA-controlling algorithms often tend to yield slightly lower conditional coverage probabilities with confidence sets of comparable sizes. Also, conditional and unconditional coverage rates are identical when PFA-controlling detection algorithms are employed, due to the very low probability of false detections, while the ARL controlling algorithms lead to lower unconditional conditional coverage rates for $T=500$, supporting our theory. Moreover, we observe that the average absolute deviations of the point estimates are barely smaller with the PFA controlling algorithms.
Overall, the major differences between these two types lie in detection delay and unconditional coverage properties. We do not observe significant differences in the size and conditional coverage properties of the confidence sets nor in the absolute deviation of the point estimates, across these two types of detection algorithms.
6 Conclusion
We developed a framework to construct a confidence set for the unknown changepoint $T$, providing conditional coverage guarantees (conditioned on the event that the stopping occurs after the changepoint) when the pre- and post-change distributions are known. We also extended our framework to handle composite pre- and post-change distributions and provide conditional coverage guarantees, demonstrating its good empirical performance. We impose no specific conditions on the change detection algorithms; they can function as entirely black-box or heuristic methods; we only need to be able to run the algorithm on simulated data sequences. By filling this significant gap in post-detection analysis, our work provides crucial tools for precise and timely changepoint localization in real-time data streams. Our framework offers important implications for applications where rapid and accurate inference is critical, paving the way for more effective changepoint localization in sequential settings. While our method is designed for parametric pre- and post-change models in the partitioned setting, future work can explore its extensions to the non-partitioned setting and nonparametric models.
Appendix A Omitted proofs
Here, we provide the formal mathematical details that were omitted in the main paper.
A.1 Omitted proofs from Section 2
Proof of Theorem 2.2.
First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d. under $H_{0,t}$.
Hence, $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})\iff M_{t}$ is one among the $\lceil(1-\alpha)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$. So, for any fixed $\alpha\in(0,1)$,
$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-\alpha)(B+1)\rceil}{B+1}\geq 1-\alpha,\text{ for all }t.$ (A.1)
For $L\in\mathbb{N}\cup\{\infty\},t\in\mathbb{N}$, define $C_{t,L}=\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B}))$. Note that
$\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\},\tau<T)\leq\mathbb{P}_{T}[\tau<T]=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$
the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$.
Also, observe that for any $L\in\mathbb{N}\cup\{\infty\}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$ and hence, $C_{t,L}\geq C_{t,\infty}$, for all $t\in\mathbb{N}$.
$1-\alpha$ $\leq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})$
$\leq\mathbb{P}_{T}(M_{T}\leq C_{t,L})$
$=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})$
$=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\},\tau\geq T)+\mathbb{P}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\},\tau<T)$
$\leq\mathbb{P}_{T}(T\in\mathcal{C})+\mathbb{\mathbb{P}}_{\infty}(\tau<T).$
Therefore, $\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<T)$.
Now, if $\mathcal{A}$ controls PFA at level $\delta$, i.e., $\mathbb{P}_{\infty}(\tau<\infty)\leq\delta$, we have
$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<T)\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<\infty)\geq 1-\alpha-\delta.$
Proof of Proposition 2.4.
For the sake of contradiction, assume that $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$ for some $\mathcal{C}^{*}\subseteq\{1,\cdots,\tau\}$. Now we observe that
$\mathbb{P}_{\infty}(\tau\geq T)=\mathbb{P}_{T}(\tau\geq T)\geq\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha.$
Threfore, $\mathbb{P}_{\infty}(\tau=\infty)=1-\mathbb{P}_{\infty}(\cup_{n=1}^{\infty}[\tau<n])=1-\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau<n)=\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau\geq n)\geq 1-\alpha.$
Proof of Proposition 2.5.
Note that $\mathcal{C}^{*}=\{1\leq t\leq\tau:M_{t}\leq C_{t,L}\}\cup\{\tau+1,\tau+2,\cdots\}=\{t\geq 1:M_{t}\leq C_{t,L}\}$, since we defined $M_{t}=-\infty\leq C_{t,L}$, for all $t>\tau.$
Therefore, we have
$\mathbb{P}_{T}(T\in\mathcal{C}^{*})=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})=\mathbb{P}_{T}(M_{T}\leq C_{t,L})\geq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})\geq 1-\alpha,$ (A.2)
where the last inequality follows from (A.1). ∎
Proof of Corollary 2.6.
From Theorem 2.2, we get $\mathbb{P}_{T}(T\in\mathcal{C})\geq\mathbb{P}_{\infty}(\tau\geq T)-\alpha.$
Also, $\mathbb{P}_{T}(\tau\geq T)=\mathbb{P}_{\infty}(\tau\geq T)$ because of the fact the event $[\tau\geq T]$ does not depend $X_{i}$ for $i\geq T$. Therefore,
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)=\frac{\mathbb{P}_{T}(T\in\mathcal{C})}{\mathbb{P}_{T}(\tau\geq T)}\geq 1-\frac{\alpha}{\mathbb{P}_{\infty}(\tau\geq T)}.$
Therefore, if $\mathcal{A}$ controls PFA at level $\delta$, we have
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{\mathbb{P}_{\infty}(\tau\geq T)}\geq 1-\frac{\alpha}{\mathbb{P}_{\infty}(\tau=\infty)}\geq 1-\frac{\alpha}{1-\gamma}.$
A.2 Omitted proofs from Section 3
Proof of Theorem 3.1.
First, observe that $M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}$ are i.i.d. under $\mathbb{P}_{\theta_{0},t,\theta_{1}}$. So, $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-\alpha)(B+1)\rceil$ smallest of $M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}\}$. Hence, for any fixed $c\in(0,1)$,
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (A.3)
Define, $C_{t,L}^{\alpha}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$. Note that $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}\}),$ for all $t\in\mathbb{N},\alpha\in(0,1)$. Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
Then, we have
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}})$
$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B})\mid\hat{\alpha}_{t}\right)\right)+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$\leq\mathbb{E}(\hat{\alpha}_{t})+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$=\alpha\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)\leq(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
Similarly, we have for fixed $\alpha\in(0,1),$ with $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverage $1-\beta$
$P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta.$ (A.4)
We also have that
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha r_{t}})$
$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\})$
$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\},\tau<T)$
$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
Similarly, we have
$P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$ (A.5)
Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha r_{t}})\geq 1-(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta$ from the above inequality.
If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\geq 1-\alpha-\beta$, we obtain
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha+\beta}{\mathbb{P}_{\theta_{0},\infty}(\tau=\infty)}\geq 1-\frac{\alpha+\beta}{1-\delta}.$
Similarly,
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)-(\alpha+\beta)\geq 1-\alpha-\beta-\delta.$
A.3 Omitted proofs from Section 4
Proof of Theorem 4.2.
First, observe that $M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}$ are i.i.d. under $\mathbb{P}_{\theta_{0},t,\theta_{1}}$.
So, $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-\alpha)(B+1)\rceil$ smallest of $M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}\}$. Hence, for any fixed $c\in(0,1)$,
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (A.6)
Define, $\hat{\alpha}_{t}:=\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)$ and $C_{t,L}^{\alpha}=\operatorname{Quantile}(1-\alpha;M_{t},\{\max_{\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. Note that $\theta_{0}\in\mathcal{S}_{t-1}$ and $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}\}),$ for any $\alpha\in(0,1).$ By Assumption 4.1, we have that $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$.
Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
And similarly,
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\in\mathcal{S}_{t-1})=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})\mid r_{t}))\geq\mathbb{E}(1-\gamma r_{t})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
Then, we have
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\hat{\alpha}_{t}})$
$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\hat{\alpha}_{t}},\theta_{0}\in\mathcal{S}_{t-1},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\notin\mathcal{S}_{t-1}))+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))+(\beta+\gamma)\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$
$\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})\mid\hat{\alpha}_{t}\right)\right)+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$\leq\mathbb{E}(\hat{\alpha}_{t})+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$=\alpha\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
$\leq(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
Similarly, we have for fixed $\alpha\in(0,1),$ with $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverages $1-\beta$ and $1-\gamma$ respectively,
$P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta+\gamma.$ (A.7)
We also have that
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\hat{\alpha}_{t}})$
$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\})$
$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\},\tau<T)$
$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L})\geq 1-(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma$ from the above inequality. Similarly, we have
$P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$ (A.8)
If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\geq 1-\alpha-\beta$, we obtain
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta+\gamma}{\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha+\beta+\gamma}{\mathbb{P}_{\theta_{0},\infty}(\tau=\infty)}\geq 1-\frac{\alpha+\beta+\gamma}{1-\delta}.$
Similarly,
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)-(\alpha+\beta+\gamma)\geq 1-\alpha-\beta-\gamma-\delta.$
Appendix B Simulations for point estimates
In this section, we study the empirical performances of our point estimates of the changepoint defined in Equations 2.7, 3.5 and 4.6, which are the sequential analogs of the maximum likelihood estimators in the offline setting. Our empirical findings indicate that these estimators perform well in practice, even though we do not have theoretical guarantees on their properties.
B.1 Known pre- and post-change distributions
In this experiment, $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$ For detecting the change, we use the usual cumulative sum (CUSUM) and Shiryaev-Roberts (SR) procedures Tartakovsky et al., (2014) as defined below:
$\tau_{\text{cusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak$ (B.1)
$\tau_{\text{sr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (B.2)
where $A$ is some fixed threshold. For our experiments, we set the value of $A$ at $10000$. Once the change is detected, we obtain $\hat{T}$, as defined in (2.7), based on observations up to the detection time. Fig. B.1 shows histograms of the values taken by $\hat{T}$ over $10000$ independent experiments.
Figure B.1: Histogram of $\hat{T}$, defined in (2.7), for known pre- and post-change over $10000$ independent experiments with $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$
B.2 Known pre-change and composite post-change distributions
In this experiment, $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta>0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
Here, we use the weighted CUSUM/SR detectors defined in (3.16) and (3.17). We considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. Fig. B.2 shows histograms of the values taken by $\hat{T}$ defined in (3.5) over $10000$ independent experiments.
Figure B.2: Histogram of $\hat{T}$, defined in (3.5), for known pre-change and unknown post-change over $10000$ independent experiments with $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta\neq 0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
B.3 Composite pre- and post-change distributions
In this experiment, $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively. We use the weighted CUSM/SR-type change detectors defined in (4.17) and (4.18).
In this experiment, we considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=0.6,d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ Fig. B.3 shows histograms of the values taken by $\hat{T}$ defined in (4.6) over $10000$ independent experiments. It is crucial to note that the RIPr-based estimator has a higher bias but lower variance.
Figure B.3: Histogram of $\hat{T}$, defined in (4.6), for unknown pre- and post-change over $10000$ independent experiments with $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively.
Appendix C A heuristic method for composite pre- or post-change
This section provides an alternative method for constructing confidence intervals for composite pre/post-change distributions. While this approach lacks theoretical guarantees, it performs slightly better in our experiments than our previous approach. Specifically, we propose a heuristic bootstrap-based approach that estimates the unknown distributions from the observed data and subsequently generates simulations from the estimated distributions.
C.1 Known pre-change and unknown post-change distributions
Here we assume that the pre-change distribution $F_{0}$ is known and the composite post-change model is $\mathcal{P}_{1}$. Building on the approach from previous sections, we define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},\text{ for some (unknown) }F_{1}\in\mathcal{P}_{1}.$
For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined in (2.8)). Since the pre-change $F_{0}$ is known, we can generate $N$ many i.i.d. sequences under $F_{0}$ until a change is detected, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$. Now, the challenge in conducting a level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq t)$ test for each $t$ lies in the composite nature of $\mathcal{H}_{0,t}$, which prevents us from straightforwardly simulating data to determine a rejection threshold as we did in Algorithm 1. To address this, we employ a simple bootstrap-based approach. The main idea is to first obtain some point estimate of the changepoint (for example, a likelihood ratio based estimate, as defined in (3.5)). We denote the estimate as $\hat{T}$. Then we use the observations from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with known pre-change distribution $F_{0}$ and estimated post-change distribution $\hat{F}_{1}$ having changepoint at $t$.
For each sampled stream $j$, we compute the same test statistic $M_{t}^{j}$ under the null hypothesis, resulting in statistics $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if
$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 5 contains an overview of the method.
Input :  $\alpha$, $N,B$, change detection algorithm $\mathcal{A}$ and data sequence $X_{1},X_{2},\cdots$ until change is detected at $\tau$ using $\mathcal{A}$
output : A confidence set $\mathcal{C}$ for changepoint $T$
1 Compute a point estimator $\hat{T}$ for $T$, (eg., as defined in (3.5))
2 Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
3 for $i=1,\cdots,N$ do
4       Simulate $X_{1}^{i},X_{2}^{i}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first.
5
6 end for
[…]
8 for $t=1,\cdots,\tau$ do
9       Compute some test statistics $M_{t}$ (eg., as defined in (3.7)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
10       for $j=1,\cdots,B$ do
11             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
12             Compute the same test statistics $M_{t}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}$
13
14       end for
15      if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
16            $\mathcal{C}=\mathcal{C}\cup\{t\}$
17       end if
18
19 end for
Algorithm 5 CI for known pre-change $F_{0}$ and composite post-change $\mathcal{P}_{1}$
C.2 Composite pre- and post-change distributions
We now consider the general case where both pre and post-change distributions follow some composite models $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively. Following the approach from the previous sections, we define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0},\text{ for some }F_{0}\in\mathcal{P}_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},\text{ for some }F_{1}\in\mathcal{P}_{1}.$
For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined in (4.8)).
As in Algorithm 5, we employ a simple bootstrap-based approach, but here we need to estimate both pre- and post-change distributions. We first obtain some point estimate $\hat{T}$ of the changepoint $T$ (eg., the likelihhod-ratio-based estimate as defined in (4.6)). Then we use the observations from $1$ to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. For some $N\in\mathbb{N}$, we generate $N$ many i.i.d. sequences under $\hat{F}_{0}$ until a change is detected and denote the detection times as $\tau_{1},\cdots,\tau_{N}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with the estimated pre- and post-change distributions $\hat{F}_{0}$ and $\hat{F}_{1}$ respectively, having changepoint at $t$. For each sampled stream $j$, we calculate the test statistic $M_{t}^{j}$ to obtain $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if we have
$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 6 contains an overview of the method.
Input :  $\alpha$, $N,B$, change detection algorithm $\mathcal{A}$ and data sequence $X_{1},X_{2},\cdots$ until change is detected at $\tau$ using $\mathcal{A}$
output : A confidence set $\mathcal{C}$ for changepoint $T$
1 Compute some point estimator $\hat{T}$ for $T$ (eg., as defined in (4.6))
2 Obtain an estimate $\hat{F}_{0}$ of the (unknown) pre-change $F_{0}\in\mathcal{P}_{0}$ based on $X_{1},\cdots,X_{\hat{T}-1}$
3 Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
4 for $i=1,\cdots,N$ do
5       Simulate $X_{1}^{i},X_{2}^{i}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first.
6
7 end for
[…]
9 for $t=1,\cdots,\tau$ do
10       Compute some test statistics $M_{t}$ (eg., as defined in (4.8)) based on $X_{1},\cdots,X_{\tau}$
11       for $j=1,\cdots,B$ do
12             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
13             Compute the same test statistics $M_{t}^{j}$ based on $X_{1}^{j},\cdots,X_{\tau_{j}}^{j}$
14
15       end for
16      if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
17            $\mathcal{C}=\mathcal{C}\cup\{t\}$
18       end if
19
20 end for
Algorithm 6 CI for unknown pre- and post-change
C.3 Experimental Results
C.3.1 Known pre-change and composite post-change (Setting II)
We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$. For detection, we employ a weighted CUSUM detector (3.16) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 5 to construct confidence sets. Table C.1 summarizes the results, including the average size and the conditional and unconditional coverage across $100$ independent runs. For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
Table C.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.16) with threshold $A=1000$.
T Pre-change
Composite
post-change
True (unknown)
post-change
Conditional
coverage
Unconditional
coverage
Size
100 $N(0,1)$$\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.94 0.94 14.76
100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.95 20.95
500 $N(0,1)$$\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.95 0.86 14.06
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.85 20.83
C.3.2 Composite pre- and post-change (Setting III)
Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$. We use a weighted CUSUM detector with RIPr defined in (4.17) for change detection. The thresholds are set at $A=1000$ for both $T=100$ and $T=500$. We considered the same discrete weight distribution as above, taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=0.8,d=0.2$ and for Poisson distribution $\theta=1.8,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 6 to construct confidence sets. We report the average size and the conditional and unconditional coverage across $100$ independent runs in Table C.2. The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
Table C.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 6 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
T
Composite
post-change
Composite
post-change
True
pre-change
True
post-change
Conditional
coverage
Unconditional
coverage
Size
100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 19.52
100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.94 0.94 23.32
500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 18.56
500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 23.04
Appendix D Another heuristic method for composite pre- or post-change
In this section, we provide another heuristic method for dealing with composite pre- or post-change distributions. We use our algorithm for known pre- and post-change (Algorithm 2) with $F_{0}$ and $F_{1}$ being the “closest” pair of distributions in the composite pre and post-change models. For instance, when the pre-change model is given by $\{F_{\theta}:\theta\leq a\}$ and the post-change model is $\{F_{\theta}:\theta\geq b\}$, we set $F_{0}=F_{a}$ and $F_{1}=F_{b}$ in Algorithm 2. The intuition behind this choice is that the least separated pair of distributions might form a “worst-case” scenario. Nonetheless, we do not have any theoretical guarantee for coverage for this method.
For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section C.3.1 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.1. Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section C.3.2 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.1. From Table D.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
Table D.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.16) with threshold $A=1000$.
T Pre-change
Composite
post-change
True (unknown)
post-change
Conditional
coverage
Unconditional
coverage
Size
100 $N(0,1)$$\{N(\mu,1):\mu>0.75\}$ $N(1,1)$ 0.95 0.95 23.16
100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.94 0.94 25.83
500 $N(0,1)$$\{N(\mu,1):\mu>0.75\}$ $N(1,1)$ 0.95 0.88 23.82
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.86 27.17
Table D.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
T
Composite
post-change
Composite
post-change
True
pre-change
True
post-change
Conditional
coverage
Unconditional
coverage
Size
100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 26.03
100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 29.64
500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 26.52
500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 30.49
While we do not have any theoretical guarantees for this method, it is computationally more efficient than the previous approaches with theoretical guarantees, as demonstrated in Table E.1.
Appendix E Summary of comparisons of heuristic and non-heuristic methods
We first compare the execution times of the three types of algorithms and present the results in Table E.1. Algorithm 3 and Algorithm 4 are collectively referred to as Proposed Method, which comes with theoretical guarantees but requires more computational time compared to the heuristic approaches. The first heuristic approach, described in Algorithm 5 and Algorithm 6, is referred to as Heuristic 1. Finally, the heuristic introduced in this section (which is Algorithm 2, with $F_{0}$ and $F_{1}$ being the “closest” pair of distributions in the composite pre- and post-change models) is labeled as Heuristic 2.
Table E.1: Average (of $10$ independent runs) execution times (in seconds) taken by different methods, with $N=B=100$, $\alpha=0.05$, and for the proposed method, $\beta=\gamma=0.025$.
T A
Pre-change model
Post-change model
Proposed Method Heuristic 1 Heuristic 2
100 1000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 41.25 10.43 9.72
100 10000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 43.14 11.38 10.97
100 1000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 66.13 11.49 11.76
100 10000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 69.47 11.86 11.52
500 1000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 364.18 143.98 148.08
500 10000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 380.46 155.89 144.75
500 1000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 535.82 171.13 167.46
500 10000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 561.74 177.28 169.41
From Table E.1, we observe that the computational times are lower for the heuristic algorithms. Moreover, comparing the corresponding tables, we observe that the theoretically valid method (Proposed Method) gives higher coverage, also higher size of the confidence set, as compared to both heuristic methods. This suggests that while the heuristic approaches offer practical benefits in terms of speed, they may lead to less conservative inference, despite not having theoretical guarantees. This points at a natural direction for future work: to establish guarantees for our heuristic algorithms, or demonstrate settings where they empirically do not have coverage.

### v2

@@
 meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$.
-We show how to construct such sets $\mathcal{C}$ for both types of algorithms (those with bounded ARL or PFA).
+We show how to construct such sets $\mathcal{C}$ for both types of algorithms (bounded ARL or PFA).
 Sequential (also known as the online or quickest) changepoint detection is a well-studied field, where the goal is to detect that a change has occurred as soon as possible, is crucial in applications requiring timely responses to distributional shifts, such as quality control, cybersecurity, medical diagnostics, and so on.
 The field traces back to early work on statistical process control.
-Shewhart, (1925) introduced the concept of a control chart, which computes a test statistic for each sample, with values within a threshold indicating the process remains in control.
-Page, (1954) proposed the CUSUM (cumulative sum) procedure, which computes the cumulative log-likelihood ratio under the assumption of two exactly known distributions.
-Another influential approach is the Shiryaev-Roberts (SR) procedure (Shiryaev,, 1963; Roberts,, 1966).
+Shewhart [19] introduced the concept of a control chart, which computes a test statistic for each sample, with values within a threshold indicating the process remains in control.
+Page [16] proposed the CUSUM (cumulative sum) procedure, which computes the cumulative log-likelihood ratio under the assumption of two exactly known distributions.
+Another influential approach is the Shiryaev-Roberts (SR) procedure [21, 18].
 These early methods, however, assume the pre- and post-change distributions to be fully specified.
-To relax the assumption of known post-change distributions, the Generalized Likelihood Ratio (GLR) procedure was introduced (Lorden,, 1971; Siegmund and Venkatraman,, 1995).
-Tartakovsky et al., (2014) provide comprehensive overviews of these classical methods, including CUSUM, SR, and GLR, and their numerous variants.
-Several nonparametric change-point detection algorithms have been developed, such as kernel-based methods (Harchaoui et al.,, 2008; Li et al.,, 2015) and e-detectors (Shin et al.,, 2023).
+To relax the assumption of known post-change distributions, the Generalized Likelihood Ratio (GLR) procedure was introduced [15, 22].
+Tartakovsky et al.
+[23] provide comprehensive overviews of these classical methods, including CUSUM, SR, and GLR, and their numerous variants.
+Several nonparametric change-point detection algorithms have been developed, such as kernel-based methods [7, 14] and e-detectors [20].
 Localizing the changepoint is another fundamental and crucial problem.
 As far as we are aware, the existing approaches to changepoint localization focus on offline or batch settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times.
-Hinkley, (1970) shows some asymptotic results about the maximum likelihood estimator (MLE) of the changepoint and also provides a confidence set based on likelihood ratio statistics.
-Yao, (1987) provides an approximation of the distribution of the MLE of the changepoint.
-Darkhovskh, (1976) presents a nonparametric changepoint estimator based on Mann-Whitney statistics.
-Carlstein, (1988) proposes a class of strongly consistent estimators for the changepoint in the nonparametric setting.
-Worsley, (1986) develops likelihood ratio-based confidence regions and tests for a changepoint for exponential family random variables.
-Dumbgen, (1991) constructs nonparametric confidence regions for changepoint by testing for each candidate time-point $t$ at level $\alpha$ using bootstrapping, achieving asymptotic coverage of $1-\alpha$, as the sample size $n\to\infty$.
-A more recent method by Jang and Walther, (2024) constructs confidence sets by evaluating a local test statistic on a triplet of time points and combining the results of the local tests with a weighted Bonferroni correction.
+Hinkley [8] shows some asymptotic results about the maximum likelihood estimator (MLE) of the changepoint and also provides a confidence set based on likelihood ratio statistics.
+Yao [26] provides an approximation of the distribution of the MLE of the changepoint.
+Darkhovskh [3] presents a nonparametric changepoint estimator based on Mann-Whitney statistics.
+Carlstein [1] proposes a class of strongly consistent estimators for the changepoint in the nonparametric setting.
+Worsley [25] develops likelihood ratio-based confidence regions and tests for a changepoint for exponential family random variables.
+Dümbgen [5] constructs nonparametric confidence regions for changepoint by testing for each candidate time-point $t$ at level $\alpha$ using bootstrapping, achieving asymptotic coverage of $1-\alpha$, as the sample size $n\to\infty$.
+A more recent method [10] constructs confidence sets by evaluating a local test statistic on a triplet of time points and combining the results of the local tests with a weighted Bonferroni correction.
 To re-emphasize, all these methods are for the offline setting, while our main focus is on the sequential problem of constructing confidence sets after stopping, which has been largely overlooked in the literature.
@@
 The key idea behind constructing the confidence interval is simple: for each candidate $t\in\{1,\cdots,\tau\}$, we will test whether $t$ is the true changepoint and form the confidence set by inverting these tests.
-Our approach is conceptually similar to Dumbgen, (1991), it adapts those ideas to the sequential setting in a nontrivial manner.
+Our approach is conceptually similar to [5], it adapts those ideas to the sequential setting in a nontrivial manner.
 For example, we show that for change detection algorithms that have finite average run length (ARL), it is impossible to obtain finite length confidence intervals with marginal coverage guarantees.
@@
 Our method is designed for the partitioned setting.
-While extending it to the non-partitioned case is an interesting direction, it is left for future work.
-Our contributions.
+Extensions to the non-partitioned case are left for future work.
+1.1 Our contributions
 Here we summarize the key contributions of the paper.
-  • For detection algorithms that control PFA, we propose Algorithm 1 for constructing a confidence set for the changepoint and we establish its unconditional and conditional coverage guarantees in Theorem 2.2 and Corollary 2.6, assuming both pre- and post-change distributions to be known.
+  • For detection algorithms that control PFA, we propose Algorithm 1 for constructing a confidence set for the changepoint and establish its unconditional and conditional coverage (Theorem 2.2, Corollary 2.6), assuming known pre- and post-change distributions.
   • For detection algorithms with bounded ARL, we show in Proposition 2.4 that we cannot have an unconditional coverage guarantee and propose Algorithm 2, which is a modification of Algorithm 1.
@@
   • We extend our framework to composite pre- and post-change classes in Algorithm 3 and Algorithm 4 and provide the coverage guarantees in Theorem 3.1 and Theorem 4.2.
-  • We demonstrate that the approach is not overly conservative (and thus practical) by examining the achieved coverage and size properties of our constructed confidence sets in extensive simulations.
-We point out again that for obtaining the aforementioned guarantees, we just need to have black-box access to the detection algorithm $\mathcal{A}$, meaning that we do not need to know its inner workings (e.g., it can be a heuristic algorithm), and mainly need to be able to run that algorithm on any stream of data.
-Paper outline.
-The rest of the paper is organized as follows.
+  • We demonstrate that the approach is not overly conservative (and thus practical) by examining the achieved coverage and size of constructed confidence sets in extensive simulations.
+We point out again that for obtaining the aforementioned guarantees, we only need to have black-box access to the detection algorithm $\mathcal{A}$.
+It need not have formal guarantees, we do not need to know its inner workings, and just need to be able to run $\mathcal{A}$ on any stream of data.
+1.2 Paper outline
 In Section 2, we construct our confidence sets for the changepoint for known pre- and post-change and study its unconditional and conditional coverage probabilities.
@@
 Omitted proofs can be found in Appendix A.
-Some additional experimental results are provided in Appendix B and some heuristic algorithms are presented in Appendices C and D, which are computationally more efficient.
+An additional illustrative example using the Poisson distribution is given in Appendix B.
+Appendix B.
+Some additional experimental results are provided in Appendix C and some heuristic algorithms are presented in Appendices D and E, which are computationally more efficient.
 A summary of comparisons of the heuristic and non-heuristic methods is provided in Appendix E, pointing at open avenues for further theoretical work.
@@
 The other one is the probability of false alarm (PFA), that is, $\mathbb{P}_{\infty}(\tau<\infty)$.
-We will provide solutions for algorithms $\mathcal{A}$ that control either metric.
+We provide solutions for algorithms $\mathcal{A}$ that control either metric.
 2.1 A first confidence set
-Our main idea is to test whether there is a changepoint at $t$, for each $t\in\{1,\cdots,\tau\}$ and then construct the confidence set consisting of all time points $t$ for which we fail to reject the tests.
-Formally, we test the following null hypothesis at each candidate time point $t$:
+We will test whether there is a changepoint at $t$, for each $t\in\{1,\cdots,\tau\}$ and then construct the confidence set consisting of all time points $t$ for which we fail to reject.
+Formally, we test the following null hypothesis at each candidate time $t$:
 $H_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\text{and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}.$
-We consider some test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ such that for each $t\in\mathbb{N}$,
+Consider some test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ such that for each $t\in\mathbb{N}$,
 $M_{t}:=\begin{cases}M_{t}^{(\tau)}(X_{1},\cdots,X_{\tau}),&\text{if }t\leq\tau<\infty,\\ -\infty&\text{if }t>\tau\text{ or }\tau=\infty,\end{cases}$ (2.1)
@@
 Output: Confidence set $\mathcal{C}$ for the changepoint $T$
-1 Detect change at $\tau$ using $\mathcal{A}$
-2 $\mathcal{C}=\varnothing$
+1 Detect change at $\tau$ using $\mathcal{A}$; Set $\mathcal{C}=\varnothing$
 2 for $t=1,\cdots,\tau$ do
@@
 6             Subroutine $\mathcal{L}$: Compute $M_{t,L}^{j}$ (defined in (2.3),(2.4)) based on $\tau_{j}^{\prime}$, $L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$
-8
+7
 8       end for
@@
 11       end if
-13
+12
 13 end for
 Algorithm 1 CI for known pre-change $F_{0}$ and post-change $F_{1}$
-We now provide more details about Subroutine $\mathcal{L}$.
 2.2 Subtleties regarding Subroutine $\mathcal{L}$
 The problem with $L=\infty$ is that the stopping time may never be reached for the $j$-th sampled sequence, i.e., we may have $\tau^{\prime}_{j}=\infty$.
-One can safely set $L=\infty$ if the $\mathcal{A}$ (i.e., its stopping time) satisfies the condition $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$, meaning that if there is a change, $\mathcal{A}$ stops almost surely.
-This is a very mild assumption because any reasonable change detector should always eventually raise an alarm if there is a change.
+One can set $L=\infty$ if $\mathcal{A}$ (i.e., its stopping time) satisfies the condition $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely).
+This is a very mild assumption because any reasonable change detector will always eventually raise an alarm if there is a change.
 Indeed, all algorithms with finite ARL satisfy the condition, and indeed typically so do algorithms that control PFA.
 To avoid making even this mild assumption, one can simply pick $L<\infty$.
-If $L<\infty$ and the $j$th sequence stopped because the algorithm $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual, but if $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it).
+If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual, but if $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it).
 Therefore, for $L<\infty$,
@@
 The rationale behind setting $M^{j}_{t,\infty}=-\infty$ when $\tau^{\prime}_{j}=\infty$ is the following.
-If for all $j=1,\cdots,B$, we have $\tau_{j}=\infty$, then we want to reject $H_{0,t}$, which is achieved by setting $M^{j}_{t,\infty}=-\infty$.
+If $\tau_{j}=\infty$, for all $j=1,\cdots,B$, then we want to reject $H_{0,t}$, which is achieved by setting $M^{j}_{t,\infty}=-\infty$.
 Remark 2.1.
@@
 $L_{T}^{n}=\prod_{i=1}^{T-1}f_{0}(X_{i})\prod_{i=T}^{n}f_{1}(X_{i}),$ (2.5)
-and the maximum likelihood estimator of $T$ (Hinkley,, 1970; Yao,, 1987) is given by
+and the maximum likelihood estimator of $T$ [8, 26] is given by
 $\hat{T}_{n}=\operatorname*{arg\,max}_{1\leq j\leq n}L_{j}^{n}=\operatorname*{arg\,max}_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ (2.6)
@@
 Remark 2.3.
-When $T\ll$ ARL, we typically have $\mathbb{\mathbb{P}}_{\infty}(\tau<T)\approx 0$ and so $\mathbb{P}_{T}(T\in\mathcal{C})\gtrapprox 1-\alpha$.
+When $T\ll$ ARL, we typically have $\mathbb{\mathbb{P}}_{\infty}(\tau<T)\approx 0$ and so $\mathbb{P}_{T}(T\in\mathcal{C})\gtrsim 1-\alpha$.
 The confidence set constructed using our method will approximately achieve the desired unconditional coverage level $1-\alpha,$ in situations where the probability of false detections before the true changepoint is negligible.
@@
 2.5 Unconditional coverage when $\mathcal{A}$ has finite ARL
-If the detection algorithm has finite ARL (i.e., $\tau$ is finite almost surely), then it is impossible for any confidence set $\mathcal{C}^{*}$, which is a subset of $\{1,\cdots,\tau\}$, to produce an unconditional coverage guarantee at a predetermined level $1-\alpha$.
-The intuitive reason is that if $\tau$ is random and finite almost surely, it is possibly a false detection, i.e., $\tau<T$, in which case the set $\{1,\cdots,\tau\}$ itself does not cover $T$.
-The next proposition shows that.
+If the detection algorithm has finite ARL (i.e., $\tau$ is finite almost surely even if there is no change), then it is impossible for any confidence set $\mathcal{C}^{*}$, which is a subset of $\{1,\cdots,\tau\}$, to produce an unconditional coverage guarantee at a predetermined level $1-\alpha$.
+The intuitive reason is that if $\tau$ is finite almost surely, it is possibly a false detection, i.e., $\tau<T$, in which case the set $\{1,\cdots,\tau\}$ itself does not cover $T$.
+The next proposition formalizes this fact.
 Proposition 2.4.
-If a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies an unconditional coverage guarantee $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$, for some $\alpha\in(0,1)$, then $\mathbb{P}_{\infty}(\tau=\infty)\geq 1-\alpha$ (thus, the PFA of the algorithm is at most $\alpha$ and thus the ARL of the algorithm is infinite).
+If a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies an unconditional coverage guarantee $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$, for some $\alpha\in(0,1)$, then $\mathbb{P}_{\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL of the algorithm is infinite).
 Said differently, for any algorithm with finite ARL, it is impossible to construct a valid confidence set that is only a subset of $\{1,\dots,\tau\}$.
-The proposition says that if we can have a confidence set with high unconditional coverage, then the detection algorithm is such that it never stops with high probability if there is no change.
+The proposition says that if we have a confidence set with high unconditional coverage, then the detection algorithm never stops with high probability if there is no change.
 Therefore, in order to achieve an unconditional coverage guarantee with $\mathcal{A}$ having finite ARL, we must extend our confidence set beyond the stopping time $\tau$.
@@
 $\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (2.13)
-where $\hat{\alpha}_{t}=\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$. Algorithm 2 contains a structured overview of the method.
+where $\hat{\alpha}_{t}=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$. Algorithm 2 contains a structured overview of the method.
 Input: $\alpha\in(0,1)$, $N\in\mathbb{N},B\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
@@
 6for $t=1,\cdots,\tau$ do
-7       Compute some test statistic $M_{t}$ (e.g., as defined in (2.9)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
+7       Compute some test statistic $M_{t}$ (e.g. (2.9)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
 8       for $j=1,\cdots,B$ do
@@
 12       end for
-13      Compute $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$
+13      Compute $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$
 14       if $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$ then
@@
 Proof.
-Step 1 [Establish the relationship between the level of testing ${H}_{0,T}$ and the conditional coverage]: For each $t$, define $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$.
-Then, the level of our test for ${H}_{0,T}$ is $P_{T}(M_{T}>C_{t,L}).$ Now,
-$\mathbb{P}_{T}(M_{T}\leq C_{t,L})$ $=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\})$
+First note the fact that $\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
+Step 1 [Relate the level of testing ${H}_{0,T}$ and the conditional coverage]: Defining $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, the level of testing ${H}_{0,T}$ is $P_{T}(M_{T}>C_{t,L}).$ Thus,
+$\mathbb{P}_{T}(M_{T}\leq C_{t,L})$
+$=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\})$
 $=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau\geq T)+\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau<T)$
-$\leq\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\mid\tau\geq T)\mathbb{P}_{T}(\tau\geq T)+\mathbb{\mathbb{P}}_{\infty}(\tau<T)$
-$=\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{\infty}(\tau\geq T).$
-The inequality in the third line holds because
-$\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t}^{N}\},\tau<T)\leq\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$
-where the last equality holds since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$ and by the same argument, $\mathbb{P}_{t}(\tau\geq t)=\mathbb{\mathbb{P}}_{\infty}(\tau\geq t)$ for all $t$.
-So, we obtain
+$\leq\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\mid\tau\geq T)\mathbb{P}_{T}(\tau\geq T)+\mathbb{\mathbb{P}}_{T}(\tau<T)$
+$=\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{\infty}(\tau\geq T),$
+where we twice invoked $\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T)$ as explained earlier. So, we obtain
 $\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{t,L})}{P_{\infty}(\tau\geq T)}.$ (2.15)
@@
 So, for any fixed $c\in(0,1)$,
-$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c\text{ for all }t.$ (2.16)
-Then, it follows that
-$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t})\geq\frac{\lceil(1-\hat{\alpha}_{t})(B+1)\rceil}{B+1}\geq 1-\hat{\alpha}_{t},\text{ almost surelyfor all }t.$ (2.17)
+$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c\text{ for all }t.$
+Since $\hat{\alpha}_{t}$ is independent of $M_{t}$, it follows that
+$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t})\geq\frac{\lceil(1-\hat{\alpha}_{t})(B+1)\rceil}{B+1}\geq 1-\hat{\alpha}_{t},\text{ almost surelyfor all }t.$
 Therefore, we have
-$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})=\mathbb{E}_{t}(\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t}))\geq\mathbb{E}_{t}(1-\hat{\alpha}_{t})=1-\alpha\mathbb{P}_{t}(\tau\geq t)=1-\alpha\mathbb{P}_{\infty}(\tau\geq t).$ (2.18)
-Also, for all $L\in\mathbb{N}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$ and hence, $C_{t,L}\geq C_{t,\infty}$.
-Thus, we obtain $\mathbb{P}_{t}(M_{t}>C_{t,L})\leq\mathbb{P}_{t}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{\infty}(\tau\geq t)$, for all $t$.
-Therefore, from (2.15), we have $\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
+$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})=\mathbb{E}_{t}(\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t}))\geq\mathbb{E}_{t}(1-\hat{\alpha}_{t})=1-\alpha\mathbb{P}_{t}(\tau\geq t).$
+As noted earlier, $\mathbb{P}_{t}(\tau\geq t)=\mathbb{P}_{\infty}(\tau\geq t)$ and hence, $\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})\geq 1-\alpha\mathbb{P}_{\infty}(\tau\geq t).$ Also, for all $L\in\mathbb{N}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$.
+Therefore, $C_{t,L}\geq C_{t,\infty}$.
+Thus, for all $t$,
+$\mathbb{P}_{t}(M_{t}>C_{t,L})\leq\mathbb{P}_{t}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{\infty}(\tau\geq t)$ (2.16)
+Therefore, combining (2.15) and (2.16), we obtain $\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
 Remark 2.8.
@@
 Notably, conventional confidence intervals are invalid in this context, since $\tau$ depends on the data.
-Therefore, we rely on confidence sequences (Darling and Robbins,, 1967; Jennison and Turnbull,, 1989; Howard et al.,, 2021).
+Therefore, we rely on confidence sequences [4, 11, 9, 2].
 A confidence sequence having coverage $1-c$ is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform coverage guarantee: $\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage.
-Otherwise, we construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage.
+Otherwise, we construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage.
 Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS has coverage probability $\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; we formalize this more explicitly in the proof of the Theorem 3.1.
@@
 $\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\},$ (3.2)
-where $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$.
+where $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$.
 If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$, we can have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
@@
 Algorithm 3 contains an overview of the method.
-Note that, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set, it can be computationally challenging to approximate the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$.
-However, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient.
-We provide details in Section 3.2 and a concrete Gaussian mean shift example in Section 3.3.
+We note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, one can approximate the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ by discretizing the parameter space into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum.
+Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient, and exact (i.e., does not require any approximation).
+We provide the details in Section 3.2.
+Additionally, we illustrate concrete examples of the Gaussian mean shift problem in Section 3.3 and the Poisson rate change problem in Appendix B.
 Input: $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N},L\in\mathbb{N}\cup\{\infty\},$ a confidence sequence procedure $\operatorname{CS(data;coverage)}$, a change detection algorithm $\mathcal{A}$, and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
@@
 6       end for
-7      $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, for $t=1,\cdots,\tau$
+7      $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, for $t=1,\cdots,\tau$
 8 end if
@@
 If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, then for any $\theta_{1}\in\Theta_{1}$,
-$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\delta\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{1-\delta},\text{ when }\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0.$
+$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\delta\leavevmode\nobreak\ \text{ , and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{1-\delta},$
+where the latter guarantee holds if $\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0.$
 3.1 Prototypical choice of test statistic $M_{t}$
@@
 We define $\hat{\theta}_{1,j:n}$ to be a maximum likelihood estimator (MLE) of the post-change parameter $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$.
-The point estimate of $T$ can be naturally extended to the following form:
+The point estimate of $T$ can be naturally extended to:
 $\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (3.5)
@@
 Note that the test statistic $M_{t}$ for testing $\mathcal{H}_{0,t}$ reduces to
-$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{ if }t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (3.7)
+$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (3.7)
 3.2 Computationally and statistically efficient subroutines
@@
 In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$.
+In the case of discrete distributions, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta)\}_{n}$ across all $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Section B.1 for the Poisson distribution.
 Let us now fix $t\in\{1,\cdots,\tau\}$.
@@
 Again, we bound $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)$ by $V_{t,L}^{j}$, where for $L<\infty$,
-$V_{t,L}^{j}:=\begin{cases}-\infty&\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{ if }t\leq t_{2}^{j}\leq L\\ \infty&\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}$ (3.9)
+$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (3.9)
 And for $L=\infty$,
-$V_{t,\infty}^{j}:=\begin{cases}\displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{ if }t\leq t_{2}^{j}\text{ and }t_{1}^{j}<\infty\\ -\infty&\text{ if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \end{cases}$ (3.10)
-which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, as we explain in Section 3.3). We define the confidence interval as
-$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (3.11)
+$V_{t,\infty}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{otherwise,}\\ \end{cases}$ (3.10)
+which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section 3.3). We define the confidence set as
+$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (3.11)
 $\mathcal{C}^{\prime}_{2}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}.$ (3.12)
 We observe that $\mathcal{C}^{\prime}\supseteq\mathcal{C}$, where $\mathcal{C}$ is the set produced by Algorithm 3, since
-$\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t);M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$ $\geq\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t);M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
-$\geq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t);M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}).$
+$\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$
+$\geq\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
+$\geq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}).$
 Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 3.1.
@@
 $\mathcal{C}^{\prime}_{1}$ and $\mathcal{C}^{\prime}_{2}$ defined in (3.11) and (3.12) respectively, satisfy the same coverage guarantees as $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ in Theorem 3.1.
+In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with Poisson distribution in Appendix B.
 3.3 A concrete example: Gaussian mean shift problem
-We consider the most common changepoint model, where the pre-change distribution is $F_{0}$ is $N(\theta_{0},1)$ and the post-change model is $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
-In this setting, we have the following example of a $1-\beta$ confidence sequence (Howard et al.,, 2021) for $\theta_{1}$
+Let the pre-change distribution $F_{0}$ be $N(\theta_{0},1)$ and the post-change model be $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
+In this setting, we have the following example of a $1-\beta$ confidence sequence [9] for $\theta_{1}$
 $\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (3.13)
@@
 Note that we can consider $\theta^{\prime\prime}=\theta_{0}$ and $G(x,\theta)=x+\theta-\theta_{0}$.
-Therefore, only need to generate the sequence $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(\theta_{0},1)$ for $j=1,\cdots,B$.
+Therefore, only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(\theta_{0},1)$ for $j=1,\cdots,B$.
 For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
-$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (3.14)
+$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t,\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (3.14)
 where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$.
 We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (2.8), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
-Let, $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$, where $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$.
+Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$.
 Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark 3.3) and they are finite.
 So, we can have $L=\infty$.
-If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,V_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
-$V_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}{L_{t}(\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j}+\theta;Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}$
+If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t$, it can be simplified in our setting as below:
+$V_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j}}{t^{\prime}-t+1}+\theta;Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}$
 $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
-$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta_{t^{\prime}}^{i});Y_{1}^{t}(X_{1}^{j},\theta_{t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta_{t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$ (3.15)
+$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta_{t^{\prime}}^{i});Y_{1}^{t}(X_{1}^{j},\theta_{t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta_{t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
 where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta)=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta),$ and $\theta_{t^{\prime}}^{i}$ is the unique element that maximizes $L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)$, or equivalently minimizes
 $-2\log L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)=$
-$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\text{constant (wrt }\theta),\text{ if }i<t,\\ \text{constant (wrt }\theta),\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{constant (wrt }\theta),\text{ if }i>t,\end{cases}$
+$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{(t-i)(\theta-\theta_{0})}{t^{\prime}-i+1}\right)^{2}+\text{constant},\text{ if }i<t,\\ \text{constant},\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{ constant},\text{ if }i>t,\end{cases}$
 over $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$.
+Here, constant refers to the terms that do not involve $\theta$.
 Note that $\theta_{t^{\prime}}^{i}$ has a simple closed-form expression since the above equation is a quadratic function of $\theta$.
-And hence, $V_{t}^{j}$ (3.3) can be computed easily.
+And hence, $V_{t,\infty}^{j}$ can be computed easily.
 Remark 3.3.
@@
 For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$.
-Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{\tau}(2X_{i}-\theta_{0}-\theta)\}$ increses as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases.
+Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{\tau}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases.
 Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
@@
 Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
-One can easily check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section 5.2.
+One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section 5.2.
 The Gaussian mean shift example we provided serves as a concrete and illustrative case to clarify the methodology.
@@
   1. 1.
-    The coupling function $G$ is available (which is always the case when distributions are continuous).
+    The coupling function $G$ is available (which is always the case when distributions are continuous and in the case of discrete distributions, even if a suitable $G$ is not directly available, our method may still be implementable, as demonstrated in Section B.1 for the Poisson distribution).
   2. 2.
@@
   4. 4.
-    If we employ the likelihood-rato-based test statistic (2.8), then we have to be able to maximize the function $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
+    If we employ the likelihood-ratio-based test statistic (2.8), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
 For instance, one can easily verify that this approach applies to Gaussian, Laplace, or exponential scale change problems.
@@
 sequences under $F_{\theta_{0}^{*}}$ until a change is detected using $\mathcal{A}$, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$, that will help us estimate $\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$, which is a lower bound on $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$.
-If $\mathcal{A}$ controls PFA at a known level $\delta\in(0,1)$, we can avoid this step and obtain a conditional/unconditional coverage guarantee without any such assumption.
+If $\mathcal{A}$ controls PFA at a known level $\delta\in(0,1)$, we can avoid this step and still obtain conditional or unconditional coverage.
 Since both pre- and post-change parameters are unknown, for each $t$, we estimate them based on $X_{1},\cdots,X_{t-1}$ and $X_{t},\cdots,X_{\tau}$ respectively.
-However, as noted earlier, the sample size used for the estimates might be too small to yield reliable estimates.
-Therefore, instead of point estimates, we rely on interval estimates in order to achieve a predetermined control over the error.
+Instead of point estimates, we rely on interval estimates in order to achieve a predetermined control over the error.
 Suppose that we have some confidence interval procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$, which constructs $1-c$ confidence interval using $X_{1},\cdots,X_{n}$ for $\theta_{0}$ and a confidence sequence procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, which constructs $1-c$ confidence sequence using $\{X_{n}\}_{n}$ for $\theta_{1}$.
 If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage and $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma)$ for $\theta_{0}$ having $1-\gamma$ coverage.
-Otherwise, we construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$.
+Otherwise, we construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$.
 Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS and CI have coverage probabilities $1-\beta\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ and $1-\gamma\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ respectively; we formalize this more explicitly in the proof of Theorem 4.2.
@@
 $M^{j}_{t,\infty}(\theta,\theta^{\prime}):=\begin{cases}M_{t}(X_{1}^{j}(\theta,t,\theta^{\prime}),\cdots,X_{\tau^{j,t}_{\theta,\theta^{\prime}}}^{j}(\theta,t,\theta^{\prime})),&\text{if }t\leq\tau^{j,t}_{\theta,\theta^{\prime}}<\infty,\\ -\infty,&\text{if }\tau^{j,t}_{\theta,\theta^{\prime}}=\infty\text{ or }t>\tau^{j,t}_{\theta,\theta^{\prime}}.\\ \end{cases}$ (4.2)
-Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$ at level $\alpha\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$,
+Our confidence set is the set of $t\leq\tau$ for which we fail to reject $\mathcal{H}_{0,t}$ at level $\alpha\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$,
 $\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\},$ (4.3)
-where $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$.
+where $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$.
 If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$ , we have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
 $\mathcal{C}_{2}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}.$ (4.4)
-Algorithm 4 contains an overview of the method.
-Note that, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are finite sets, it can be computationally challenging to approximate the threshold $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$.
-However, in many cases, we can derive a suitable upper bound of this threshold, which is computationally and statistically efficient.
+Algorithm 4 contains an overview.
+As before, even if $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are uncountable, one can approximate the threshold $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ by discretizing the parameter space into a grid and evaluating the quantile for values of $(\theta,\theta^{\prime})$ on this grid, taking their maximum.
+Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient, and exact (i.e., does not require any approximation).
 We provide the details in Section 4.2.
-We offer a concrete example of the Gaussian mean shift problem in Section 4.3.
+Additionally, we illustrate concrete examples of the Gaussian mean shift problem in Section 4.3 and the Poisson rate change problem in Appendix B.
 Input: $\theta_{0}^{*},\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N},L\in\mathbb{N}\cup\{\infty\}$, change detection algorithm $\mathcal{A}$, a confidence interval procedure $\operatorname{CI(data;coverage)},$ a confidence sequence procedure $\operatorname{CS(data;coverage)}$, and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
@@
 3       for $i=1,\cdots,N$ do
-4             Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}^{*}}$, define $\tau_{i}$ as the time when $\mathcal{A}$ detects a change, or $\tau$, whichever occurs first
+4             Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}^{*}}$, define $\tau_{i}$ as the time when $\mathcal{A}$ detects a change, or $\tau$, whichever is first
 5
 6       end for
-7      Define $r_{t}=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$ for $t=1,\dots,\tau$;
+7      Define $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$ for $t=1,\dots,\tau$;
 8 end if
@@
 $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
-If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, $\mathcal{C}_{2}$ satisfies the following guarantee for any $\theta_{i}\in\Theta_{i}(i=0,1)$:
-$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\gamma-\delta,\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta+\gamma}{1-\delta},\text{ when }\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0.$
+If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, we have that for any $\theta_{i}\in\Theta_{i}(i=0,1)$:
+$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\gamma-\delta,\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta+\gamma}{1-\delta}$
+[…]
 Remark 4.3.
@@
 $Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (4.9)
+Note that, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta,\theta^{\prime})\}_{n}$ across all $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Section B.2 for the Poisson distribution.
 Let us now fix $t\in\{1,\cdots,\tau\}$.
@@
 However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$ and $t_{2}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$.
-Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient.
+Also, we can often derive a suitable upper bound on the threshold, which does not require any approximation and is computationally and statistically efficient.
 E.g., we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$.
 Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where for $L<\infty$,
-$U_{t,L}^{j}:=\begin{cases}-\infty&\text{if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),&\text{if}t\leq t_{2}^{j}\leq L\\ \infty&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (4.10)
+$U_{t,L}^{j}:=\begin{cases}-\infty,\leavevmode\nobreak\ \leavevmode\nobreak\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (4.10)
 And for $L=\infty$,
-$U_{t,\infty}^{j}:=\begin{cases}\displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),&\text{if }t\leq t_{2}^{j}\text{ and }t_{1}^{j}<\infty\\ -\infty&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty.\end{cases}$ (4.11)
+$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (4.11)
 We consider the following confidence sets:
-$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (4.12)
+$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (4.12)
 $\mathcal{C}^{\prime\prime}_{2}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}.$ (4.13)
@@
 $\mathcal{C}^{\prime\prime}_{1}$ and $\mathcal{C}^{\prime\prime}_{2}$ defined in (4.12) and (4.13) respectively satisfy the same coverage guarantees as $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ in Theorem 4.2.
+In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with Poisson distribution in Appendix B.
 4.3 A concrete example: Gaussian mean shift problem
 We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$.
-In this setting, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$
+Then, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$:
 $\mathcal{S}_{t-1}=\left[\bar{X}_{1:t-1}-\frac{q_{\gamma/2}}{\sqrt{t-1}},\bar{X}_{1:t-1}+\frac{q_{\gamma/2}}{\sqrt{{t-1}}}\right]\cap\Theta_{0},$ (4.14)
@@
 Note that we can consider $\theta^{\prime\prime}=0$ and $G(x,\theta)=x+\theta$.
-Therefore, only need to generate the sequence $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$.
-For $\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}$, we define
+Therefore, only need to generate $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$.
+For $\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}$, define
 $Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (4.15)
 We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (4.8), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
-Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms as we illustrate in Remark 4.5.
+Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark 4.5.
 For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
-$U_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j}+\theta_{0},\bar{\epsilon}_{t:t^{\prime}}^{j}+\theta_{1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}$
+$\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j}+\theta_{0},\bar{\epsilon}_{t:t^{\prime}}^{j}+\theta_{1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}$
 $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
@@
 where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime})=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime}),$ and $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ is the one such pair that maximizes $L_{i}\left(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)$, or equivalently minimizes
-$-2\log L_{i}\left(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)=$
-$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{ constant },\text{ if }i<t,\\ \text{ constant},\text{ if }i=t,\\ \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{ constant },\text{ if }i>t,\end{cases}$ (4.16)
+$-2\log L_{i}\Big{(}\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\dots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\Big{)}=$
+$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (4.16)
 over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
+Here, constant means that it does not involve $\theta,\theta^{\prime}$.
 Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (4.3) is a quadratic function of $\theta_{1}-\theta_{0}$.
@@
 […]
-$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\}$ (4.17)
-[…]
+$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (4.17)
+$\tau_{\text{wsr-ripr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (4.18)
 where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$.
@@
 Therefore,
-$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
-Similarly, it is easy to check that for $\mathcal{P}_{0}=\{N(\theta,1):\theta\geq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\leq b\}$, for some $b\leq a$,
-$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
+$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
+$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
+Similarly, one can check that for $\mathcal{P}_{0}=\{N(\theta,1):\theta\geq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\leq b\}$, for $b\leq a$,
+$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
+$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
 Remark 4.6.
@@
 Then,
-$\mathbb{P}_{\theta_{0},\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}_{\theta_{0},\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)<A\right]$
+$\mathbb{P}_{\theta_{0},\infty}[\tau_{\text{wcs-ripr}}\geq t]$
+$=\mathbb{P}_{\theta_{0},\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)<A\right]$
 $=\mathbb{P}_{\theta_{0},\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta}(X_{i})}{f_{a}(X_{i})}dW(\theta)<A\right]$
@@
 The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology.
-It is worth noting that $\mathcal{C}^{\prime\prime}$ in (4.12) is computable for parametric families if the following conditions are met:
+Note that $\mathcal{C}^{\prime\prime}$ in (4.12) is computable if the following conditions are met:
   1. 1.
     The coupling function $G$ is available (which is always the case when distributions are continuous).
+For discrete distributions, even if a suitable $G$ is not available, our method may still be implementable, as demonstrated in Section B.2 for the Poisson case).
   2. 2.
-    We can construct a confidence interval (and hence confidence sequence) for the parameter of interest.
+    We can construct a confidence interval (and hence confidence sequence) for $\theta$.
   3. 3.
@@
     If the detection algorithm $\mathcal{A}$ controls ARL (or is a black box), we need to satisfy Assumption 4.1. If $\mathcal{A}$ controls PFA, we do not need this assumption.
-For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions.
+For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions. Another concrete example with the Poisson distribution is in Appendix B.
 5 Experiments
-We now experimentally validate our algorithms for constructing confidence sets in order to evaluate their performance across various settings and demonstrate their practical utility.
+We validate our algorithms across various settings and demonstrate their practicality. Experiments below are for Gaussians; Poisson is handled in Appendix B.
 5.1 Experiments with detection algorithms that control ARL
 5.1.1 Known pre- and post-change (Setting I)
-We consider both Normal and Poisson mean-change scenarios having the true changepoint at $T=100,500$, using the CUSUM change detection method defined in (B.1) with thresholds at $A=1000,10000$ (which controls ARL at level $A$ (Tartakovsky et al.,, 2014)).
-After detection, we employ Algorithm 2 to construct confidence sets for $T$.
-Figures 5.1 and 5.2 provide visualizations of the confidence sets and point estimates (2.7) of the changepoints across $5$ runs.
+We consider Gaussian mean-change from $N(0,1)$ (pre-change) to $N(1,1)$ (post-change) having the true (unknown) changepoint at $T=100,500$.
+We employ the CUSUM change detection method (C.1) with thresholds at $A=1000,10000$ (which controls ARL at level $A$ [23]).
+After detection, we employ Algorithm 2 to construct confidence sets for $T$, with $B=N=100$, $\alpha=0.05$.
+Fig.
+5.1 visualizes the confidence sets and point estimates (2.7) across $5$ runs.
+Table 5.2 has the average size and the coverages across $100$ independent runs.
 Figure 5.1: Setting I (known pre- and post-change Gaussian distributions): The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$ for $T=100$ (top row) and $T=500$ (bottom row).
-Change is detected using the CUSUM detector (B.1) with $A=10000$ (top row) or $A=1000$ (bottom row).
-The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points.
-$B=N=100$, $\alpha=0.05,L=\infty$.
+Change is detected using the CUSUM detector (C.1) with $A=10000$ (top row) or $A=1000$ (bottom row).
+The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$.
 Results of $5$ independent simulations are shown.
-Figure 5.2: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois$(1)$ and the rest from Pois$(2)$, with $T=100$ (top row) or $T=500$ (bottom row).
-Change is detected using CUSUM detector (B.1) with $A=10000$ (top row) or $A=1000$ (bottom row).
-The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points.
-$B=N=100$, $\alpha=0.05,L=\infty$.
-Results of $5$ independent simulations are shown.
-We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.3.
-Table 5.3: Setting I: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (B.1) with threshold $A=1000$, while varying $T$.
-$T$ $F_{0}$ $F_{1}$
+Table 5.2: Known pre- and post-change (Setting I, finite ARL)
+$T$ A
 Conditional
@@
 Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 $N(0,1)$ $N(1,1)$ 0.96 0.94 12.34 2.85 13.97
-100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.95 17.45 3.98 15.72
-500 $N(0,1)$ $N(1,1)$ 0.95 0.86 12.57 2.62 13.22
-500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.83 18.72 3.96 16.14
+100 10000 0.96 0.96 12.21 2.97 17.63
+100 1000 0.96 0.94 12.34 2.85 13.97
+500 10000 0.95 0.95 12.88 2.59 17.26
+500 1000 0.95 0.90 12.57 2.62 13.22
 For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible.
@@
 5.1.2 Known pre-change and composite post-change (Setting II)
-We perform experiments for Gaussian mean-change scenarios having pre-change distribution as $N(0,1)$, and the true (unknown) post-change data distribution as $N(1,1)$, with the changepoint at $T=100,500.$ For detection, we employ a weighted CUSUM detector (3.16) with thresholds at $A=1000,10000$ (which controls ARL at $A$).
+We perform experiments for Gaussian mean-change scenarios having pre-change distribution as $N(0,1)$, and the true (unknown) post-change data distribution as $N(1,1)$, with the changepoint at $T=100,500.$ For detection, we employ a weighted CUSUM detector (defined in (3.15)) with thresholds at $A=1000$ (which is known to control ARL at level $A$).
 For weighted CUSUM, we employ a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.75\}$ and $\theta=0.9$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
 After detection, we construct confidence sets (as defined in (3.11)) for $T$.
+We report the average size and coverages across $100$ independent runs in Table 5.4.
 Fig.
@@
 Results of $5$ independent simulations are shown.
-We report the average size and the conditional and unconditional coverage across $100$ independent runs in (5.5).
-For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible.
-However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage.
-However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.3).
-We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
-We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
-Table 5.5: Setting II: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.11) with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ (3.5) and average detection delay (given no false detection) of the change detection method used (3.16) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and composite post-change model.
+Table 5.4: Known pre-change but unknown post-change (Setting II, finite ARL)
 T
@@
 Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 23.25 3.95 16.87
-100 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 20.11 3.67 16.21
-500 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.87 30.83 3.45 17.63
-500 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.86 28.56 3.48 15.81
+100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 26.25 3.95 16.87
+100 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 24.11 3.67 16.21
+500 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.89 30.83 3.45 17.63
+500 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.91 28.56 3.48 15.81
+For $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in much lower unconditional coverage.
+However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments).
+It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.2).
+We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
+We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
 5.1.3 Composite pre- and post-change (Setting III)
 Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, with the changepoint at $T=100$ and $T=500$.
-For change detection, we employ a weighted CUSUM-type detector (defined in (4.17)), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in setting II is considered.
+For change detection, we employ a weighted CUSUM-type detector (4.17), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in setting II is considered.
 We set the thresholds at $A=1000,10000$.
-This is a special case of “e-detector” and is known to control ARL at level $A$ (Shin et al.,, 2023).
-After detection, we construct confidence sets (as defined in (4.12)) for $T$ (assuming $T\neq 1$).
-We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs in Table 5.6.
-Table 5.6: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ (4.6) and average detection delay (given no false detection) of the change detection method used (4.17) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and pre- and post-change model.
+This is a special case of “e-detector” and is known to control ARL at level $A$ [20].
+After detection, we construct confidence sets (4.12) for $T$ (assuming $T\neq 1$), with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$.
+Fig.
+5.5 visualizes the confidence sets (4.12) and point estimates (4.6) of the changepoint across five random runs.
+Figure 5.5: Setting III: The first $T-1$ observations are drawn from $N(0,1)$ and the remaining observations from $N(1,1)$ with $T=100$ (top row) or $T=500$ (bottom row).
+Change is detected using the CUSUM detector defined in (4.17) with $A=1000$.
+The point estimate (4.6) is shown in vertical red dashed line and the confidence set (3.11) is shown in red points.
+$\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.1\}$, $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, $N=B=100$, $\alpha=0.05,\beta=\gamma=0.025$ and $L=\infty$.
+Results of $5$ independent simulations are shown.
+Table 5.6 contains the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs.
+The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments).
+It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.2).
+We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table 5.2), exhibiting both higher coverage and larger size.
+Table 5.6: Composite pre- and post-change (Setting III, finite ARL)
 T
@@
 Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 25.51 4.36 25.81
-100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 22.97 4.19 23.15
-500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.95 0.95 30.74 4.03 24.48
-500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 29.16 4.07 23.07
-However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.3).
-We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table 5.3), exhibiting both higher coverage and larger size.
-Fig. 5.7 visualizes the confidence sets (4.12) and point estimates (4.6) of the changepoint across $5$ random runs.
-Figure 5.7: Setting III: The first $T-1$ observations are drawn from $N(0,1)$ and the remaining observations from $N(1,1)$ with $T=100$ (top row) or $T=500$ (bottom row).
-Change is detected using the CUSUM detector defined in (4.17) with $A=1000$.
-The point estimate (4.6) is shown in vertical red dashed line and the confidence set (3.11) is shown in red points.
-$\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.1\}$, $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, $N=B=100$, $\alpha=0.05,\beta=\gamma=0.025$ and $L=\infty$.
-Results of $5$ independent simulations are shown.
+100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 27.84 4.36 25.81
+100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 25.71 4.19 23.15
+500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 31.45 4.03 24.48
+500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 30.66 4.07 23.07
 5.2 Experiments with detection algorithms that control PFA
@@
 with thresholds at $A=1000$.
-Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality (Ville,, 1939), the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
-After detection, we employ Algorithm 1 to construct confidence sets for $T$.
+Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality [24], the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
+After detection, we employ Algorithm 1 to construct confidence sets for $T$, with $B=100$ and $\alpha=0.05$.
 We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.7.
 We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
-Table 5.8: Setting A: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 1 with $B=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (defined in (5.1)) with $A=1000$, while varying the changepoint $T$.
+Table 5.7: Known pre- and post-change distributions (Setting A, PFA $\leq 0.001$)
 $T$ $F_{0}$ $F_{1}$
@@
 After detection, we construct confidence sets for $T$ as defined in (3.12).
-We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.9.
-We also report the average absolute deviation of $\hat{T}$ (defined in (3.5)) and delay (given no false detection) in the table.
-Table 5.9: Setting B: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.12) with $B=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation of $\hat{T}$ (defined in (3.5)) and average detection delay (given no false detection) of the change detector used (as defined in (5.2)) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and composite post-change model.
+We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.8, with $B=100$ and $\alpha=0.05,\beta=0.025,L=\infty$.
+We also report the average absolute deviation of $\hat{T}$ in (3.5) and delay (given no false detection) in the table.
+Table 5.8: Known pre-change but unknown post-change (Setting B, PFA $\leq 0.001$)
 T
@@
 5.2.3 Composite pre- and post-change (Setting C)
-We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) Larsson et al., (2025); Grünwald et al., (2024); Lardy et al., (2024) being the representative of the pre-change class) defined as
+We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) [13, 6, 12] being the representative of the pre-change class) defined as
 $\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (5.3)
 with $f^{*}_{\theta}$ being the RIPr of $f_{\theta}$ on the pre-change class $\mathcal{P}_{0}$ and with thresholds at $A=1000$.
-This is a stop time of an ”e-process” (Ramdas et al.,, 2023) and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
+This is a stop time of an ”e-process” [17] and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
 The weights are chosen just as described in Section 5.1.3.
 After detection, we construct confidence sets for $T$ as defined in (4.13) (assuming $T\neq 1$).
-We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.10.
-We also report the average absolute deviation of $\hat{T}$ (defined in (4.6)) and delay (given no false detection) in the table.
-Table 5.10: Setting C: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $B=100$ and $\alpha=0.05,\beta=\gamma=0.025,L=\infty$, the average absolute deviation of $\hat{T}$ (defined in (4.6)) and average detection delay (given no false detection) of the change detector used (as defined in (5.3)) with threshold $A=1000$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(0,1)$ respectively, while we vary the changepoint $T$ and pre- and post-change models.
+The pre- and post-change data-generating distributions are $N(0,1)$ and $N(0,1)$ respectively, but neither of which is assumed known, and we vary the pre- and post-change classes.
+We use $B=100$ and $\alpha=0.05,\beta=\gamma=0.025,L=\infty$.
+We report the average size and coverage rates across $100$ independent runs in Table 5.9.
+We also report the average absolute deviation of $\hat{T}$ in (4.6) and delay (given no false detection).
+The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
+Table 5.9: Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
 T
@@
 6 Conclusion
-We developed a framework to construct a confidence set for the unknown changepoint $T$, providing conditional coverage guarantees (conditioned on the event that the stopping occurs after the changepoint) when the pre- and post-change distributions are known.
+We developed a framework to construct a confidence set for the unknown changepoint $T$ after a sequential detection algorithm declares a change at a stopping time, providing conditional coverage guarantees (conditioned on the event that the stopping occurs after the changepoint) when the pre- and post-change distributions are known.
 We also extended our framework to handle composite pre- and post-change distributions and provide conditional coverage guarantees, demonstrating its good empirical performance.
@@
 By filling this significant gap in post-detection analysis, our work provides crucial tools for precise and timely changepoint localization in real-time data streams.
-Our framework offers important implications for applications where rapid and accurate inference is critical, paving the way for more effective changepoint localization in sequential settings.
-While our method is designed for parametric pre- and post-change models in the partitioned setting, future work can explore its extensions to the non-partitioned setting and nonparametric models.
 Appendix A Omitted proofs
@@
 First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d. under $H_{0,t}$.
-Hence, $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})\iff M_{t}$ is one among the $\lceil(1-\alpha)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$.
-So, for any fixed $\alpha\in(0,1)$,
-$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-\alpha)(B+1)\rceil}{B+1}\geq 1-\alpha,\text{ for all }t.$ (A.1)
+[…]
+$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-\alpha)(B+1)\rceil}{B+1}\geq 1-\alpha.$ (A.1)
 For $L\in\mathbb{N}\cup\{\infty\},t\in\mathbb{N}$, define $C_{t,L}=\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B}))$. Note that
@@
 Proof of Proposition 2.5.
-Note that $\mathcal{C}^{*}=\{1\leq t\leq\tau:M_{t}\leq C_{t,L}\}\cup\{\tau+1,\tau+2,\cdots\}=\{t\geq 1:M_{t}\leq C_{t,L}\}$, since we defined $M_{t}=-\infty\leq C_{t,L}$, for all $t>\tau.$
+Note that $\mathcal{C}^{*}=\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\cup\{\tau+1,\tau+2,\cdots\}=\{t\geq 1:M_{t}\leq C_{t,L}\}$, since we defined $M_{t}=-\infty\leq C_{t,L}$, for $t>\tau.$
 Therefore, we have
-$\mathbb{P}_{T}(T\in\mathcal{C}^{*})=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})=\mathbb{P}_{T}(M_{T}\leq C_{t,L})\geq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})\geq 1-\alpha,$ (A.2)
+$\mathbb{P}_{T}(T\in\mathcal{C}^{*})=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})=\mathbb{P}_{T}(M_{T}\leq C_{t,L})\geq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})\geq 1-\alpha,$
 where the last inequality follows from (A.1). ∎
@@
 Note that $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}\}),$ for all $t\in\mathbb{N},\alpha\in(0,1)$.
-Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
-$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
+  1. (a)
+    For $\mathcal{C}_{1}$: For all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
+    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))$
+    $\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
     Then, we have
@@
     $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
-$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
-$\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B})\mid\hat{\alpha}_{t}\right)\right)+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
-$\leq\mathbb{E}(\hat{\alpha}_{t})+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
-$=\alpha\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)+\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)\leq(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
-Similarly, we have for fixed $\alpha\in(0,1),$ with $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverage $1-\beta$
-$P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta.$ (A.4)
+    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))+\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
+    $\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B})\mid r_{t}\right)\right)+\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
+    $\leq\mathbb{E}(\alpha r_{t})+\beta\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
+    $=(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
     We also have that
@@
     $=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\})$
-$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\},\tau<T)$
-$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
-Similarly, we have
-$P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$ (A.5)
+    $\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)$
+    $=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
     Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha r_{t}})\geq 1-(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta$ from the above inequality.
-If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\geq 1-\alpha-\beta$, we obtain
+  2. (b)
+    For $\mathcal{C}_{2}$: Here, we have $r_{t}=1$.
+The proof for $\mathcal{C}_{2}$ is similar.
+For fixed $\alpha\in(0,1),$ with $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverage $1-\beta$ we show that
+    $P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta.$
+    Then, we have
+    $P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
+    If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, we obtain
     $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha+\beta}{\mathbb{P}_{\theta_{0},\infty}(\tau=\infty)}\geq 1-\frac{\alpha+\beta}{1-\delta}.$
-Similarly,
+    And similarly,
     $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)-(\alpha+\beta)\geq 1-\alpha-\beta-\delta.$
@@
 $\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (A.3)
-Define, $\hat{\alpha}_{t}:=\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)$ and $C_{t,L}^{\alpha}=\operatorname{Quantile}(1-\alpha;M_{t},\{\max_{\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$.
+Define, $C_{t,L}^{\alpha}=\operatorname{Quantile}(1-\alpha;M_{t},\{\max_{\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$.
 Note that $\theta_{0}\in\mathcal{S}_{t-1}$ and $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}\}),$ for any $\alpha\in(0,1).$ By Assumption 4.1, we have that $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$.
-Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
-$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
+  1. (a)
+    For $\mathcal{C}_{1}$: Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
+    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))$
+    $\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
     And similarly,
-$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\in\mathcal{S}_{t-1})=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})\mid r_{t}))\geq\mathbb{E}(1-\gamma r_{t})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
+    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{o}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})\mid r_{t}))$
+    $\geq\mathbb{E}(1-\gamma r_{t})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
     Then, we have
-$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\hat{\alpha}_{t}})$
-$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\hat{\alpha}_{t}},\theta_{0}\in\mathcal{S}_{t-1},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\notin\mathcal{S}_{t-1}))+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
-$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))+(\beta+\gamma)\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$
-$\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})\mid\hat{\alpha}_{t}\right)\right)+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
-$\leq\mathbb{E}(\hat{\alpha}_{t})+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
+    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}})$
+    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}},\theta_{0}\in\mathcal{S}_{t-1},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\notin\mathcal{S}_{t-1}))+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
+    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))+(\beta+\gamma)\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$
+    $\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})\mid r_{t}\right)\right)+(\beta+\gamma)\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
+    $\leq\mathbb{E}(\alpha r_{t})+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
     $=\alpha\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
     $\leq(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t).$
-Similarly, we have for fixed $\alpha\in(0,1),$ with $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverages $1-\beta$ and $1-\gamma$ respectively,
-$P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta+\gamma.$ (A.7)
     We also have that
-$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
+    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
     the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
-$\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\hat{\alpha}_{t}})$
-$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\})$
-$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\hat{\alpha}_{t}}\},\tau<T)$
+    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\alpha r_{t}})$
+    $=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\})$
+    $=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)$
     $\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
     Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L})\geq 1-(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma$ from the above inequality.
+  2. (b)
+    For $\mathcal{C}_{2}$: Proof is similar to part (a).
+Here we have for fixed $r_{t}=1$ with $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverages $1-\beta$ and $1-\gamma$ respectively,
+    $P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta+\gamma.$
     Similarly, we have
-$P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$ (A.8)
+    $P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
     If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\geq 1-\alpha-\beta$, we obtain
@@
     $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)-(\alpha+\beta+\gamma)\geq 1-\alpha-\beta-\gamma-\delta.$
-Appendix B Simulations for point estimates
+Appendix B Another concrete example with Poisson distribution
+B.1 Known pre-change and composite post-change
+Here is $F_{0}$ is $\text{Pois}(\theta_{0})$ and the post-change model is $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\in\Theta_{1}\}$.
+In this setting, we have the following example of a $1-\beta$ confidence sequence [2] for $\theta_{1}$
+$\mathcal{S}^{\prime}_{n}=\{\theta:n\theta-\sum_{i=t}^{t+n-1}X_{i}\log\theta\leq\log\left(\frac{1}{\beta}\right)+\log I(c,c\theta)-\log I(n+c,\sum_{i=t}^{t+n-1}X_{i}+c\theta)\}\cap\Theta_{1},$ (B.1)
+[…]
+We fix $t\in\{1,\cdots,\tau\}$.
+Let $\lambda_{t}=\max\{\theta_{0},\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$.
+We only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Pois}(\lambda_{t})$ and $U_{n,1}^{j},\cdots,U_{n,X_{n}^{j}}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Uniform}(0,1)$, independently for each $n\in\mathbb{N},j=1,\cdots,B$.
+For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
+$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta_{0}/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ otherwise.
+}\end{cases}$ (B.2)
+Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta_{0})$, for $n<t$ and $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta)$, for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1},n\geq t$.
+We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (2.8), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
+Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$.
+Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.1) and they are finite.
+So, we can choose $L=\infty$.
+If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,V_{t,\infty}^{j}$ can be simplified as below:
+$V_{t,\infty}^{j}$ $=\max_{\begin{subarray}{c}i,t^{\prime}\in\mathbb{N}:1\leq i\leq t^{\prime},\\ \max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}\end{subarray}}\max_{\theta\in{S}^{j}_{t,t^{\prime}}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-t+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}.$
+Here ${S}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},t\leq n\leq t^{\prime},\inf\mathcal{S}^{\prime}_{\tau-t+1}\leq\lambda_{t}U_{n,i}^{j}\leq\sup\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\inf\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$ is the set consisting of one candidate element from is a finite set with at most $2+\sum_{n=t}^{t^{\prime}}X_{n}^{j}$ many elements.
+All the maximum are taken over a finite set; hence, $V_{t}^{j}$ can be computed easily.
+Remark B.1.
+For weighted CUSUM or SR detectors (defined in (3.15),(3.16)), finding $t_{1}^{j},t_{2}^{j}$ is often straightforward.
+For example, consider $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$.
+Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{(\tau-j+1)\times(\theta-\theta_{0})\}(\theta/\theta_{0})^{\sum_{i=j}^{\tau}X_{i}}$ increases as $X_{i}$’s increases and hence both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases.
+Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
+Similarly, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
+Also, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\neq\theta_{0}\}$, if we employ Gamma weights (it is a conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
+B.2 Composite pre- and post-change
+Here we assume the pre and post-change model to be $\mathcal{P}_{i}=\{\text{Pois}(\theta):\theta\in\Theta_{i}\}$, for $i=0,1$ respectively.
+In this setting, we have the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (B.1) and the same method can be employed for constructing $1-\gamma$ confidence interval for $\theta_{0}$ as well.
+We fix $t\in\{1,\cdots,\tau\}$.
+Let $\lambda_{t}=\max\{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$.
+We only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Pois}(\lambda_{t})$ and $U_{n,1}^{j},\cdots,U_{n,X_{n}^{j}}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Uniform}(0,1)$, independently for each $n\in\mathbb{N},j=1,\cdots,B$.
+For $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$,
+$Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise.
+}\end{cases}$ (B.3)
+Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta)$, for $n<t$, $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta^{\prime})$, for $n\geq t$ and $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
+We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (4.8), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
+For Poisson distribution as well, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly for the same detection algorithms, similar to the Gaussian setting discussed in Remark 4.5.
+If $t_{2}^{j}<t$, we have $U_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified as below:
+$U_{t,\infty}^{j}$ $=\max_{\begin{subarray}{c}i,t^{\prime}\in\mathbb{N}:1\leq i\leq t^{\prime},\\ \max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}\end{subarray}}\max_{\begin{subarray}{c}\theta\in{A}^{i,j}_{t,t^{\prime}}\\ \theta^{\prime}\in{B}^{i,j}_{t,t^{\prime}}\end{subarray}}\frac{L_{i}\left(\frac{\sum_{k=1}^{i-1}Y_{k}^{j}(\theta,\theta^{\prime})}{i-1},\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime})}{t^{\prime}-i+1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)}{L_{t}\left(\frac{\sum_{k=1}^{t-1}Y_{k}^{j}(\theta,\theta^{\prime})}{t-1},\frac{\sum_{k=t}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime})}{t^{\prime}-t+1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)}.$
+Here ${A}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},n\leq t-1,\inf\mathcal{S}_{t-1}\leq\lambda_{t}U_{n,k}^{j}\leq\sup\mathcal{S}_{t-1}\}\cup\{\inf\mathcal{S}_{t-1}\}\cup\{\sup\mathcal{S}_{t-1}\}$ and ${B}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},t\leq n\leq t^{\prime},\inf\mathcal{S}^{\prime}_{\tau-t+1}\leq\lambda_{t}U_{n,k}^{j}\leq\sup\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\inf\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$ are finite sets with at most $(2+\sum_{n=1}^{t^{\prime}}X_{n}^{j})$ many elements.
+All the maximum are taken over a finite set; hence, $U_{t}^{j}$ can be computed easily.
+B.3 Experimental results
+B.3.1 Known pre- and post-change (Setting I)
+We consider Poisson the mean-change problem: $\text{Pois}(1)$ (pre-change) vs $\text{Pois}(2)$ (post-change) with the true (unknown) changepoint at $T=100,500$.
+We employ the CUSUM change detection method defined in (C.1) with thresholds at $A=1000,10000$.
+After detection, we employ Algorithm 2 to construct confidence sets for $T$.
+5.1 and B.1 provide visualizations of the confidence sets and point estimates (2.7) of the changepoints across $5$ runs.
+Figure B.1: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois$(1)$ and the rest from Pois$(2)$, with $T=100$ (top row) or $T=500$ (bottom row).
+Change is detected using CUSUM detector (C.1) with $A=10000$ (top row) or $A=1000$ (bottom row).
+The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points.
+$B=N=100$, $\alpha=0.05,L=\infty$.
+Results of $5$ independent simulations are shown.
+We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table B.2.
+Table B.2: Setting I: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (C.1) with threshold $A$, while varying $T$ and $A$.
+$T$ $A$
+Conditional
+coverage
+Unconditional
+coverage
+Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
+100 10000 0.94 0.94 18.04 4.02 22.15
+100 1000 0.94 0.94 17.45 3.98 15.72
+500 10000 0.95 0.95 18.53 4.30 22.10
+500 1000 0.96 0.89 18.72 3.96 16.14
+For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible.
+However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage.
+Nevertheless, the empirical conditional coverage rates meet the desired level, supporting our theory.
+We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
+B.3.2 Known pre-change and composite post-change (Setting II)
+We perform experiments for Poisson rate-change scenarios having pre-change distribution as $\text{Pois}(1)$, and the true (unknown) post-change data distribution as $\text{Pois}(2)$, with the changepoint at $T=100,200,$ assuming the composite post-change model to be $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$.
+For detection, we employ a weighted CUSUM detector (3.15) with thresholds at $A=1000,10000$.
+For weighted CUSUM, we employ a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, and $\theta=1.9$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
+After detection, we construct confidence sets (as defined in (3.11)) for $T$.
+We report the average size and the conditional and unconditional coverage across $100$ independent runs in (5.4).
+The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments).
+It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table B.2).
+We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
+We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
+Table B.3: Setting II: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.11) with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ (3.5) and average detection delay (given no false detection) of the change detection method used (3.15) with threshold $A$, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
+T A
+Conditional
+coverage
+Unconditional
+coverage
+Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
+100 10000 0.97 0.97 33.13 6.11 27.68
+100 1000 0.97 0.96 33.85 5.59 19.14
+200 10000 0.98 0.98 36.08 5.98 26.81
+200 1000 0.97 0.93 35.24 5.74 21.27
+B.3.3 Composite pre- and post-change (Setting III)
+Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $200$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively.
+For change detection, we employ a weighted CUSUM-type detector (defined in (4.17)), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in Section B.3.2 is considered.
+After detection, we construct confidence sets (as defined in (4.12)) for $T$ (assuming $T\neq 1$).
+We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs in Table B.4.
+Table B.4: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ (4.6) and average detection delay (given no false detection) of the change detection method used (4.17) with threshold $A$, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
+T A
+Conditional
+coverage
+Unconditional
+coverage
+Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
+100 10000 0.97 0.97 35.04 6.34 31.56
+100 1000 0.97 0.97 35.19 6.27 26.33
+200 10000 0.98 0.98 37.54 5.92 30.45
+200 1000 0.98 0.97 37.86 6.18 25.82
+However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments).
+It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table B.2).
+We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table B.2), exhibiting both higher coverage and larger size.
+Appendix C Simulations for point estimates
 In this section, we study the empirical performances of our point estimates of the changepoint defined in Equations 2.7, 3.5 and 4.6, which are the sequential analogs of the maximum likelihood estimators in the offline setting.
 Our empirical findings indicate that these estimators perform well in practice, even though we do not have theoretical guarantees on their properties.
-B.1 Known pre- and post-change distributions
-In this experiment, $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$ For detecting the change, we use the usual cumulative sum (CUSUM) and Shiryaev-Roberts (SR) procedures Tartakovsky et al., (2014) as defined below:
-$\tau_{\text{cusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak$ (B.1)
-$\tau_{\text{sr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (B.2)
+C.1 Known pre- and post-change distributions
+In this experiment, $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$ For detecting the change, we use the usual cumulative sum (CUSUM) and Shiryaev-Roberts (SR) procedures [23] as defined below:
+$\tau_{\text{cusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak$ (C.1)
+$\tau_{\text{sr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (C.2)
 where $A$ is some fixed threshold.
@@
 Fig.
-B.1 shows histograms of the values taken by $\hat{T}$ over $10000$ independent experiments.
-Figure B.1: Histogram of $\hat{T}$, defined in (2.7), for known pre- and post-change over $10000$ independent experiments with $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$
-B.2 Known pre-change and composite post-change distributions
+C.1 shows histograms of the values taken by $\hat{T}$ over $10000$ independent experiments.
+Figure C.1: Histogram of $\hat{T}$, defined in (2.7), for known pre- and post-change over $10000$ independent experiments with $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$
+C.2 Known pre-change and composite post-change distributions
 In this experiment, $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta>0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
@@
 Fig.
-B.2 shows histograms of the values taken by $\hat{T}$ defined in (3.5) over $10000$ independent experiments.
-Figure B.2: Histogram of $\hat{T}$, defined in (3.5), for known pre-change and unknown post-change over $10000$ independent experiments with $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta\neq 0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
-B.3 Composite pre- and post-change distributions
+C.2 shows histograms of the values taken by $\hat{T}$ defined in (3.5) over $10000$ independent experiments.
+Figure C.2: Histogram of $\hat{T}$, defined in (3.5), for known pre-change and unknown post-change over $10000$ independent experiments with $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta\neq 0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
+C.3 Composite pre- and post-change distributions
 In this experiment, $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively.
@@
 In this experiment, we considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=0.6,d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ Fig.
-B.3 shows histograms of the values taken by $\hat{T}$ defined in (4.6) over $10000$ independent experiments.
+C.3 shows histograms of the values taken by $\hat{T}$ defined in (4.6) over $10000$ independent experiments.
 It is crucial to note that the RIPr-based estimator has a higher bias but lower variance.
-Figure B.3: Histogram of $\hat{T}$, defined in (4.6), for unknown pre- and post-change over $10000$ independent experiments with $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively.
-Appendix C A heuristic method for composite pre- or post-change
+Figure C.3: Histogram of $\hat{T}$, defined in (4.6), for unknown pre- and post-change over $10000$ independent experiments with $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively.
+Appendix D A heuristic method for composite pre- or post-change
 This section provides an alternative method for constructing confidence intervals for composite pre/post-change distributions.
@@
 Specifically, we propose a heuristic bootstrap-based approach that estimates the unknown distributions from the observed data and subsequently generates simulations from the estimated distributions.
-C.1 Known pre-change and unknown post-change distributions
+D.1 Known pre-change and unknown post-change distributions
 Here we assume that the pre-change distribution $F_{0}$ is known and the composite post-change model is $\mathcal{P}_{1}$.
@@
 For each sampled stream $j$, we compute the same test statistic $M_{t}^{j}$ under the null hypothesis, resulting in statistics $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if
-$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
+$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
 Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 5 contains an overview of the method.
@@
 14       end for
-15      if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
+15      if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
 16            $\mathcal{C}=\mathcal{C}\cup\{t\}$
@@
 Algorithm 5 CI for known pre-change $F_{0}$ and composite post-change $\mathcal{P}_{1}$
-C.2 Composite pre- and post-change distributions
+D.2 Composite pre- and post-change distributions
 We now consider the general case where both pre and post-change distributions follow some composite models $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively.
@@
 We reject $\mathcal{H}_{0,t}$ if we have
-$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
+$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
 Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 6 contains an overview of the method.
@@
 Algorithm 6 CI for unknown pre- and post-change
-C.3 Experimental Results
-C.3.1 Known pre-change and composite post-change (Setting II)
+D.3 Experimental Results
+D.3.1 Known pre-change and composite post-change (Setting II)
 We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$.
@@
 After detection, we employ Algorithm 5 to construct confidence sets.
-Table C.1 summarizes the results, including the average size and the conditional and unconditional coverage across $100$ independent runs.
+Table D.1 summarizes the results, including the average size and the conditional and unconditional coverage across $100$ independent runs.
 For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible.
@@
 Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
-Table C.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.16) with threshold $A=1000$.
+Table D.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.15) with threshold $A=1000$.
 T Pre-change
@@
 100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.95 20.95
-500 $N(0,1)$$\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.95 0.86 14.06
-500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.85 20.83
-C.3.2 Composite pre- and post-change (Setting III)
+500 $N(0,1)$$\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.95 0.91 14.06
+500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.89 20.83
+D.3.2 Composite pre- and post-change (Setting III)
 Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$.
@@
 After detection, we employ Algorithm 6 to construct confidence sets.
-We report the average size and the conditional and unconditional coverage across $100$ independent runs in Table C.2.
+We report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.2.
 The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
-Table C.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 6 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
+Table D.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 6 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
 T
@@
 500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 23.04
-Appendix D Another heuristic method for composite pre- or post-change
+Appendix E Another heuristic method for composite pre- or post-change
 In this section, we provide another heuristic method for dealing with composite pre- or post-change distributions.
@@
 Nonetheless, we do not have any theoretical guarantee for coverage for this method.
-For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section C.3.1 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.1.
-Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section C.3.2 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.1.
-From Table D.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
-Table D.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.16) with threshold $A=1000$.
+For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section D.3.1 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table E.1.
+Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section D.3.2 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table E.1.
+From Table E.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
+Table E.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.15) with threshold $A=1000$.
 T Pre-change
@@
 500 $N(0,1)$$\{N(\mu,1):\mu>0.75\}$ $N(1,1)$ 0.95 0.88 23.82
-500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.86 27.17
-Table D.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
+500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.90 27.17
+Table E.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
 T
@@
 While we do not have any theoretical guarantees for this method, it is computationally more efficient than the previous approaches with theoretical guarantees, as demonstrated in Table E.3.
-Appendix E Summary of comparisons of heuristic and non-heuristic methods
-We first compare the execution times of the three types of algorithms and present the results in Table E.1.
+Summary of comparisons of heuristic and non-heuristic methods
+We first compare the execution times of the three types of algorithms and present the results in Table E.3 for the Gaussian mean shift setting.
 Algorithm 3 and Algorithm 4 are collectively referred to as Proposed Method, which comes with theoretical guarantees but requires more computational time compared to the heuristic approaches.
@@
 Finally, the heuristic introduced in this section (which is Algorithm 2, with $F_{0}$ and $F_{1}$ being the “closest” pair of distributions in the composite pre- and post-change models) is labeled as Heuristic 2.
-Table E.1: Average (of $10$ independent runs) execution times (in seconds) taken by different methods, with $N=B=100$, $\alpha=0.05$, and for the proposed method, $\beta=\gamma=0.025$.
+Table E.3: Average (of $10$ independent runs) execution times (in seconds) taken by different methods, with $N=B=100$, $\alpha=0.05$, and for the proposed method, $\beta=\gamma=0.05$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(1,1)$ respectively.
 T A

### v3

@@
 This paper addresses a fundamental but largely unexplored challenge in sequential changepoint analysis: conducting inference following a detected change.
-We study the problem of localizing the changepoint using only the data observed up to a data-dependent stopping time at which a sequential detection algorithm $\mathcal{A}$ declares a change.
-We first construct confidence sets for the unknown changepoint when pre- and post-change distributions are assumed to be known.
-We then extend our framework to composite pre- and post-change scenarios.
-We impose no conditions on the observation space or on $\mathcal{A}$ — we only need to be able to run $\mathcal{A}$ on simulated data sequences.
-In summary, this work offers both theoretically sound and practically effective tools for sequential changepoint localization.
+We construct confidence sets for the unknown changepoint using only the data observed up to a data-dependent stopping time at which an arbitrary sequential detection algorithm $\mathcal{A}$ declares a change.
+Our framework is nonparametric, makes no assumption on the composite post-change class.
+We extend it to the composite pre-change class under a suitable assumption.
+We impose no conditions on the observation space or on $\mathcal{A}$.
+We thus present the first general tool for sequential changepoint localization, which is theoretically sound and practically effective.
 1 Introduction
-We consider the following problem of sequential change analysis in a general space $\mathcal{X}$: a sequence of $\mathcal{X}$-valued observations $X_{1},X_{2},\cdots$ arrives sequentially, such that for some unknown time $T\in\mathbb{N}\cup\{\infty\}$ (the changepoint),
-$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},$ (1.1)
-for some pre-change distribution $F_{0}\in\mathcal{P}$ and some post-change distribution $F_{1}\in\mathcal{P}$, where $\mathcal{P}$ denotes a class of probability distributions on $\mathcal{X}.$ We start by assuming $F_{0}$ and $F_{1}$ are known, and generalize our methods to encompass settings where $F_{0}$ and $F_{1}$ are only known to lie in some classes, but are not known exactly.
-A sequential changepoint detection algorithm $\mathcal{A}$ is simply a data-dependent stopping rule $\tau$ to raise an alarm that a change has previously occurred.
-Our aim in this paper is to construct confidence sets for the unknown changepoint $T$ after stopping at $\tau$, assuming only black-box access to $\mathcal{A}$ (and ideally no further assumptions or restrictions).
-In other words, we hope to obtain a set $\mathcal{C}\subseteq\{1,\dots,\tau\}$ such that
-$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha,$
-where $\mathbb{P}_{T}$ is the data distribution in (1.1), i.e.
-having the true changepoint at $T$.
-We will show for algorithms $\mathcal{A}$ with a bounded average run length (ARL), no set $\mathcal{C}$ can satisfy such a guarantee, but if $\mathcal{A}$ has probability of false alarm (PFA) bounded above by $\alpha$, such a guarantee is achievable.
-In either case, another natural criterion for $\mathcal{C}$ is for it to satisfy
-$\mathbb{P}_{T}(T\in\mathcal{C}|\tau\geq T)\geq 1-\alpha,$
+Consider the following general problem of sequential change analysis in an arbitrary space $\mathcal{X}$ (which could be the Euclidean space $\mathbb{R}^{d}$, or images, or anything else).
+A sequence of $\mathcal{X}$-valued observations ${\bf X}=X_{1},X_{2},\cdots$ arrives sequentially, such that for some unknown time ${}T\in\mathbb{N}\cup\{\infty\}$ (the changepoint),
+$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1.1)
+for some $F_{0}$ in a pre-change class $\mathcal{P}_{0}$ and $F_{1}$ in a post-change class $\mathcal{P}_{1}$, where $\mathcal{P}_{0}$ and $\mathcal{P}_{0}$ denote some non-intersecting classes of probability distributions.
+Initially, we shall assume all observations to be independent, although our methodology later extends to handling dependent data as well.
+Suppose the user has already decided to use a sequential changepoint detection algorithm $\mathcal{A}$ for this task, which is a mapping from the data sequence $\bf X$ to a stopping time [^1] While sequential change detection methods are commonly described solely by stopping times, we describe them by algorithms $\mathcal{A}$ because our method will feed in simulated (synthetically generated) data sequences $\bf X^{\prime}$ into $\mathcal{A}$; we find it easier to describe our methodology when making $\mathcal{A}$ explicit.
+$\tau=\mathcal{A}({\bf X})$ at which it raises an alarm that a change may have previously occurred.
+When $\mathcal{A}$ raises an alarm at time $\tau$, we do not know if it is a false alarm, and if it is not, then we only know that there has been some possible change before $\tau$, but not when that change occurred.
+Our aim here is to construct confidence sets for the unknown changepoint $T$ after stopping at $\tau$, assuming only black-box access to $\mathcal{A}$ (and ideally no further assumptions or restrictions).
+By “black-box”, we mean that we can run $\mathcal{A}$ on any input data sequences we like, but we do not require any further insight into its inner workings.
+The fact that we do not require knowledge of internal details of $\mathcal{A}$ is a particularly important feature; it implies that our changepoint localization method can wrap around any existing workflow for changepoint detection — it could be employed both with simple classical statistical algorithms (like CUSUM), and also with complex modern pipelines in the information technology industry that may involve heuristic machine learning algorithms (to deal with complex data types).
+As an initial goal, one may hope to obtain a set $\mathcal{C}\subseteq\{1,\dots,\tau\}$ such that
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha,$ (1.2)
+where $\mathbb{P}_{F_{0},T,F_{1}}$ is the data distribution in (1.1), i.e. having the true changepoint at $T$. But this goal is harder to achieve than may initally appear; we elaborate below.
+There are two classes of algorithms $\mathcal{A}$ in the literature: (i) algorithms with probability of false alarm (PFA) bounded by $\delta<1$; we call these “bounded PFA” for short, and (ii) algorithms with PFA equal to one, but have a bounded average run length to false alarm (ARL2FA, or just ARL for short); we call these “bounded ARL”.
+Since algorithms with bounded ARL have PFA equal to one, false detections occur with a nontrivial frequency.
+But the flipside is that bounded ARL algorithms typically have smaller detection delay than bounded PFA algorithms.
+Typically, detection delays of bounded ARL algorithms can be bounded by quantities independent of the changepoint $T$, but this guarantee is impossible to achieve with bounded PFA algorithms.
+In this paper, we remain agnostic to the choice of $\mathcal{A}$.
+For algorithms with a bounded ARL, we prove that no set $\mathcal{C}$ can satisfy guarantee (1.2) (for algorithms with bounded PFA, we show how to achieve (1.2) in the Supplementary).
+This impossibility result (which we prove formally later) intuitively holds because it is meaningless to ask for coverage if the algorithm stopped before the changepoint; indeed, on the event $\{\tau<T\}$, we have no information to rule out the possibility that $T=\infty$, and thus the only way to cover $T$ is to output the set of all natural numbers $\mathbb{N}$.
+Thus, an arguably more sensible goal is to have
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,$ (1.3)
 meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$.
-We show how to construct such sets $\mathcal{C}$ for both types of algorithms (bounded ARL or PFA).
-Sequential (also known as the online or quickest) changepoint detection is a well-studied field, where the goal is to detect that a change has occurred as soon as possible, is crucial in applications requiring timely responses to distributional shifts, such as quality control, cybersecurity, medical diagnostics, and so on.
-The field traces back to early work on statistical process control.
-Shewhart [19] introduced the concept of a control chart, which computes a test statistic for each sample, with values within a threshold indicating the process remains in control.
-Page [16] proposed the CUSUM (cumulative sum) procedure, which computes the cumulative log-likelihood ratio under the assumption of two exactly known distributions.
-Another influential approach is the Shiryaev-Roberts (SR) procedure [21, 18].
-These early methods, however, assume the pre- and post-change distributions to be fully specified.
-To relax the assumption of known post-change distributions, the Generalized Likelihood Ratio (GLR) procedure was introduced [15, 22].
-Tartakovsky et al.
-[23] provide comprehensive overviews of these classical methods, including CUSUM, SR, and GLR, and their numerous variants.
-Several nonparametric change-point detection algorithms have been developed, such as kernel-based methods [7, 14] and e-detectors [20].
-Localizing the changepoint is another fundamental and crucial problem.
-As far as we are aware, the existing approaches to changepoint localization focus on offline or batch settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times.
-Hinkley [8] shows some asymptotic results about the maximum likelihood estimator (MLE) of the changepoint and also provides a confidence set based on likelihood ratio statistics.
-Yao [26] provides an approximation of the distribution of the MLE of the changepoint.
-Darkhovskh [3] presents a nonparametric changepoint estimator based on Mann-Whitney statistics.
-Carlstein [1] proposes a class of strongly consistent estimators for the changepoint in the nonparametric setting.
-Worsley [25] develops likelihood ratio-based confidence regions and tests for a changepoint for exponential family random variables.
-Dümbgen [5] constructs nonparametric confidence regions for changepoint by testing for each candidate time-point $t$ at level $\alpha$ using bootstrapping, achieving asymptotic coverage of $1-\alpha$, as the sample size $n\to\infty$.
-A more recent method [10] constructs confidence sets by evaluating a local test statistic on a triplet of time points and combining the results of the local tests with a weighted Bonferroni correction.
-To re-emphasize, all these methods are for the offline setting, while our main focus is on the sequential problem of constructing confidence sets after stopping, which has been largely overlooked in the literature.
-This gap is significant because in most real-world applications, data arrives sequentially, and making timely inferences about the changepoint immediately upon detecting a change is essential.
-Thus, developing methodologies for post-detection inference in a sequential framework is a significant step forward in changepoint analysis.
-The key idea behind constructing the confidence interval is simple: for each candidate $t\in\{1,\cdots,\tau\}$, we will test whether $t$ is the true changepoint and form the confidence set by inverting these tests.
-Our approach is conceptually similar to [5], it adapts those ideas to the sequential setting in a nontrivial manner.
-For example, we show that for change detection algorithms that have finite average run length (ARL), it is impossible to obtain finite length confidence intervals with marginal coverage guarantees.
-Further, for $1-\alpha$ conditional coverage, we show that we need to test the $t$-th hypothesis at level $\alpha\times\mathbb{P}_{\infty}(\tau\geq t)$ (a quantity we do not know apriori), whereas testing for each $t$ at level $\alpha$ works in the offline setting.
-We develop simulation based approaches to get around these hurdles.
-There are two settings for the changepoint problem: non-partitioned and partitioned.
-The non-partitioned changepoint problem assumes that the data-generating distribution changes from one unknown distribution $F_{0}\in\mathcal{P}$ to another unknown distribution $F_{1}\in\mathcal{P}$, without requiring any pre-specified partitioning of $\mathcal{P}$ into pre- and post-change distribution classes (eg: detecting a change in a Gaussian mean from any value to any other value).
-In contrast, the partitioned setting assumes that we know some partition of $\mathcal{P}$ into $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ such that $F_{0}$ and $F_{1}$ are known to lie in $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively.
-This formulation is useful when changes within $\mathcal{P}_{0}$ are not of interest, and we only seek to detect shifts from $\mathcal{P}_{0}$ to $\mathcal{P}_{1}$ (eg: detecting a change in Gaussian mean from negative to positive).
-Our method is designed for the partitioned setting.
-Extensions to the non-partitioned case are left for future work.
-1.1 Our contributions
-Here we summarize the key contributions of the paper.
-  • For detection algorithms that control PFA, we propose Algorithm 1 for constructing a confidence set for the changepoint and establish its unconditional and conditional coverage (Theorem 2.2, Corollary 2.6), assuming known pre- and post-change distributions.
-  • For detection algorithms with bounded ARL, we show in Proposition 2.4 that we cannot have an unconditional coverage guarantee and propose Algorithm 2, which is a modification of Algorithm 1.
-We provide its conditional coverage guarantee in Theorem 2.7.
-  • We extend our framework to composite pre- and post-change classes in Algorithm 3 and Algorithm 4 and provide the coverage guarantees in Theorem 3.1 and Theorem 4.2.
-  • We demonstrate that the approach is not overly conservative (and thus practical) by examining the achieved coverage and size of constructed confidence sets in extensive simulations.
-We point out again that for obtaining the aforementioned guarantees, we only need to have black-box access to the detection algorithm $\mathcal{A}$.
-It need not have formal guarantees, we do not need to know its inner workings, and just need to be able to run $\mathcal{A}$ on any stream of data.
-1.2 Paper outline
-In Section 2, we construct our confidence sets for the changepoint for known pre- and post-change and study its unconditional and conditional coverage probabilities.
-Section 3 extend this approach to handle composite post-change, assuming the pre-change distribution to be known.
-In Section 4, we finally extend this approach to handle composite pre- and post-change setting.
-Section 5.1 and Section 5.2 present a comprehensive set of simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach.
+This definition has been considered by, for example, [5] and [39, 36], who achieved it (in a certain asymptotic sense) in much more restricted settings than the current investigation.
+The aforesaid impossibility result no longer applies to the new goal (1.3).
+It is important to note that we do not prescribe which $\mathcal{A}$ to use in which situation, but simply enable post-hoc inference on $T$ for any user-chosen $\mathcal{A}$.
+One can view our techniques as providing a wrapper to convert $\mathcal{A}$ from producing only a stopping time $\tau$ to also producing a confidence set for $T$  [^2] In some sense, our goal can be viewed in an analogous way to conformal inference, which wraps around an arbitrary prediction algorithm to convert a “point prediction” into a prediction set, by running the same algorithm on many hypothetical datasets (for full conformal prediction).
+However, our methods are not distribution-free, and rely on smartly using simulations to accomplish the goal..
+As we survey below, there is a rich literature on designing change detection algorithms $\mathcal{A}$, but a substantial paucity of inference methodology for localizing the change after detection; our focus is on the latter goal, assuming the user has chosen a suitable algorithm $\mathcal{A}$ for detection.
+While inference for $T$ is intrinsically interesting and important, a side benefit of confidence sets for $T$ is that one can use these to derive post-hoc confidence sets for unknown parameters of $F_{0}$ and $F_{1}$.
+To clarify, suppose $\mathcal{P}_{i}=\{F_{\theta}\}_{\theta\in\Theta_{i}};i=0,1$.
+Our method that delivers (1.3) also delivers confidence sets $\mathcal{C}^{\prime}\subseteq\Theta_{0}\times\Theta_{1}$ such that $\mathbb{P}_{\theta_{0},T,\theta_{1}}((\theta_{0},\theta_{1})\in\mathcal{C}^{\prime}\mid\tau\geq T)\geq 1-\alpha,\forall\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},$ where $\mathbb{P}_{\theta_{0},T,\theta_{1}}$ is the distribution of the sequence $X_{1},\dots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{0}}$, $X_{T},X_{T+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{1}}$.
+We stress that our confidence sets do not rely on asymptotics and are valid in finite samples.
+1.1 Related work
+Prior work on post-detection confidence sets for $T$.
+The problem of constructing confidence sets for the changepoint after stopping has been largely overlooked in the literature.
+This gap is significant because in most real-world applications, data arrives sequentially, and it can be essential to make timely inferences about the changepoint immediately upon detecting a change.
+[…]
+In sharp contrast, our proposed confidence sets for $T$ are very general: We make no assumptions on the detection procedure $\mathcal{A}$, and our framework acts as a versatile wrapper that can be layered on top of any sequential change detection algorithm.
+It is not restricted to exponential families — it works for any known $F_{0}$ and any nonparametric post-change class $\mathcal{P}_{1}$ on any general space (e.g., $\mathbb{R}^{d}$) and offers non-asymptotic coverage guarantees.
+It also extends to composite pre-change classes under some assumptions.
+Thus, our framework is the first nonparametric, theoretically sound, and widely applicable practical tool for sequential change localization tasks.
+We summarize the comparison in Table 1.1.
+Table 1.1: Comparison of existing methods and our proposed approach
+Pre-, Post- Existing Methods [5, 36] Our Proposed Method
+change distributions Assumptions Guarantee Assumptions Guarantee
+Known, Known $\mathcal{A}=$ CUSUM. $F_{0},F_{1}$ Asymptotic No assumption Nonasymptotic
+[…]
+Known, Unknown No existing work No assumption Nonasymptotic
+($F_{0},\mathcal{P}_{1}$) (any $\mathcal{A},F_{0},\mathcal{P}_{1}$) (Theorem 4.4)
+Unknown, Unknown No existing work Assumption 3.3 Nonasymptotic
+($\mathcal{P}_{0},\mathcal{P}_{1}$) on $\mathcal{A},\mathcal{P}_{0}$ (any $\mathcal{P}_{1}$) (Theorem 4.5)
+Literature on sequential change detection.
+In sequential change detection, the goal is to detect that a change has occurred as soon as possible, but not to provide inferential tools for when the change occurred.
+This methodology has proved crucial in many applications, such as quality control, cybersecurity, medical diagnostics, and so on.
+Classical methods such as Shewhart control charts [25], the CUSUM procedure [18], and the Shiryaev-Roberts approach [27, 22] assume known pre- and post-change distributions.
+To address scenarios with unknown post-change distributions, generalized likelihood ratio (GLR) procedures [17, 28] were developed.
+Comprehensive reviews of these methods are provided by [30].
+More recently, kernel-based techniques [9], e-detectors [26], etc., have emerged to handle nonparametric settings.
+Offline vs.
+online change localization.
+As far as we are aware, most existing approaches to changepoint localization focus on offline settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times.
+Classical works, e.g., [10] study the asymptotic behavior of the MLE for the changepoint and construct likelihood-ratio-based confidence sets.
+[3] proposed some nonparametric estimators.
+[35] developed likelihood ratio-based confidence sets for exponential family distributions, while [6] used bootstrap-based tests for each candidate time to construct nonparametric confidence sets with asymptotic coverage.
+Recently, [31] studies optimal rates for both changepoint detection and localization.
+[13] constructs confidence sets by evaluating a local test statistic.
+To re-emphasize, all these methods are for the offline setting.
+In the sequential setting, [29, 7, 37, 1] develop point estimators for $T$; but very few works construct confidence sets [5, 36], whose drawbacks were discussed earlier.
+Inference on $\theta_{0},\theta_{1}$ after stopping.
+As with the problem of post-detection inference on the changepoint $T$, there is limited literature on valid estimation or inference for the pre- and post-change parameters $\theta_{0}$ and $\theta_{1}$ post-stopping.
+[39] constructs asymptotic bias-corrected point estimates and confidence intervals for the post-change mean after a mean change is detected by a CUSUM procedure.
+[38] generalizes these results for the case when there is a possible change in the variance as well.
+These methods are also discussed in [36].
+However, these approaches share several limitations with the existing post-detection changepoint inference methods, as mentioned earlier: they rely on asymptotics, assume specific parametric forms, and apply only to the CUSUM procedure.
+1.2 Contributions and paper outline
+  • We propose a general framework for constructing a confidence set for the changepoint based on inverting tests obtained using a universal threshold, assuming all observations to be independent.
+We show that it provides $1-\alpha$ conditional coverage for known pre-change and arbitrary (nonparametric) post-change classes under no additional assumption in Theorem 3.2.
+We also extend it for composite pre-change under a suitable assumption in Theorem 3.5.
+  • We propose an alternative method of constructing the tests using adaptive thresholds (which is often tighter than the universal threshold), assuming pre- and post-change distributions to be known, in Algorithm 1.
+We extend this framework to handle parametric pre- and/or post-change classes in Algorithms 2 and 3.
+These adaptive methods allow the observations to be dependent.
+We establish corresponding coverage guarantees in Theorems 4.1, 4.4 and 4.5.
+  • As a byproduct of our methods, we derive confidence sets for the unknown pre- and post-change parameters and show their coverage guarantees in Theorem 5.1.
+In short, we present the first nonparametric, theoretically grounded, and widely applicable post-detection confidence sets that are agnostic to the detection method.
+The rest of the paper is organized as follows.
+In Section 2, we establish some fundamental results showing the impossibility of unconditional coverage and linking conditional confidence sets to hypothesis tests.
+Section 3 develops a general framework for constructing a confidence set for $T$ based on inverting tests obtained using a universal threshold.
+In Section 4, we introduce a simulation-based approach for constructing confidence sets using an adaptive threshold in the parametric setting and study their conditional coverage probabilities.
+In Section 5, we derive confidence sets for the changepoint, as well as the pre- and post-change parameters.
+Section 6 presents an extensive simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach.
 This article is concluded in Section 7.
-Omitted proofs can be found in Appendix A.
-An additional illustrative example using the Poisson distribution is given in Appendix B.
-Appendix B.
-Some additional experimental results are provided in Appendix C and some heuristic algorithms are presented in Appendices D and E, which are computationally more efficient.
-A summary of comparisons of the heuristic and non-heuristic methods is provided in Appendix E, pointing at open avenues for further theoretical work.
-2 Known pre- and post-change distributions
-We now address the question of constructing confidence sets for the changepoint, assuming that our change detection procedure stopped after a changepoint.
-We start from the simplest case when both pre- and post-change distributions are known, addressing more general cases in later sections.
+All the mathematical proofs, implementation details of some algorithms in the paper, and additional experiments are presented in the Supplementary Material.
+2 Some Fundamental Results
+In this section, we highlight some key observations that shape our methodology of post-detection inference for the change-point $T$.
+We begin by clarifying some notational conventions that will be used throughout the paper.
 Notation.
-For $t\in\mathbb{N}$, let $\mathbb{P}_{t}$ denote the joint distribution of the sequence of independent random variables $X_{1},X_{2},\dots$ when there is a changepoint at time t, meaning that $X_{1},\dots,X_{t-1}$ are from $F_{0}$ (with density $f_{0}$) and $X_{t},X_{t+1},\dots$ are from $F_{1}$ (with density $f_{1}$).
-As corner cases, let $\mathbb{P}_{1}$ denotes the joint distribution where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}$ and $\mathbb{P}_{\infty}$ denotes the case where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}$.
-Let $\mathbb{E}_{t}$ denote the expectation under $\mathbb{P}_{t}$, for $t\in\mathbb{N}\cup\{\infty\}$.
-Suppose we have a data sequence $\{X_{n}\}\sim\mathbb{P}_{T}$ for some unknown $T\in\mathbb{N}\cup\{\infty\}$ and some sequential change detection algorithm $\mathcal{A}$ that raises an alarm at the stopping time $\tau$, which is finite for the observed sequence.
-In sequential change detection, there are typically two types of Type I error metrics.
-One is the average run length (ARL), which is defined as $\mathbb{E}_{\infty}[\tau]$, the expected time until a false alarm occurs under $T=\infty$ (i.e., no change).
-The other one is the probability of false alarm (PFA), that is, $\mathbb{P}_{\infty}(\tau<\infty)$.
-We provide solutions for algorithms $\mathcal{A}$ that control either metric.
-2.1 A first confidence set
-We will test whether there is a changepoint at $t$, for each $t\in\{1,\cdots,\tau\}$ and then construct the confidence set consisting of all time points $t$ for which we fail to reject.
-Formally, we test the following null hypothesis at each candidate time $t$:
-$H_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\text{and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}.$
-Consider some test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ such that for each $t\in\mathbb{N}$,
-$M_{t}:=\begin{cases}M_{t}^{(\tau)}(X_{1},\cdots,X_{\tau}),&\text{if }t\leq\tau<\infty,\\ -\infty&\text{if }t>\tau\text{ or }\tau=\infty,\end{cases}$ (2.1)
-where $M_{t}^{(n)}$ is some test statistic based on $n$ many samples, for each $t,n\in\mathbb{N},t\leq n$.
-While our theoretical guarantees will not depend on any particular choice of $M_{t}$, we provide a prototypical choice in a later subsection that we use for our experiments.
-To test $H_{0,t}$, we draw $B$ many independent streams of data under $H_{0,t}$.
-For the $j$th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at the stopping time $\tau_{j}^{\prime}$ using $\mathcal{A}$ or until time $L$, whichever occurs first.
-Then, we compute the same test statistics truncated at $L$ (which we discuss in detail in the next subsection), based on $\tau_{j}^{\prime},L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$, which is denoted as $M^{j}_{t,L}$.
-We reject $H_{0,t}$ if $M_{t}>\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, and the confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$, i.e,
-$\mathcal{C}=\Big{\{}t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\Big{\}}.$ (2.2)
-Algorithm 1 provides an overview of the method.
-Input: $\alpha\in(0,1)$, $B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, algorithm $\mathcal{A}$ and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
-Output: Confidence set $\mathcal{C}$ for the changepoint $T$
-1 Detect change at $\tau$ using $\mathcal{A}$; Set $\mathcal{C}=\varnothing$
-2 for $t=1,\cdots,\tau$ do
-3       Compute some test statistic $M_{t}$ (e.g., as defined in (2.9)) based on $X_{1},\cdots,X_{\tau}$
-4       for $j=1,\cdots,B$ do
-5             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$, or until time $L$, whichever happens first.
-6             Subroutine $\mathcal{L}$: Compute $M_{t,L}^{j}$ (defined in (2.3),(2.4)) based on $\tau_{j}^{\prime}$, $L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$
-7
-8       end for
-9      if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$ then
-10            $\mathcal{C}=\mathcal{C}\cup\{t\}$
-11       end if
-12
-13 end for
-Algorithm 1 CI for known pre-change $F_{0}$ and post-change $F_{1}$
-2.2 Subtleties regarding Subroutine $\mathcal{L}$
-The problem with $L=\infty$ is that the stopping time may never be reached for the $j$-th sampled sequence, i.e., we may have $\tau^{\prime}_{j}=\infty$.
-One can set $L=\infty$ if $\mathcal{A}$ (i.e., its stopping time) satisfies the condition $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely).
-This is a very mild assumption because any reasonable change detector will always eventually raise an alarm if there is a change.
-Indeed, all algorithms with finite ARL satisfy the condition, and indeed typically so do algorithms that control PFA.
-To avoid making even this mild assumption, one can simply pick $L<\infty$.
-If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual, but if $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it).
-Therefore, for $L<\infty$,
-$M^{j}_{t,L}:=\begin{cases}-\infty,&\text{if }t>\tau_{j}^{\prime},\\ M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),&\text{if }t\leq\tau_{j}^{\prime}\leq L,\\ \infty,&\text{if }t\leq\tau_{j}^{\prime}\text{ and }\tau_{j}^{\prime}>L.\\ \end{cases}$ (2.3)
-For the proof of the upcoming theorem, we still need a definition of $M^{j}_{t,\infty}$ in the case that $L=\infty$ and $\tau_{j}=\infty$. We define
-$M^{j}_{t,\infty}:=\begin{cases}M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),&\text{if }t\leq\tau_{j}^{\prime}<\infty,\\ -\infty,&\text{if }t>\tau_{j}^{\prime}\text{ or }\tau_{j}^{\prime}=\infty.\\ \end{cases}$ (2.4)
-The rationale behind setting $M^{j}_{t,\infty}=-\infty$ when $\tau^{\prime}_{j}=\infty$ is the following.
-If $\tau_{j}=\infty$, for all $j=1,\cdots,B$, then we want to reject $H_{0,t}$, which is achieved by setting $M^{j}_{t,\infty}=-\infty$.
-Remark 2.1.
-The value of $L$ does not need to be a fixed constant; it can be set adaptively depending on the observed data sequence, for instance, one can choose $L$ to be $2\tau$.
-The proof of coverage works by arguing validity when $L=\infty$, and then pointing out that any smaller choice of $L$ only leads to more conservatism.
-2.3 Prototypical choice of test statistic $M_{t}$
-In the offline setting, where data is observed up to some fixed time $n$, the likelihood of data is
-$L_{T}^{n}=\prod_{i=1}^{T-1}f_{0}(X_{i})\prod_{i=T}^{n}f_{1}(X_{i}),$ (2.5)
-and the maximum likelihood estimator of $T$ [8, 26] is given by
-$\hat{T}_{n}=\operatorname*{arg\,max}_{1\leq j\leq n}L_{j}^{n}=\operatorname*{arg\,max}_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ (2.6)
-In our sequential setting, where the data is observed up to a data-dependent stopping time $\tau$, we adapt this estimator by replacing $n$ with $\tau$ and define the estimator of the changepoint as:
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ (2.7)
-The prototypical test statistic that we employ in our simulations is the likelihood-ratio-based test statistic $M_{t}^{(n)}$, where for all $n\in\mathbb{N},t\leq n$,
-$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}^{n}}{L_{t}^{n}}.$ (2.8)
-Note that for all $t\in\mathbb{N}$, the test statistic $M_{t}$ for testing $H_{0,t}$ simplifies to
-$M_{t}=\begin{cases}L_{\hat{T}}^{\tau}/L_{t}^{\tau},&\text{ if }t\leq\tau<\infty\\ -\infty&\text{if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (2.9)
-2.4 Unconditional coverage when $\mathcal{A}$ controls PFA
-We now investigate the coverage properties and provide a bound on the coverage of the confidence set produced by Algorithm 1.
-Theorem 2.2.
-Let $\mathcal{C}$ be the confidence set for changepoint $T\in\mathbb{N}$ produced by Algorithm 1, for some $\alpha\in(0,1)$, $B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$ and change detection algorithm $\mathcal{A}$.
-Then,
-$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<T).$
-Therefore, if $\mathcal{A}$ controls PFA at level $\delta$,
-$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\delta.$
-Remark 2.3.
-When $T\ll$ ARL, we typically have $\mathbb{\mathbb{P}}_{\infty}(\tau<T)\approx 0$ and so $\mathbb{P}_{T}(T\in\mathcal{C})\gtrsim 1-\alpha$.
-The confidence set constructed using our method will approximately achieve the desired unconditional coverage level $1-\alpha,$ in situations where the probability of false detections before the true changepoint is negligible.
-However, $T$ is not known to us, and hence, we are unable to produce a suitable bound on the unconditional coverage for the detection algorithms that almost surely stop at a finite time, even when there is no change.
-2.5 Unconditional coverage when $\mathcal{A}$ has finite ARL
-If the detection algorithm has finite ARL (i.e., $\tau$ is finite almost surely even if there is no change), then it is impossible for any confidence set $\mathcal{C}^{*}$, which is a subset of $\{1,\cdots,\tau\}$, to produce an unconditional coverage guarantee at a predetermined level $1-\alpha$.
+For $t\in\mathbb{N}$, let $\mathbb{P}_{F_{0},t,F_{1}}$ denote joint distribution of the sequence $\{X_{n}\}_{n}$ when $X_{1},\dots,X_{t-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$ and $X_{t},X_{t+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{1}$.
+333 To understand the key ideas in the paper, it will suffice for the reader to focus on the case when all $X_{i}$’s are all independent, and $F_{0}$ and $F_{1}$ denote the pre- and post-change marginal distributions.
+However, our framework in Section 4 can actually handle dependent data.
+Under dependence, $F_{0}$ and $F_{1}$ denotes the joint distributions of $X_{1},\dots,X_{t-1}$ and $X_{t},X_{t+1},\dots$ respectively, and is in fact ambivalent to whether $X_{t}$ is drawn independently of $X_{1},\dots,X_{t-1}$ or conditional on it, so we do not emphasize the distinction in our notation.
+As a corner cases, let $\mathbb{P}_{F_{0},\infty}$ be the joint distribution when $X_{1},X_{2},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{0}$.
+In parametric settings, we write $\mathbb{P}_{\theta_{0},t,\theta_{1}}$ and $\mathbb{P}_{\theta_{0},\infty}$ as shorthand for $\mathbb{P}_{F_{\theta_{0}},t,F_{\theta_{1}}}$ and $\mathbb{P}_{F_{\theta_{0}},\infty}$ respectively.
+It is impossible for any set $\mathcal{C}^{*}\subseteq\{1,\cdots,\tau\}$, to produce an unconditional coverage guarantee at a predetermined level $1-\alpha$, if the detection algorithm has finite ARL (i.e., $\tau$ is finite almost surely even if there is no change).
 The intuitive reason is that if $\tau$ is finite almost surely, it is possibly a false detection, i.e., $\tau<T$, in which case the set $\{1,\cdots,\tau\}$ itself does not cover $T$.
@@
 Proposition 2.1.
-If a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies an unconditional coverage guarantee $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$, for some $\alpha\in(0,1)$, then $\mathbb{P}_{\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL of the algorithm is infinite).
-Said differently, for any algorithm with finite ARL, it is impossible to construct a valid confidence set that is only a subset of $\{1,\dots,\tau\}$.
-The proposition says that if we have a confidence set with high unconditional coverage, then the detection algorithm never stops with high probability if there is no change.
-Therefore, in order to achieve an unconditional coverage guarantee with $\mathcal{A}$ having finite ARL, we must extend our confidence set beyond the stopping time $\tau$.
-So, we consider the union of the confidence set constructed using Algorithm 1 and all integers greater than $\tau$.
-The following proposition establishes that this augmented confidence set achieves the desired coverage guarantee of $1-\alpha$.
+If, for some $\alpha\in(0,1)$, a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}^{*})\geq 1-\alpha$, $\forall T\in\mathbb{N}$, then $\mathbb{P}_{F_{0},\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL is infinite).
+Said differently, for any algorithm with finite ARL, it is impossible to construct a $1-\alpha$ confidence set that is only a subset of $\{1,\dots,\tau\}$.
+Thus, for post-detection inference, the usual notion of coverage is not an appropriate one.
+Instead, we focus on conditional coverage (1.3), a more reasonable criterion, which was also considered in prior work [5, 36].
+The key idea behind constructing our confidence intervals is simple, but several details are subtle.
+For each candidate $t\in\{1,\cdots,\tau\}$, we test whether $t$ is the true changepoint from some pre-change class $\mathcal{P}_{0}$ to some post-change class $\mathcal{P}_{1}$:
+$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (2.1)
+Let $\phi_{\alpha^{*}}^{(t)}=\mathds{1}(H_{0,t}\text{ is rejected})$ denote some test for the null $H_{0,t}$ such that
+$\mathbb{P}_{F_{0},t,F_{1}}(\phi_{\alpha^{*}}^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha^{*},\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$ (2.2)
+and form the confidence set by inverting these tests:
+$\mathcal{C}:=\{t\in\mathbb{N}:t\leq\tau,H_{0,t}\text{ is not rejected}\}$ (2.3)
+[…]
 Proposition 2.2.
-Let $\mathcal{C}$ be the confidence set for changepoint $T\in\mathbb{N}$ produced by Algorithm 1, for some $\alpha\in(0,1)$, $B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$ and change detection algorithm $\mathcal{A}$.
-Then, $\mathcal{C}^{*}:=\mathcal{C}\cup\{\tau+1,\tau+2,\cdots\}$ satisfies
-$\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha.$ (2.10)
-So, the confidence set constructed using Algorithm 1 can be interpreted as the pre-detection part of the augmented confidence set, which has the desired coverage guarantee of $1-\alpha$.
-2.6 Conditional coverage when $\mathcal{A}$ has finite ARL
-In the previous subsection, we established that the conventional notion of coverage is unsuitable for detection algorithms with finite ARL, where the confidence set is a subset of the random set $\{1,\cdots,\tau\}$.
-Intuitively, it is more appropriate to consider the conditional coverage, given that no false detection occurs, i.e., $T\leq\tau$.
-This adjustment is crucial because discussing coverage becomes meaningless when $\tau<T$, meaning that the detection occurs before the true changepoint — the target point we wish to cover by our confidence set.
-The following is an immediate corollary of Theorem 2.2, where we show a bound on the conditional coverage of the confidence set produced by Algorithm 1.
-Corollary 2.6.
-For any $\alpha\in(0,1)$, $B\in\mathbb{N}$ and $T\in\mathbb{N}$, and change detection algorithm $\mathcal{A}$ satisfying $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ produced by Algorithm 1 satisfies
-$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)}.$ (2.11)
-Therefore, if $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$,
-$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{1-\delta}.$ (2.12)
-The restriction that $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$ is very weak and clearly necessary for conditional coverage.
-Since $T$ is arbitrary, the condition means that there is no $S<\infty$ satisfying $\mathbb{P}_{\infty}(\tau<S)=1$.
-If the algorithm always stops under $\mathbb{P}_{\infty}$ before $S$, then it can never detect any change that occurs after $S$.
-This is because the information that distinguishes $\mathbb{P}_{\infty}$ from $\mathbb{P}_{T}$ for $T>S$ occurs after $S$, but the algorithm has halted before $S$.
-When $\tau$ is finite almost surely (i.e., PFA is $1$), we cannot have a positive lower bound on $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)$ uniformly over $T\in\mathbb{N}$.
-Therefore, Algorithm 1 does not guarantee conditional coverage at some prespecified level for detection algorithms with finite ARL (which implies that PFA is 1, although the converse is not true).
-Furthermore, if $\mathcal{A}$ is completely black-box — meaning its PFA is unknown — it becomes infeasible to construct a confidence set with the desired conditional or unconditional coverage using this approach.
-We propose a simple modification to address the issue in the following subsection, which applies to any black-box detection algorithm.
-2.7 Modified confidence set when $\mathcal{A}$ has finite ARL (or $\mathcal{A}$ is a blackbox)
-Equation 2.11 indicates that to ensure a conditional coverage guarantee of $1-\alpha$, we observe that we need to test $H_{0,T}$ at level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq T),$ rather than simply $\alpha$.
-To achieve that, for each $t\in\{1,\cdots,\tau\}$, we test $H_{0,t}$ at level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq t).$ For each $i\in\{1,\cdots,N\}$, we generate a sequence $\{Z_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$ until a change is detected, recording the detection times as $\tau_{i}$, which help to estimate the probability $\mathbb{P}_{\infty}(\tau\geq t)$, for $t=1,\cdots,\tau$.
-Therefore, it suffices to draw data points only up to $\min\{\tau,\tau_{j}\}$, for $j=1,\cdots,N.$
-Then, we draw $B$ many independent streams of data under $H_{0,t}$.
-For the $j$th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at the stopping time $\tau_{j}^{\prime}$, using the same change detection method $\mathcal{A}$.
-Then, we compute the test statistic $M_{t,L}^{j}$ as described in Section 2.2, based on $\tau_{j}^{\prime}\wedge L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$.
-Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$, i.e.,
-$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (2.13)
-where $\hat{\alpha}_{t}=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$. Algorithm 2 contains a structured overview of the method.
-Input: $\alpha\in(0,1)$, $N\in\mathbb{N},B\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
-Output: A confidence set $\mathcal{C}$ for the changepoint $T$
-1 $\mathcal{C}=\varnothing$
-2 for $i=1,\cdots,N$ do
-3       Simulate $\{Z_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until $\mathcal{A}$ detects a change at $\tau_{j}$, or until $\tau$, whichever happens first.
-4
-5 end for
-6for $t=1,\cdots,\tau$ do
-7       Compute some test statistic $M_{t}$ (e.g. (2.9)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
-8       for $j=1,\cdots,B$ do
-9             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{1}$, until $\mathcal{A}$ detects a change at $\tau_{j}^{\prime}$, or time $L$, whichever happens first.
-10             Subroutine $\mathcal{L}$: Compute $M_{t,L}^{j}$ (defined in (2.3),(2.4)) based on $\tau_{j}^{\prime}$, $L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$
-11
-12       end for
-13      Compute $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$
-14       if $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$ then
-15            $\mathcal{C}=\mathcal{C}\cup\{t\}$
-16       end if
-17
-18 end for
+Let $\tau$ be a stopping time and $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying (2.5).
+Then, the test defined by $\phi^{(t)}(X_{1},\cdots,X_{\tau})=\mathds{1}(t\notin\mathcal{C})$ satisfies
+$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall t\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.4)
+Conversely, given any test $\phi^{(t)}$ for $H_{0,t}$ using data up to $\tau$ satisfying (2.4), the inverted confidence set $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\leavevmode\nobreak\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.5)
+The above result establishes the duality between hypothesis tests and a post-detection confidence set.
+In the next section, we propose a general framework for constructing confidence sets for $T$ by inverting tests based on the above.
+3 A general framework for nonparametric post-change
+Let $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ denote the pre- and post-change classes respectively, $\mathcal{P}_{0}\cap\mathcal{P}_{1}=\emptyset$.
+Our goal is to construct post-detection confidence sets for the changepoint $T$ by inverting tests for $H_{0,t}$, as formalized in Proposition 2.2.
+To motivate the general construction, we begin with the simplest case when both pre- and post-change distributions $F_{0}$ and $F_{1}$ (with densities $f_{0}$ and $f_{1}$) are known.
+Then, the likelihood of the data $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}$, $X_{T},\cdots,X_{n}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}$ is given by $L_{T}^{n}:=\prod_{i=1}^{T-1}f_{0}(X_{i})\prod_{i=T}^{n}f_{1}(X_{i})$, and the maximum likelihood estimator of $T$ [10] can be written as $\hat{T}_{n}=\operatorname*{arg\,max}_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ In our setting, having data is observed up to a stopping time $\tau$, the analogous estimator is obtained by replacing $n$ with $\tau$ [37, 36]:
+$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ (3.1)
+Now, a prototypical likelihood-ratio-based test statistic is the following: $M_{t}=\displaystyle\max_{1\leq i\leq\tau}{L_{i}^{\tau}}/{L_{t}^{\tau}},\text{ if }t\leq\tau<\infty$ and $M_{t}=-\infty\text{ if }t>\tau\text{ or }\tau=\infty$.
+It simplifies to
+$M_{t}=\begin{cases}\prod_{i=t}^{\hat{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})},&\text{ if }t<\hat{T}\leq\tau<\infty\\ 1,&\text{ if }t=\hat{T}\leq\tau<\infty\\ \prod_{i=\hat{T}}^{t-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})},&\text{ if }\hat{T}<t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (3.2)
+One natural way to extend this test statistic for the general nonparametric setting is to replace the likelihood ratio processes by e-processes [19, 8], which have emerged as a central tool in anytime-valid inference.
+While e-processes have been used for sequential testing [33, 20], estimation [34], and change detection [26], we employ them for the first time for changepoint localization.
+For $t\in\mathbb{N}$, we define two different types of e-processes:
+  • Forward $t$-delay e-process under $\mathcal{P}_{1}$ (against $\mathcal{P}_{0}$).
+This is a stochastic process $\{R^{(t)}_{n}\}_{n\geq t}$ adapted to the forward filtration $\{\mathcal{F}^{(t)}_{n}\}_{n=t,t+1,\cdots}$, where $\mathcal{F}^{(t)}_{n}:=\sigma(X_{t},X_{t+1},\dots,X_{n}),$ such that for all stopping times $\kappa$ with respect to $\{\mathcal{F}^{(t)}_{n}\}_{n\geq t}$, $\sup_{F_{1}\in\mathcal{P}_{1}}\mathbb{E}_{F_{1}}[R^{(t)}_{\kappa}]\leq 1.$ Ideally, this process should have power against $\mathcal{P}_{0}$, i.e., $R^{(t)}_{n}$ should grow to infinity (as $n\to\infty$) under $\mathcal{P}_{1}$.
+  • Backward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$).
+This is a stochastic process $\{S^{(t)}_{n}\}_{n<t}$ adapted to the backward filtration $\{\mathcal{G}^{(t)}_{n}\}_{n=t-1,t-2,\cdots}$, where $\mathcal{G}^{(t)}_{n}:=\sigma(X_{n},X_{n+1},\cdots,X_{t-1}),$ such that for all stopping times $\kappa$ with respect to $\{\mathcal{G}^{(t)}_{n}\}_{n<t}$, $\sup_{F_{0}\in\mathcal{P}_{0}}\mathbb{E}_{F_{0}}[S^{(t)}_{\kappa}]\leq 1.$ Ideally, this process should have power against $\mathcal{P}_{1}$, i.e., $S^{(t)}_{n}$ should grow to infinity (as $n\to\infty$) under $\mathcal{P}_{0}$.
+Such powered e-processes for $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ (and vice versa) exist if the total variation distance between the two convex hulls of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ is positive (see e.g., Section 3.4 of [21]).
+We require all $X_{i}$’s to be independent for both powered (nontrivial) e-processes to be adapted to their respective filtrations.
+The final ingredient we need is $\hat{T}$, which is a preliminary point estimate of $T$.
+[^4] Here, we again note that our confidence set based on a point estimate $\hat{T}$ can be viewed analogously to the conformal prediction set, which is constructed around a preliminary “point prediction”.
+For example, one can construct such an estimator based on e-processes under $\mathcal{P}_{0}$.
+To elaborate, let $\{E^{(t)}_{n}\}_{n\geq t}$ be some forward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$).
+Then, one reasonable estimate is
+$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau}E^{(t)}_{\tau}.$ (3.3)
+This can be viewed as a nonparametric generalization of the maximum likelihood estimator for known pre- and post-change defined in (3.1).
+With the above ingredients in place, we construct a test statistic for $H_{0,t}$ by comparing it against a competing hypothesis $H_{0,T}$.
+Suppose $t<T$.
+Then under $H_{0,t}$ (which is false), the segment $X_{t},\dots,X_{T-1}$ follows a post-change distribution from $\mathcal{P}_{1}$ (because under $H_{0,t}$, the change occurred at $t$), whereas under $H_{0,T}$ (which is true), it follows a pre-change distribution from $\mathcal{P}_{0}$.
+Outside this segment $X_{t},\dots,X_{T-1}$, both hypotheses agree on the distribution of the observations, and hence those are non-informative for testing between $H_{0,t}$ and $H_{0,T}$.
+Therefore, we use a forward e-process $R^{(t)}_{T-1}$ that accumulates evidence in favor of $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ over this segment.
+Similarly, if $t>T$, we employ a backward e-process $S^{(t)}_{T}$ to accumulate evidence for $\mathcal{P}_{0}$ against $\mathcal{P}_{1}$.
+Since the true change-point $T$ is unknown, we replace it with an estimated change-point $\hat{T}$.
+This naturally leads us to define the following:
+$M_{t}=\begin{cases}R^{(t)}_{\hat{T}-1},&\text{ if }t<\hat{T}\leq\tau<\infty\\ 1,&\text{ if }t=\hat{T}\leq\tau<\infty\\ S^{(t)}_{\hat{T}},&\text{ if }\hat{T}<t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (3.4)
+[…]
+Theorem 3.1.
+$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha]\leq\alpha,$ for any $t\in\mathbb{N},F_{0}\in\mathcal{P}_{0},$ and $F_{1}\in\mathcal{P}_{1}.$
+While this bound may be conservative, it is remarkably universal: it holds without any assumptions on the distribution classes $\mathcal{P}_{0}$, $\mathcal{P}_{1}$, or on the detection method $\mathcal{A}$.
+For constructing a valid confidence set, Proposition 2.2 implies that we require tests of level $\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t)$, whose value is of course not known to us.
+To circumvent this issue, we first assume that the pre-change distribution $F_{0}$ is known.
+This is a common assumption in sequential changepoint literature.
+3.1 Confidence set for $T$ with known pre-change distribution
+Fix some $N\in\mathbb{N}$.
+For each $j=1,\dots,N$, simulate a sequence of i.i.d.
+data points $Z_{1}^{(j)},Z_{2}^{(j)},\dots\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}$ until either the detection algorithm triggers or the sample size reaches $\tau$—whichever occurs first.
+Let $\tau_{j}:=\min\{\tau,\mathcal{A}(\{Z_{n}^{(j)}\}_{n})\}$ denote the stopping time for the $j$-th simulated sequence.
+We then estimate $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$ using
+$r_{t}=\frac{1}{N}\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t).$ (3.5)
+Define the confidence set as
+$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}}\right\}.$ (3.6)
+The next theorem proves the validity of the above confidence set.
+Theorem 3.2.
+Suppose that the pre-change distribution is known, i.e., $\mathcal{P}_{0}=F_{0}$.
+Then, for any $\alpha\in(0,1),T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfying $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (3.6) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{1}\in\mathcal{P}_{1}.$
+Next, we extend it to composite pre-change under an additional assumption.
+3.2 A partial extension to composite pre-change
+To extend our method to composite pre-change classes $\mathcal{P}_{0}$, we require the following assumption on the change detection method and $\mathcal{P}_{0}$:
+Assumption 3.3.
+$\exists F_{0}^{*}\in\mathcal{P}_{0}$ such that $\inf_{F_{0}\in\mathcal{P}_{0}}\mathbb{P}_{F_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t)$, $\forall t\in\mathbb{N}$.
+This assumption posits the existence of a “least favorable” distribution $F_{0}^{*}$, in the sense that the detection algorithm is most prone to false alarms under $F_{0}^{*}$.
+We show that it holds for monotone likelihood ratio families with CUSUM/SR type change detectors when the pre- and post-change parameter spaces, $\Theta_{0}$ and $\Theta_{1}$, are ordered, e.g., $\Theta_{0}=(-\infty,a],\Theta_{1}=(b,\infty)$, for some $a<b$.
+[^5] For detection algorithms with bounded PFA, one can avoid Assumption 3.3 and obtain confidence sets with coverage guarantees without any additional assumption (i.e., for arbitrary $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$); see Section F of the Supplementary.
+Proposition 3.4.
+Suppose that $\{F_{\theta}:\theta\in\Theta\}$ is some monotone likelihood ratio family with respect to some sufficient statistic for the parameter $\theta$ and $\sup\Theta_{0}\leq\inf\Theta_{1}$.
+For the CUSUM type stopping time $\tau$ defined in (C.12), and for any $\theta\leq\theta^{\prime}\leq\sup\Theta_{0}$, $\mathbb{P}_{\theta,\infty}(\tau\geq t)\geq\mathbb{P}_{\theta^{\prime},\infty}(\tau\geq t),$ and hence, $\theta_{0}^{*}=\sup\Theta_{0}$ satisfies Assumption 3.3.
+Similarly, if $\inf\Theta_{0}\geq\sup\Theta_{1}$, then $\theta_{0}^{*}=\inf\Theta_{0}$ satisfies Assumption 3.3.
+For each $j=1,\dots,N$, we now simulate a sequence of i.i.d.
+data points $Z_{1}^{(j)},Z_{2}^{(j)},\dots\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}^{*}$ until either the detection algorithm triggers or the sample size reaches $\tau$—whichever occurs first.
+Let $\tau_{j}^{*}:=\min\{\tau,\mathcal{A}(\{Z_{n}^{(j)}\}_{n})\}$ denote the stopping time for $\{Z_{n}^{(j)}\}_{n}$.
+We estimate the probability $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$ using
+$r_{t}^{*}=\frac{1}{N}\sum_{j=1}^{N}\mathds{1}(\tau_{j}^{*}\geq t).$ (3.7)
+Define the confidence set as
+$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}^{*}}\right\}.$ (3.8)
+The next theorem proves the validity of the above confidence set.
+Theorem 3.5.
+Suppose that the pre-change class $\mathcal{P}_{0}$ and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfy Assumption 3.3 and $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,\forall F_{0}\in\mathcal{P}_{0}$.
+Then, for any $\alpha\in(0,1),$ $T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, the confidence set $\mathcal{C}$ in (3.8) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
+Next, we explore how, in parametric settings, one can obtain adaptive thresholds via simulation.
+Although such adaptive thresholds can sometimes be tighter, our simulations indicate that the resulting confidence sets from the universal threshold are not significantly wider, and are useful for their simplicity and generality.
+4 A simulation-based approach for parametric pre- and post-change settings
+In this section, we introduce an alternative approach that, while restricted to the parametric setting, allows observations to be dependent, and often yields tighter confidence sets than the method proposed in the previous section, as shown by our simulations later.
+We begin with the most basic case, where the pre- and post-change distributions are known, and then extend it to more general parametric settings subsequently.
+4.1 Known pre- and post-change distributions
+We develop a simulation-based scheme to approximate the exact $\alpha$-th quantile of the distribution of the test statistic $M_{t}$.
+Here, $M_{t}$ need not be of the form (3.4), it can be any test statistic based on $X_{1},\cdots,X_{\tau}$.
+Assuming known pre- and post-change distributions, we draw $B$ many independent streams of data under $H_{0,t}$.
+For the $j$-th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at $\tau_{j}^{\prime}=\mathcal{A}(\{X_{n}^{j}\}_{n})$ or until a fixed time $L$, whichever happens first.
+Then, we compute the test statistic $M_{t,L}^{j}$ as described in Section 4.1.1, based on $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$.
+Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$:
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (4.1)
+where $r_{t}$ is as defined in (3.7). Algorithm 1 contains an overview of the method.
 Algorithm 1 CI for known pre-change $F_{0}$ and post-change $F_{1}$
-By suitably adjusting the test levels for each $t$, this approach ensures that the desired conditional coverage guarantee is achieved.
-Notably, this guarantee does not depend on any specific property of $\mathcal{A}$ (e.g., its PFA or ARL).
-We prove it in the following theorem.
+[…]
+2:A confidence set $\mathcal{C}$ for the changepoint $T$
+[…]
+4:for $i=1$ to $N$ do
+5:  Draw $\{Z_{n}^{i}\}_{n}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$, until $\mathcal{A}$ detects at $\tau_{i}$, or until $\tau$, whichever happens first
+6:end for
+7:for $t=1$ to $\tau$ do
+8:  Compute test statistic $M_{t}$ (e.g. (3.2)) based on $\tau$ and $X_{1},\dots,X_{\tau}$
+9:  for $j=1$ to $B$ do
+10:   Simulate $X_{1}^{j},\dots,X_{t-1}^{j}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$; $X_{t}^{j},X_{t+1}^{j},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{1}$, until $\mathcal{A}$ detects a change at $\tau_{j}^{\prime}$, or time $L$, whichever happens first
+11:   Subroutine $\mathcal{L}$: Compute $M_{t,L}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\dots,X_{\min(\tau_{j}^{\prime},L)}^{j}$
+12:  end for
+13:  if $M_{t}\leq\operatorname{Quantile}\left(1-\frac{\alpha}{N}\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);\,M_{t},M_{t,L}^{1},\dots,M_{t,L}^{B}\right)$ then
+14:   $\mathcal{C}\leftarrow\mathcal{C}\cup\{t\}$
+15:  end if
+16:end for
 Theorem 4.1.
-For any fixed $\alpha\in(0,1)$, $B,N\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, $T\in\mathbb{N}$ and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (2.13) satisfies
-$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-{\alpha}.$ (2.14)
-Proof.
-First note the fact that $\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
-Step 1 [Relate the level of testing ${H}_{0,T}$ and the conditional coverage]: Defining $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, the level of testing ${H}_{0,T}$ is $P_{T}(M_{T}>C_{t,L}).$ Thus,
-$\mathbb{P}_{T}(M_{T}\leq C_{t,L})$
-$=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\})$
-$=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau\geq T)+\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau<T)$
-$\leq\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\mid\tau\geq T)\mathbb{P}_{T}(\tau\geq T)+\mathbb{\mathbb{P}}_{T}(\tau<T)$
-$=\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{\infty}(\tau\geq T),$
-where we twice invoked $\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T)$ as explained earlier. So, we obtain
-$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{t,L})}{P_{\infty}(\tau\geq T)}.$ (2.15)
-Step 2 [Establish that our test for each ${H}_{0,t}$ has level $\alpha\times\mathbb{P}_{\infty}(\tau\geq t)$]: First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d.
-under $H_{0,t}$.
-Hence, $M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-c)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$.
-So, for any fixed $c\in(0,1)$,
-$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c\text{ for all }t.$
-Since $\hat{\alpha}_{t}$ is independent of $M_{t}$, it follows that
-$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t})\geq\frac{\lceil(1-\hat{\alpha}_{t})(B+1)\rceil}{B+1}\geq 1-\hat{\alpha}_{t},\text{ almost surelyfor all }t.$
-Therefore, we have
-$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})=\mathbb{E}_{t}(\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t}))\geq\mathbb{E}_{t}(1-\hat{\alpha}_{t})=1-\alpha\mathbb{P}_{t}(\tau\geq t).$
-As noted earlier, $\mathbb{P}_{t}(\tau\geq t)=\mathbb{P}_{\infty}(\tau\geq t)$ and hence, $\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})\geq 1-\alpha\mathbb{P}_{\infty}(\tau\geq t).$ Also, for all $L\in\mathbb{N}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$.
-Therefore, $C_{t,L}\geq C_{t,\infty}$.
-Thus, for all $t$,
-$\mathbb{P}_{t}(M_{t}>C_{t,L})\leq\mathbb{P}_{t}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{\infty}(\tau\geq t)$ (2.16)
-Therefore, combining (2.15) and (2.16), we obtain $\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
+For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (4.1) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
 Remark 4.2.
-As mentioned before, the assumption $\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ is essential; otherwise, conditioning the coverage probability on $\{\tau\geq T\}$ becomes undefined.
-If this assumption does not hold, we have $\mathbb{P}_{\theta_{0},\infty}(\tau<T)=\mathbb{P}_{T}(\tau<T)=1$, implying that the detection is almost surely false.
-In such a scenario, constructing a confidence interval would not be meaningful.
+Notably, this method allows arbitrary (known) dependence among observations and is valid for any choice of $B,N,L$, test statistic $M_{t}$, and detection algorithm $\mathcal{A}$, which can function as an entirely black-box or heuristic method.
+The only necessity is a stopping rule that can be applied to generated data sequences.
+4.1.1 Subtleties regarding Subroutine $\mathcal{L}$
+The problem with $L=\infty$ is that the stopping time may never be reached for some sampled sequence.
+One can set $L=\infty$ if $\mathcal{A}$ is known to satisfy $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely).
+This is a very mild assumption because any reasonable change detector should eventually raise an alarm if there is a change; indeed, for example, this holds for all algorithms with finite ARL.
+If $\mathcal{A}$ is not known to satisfy this condition, one can simply pick $L<\infty$.
+If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual.
+If $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it).
+For $L<\infty$,
+$M^{j}_{t,L}:=\begin{cases}-\infty,&\text{if }t>\tau_{j}^{\prime},\\ M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),&\text{if }t\leq\tau_{j}^{\prime}\leq L,\\ \infty,&\text{if }t\leq\tau_{j}^{\prime}\text{ and }\tau_{j}^{\prime}>L.\\ \end{cases}$ (4.2)
+We still need a definition of $M^{j}_{t,\infty}$ in the case that $L=\infty$ and $\tau_{j}=\infty$.
+We define $M^{j}_{t,\infty}:=M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),\text{if }t\leq\tau_{j}^{\prime}<\infty,$ and $M^{j}_{t,\infty}:=-\infty,\text{if }t>\tau_{j}^{\prime}\text{ or }\tau_{j}^{\prime}=\infty.$
 Remark 4.3.
-We do not require any specific property of the sequential change detection algorithm; it can function as an entirely black-box or heuristic method.
-The only necessity is a stopping rule that can be applied to generated data sequences.
-Remark 2.10.
-For a fixed $t$, the sampled sequences $\{Z^{j}_{n}\}_{n}$ must be independent across all $j$ to ensure the validity of the above theorem.
-However, for testing the $t$-th hypothesis ($t\geq 3$), one can reuse the first $t-2$ data points from the sequences generated for testing the $(t-1)$-th hypothesis, thereby reducing computational redundancy.
-Nevertheless, in our experiments, we sampled the sequences independently across $t$.
-3 Known pre-change but unknown post-change distribution
-In this section, we assume that the post-change distribution follows some parametric composite model $\mathcal{P}_{1}=\{F_{\theta}:{\theta}\in\Theta_{1}\}$ and the pre-change distribution, $F_{\theta_{0}}(\theta_{0}\notin\Theta_{1})$ is known.
-Notation.
-For $t\in\mathbb{N}$, let $\mathbb{P}_{\theta,t,\theta^{\prime}}$ denote the joint distribution of the sequence of independent random variables $\{X_{n}\}_{n\in\mathbb{N}}$, when $X_{1},\dots,X_{t-1}$ are from $F_{\theta}$ (with density $f_{\theta}$) and $X_{t},X_{t+1},\dots$ are from $F_{\theta^{\prime}}$ (with density $f_{\theta^{\prime}}$).
-As corner cases, let $\mathbb{P}_{1,\theta^{\prime}}$ denotes the joint distribution where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta^{\prime}}$ and $\mathbb{P}_{\theta,\infty}$ denotes the case where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta}$.
-Let $\mathbb{E}_{\theta,t,\theta^{\prime}}$ denote the expectation under $\mathbb{P}_{\theta,t,\theta^{\prime}}$, for $t\in\mathbb{N}\cup\{\infty\}$, for $t\in\mathbb{N}\cup\{\infty\}$ and $\mathbb{E}_{\theta,\infty}$ denote the expectation under $\mathbb{P}_{\theta,\infty}$.
-Suppose we have the data sequence $\{X_{n}\}_{n}\sim\mathbb{P}_{\theta_{0},T,\theta_{1}}$, for some unknown $T\in\mathbb{N}\cup\{\infty\},\theta_{1}\in\Theta_{1}$ and some sequential change detection algorithm $\mathcal{A}$ that raises an alarm at the stopping time $\tau$ in the sequence.
-We define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
-$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{{\theta_{0}}},\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta_{1}},\text{ for some }{\theta_{1}}\in\Theta_{1}.$
-For testing the null hypothesis $H_{0,t}$, we consider some test statistic $M_{t}:=M_{t}^{(\tau)}(X_{1},\cdots,X_{\tau})$ based on data up to time $\tau,$ if $t\leq\tau<\infty$ and $M_{t}:=-\infty$, otherwise, for each $t\in\mathbb{N}$, where $M_{t}^{(n)}$ is some test statistic based on $n$ many data points, for $n,t\in\mathbb{N},t\leq n$.
-An example is provided in Section 3.1.
-Suppose the change detection algorithm, $\mathcal{A}$ has finite ARL or $\mathcal{A}$ is possibly a blackbox (i.e., its properties or inner workings are unknown to us).
-In that case, we generate a sequence $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$ until a change is detected, recording the detection times as $\tau_{i}$.
-These will help empirically estimate the probability $\mathbb{P}_{\infty}(\tau\geq t)$, for $t=1,\cdots,\tau$.
-Therefore, it suffices to draw data points only up to $\min\{\tau,\tau_{j}\}$, for $j=1,\cdots,N$.
-However, if $\mathcal{A}$ controls PFA at a known level $\delta\in(0,1)$, we can avoid this step.
-Since the post-change parameter is unknown, for each $t$, we estimate it based on $X_{t},\cdots,X_{\tau}$.
-However, note that the sample size used for estimate might be too small to have a reliable estimate.
-To address this issue, we would use interval estimates instead of point estimates so that we can hope to achieve a predetermined control over the error.
-Notably, conventional confidence intervals are invalid in this context, since $\tau$ depends on the data.
-Therefore, we rely on confidence sequences [4, 11, 9, 2].
-A confidence sequence having coverage $1-c$ is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform coverage guarantee: $\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage.
-Otherwise, we construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage.
-Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS has coverage probability $\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; we formalize this more explicitly in the proof of the Theorem 3.1.
-For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ having joint distribution as $\mathbb{P}_{\theta_{0},t,\theta}$.
+The value of $L$ does not need to be a fixed constant; it can be set adaptively depending on the observed data sequence, for instance, one can choose $L$ to be $2\tau$.
+4.2 Adaptive confidence set for known pre-change distribution and parametric post-change class
+We now assume that the post-change distribution follows some parametric composite model $\mathcal{P}_{1}=\{F_{\theta}:{\theta}\in\Theta_{1}\}$ and the pre-change distribution, $F_{\theta_{0}}(\theta_{0}\notin\Theta_{1})$ is known.
+[…]
+$\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ (4.3)
+We construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage, where $r_{t}$ is as defined in (3.5).
+Despite $r_{t}$ being random, it turns out that marginalizing over $r_{t}$ yields that the CS has coverage $\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; we formalize this more explicitly in the proof of the Theorem 4.4.
+Fix $B\in\mathbb{N}$.
+For each $j=1,\cdots,B$, suppose we have a sequence $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ $\sim\mathbb{P}_{\theta_{0},t,\theta}$.
 Note that for a fixed $t$ and $\theta$, the sequences must be independent across $j$.
@@
 sequences $\{X^{j}_{n}(\theta_{0},\theta)\}_{n}$ are allowed to be dependent across different values of $t$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$.
-A specific method for generating these sequences is detailed in Section 3.2.
-We draw the datapoints until a change is detected at the stopping time $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until time $L$, whichever occurs first.
+A specific method for generating these sequences is detailed in Section B of the Supplementary.
+We draw the data until a change is detected at $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until a fixed time $L$, whichever occurs first.
 Then, we compute the same test statistics truncated at $L$, based on $\tau_{j,t}^{\theta}$ and $X^{j}_{1}(\theta_{0},t,\theta),\cdots,X^{j}_{\tau_{j,t}^{\theta}\wedge L}(\theta_{0},t,\theta)$, which is denoted as $M^{j}_{t,L}(\theta)$.
 For $L<\infty$,
-$M^{j}_{t,L}(\theta):=\begin{cases}-\infty,&\text{if }t>\tau_{j,t}^{\theta},\\ M_{t}^{(\tau_{j,t}^{\theta})}(X_{1}^{j}(\theta_{0},t,\theta),\cdots,X_{\tau_{j,t}^{\theta}}^{j}(\theta_{0},t,\theta)),&\text{if }t\leq\tau_{j,t}^{\theta}\leq L\\ \infty,&\text{if }t\leq\tau_{j,t}^{\theta}\text{ and }\tau_{j,t}^{\theta}>L.\end{cases}$ (3.1)
 […]
-$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\},$ (3.2)
-where $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$.
-If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$, we can have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
-$\mathcal{C}_{2}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha;M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\},$ (3.3)
-Algorithm 3 contains an overview of the method.
-We note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, one can approximate the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ by discretizing the parameter space into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum.
-Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient, and exact (i.e., does not require any approximation).
-We provide the details in Section 3.2.
-Additionally, we illustrate concrete examples of the Gaussian mean shift problem in Section 3.3 and the Poisson rate change problem in Appendix B.
-Input: $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N},L\in\mathbb{N}\cup\{\infty\},$ a confidence sequence procedure $\operatorname{CS(data;coverage)}$, a change detection algorithm $\mathcal{A}$, and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
-Output: Confidence sets $\mathcal{C}$ for changepoint $T$
-1 $r_{1}=r_{2}=\cdots=r_{\tau}=1$
-2 if $\mathcal{A}$ has finite ARL then
-3       for $i=1,\cdots,N$ do
-4             Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first.
-5
-6       end for
-7      $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, for $t=1,\cdots,\tau$
-8 end if
 […]
-10 for $t=1,\cdots,\tau$ do
-11       Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$ (e.g., as defined in (3.7))
-12       Find a confidence sequence for $\theta_{1}$: $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$
-13       for $j=1,\cdots,B$ do
-14             Subroutine 2: for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}(X_{t},\cdots,X_{\tau})$ do
-15                   Get the sequence $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ following $\mathbb{P}_{\theta_{0},t,\theta}$ (independently for all $j$), until its detection time $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until time $L$, whichever happens first.
-16                   Compute $M_{t,L}^{j}(\theta)$ based on $L,\tau_{j,t}^{\theta}$ and $X^{j}_{1}(\theta_{0},t,\theta),\cdots,X^{j}_{\tau_{j,t}^{\theta}\wedge L}(\theta_{0},t,\theta)$
-17
-18             end for
-19
-20       end for
-21      Subroutine 3: if there exists $\text{there exists }\theta\in\mathcal{S}^{\prime}_{\tau-t+1}:M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ then
-22            $\mathcal{C}=\mathcal{C}\cup\{t\}$
-23       end if
-24
-25 end for
-26
-Algorithm 3 CI for known pre-change ($F_{\theta_{0}}$) and composite post-change
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}$ (4.4)
+Algorithm 2 summarizes the method.
+Note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ can be approximated by discretizing $\mathcal{S}^{\prime}_{\tau-t+1}$ into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum.
+While this approach is conceptually straightforward, it is computationally intensive; related experiments are in Section E of the Supplementary.
+Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient and does not require any approximation.
+Details are provided in Section B of the Supplementary, along with two concrete examples of the Gaussian and Poisson mean change problems.
+Algorithm 2 CI for known pre-change ($F_{\theta_{0}}$) and parametric post-change
+[…]
+2:Confidence sets $\mathcal{C}$ for changepoint $T$
+3:for $i=1,\cdots,N$ do
+4:  Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}}$, until $\mathcal{A}$ detects at $\tau_{i}$ or until $\tau$, whichever happens first.
+5:end for
+[…]
+7:for $t=1,\cdots,\tau$ do
+8:  Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$
+9:  Compute $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$
+10:  Find a confidence sequence for $\theta_{1}$: $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$
+11:  for $j=1,\cdots,B$ do
+12:   Subroutine 2:
+13:   for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}(X_{t},\cdots,X_{\tau})$ do
+14:     Get $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ following $\mathbb{P}_{\theta_{0},t,\theta}$ (independently for all $j$), until its detection time $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until time $L$, whichever happens first.
+15:     Compute $M_{t,L}^{j}(\theta)$ based on $\tau_{j,t}^{\theta}$, and $X^{j}_{1}(\theta_{0},t,\theta),\cdots,X^{j}_{\tau_{j,t}^{\theta}\wedge L}(\theta_{0},t,\theta)$
+16:   end for
+17:  end for
+18:  Subroutine 3:
+19:  if there exists $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ then
+20:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
+21:  end if
+22:end for
 Theorem 4.4.
-Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be the confidence sets for changepoint $T\in\mathbb{N}$ as defined in (3.2) and (3.3) respectively, for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$.
-Then for any given change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$,
-$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$
-If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, then for any $\theta_{1}\in\Theta_{1}$,
-$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\delta\leavevmode\nobreak\ \text{ , and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{1-\delta},$
-where the latter guarantee holds if $\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0.$
-3.1 Prototypical choice of test statistic $M_{t}$
-Let $L_{t}(\theta_{1};X_{1},\cdots,X_{n})$ be the likelihood of data $X_{1},\cdots,X_{n}$ under $\mathcal{H}_{0,t}$. Then, for $n\geq t$,
-$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (3.4)
-We define $\hat{\theta}_{1,j:n}$ to be a maximum likelihood estimator (MLE) of the post-change parameter $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$.
-The point estimate of $T$ can be naturally extended to:
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (3.5)
-The prototypical test statistic that we employ in our simulations is the likelihood-ratio-based test statistic $M_{t}^{(n)}$, where for $t,n\in\mathbb{N},t\leq n$,
-$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (3.6)
-Note that the test statistic $M_{t}$ for testing $\mathcal{H}_{0,t}$ reduces to
-$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (3.7)
-3.2 Computationally and statistically efficient subroutines
-For computational and statistical efficiency, we generate sequences from all $t\in\{1,\cdots,\tau\}$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ simultaneously using a single shared source of randomness.
-To achieve this, we assume that we have a fixed element $\theta^{\prime\prime}\in\Theta$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta$.
-Therefore, for each $j=1,\cdots,B$, we generate only one data sequence $\{X_{n}^{j}\}_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{\theta^{\prime\prime}}$, and then for each $t\in\{1,\cdots,\tau\}$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ (having joint distribution $\mathbb{P}_{\theta_{0},t,\theta}$) by applying the function $G(.,\theta_{0})$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta)$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
-$Y_{n}^{t}(x,\theta):=\begin{cases}G(x,\theta_{0}),&\text{if }n<t\\ G(x,\theta),&\text{ otherwise. }\end{cases}$ (3.8)
-Let us provide a few examples of the function $G$.
-If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}}(x))$ works.
-In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$.
-In the case of discrete distributions, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta)\}_{n}$ across all $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Section B.1 for the Poisson distribution.
-Let us now fix $t\in\{1,\cdots,\tau\}$.
-Computing the detection time $\tau_{j,t}^{\theta}$ for each $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether there exists $\text{there exists }\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set.
-However, it is often straightforward to compute $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$.
-Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient.
-For example, we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$.
-Again, we bound $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)$ by $V_{t,L}^{j}$, where for $L<\infty$,
-$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (3.9)
-And for $L=\infty$,
-$V_{t,\infty}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{otherwise,}\\ \end{cases}$ (3.10)
-which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section 3.3). We define the confidence set as
-$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (3.11)
-$\mathcal{C}^{\prime}_{2}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}.$ (3.12)
-We observe that $\mathcal{C}^{\prime}\supseteq\mathcal{C}$, where $\mathcal{C}$ is the set produced by Algorithm 3, since
-$\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$
-$\geq\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
-$\geq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}).$
-Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 3.1.
-Corollary 3.2.
-$\mathcal{C}^{\prime}_{1}$ and $\mathcal{C}^{\prime}_{2}$ defined in (3.11) and (3.12) respectively, satisfy the same coverage guarantees as $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ in Theorem 3.1.
-In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with Poisson distribution in Appendix B.
-3.3 A concrete example: Gaussian mean shift problem
-Let the pre-change distribution $F_{0}$ be $N(\theta_{0},1)$ and the post-change model be $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
-In this setting, we have the following example of a $1-\beta$ confidence sequence [9] for $\theta_{1}$
-$\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (3.13)
-where $\bar{X}_{t:\tau}=\frac{1}{\tau-t+1}\sum_{i=t}^{\tau}X_{i}$ and $s_{n}(\beta)=\sqrt{\log\log(2n)+0.72\log(10.4/\beta)}$.
-We fix $t\in\{1,\cdots,\tau\}$.
-Note that we can consider $\theta^{\prime\prime}=\theta_{0}$ and $G(x,\theta)=x+\theta-\theta_{0}$.
-Therefore, only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(\theta_{0},1)$ for $j=1,\cdots,B$.
-For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
-$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t,\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (3.14)
-where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$.
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (2.8), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
-Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$.
-Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark 3.3) and they are finite.
-So, we can have $L=\infty$.
-If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t$, it can be simplified in our setting as below:
-$V_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j}}{t^{\prime}-t+1}+\theta;Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}$
-$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
-$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta_{t^{\prime}}^{i});Y_{1}^{t}(X_{1}^{j},\theta_{t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta_{t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
-where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta)=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta),$ and $\theta_{t^{\prime}}^{i}$ is the unique element that maximizes $L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)$, or equivalently minimizes
-$-2\log L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)=$
-$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{(t-i)(\theta-\theta_{0})}{t^{\prime}-i+1}\right)^{2}+\text{constant},\text{ if }i<t,\\ \text{constant},\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{ constant},\text{ if }i>t,\end{cases}$
-over $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$.
-Here, constant refers to the terms that do not involve $\theta$.
-Note that $\theta_{t^{\prime}}^{i}$ has a simple closed-form expression since the above equation is a quadratic function of $\theta$.
-And hence, $V_{t,\infty}^{j}$ can be computed easily.
-Remark 3.3.
+Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (4.4) for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$.
+Then for any given detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$
+4.3 Adaptive confidence set for parametric pre- and post-change classes
+We now assume that both pre- and post-change distributions follow some parametric composite models: $\mathcal{P}_{0}=\{F_{\theta}:\theta\in\Theta_{0}\}$ and $\mathcal{P}_{1}=\{F_{\theta}:\theta\in\Theta_{1}\}$, $\Theta_{0}\cap\Theta_{1}=\emptyset$.
 […]
-$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (3.15)
-[…]
-For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$.
-Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{\tau}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases.
-Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
-Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
-Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
-One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section 5.2.
-The Gaussian mean shift example we provided serves as a concrete and illustrative case to clarify the methodology.
-It is worth noting that $C^{\prime}$, as defined in (3.11), is practically implementable as long as the following relatively weak conditions are met:
-  1. 1.
-    The coupling function $G$ is available (which is always the case when distributions are continuous and in the case of discrete distributions, even if a suitable $G$ is not directly available, our method may still be implementable, as demonstrated in Section B.1 for the Poisson distribution).
-  2. 2.
-    We can construct a confidence sequence for the parameter of interest (which is always achievable as long as one can construct a confidence interval for the parameter).
-  3. 3.
-    The quantities $t_{j}^{1}$ and $t_{j}^{2}$ are computable (either exactly or numerically).
-  4. 4.
-    If we employ the likelihood-ratio-based test statistic (2.8), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
-For instance, one can easily verify that this approach applies to Gaussian, Laplace, or exponential scale change problems.
-4 Composite pre- and post-change
-We now assume that both pre- and post-change distributions follow some parametric composite models $\mathcal{P}_{0}=\{F_{\theta}:\theta\in\Theta_{0}\}$ and $\mathcal{P}_{1}=\{F_{\theta}:\theta\in\Theta_{1}\}$ respectively such that $\Theta_{0}\cap\Theta_{1}=\emptyset$.
-In this setting, the average run length (ARL) of the detection algorithm is defined as $\inf_{\theta\in\Theta_{0}}\mathbb{E}_{\theta,\infty}[\tau]$, which represents the least expected time until a false alarm occurs under $T=\infty$ (i.e., no change).
-On the other hand, the probability of false alarm (PFA) is defined as the worst-case probability of triggering a false alarm, that is $\sup_{\theta\in\Theta_{0}}\mathbb{P}_{\theta,\infty}(\tau<\infty)$.
-Suppose we have the data sequence $\{X_{n}\}_{n}\sim\mathbb{P}_{\theta_{0},T,\theta_{1}}$, for some unknown $T\in\mathbb{N}\cup\{\infty\},\theta_{i}\in\Theta_{i}(i=0,1)$ and some sequential change detection algorithm $\mathcal{A}$ that raises an alarm at the stopping time $\tau$ in the sequence.
-Following the approach from the previous sections, we define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
-$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta_{0}},\text{ for some }\theta_{0}\in\Theta_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta_{1}},\text{ for some }{\theta_{1}}\in\Theta_{1}.$
-We consider some test statistic $M_{t}:=M_{t}(X_{1},\cdots,X_{\tau};\tau)$ based on data up to time $\tau,$ for testing the null hypothesis $\mathcal{H}_{0,t}$. An example is provided in Section 4.1.
-Suppose the change detection algorithm, $\mathcal{A}$ has finite ARL or it is a blackbox algorithm so that its PFA is unknown.
-In that case, we have a problem in generating i.i.d.
-copies of the detection times under $T=\infty$ (i.e., no change), since the pre-change parameter $\theta_{0}$ is unknown.
-Therefore, we need to assume the following:
-Assumption 4.1.
-There exists $\theta_{0}^{*}\in\Theta$ such that $\inf_{\theta\in\Theta_{0}}\mathbb{P}_{\theta,\infty}(\tau\geq t)\geq\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$, for all $t\in\mathbb{N}$.
-This assumption holds for a variety of change detection problems.
-For instance, we illustrate this in Remark 4.6 for some Gaussian mean change problems with CUSUM and SR-type detectors.
-One can easily check that it holds for analogous Gaussian, Laplace, and exponential scale change problems as well.
-We now generate $N$ many i.i.d.
-sequences under $F_{\theta_{0}^{*}}$ until a change is detected using $\mathcal{A}$, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$, that will help us estimate $\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$, which is a lower bound on $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$.
-If $\mathcal{A}$ controls PFA at a known level $\delta\in(0,1)$, we can avoid this step and still obtain conditional or unconditional coverage.
-Since both pre- and post-change parameters are unknown, for each $t$, we estimate them based on $X_{1},\cdots,X_{t-1}$ and $X_{t},\cdots,X_{\tau}$ respectively.
-Instead of point estimates, we rely on interval estimates in order to achieve a predetermined control over the error.
-Suppose that we have some confidence interval procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$, which constructs $1-c$ confidence interval using $X_{1},\cdots,X_{n}$ for $\theta_{0}$ and a confidence sequence procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, which constructs $1-c$ confidence sequence using $\{X_{n}\}_{n}$ for $\theta_{1}$.
-If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage and $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma)$ for $\theta_{0}$ having $1-\gamma$ coverage.
-Otherwise, we construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$.
-Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS and CI have coverage probabilities $1-\beta\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ and $1-\gamma\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$ respectively; we formalize this more explicitly in the proof of Theorem 4.2.
-Note that for $t=1$, the only plausible confidence interval of $\theta_{0}$ is $\mathcal{S}_{0}=\Theta_{0}$.
-Therefore, if $\Theta_{0}$ is unbounded, then it becomes infeasible.
-To avoid this issue, one can assume that $T\in\mathbb{N}\setminus\{1\}$ and test only for $t=2,\cdots,\tau.$
 For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ having joint distribution as $\mathbb{P}_{\theta,t,\theta^{\prime}}$.
@@
 However, for any fixed $j$, the i.i.d.
-sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ are allowed to be dependent across different values of $t$ and $\theta\in\mathcal{S}_{t-1},\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$.
-A specific method for generating these sequences is detailed in Section 4.2.
-We draw the datapoints until a change is detected at the stopping time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever occurs first.
+sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ can be dependent across $t$ and $\theta\in\mathcal{S}_{t-1},\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$.
+A specific method for generating these sequences is detailed in Section C of the Supplementary.
+We draw the data until a change is detected at the stopping time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever occurs first.
 Then, we compute the same test statistics truncated at $L$, based on $\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$, which is denoted as $M^{j}_{t,L}(\theta,\theta^{\prime})$.
-For $L<\infty$,
-$M^{j}_{t,L}(\theta,\theta^{\prime}):=\begin{cases}-\infty,&\text{if }t>\tau^{j,t}_{\theta,\theta^{\prime}},\\ M_{t}(X_{1}^{j}(\theta,t,\theta^{\prime}),\cdots,X_{\tau^{j,t}_{\theta,\theta^{\prime}}}^{j}(\theta,t,\theta^{\prime})),&\text{if }t\leq\tau^{j,t}_{\theta,\theta^{\prime}}\leq L,\\ \infty,&\text{if }t\leq\tau^{j,t}_{\theta,\theta^{\prime}}\text{ and }\tau^{j,t}_{\theta,\theta^{\prime}}>L.\\ \end{cases}$ (4.1)
-And for $L=\infty$,
-$M^{j}_{t,\infty}(\theta,\theta^{\prime}):=\begin{cases}M_{t}(X_{1}^{j}(\theta,t,\theta^{\prime}),\cdots,X_{\tau^{j,t}_{\theta,\theta^{\prime}}}^{j}(\theta,t,\theta^{\prime})),&\text{if }t\leq\tau^{j,t}_{\theta,\theta^{\prime}}<\infty,\\ -\infty,&\text{if }\tau^{j,t}_{\theta,\theta^{\prime}}=\infty\text{ or }t>\tau^{j,t}_{\theta,\theta^{\prime}}.\\ \end{cases}$ (4.2)
-Our confidence set is the set of $t\leq\tau$ for which we fail to reject $\mathcal{H}_{0,t}$ at level $\alpha\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$,
-$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\},$ (4.3)
-where $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$.
-If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$ , we have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
-$\mathcal{C}_{2}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}.$ (4.4)
+Similar to (4.4), the adaptive confidence set is defined as
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha}r_{t}^{*};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (4.5)
+where $\mathcal{\mathbf{S}}^{(t)}=\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
 Algorithm 3 contains an overview.
-As before, even if $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are uncountable, one can approximate the threshold $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ by discretizing the parameter space into a grid and evaluating the quantile for values of $(\theta,\theta^{\prime})$ on this grid, taking their maximum.
-Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient, and exact (i.e., does not require any approximation).
-We provide the details in Section 4.2.
-Additionally, we illustrate concrete examples of the Gaussian mean shift problem in Section 4.3 and the Poisson rate change problem in Appendix B.
-Input: $\theta_{0}^{*},\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N},L\in\mathbb{N}\cup\{\infty\}$, change detection algorithm $\mathcal{A}$, a confidence interval procedure $\operatorname{CI(data;coverage)},$ a confidence sequence procedure $\operatorname{CS(data;coverage)}$, and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
-Output: A confidence set $\mathcal{C}$ for changepoint $T$
-1 $r_{1}=r_{2}=\cdots=r_{\tau}=1$
-2 if $\mathcal{A}$ has finite ARL then
-3       for $i=1,\cdots,N$ do
-4             Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}^{*}}$, define $\tau_{i}$ as the time when $\mathcal{A}$ detects a change, or $\tau$, whichever is first
-5
-6       end for
-7      Define $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$ for $t=1,\dots,\tau$;
-8 end if
+If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 3.3 and provides coverage guarantees, as elaborated in Section F of the Supplementary.
+As before, one can always adopt a grid-based approach, which is conceptually straightforward, but it is computationally intensive; corresponding experimental results are reported in Section E of the Supplementary.
+Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient and exact, but that can be a bit more conservative.
+We provide the details in Section C of the Supplementary, along with two concrete examples.
+Algorithm 3 CI for parametric pre-change and post-change
 […]
-10
-11for $t=1,\cdots,\tau$ do
-12       Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$ (e.g., as defined in (4.8))
-13       Find a confidence interval for $\theta_{0}$: $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})$
-14       Find a confidence sequence or $\theta_{1}$: $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$
-15       for $j=1,\cdots,B$ do
-16             Subroutine 1: for $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ do
-17                   Simulate a sequence $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ following $\mathbb{P}_{\theta,t,\theta^{\prime}}$ (independently for all $j$) until its detection time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever happens first.
-18                   Compute $M_{t,L}^{j}(\theta,\theta^{\prime})$ based on $L,\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$
-19
-20             end for
-21
-22       end for
-23      Subroutine 2: if $M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ then
-24            $\mathcal{C}=\mathcal{C}\cup\{t\}$
-25       end if
-26
-27 end for
-Algorithm 4 CI for composite pre-change and post-change
+2:A confidence set $\mathcal{C}$ for changepoint $T$
+[…]
+4:for $i=1$ to $N$ do
+5:  Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F_{\theta_{0}^{*}}$, until $\mathcal{A}$ detects at $\tau_{i}^{*}$ or until $\tau$, whichever happens first.
+6:end for
+7:for $t=1$ to $\tau$ do
+8:  Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$
+9:  Compute $r_{t}^{*}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}^{*}\geq t)/N$
+10:  Find a confidence interval for $\theta_{0}$: $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})$
+11:  Find a confidence sequence for $\theta_{1}$: $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})$
+12:  for $j=1$ to $B$ do
+13:   Subroutine 1:
+14:   for each $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ do
+15:     Simulate $\{X_{n}^{j}(\theta,t,\theta^{\prime})\}_{n}$ following $\mathbb{P}_{\theta,t,\theta^{\prime}}$ (independently for all $j$) until its detection time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever happens first
+16:     Compute $M_{t,L}^{j}(\theta,\theta^{\prime})$ based on $\tau^{j,t}_{\theta,\theta^{\prime}}$, $X_{1}^{j}(\theta,t,\theta^{\prime}),\cdots,X_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}^{j}(\theta,t,\theta^{\prime})$
+17:   end for
+18:  end for
+19:  Subroutine 2:
+20:  if $M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha r_{t}^{*};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ then
+21:   $\mathcal{C}\leftarrow\mathcal{C}\cup\{t\}$
+22:  end if
+23:end for
 Theorem 4.5.
-Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.3) and (4.4), for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$.
-For any $\theta_{i}\in\Theta_{i}(i=0,1$), and any change detection algorithm $\mathcal{A}$ satisfying Assumption 4.1 and having $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$, the following conditional coverage guarantee holds:
-$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
-If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, we have that for any $\theta_{i}\in\Theta_{i}(i=0,1)$:
-$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\gamma-\delta,\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta+\gamma}{1-\delta}$
+Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.5), for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$.
+For any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1}$, and any change detection algorithm $\mathcal{A}$ satisfying Assumption 3.3 and having $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, the following guarantee holds: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
+Remark 4.6.
+As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets, $i=0,1$, we have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$.
+So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 4.4 and 4.5 reduces to Theorem 4.1.
+5 Confidence set for pre- and post-change parameters
+Beyond localizing the changepoint, another natural inferential goal is to estimate the pre- and post-change parameters, assuming that the pre- and post-change distributions follow some parametric composite models: $\mathcal{P}_{i}=\{F_{\theta}:\theta\in\Theta_{i}\};i=0,1$.
+While our primary focus is on constructing a confidence set $\mathcal{C}$ for the changepoint $T$, we now outline how one can derive confidence sets for $\theta_{0}$ and $\theta_{1}$ using $\mathcal{C}$.
+The key idea is as follows: once a valid confidence set $\mathcal{C}\subseteq\{1,\cdots,\tau\}$ for $T$ is constructed, this implicitly identifies a subset of data that is likely to have been generated under the pre- and post-change regimes.
+For estimating the post-change parameter, one natural idea might be to consider the samples starting from $\max\mathcal{C}$ till the detection time $\tau$.
+But since both starting and ending times are random here, this strategy does not yield valid coverage guarantees.
+Instead, we fix each $t\in\mathcal{C}$ and construct a confidence sequence based on $X_{t},\cdots,X_{\tau}$ for $\theta_{1}$, denoted as $\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})$ (recall the definition in (4.3)), for some $\eta_{1}\in(0,1)$, whose value is less than the target non-coverage, $\alpha^{\prime}$.
+Finally, the union of these CS across $t\in\mathcal{C}$, yields the post-detection confidence interval for $\theta_{1}$:
+$\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*}).$ (5.1)
+There is a natural tension between how quickly $\mathcal{A}$ detects a change from $\theta_{0}$ and the tightness of inference on $\theta_{1}$ — if $\mathcal{A}$ stops very quickly, we will have very few data points with which to estimate $\theta_{1}$, leading to wider intervals.
+This is not a drawback of our methodology, but an inherent limitation of any post-detection inference method.
+Analogously, for the pre-change parameter $\theta_{0}$, we fix each $t\in\mathcal{C}$ and construct a confidence interval for $\theta_{0}$ based on the data ${X_{1},\ldots,X_{t-1}}$, denoted as $\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})$, for some $\eta_{0}\in(0,1)$, chosen to be smaller than the target (conditional) non-coverage level, $\alpha^{\prime\prime}$.
+The final confidence set for $\theta_{0}$ is the following:
+$\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$ (5.2)
+We formalize its coverage guarantee in the next theorem, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods [39, 36].
+Theorem 5.1.
+Let $\mathcal{C}$ be a confidence set, e.g., (3.8) or (4.5), such that for any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},T\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ satisfying Assumption 3.3 and $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, we have: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha$.
+Then, for any $T\in\mathbb{N}$,
+$\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{0}\in\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})\mid\tau\geq T\right)\geq 1-\alpha-\eta_{0},\text{ and }$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\in\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})\mid\tau\geq T\right)\geq 1-\alpha-\eta_{1}.$
+6 Experiments
+We validate our methods across various settings and demonstrate their practicality.
+Known pre- and post-change (Setting I): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$, with $T$ being the only unknown parameter.
+We employ the CUSUM detector, $\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ with $A=1000$.
+After detection, we construct both the proposed confidence sets (universal (3.6) and adaptive (4.1)), with $B=N=100$, $\alpha=0.05$.
+Fig.
+6.1 visualizes the confidence sets (4.1) and point estimates (3.1) across 555 runs.
+Table 6.2 has the (conditional) average size and coverages of the confidence sets, the (conditional) average absolute deviation of $\hat{T}$, and delay of detection, i.e, $\tau-T$, across 500500500 runs.
+Figure 6.1: Setting I: The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$.
+The point estimates (3.1) are shown in a vertical red dashed line, and confidence sets (adaptive (4.1)) are shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$.
+Results of 555 independent simulations are shown.
+Table 6.2: Setting I: Known pre- and post-change.
+$T$ A Conditional Coverage Conditional Size
+$|\hat{T}-T|$
+Delay
+Universal Adaptive Universal Adaptive (Conditional) (Conditional)
+100 1000 0.98 0.95 15.63 12.34 2.85 13.97
+500 1000 0.98 0.95 15.77 12.57 2.62 13.22
+Known pre- and parametric post-change (Setting II): We perform experiments with $F_{0}=N(0,1)$ (known pre-change), post-change class $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$, ; we vary $\Theta_{1}$ and $T$.
+The post-change samples are drawn from $N(1,1)$ (unknown).
+For detection, we use a weighted CUSUM detector:
+$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (6.1)
+with $A=1000$, a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized mixture over $\mathcal{P}_{1}$.
+We use
+$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})},$ (6.2)
+where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$.
+After detection, we construct both the proposed confidence sets for $T$ (universal (3.6) with $\alpha=0.075$ and adaptive (4.4) [^6] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section B.1 (and Section C.1) of the Supplementary.
+with $\alpha=0.05,\beta=0.025$) with target conditional coverage 0.9250.9250.925.
+For implementing (3.6) , we use $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})},$ where ${\theta_{1}^{*}}$ is the closest element of $\Theta_{0}$ from $\Theta_{1}$, and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w$ being the same weight mentioned above.
+In Table 6.4, we report the (conditional) average size and coverages, absolute deviation of $\hat{T}$ , and detection delay $\tau-T$ across 500500500 runs.
+Fig.
+6.3 visualizes the confidence sets (3.6) and point estimates (B.2) across 555 random runs.
+Figure 6.3: Setting II: First $T-1$ samples are drawn from $N(0,1)$ and the rest from $N(1,1)$.
+The point estimates (B.2) are shown in a vertical red dashed line, and confidence sets (universal (3.6)) are shown in red points.
+$F_{0}=N(0,1)$, $\mathcal{P}_{1}=\{N(\mu,1):\mu\geq 0.9\}$, $N=100$, $\alpha=0.075$.
+Results of 555 independent simulations are shown.
+Table 6.4: Known pre-change but unknown post-change.
+The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column.
+(Setting II)
+$T$ $\Theta_{1}$ Conditional Coverage Conditional Size
+$|\hat{T}-T|$
+Delay
+Universal Adaptive Universal Adaptive (Conditional) (Conditional)
+100 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 22.21 25.72 3.95 16.87
+100 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.97 17.85 24.11 3.67 16.21
+500 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 22.89 30.83 3.45 17.63
+500 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.98 18.14 28.56 3.48 15.81
+Composite parametric pre- and post-change (Setting III): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$.
+Now we conduct experiments for pre-change class $\mathcal{P}_{0}=\{N(\theta,1):\theta\in\Theta_{0}\}$, post-change class $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$; we vary $\Theta_{0},\Theta_{1}$ and $T$.
+For detection, we employ a weighted CUSUM-type detector
+$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta^{*}_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (6.3)
+with $A=1000$, ${\theta^{*}_{0}}$ is the “closest” element of $\Theta_{0}$ from $\Theta_{1}$, and for the post-change parameter, the same discrete weight distribution as in setting II is considered.
+We define $\hat{\theta}_{0,1:j}$ to be the MLE of $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$.
+We use
+$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})},$ (6.4)
+where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$ and $\hat{\theta}_{0,:j-1}$ is the MLE of $\theta_{0}\in\Theta_{0}$ based on $X_{1},\cdots,X_{j-1}$.
+After detection, we construct both the proposed confidence sets (universal (3.8) with $\alpha=0.1$ and adaptive (4.5) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage 0.90.90.9.
+For implementing (3.8), we use $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})}w^{\prime}(\theta)d\theta,$ with $w^{\prime}(\theta)$ being the discrete distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta-d,\theta_{3}=\theta-2d,\cdots,\theta_{10}=\theta-9d$, $d=0.2$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}w(\theta)d\theta$, with $w(\theta)$ being the same as mentioned in Setting II.
+$\theta_{i}^{*}$ is the closest element of $\Theta_{i}$ from the other space.
+Fig.
+6.5 visualizes the confidence sets (3.8) and point estimates (C.2) across 5 random runs.
+Table 6.6 reports the results averaged across 500500500 runs.
+Figure 6.5: Setting III: First $T-1$ samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1)$.
+The point estimates (C.2) are shown in a vertical red dashed line.
+Confidence sets (universal (3.8)) are shown in red points.
+$\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.1\}$, $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, $N=100$, $\alpha=0.1$.
+Results of 555 random simulations are shown.
+Table 6.6: Confidence interval for $T$ (Setting III): unknown pre- and post-change.
+The true (unknown) values of the parameters are $\theta_{0}=0$, $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns.
+$T$ $\Theta_{0}$ $\Theta_{1}$ Conditional Coverage Conditional Size
+$|\hat{T}-T|$
+Delay
+Universal Adaptive Universal Adaptive (Cond.) (Cond.)
+100 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 26.91 26.84 4.36 25.81
+100 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.97 18.63 25.71 4.19 23.15
+500 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 26.12 31.45 4.03 24.48
+500 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.98 18.86 30.66 4.07 23.07
+Known pre- and nonparametric post-change (Setting IV): Here, the pre-change samples are from known $F_{0}=\text{U}[0,1]$ (uniform distribution on [0,1][0,1][ 0 , 1 ]) and let the post-change class be $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[X]\leq\mu\right\},$ where $\mathcal{M}$ is the set of all probability measures on $([0,1],\mathcal{B}([0,1]))$, with $\mu=0.25$.
+For detection, we use an e-detector [26] $\tau_{e}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}\geq A\right\},$ where $\hat{p}_{i-1}^{\prime}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=t}^{i-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},$ for $x\in[0,1]$, with $A=1000$.
+Post-detection, we construct the confidence sets (3.6).
+The forward $t$-delay e-process $\{R^{(t)}_{n}\}_{n\geq t}$, we use numeraire e-variable for $\mathcal{P}_{1}$ against $F_{0}$, as in Section 5.2 of [15] to get $R^{(t)}_{n}=\prod_{i=t}^{n}(1+\lambda^{*}(X_{i}-\mu)),$ where $\lambda^{*}\in(0,1/\mu)$ is the solution of the equation $e^{\lambda^{*}}=\frac{1+\lambda^{*}(1-\mu)}{1-\lambda^{*}\mu}.$ For the backward e-processes $\{S^{(t)}_{n}\}_{n<t}$, we adopt the predictable plug-in method [19].
+We divide the interval [0,1][0,1][ 0 , 1 ] into 101010 equal-sized bins $B_{i}=[\frac{i-1}{10},\frac{i}{10}),i=1,\cdots,10$ and perform histogram-based estimation.
+We construct $S^{(t)}_{n}=\prod_{i=n}^{t-1}\frac{\hat{p}_{i-1}(X_{i})}{{p}_{0}(X_{i})},$ where ${p}_{0}(x)=\frac{1}{10},$ and $\hat{p}_{i-1}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=i-1}^{t-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},x\in[0,1]$.
+We use the point estimator
+$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}.$ (6.5)
+Table 6.7 contains average results across 500500500 runs and Fig. 6.8 visualizes across 555 runs.
 […]
+T
+True (unknown) post-
+change density, $f_{1}(x)$
+Coverage
+Coverage
+Size
+$|\hat{T}-T|$
+Delay
+100 $4(1-x)^{3}$ 0.98 0.97 24.14 4.06 20.65
+[…]
+500 $4(1-x)^{3}$ 0.98 0.94 24.63 4.11 19.57
+[…]
+Figure 6.8: Setting IV: First $T-1$ samples are drawn from $U(0,1)$ and the remaining observations from the density $f(x)=4(1-x)^{3},x\in[0,1]$.
+The point estimate (6.5) is shown in a vertical red dashed line, and the confidence set (3.8) is shown in red points.
+Results of 555 random simulations are shown.
+[…]
+$\tau_{e}^{\prime}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}\geq A\right\},$
+with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}-0.5.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 3.3 with $F_{0}^{*}=\text{N}(0.5,1)$.
+After detection, we construct the confidence sets (3.8).
+The forward $t$-delay e-process we use is $\{R^{(t)}_{n}\}_{n\geq t}=\prod_{i=t}^{n}e^{\lambda_{i}X_{i}-\lambda_{i}^{2}/2}$, where $\lambda_{i}=\max\{0.5,\frac{1}{i-t}\sum_{j=t}^{i-1}X_{i}\}$ and the backward e-processes is constructed similarly: $\{S^{(t)}_{n}\}_{n<t}=\prod_{i=n}^{t}e^{\mu_{i}X_{i}-\mu_{i}^{2}/2}$, where $\mu_{i}=\min\{0,\frac{1}{t-i}\sum_{j=i+1}^{t}X_{j}\}-0.5$ (see, e.g., Ch.
+6.7.5 of [21]).
+We use the following point estimator:
+$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}^{\prime}}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}.$ (6.6)
+Figure 6.9: Setting V.
+First $T-1$ samples are drawn from $N(1,1)$ and the remaining samples are from $U(-1.2,0.8)$.
+Vertical red dashed lines show the point estimates (6.6).
+The confidence set (3.8) is marked in red points.
+Results of 555 independent simulations are shown.
+Table 6.10: Setting V: Parametric pre-change, nonparametric post-change.
+T
+Coverage
+Coverage
+Size
+$|\hat{T}-T|$
+Delay
+100 0.98 0.97 27.65 3.08 27.34
+500 0.98 0.94 28.03 3.15 25.12
+Comparison with [36]: We now compare our method with Ch.
+3.5 of [36] on the Gaussian mean-change, having the (unknown) changepoint at $T=100$ or 500500500.
+We recall that, unlike our general nonasymptotic guarantees for any $\mathcal{A}$, Wu’s method only has asymptotic approximate guarantees in specific parametric settings, and only for CUSUM procedure [18], $\tau^{\prime}_{\text{cusum}}:=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}T_{j}\geq d\right\},$ where $T_{n}=\max\{0,T_{n-1}+X_{n}\}$ and $d$ is the threshold.
+We used $d=8.59$ and 7.567.567.56 for ${N}(-0.25,1)$ vs.
+${N}(0.25,1)$ and ${N}(-0.3,1)$ vs.
+${N}(0.3,1)$ respectively.
+We implement Wu’s confidence set of the form $V_{c}\cup[L_{s},\hat{\nu})$, where $\hat{\nu}=\max\{k<\tau:T_{k}=0\}$, $V_{c}=\{k\geq\hat{\nu}:T_{k}\leq c\}$, and $L_{k}$ denote the $k$-th last zero point of $T_{n}$ counted backward starting from $L_{0}:=\hat{\nu}$, with their proposed approximations (for the case when $\theta_{1}=-\theta_{0}$) $s=\log(\alpha)\times(1/(\sqrt{2}\theta_{1})+0.088)$ and $c=-\log(1-\sqrt{1-\alpha})/(2\theta_{1})-0.583$ with $\alpha=0.05$.
+Table 6.11 reports the average size and coverage rates (for target conditional coverage 0.950.950.95) across 500500500 runs.
+Table 6.11: Comparison of the confidence sets for $T$ with [36] (Setting I)
+$F_{0}$ $F_{1}$ $T$ Conditional Size Conditional Coverage Marginal Coverage
+Ours Wu (2007) Ours Wu (2007) Ours Wu (2007)
+${N}(-0.25,1)$ ${N}(0.25,1)$ 100 46.39 47.83 0.96 0.86 0.88 0.79
+${N}(-0.3,1)$ ${N}(0.3,1)$ 100 33.74 34.14 0.96 0.83 0.82 0.76
+${N}(-0.25,1)$ ${N}(0.25,1)$ 500 46.58 47.21 0.96 0.86 0.57 0.53
+${N}(-0.3,1)$ ${N}(0.3,1)$ 500 35.19 34.81 0.97 0.82 0.56 0.49
+The results show that [36] (which is only asymptotically valid and relies on approximations) exhibits noticeable undercoverage, while the sizes for both methods are comparable.
+However, our method (4.1) achieves the target conditional coverage.
 Remark 6.1.
-As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets for $i=0,1$, we can have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$.
-So, in this case, Algorithms 4 and 3 reduces to our algorithm designed for the known pre- and post-change setting and similarly, Theorems 3.1 and 4.2 reduces to Theorem 2.7.
-4.1 Prototypical choice of test statistic $M_{t}$
-Let $L_{t}^{n}(\theta_{0},\theta_{1})$ be the likelihood of data up to time $n\geq t$ under $\mathcal{H}_{0,t}$. Then,
-$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (4.5)
-We define $\hat{\theta}_{0,1:j}$ to be the maximum likelihood estimator (MLE) of the null $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$.
-Similarly, we define $\hat{\theta}_{1,j:n}$ to be the MLE of the alternative $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$.
-The point estimate of $T$ can be naturally extended to the following form:
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (4.6)
-As a prototypical example, we consider the likelihood-ratio-based test statistic $M_{t}^{(n)}$, which is defined for $t,n\in\mathbb{N},t\leq n$ as below.
-$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (4.7)
-Note that the test statistic $M_{t}$ for testing the null hypothesis $\mathcal{H}_{0,t}$ reduces to
-$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{0,1:\hat{T}-1},\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau\text{ or }\tau=\infty.\\ \end{cases}$ (4.8)
-4.2 Computationally and statistically efficient subroutines
-As in Section 3.2, we assume that we have a fixed element $\theta^{\prime\prime}$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta_{0}\cup\Theta_{1}$.
-Therefore, for each $j=1,\cdots,B$, we generate data sequence $X_{1}^{j},X_{2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta^{\prime\prime}}$ only once.
-Then, for each $t\in\{1,\cdots,\tau\}$ and $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta,\theta^{\prime})\}_{n}$ having joint distribution $\mathbb{P}_{\theta,t,\theta^{\prime}},$ by applying the function $G(.,\theta)$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta^{\prime})$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
-$Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (4.9)
-Note that, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta,\theta^{\prime})\}_{n}$ across all $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Section B.2 for the Poisson distribution.
-Let us now fix $t\in\{1,\cdots,\tau\}$.
-Computing the detection time $\tau^{j}_{\theta,\theta^{\prime}}$ for each $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether there exists ($\text{there exists }(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are a finite sets.
-However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$ and $t_{2}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$.
-Also, we can often derive a suitable upper bound on the threshold, which does not require any approximation and is computationally and statistically efficient.
-E.g., we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$.
-Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where for $L<\infty$,
-$U_{t,L}^{j}:=\begin{cases}-\infty,\leavevmode\nobreak\ \leavevmode\nobreak\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (4.10)
-And for $L=\infty$,
-$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (4.11)
-We consider the following confidence sets:
-$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (4.12)
-$\mathcal{C}^{\prime\prime}_{2}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}.$ (4.13)
-Observe that $\mathcal{C}^{\prime\prime}_{i}\supseteq\mathcal{C}_{i}$, for $i=1,2$, where $\mathcal{C}_{i}$’s are the sets produced by Algorithm 4, since
-$\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{U_{t,L}^{j}\}_{j=1}^{B})\}$ $\geq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$
-$\geq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B}).$
-Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.2.
-Corollary 4.4.
-$\mathcal{C}^{\prime\prime}_{1}$ and $\mathcal{C}^{\prime\prime}_{2}$ defined in (4.12) and (4.13) respectively satisfy the same coverage guarantees as $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ in Theorem 4.2.
-In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with Poisson distribution in Appendix B.
-4.3 A concrete example: Gaussian mean shift problem
-We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$.
-Then, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$:
-$\mathcal{S}_{t-1}=\left[\bar{X}_{1:t-1}-\frac{q_{\gamma/2}}{\sqrt{t-1}},\bar{X}_{1:t-1}+\frac{q_{\gamma/2}}{\sqrt{{t-1}}}\right]\cap\Theta_{0},$ (4.14)
-and the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (3.13).
-We fix $t\in\{2,\cdots,\tau\}$.
-Note that we can consider $\theta^{\prime\prime}=0$ and $G(x,\theta)=x+\theta$.
-Therefore, only need to generate $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$.
-For $\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}$, define
-$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (4.15)
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (4.8), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
-Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark 4.5.
-For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
-$\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j}+\theta_{0},\bar{\epsilon}_{t:t^{\prime}}^{j}+\theta_{1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}$
-$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
-$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\bar{Y}_{i:t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i});Y_{1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
-where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime})=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime}),$ and $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ is the one such pair that maximizes $L_{i}\left(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)$, or equivalently minimizes
-$-2\log L_{i}\Big{(}\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\dots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\Big{)}=$
-$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (4.16)
-over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
-Here, constant means that it does not involve $\theta,\theta^{\prime}$.
-Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (4.3) is a quadratic function of $\theta_{1}-\theta_{0}$.
-Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
-Remark 4.5.
-[…]
-$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (4.17)
-$\tau_{\text{wsr-ripr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (4.18)
-where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$.
-For example, consider $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\geq b\}$, for some $b\geq a$.
-Observe that for all $\theta\in\Theta_{1},f_{\theta}^{*}=f_{a}$, and $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}=\exp\{\frac{1}{2}(\theta-a)\times\sum_{i=j}^{\tau}(2X_{i}-a-\theta)\}$ increses as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases.
-Therefore,
-$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
-$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
-Similarly, one can check that for $\mathcal{P}_{0}=\{N(\theta,1):\theta\geq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\leq b\}$, for $b\leq a$,
-$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
-$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
-Remark 4.6.
-For both types of pre- and post-change classes discussed above, and for CUSUM and SR-type detectors defined in (4.17) and (4.18), as well as the PFA-controlling change detectors used in Section 5.2, assumption 4.1 holds with $\theta_{0}^{*}=a$.
-We provide the explanation for the CUSUM detector, and the argument for the SR detector is analogous.
-For example, consider $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\geq b\}$, for some $b\geq a$.
-Then,
-$\mathbb{P}_{\theta_{0},\infty}[\tau_{\text{wcs-ripr}}\geq t]$
-$=\mathbb{P}_{\theta_{0},\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)<A\right]$
-$=\mathbb{P}_{\theta_{0},\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta}(X_{i})}{f_{a}(X_{i})}dW(\theta)<A\right]$
-[…]
-Then, $\exp\{\frac{1}{2}(\theta-a)\times\sum_{i=j}^{\tau}(2Z_{i}+2\theta_{0}-a-\theta)\}\leq\exp\{\frac{1}{2}(\theta-a)\times\sum_{i=j}^{\tau}(2Z_{i}+a-\theta)\},\forall\theta\in\Theta_{1},\forall\theta_{0}\in\Theta_{0}.$ Therefore, $\mathbb{P}_{\theta_{0},\infty}[\tau_{\text{wcs-ripr}}\geq t]\geq\mathbb{P}_{a,\infty}[\tau_{\text{wcs-ripr}}\geq t],\forall\theta_{0}\in\Theta_{0}.$
-The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology.
-Note that $\mathcal{C}^{\prime\prime}$ in (4.12) is computable if the following conditions are met:
-  1. 1.
-    The coupling function $G$ is available (which is always the case when distributions are continuous).
-For discrete distributions, even if a suitable $G$ is not available, our method may still be implementable, as demonstrated in Section B.2 for the Poisson case).
-  2. 2.
-    We can construct a confidence interval (and hence confidence sequence) for $\theta$.
-  3. 3.
-    The quantities $t_{j}^{1}$ and $t_{j}^{2}$ are computable (either exactly or numerically).
-  4. 4.
-    If we employ the likelihood-ratio-based test statistic (4.8), then we have to be able to maximize the function $L_{i}(\hat{\theta}_{1:i-1},\hat{\theta}_{i:t^{\prime}};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))$ over $(\theta,\theta^{\prime})$ on a set (either exactly or numerically).
-  5. 5.
-    If the detection algorithm $\mathcal{A}$ controls ARL (or is a black box), we need to satisfy Assumption 4.1. If $\mathcal{A}$ controls PFA, we do not need this assumption.
-For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions. Another concrete example with the Poisson distribution is in Appendix B.
-5 Experiments
-We validate our algorithms across various settings and demonstrate their practicality. Experiments below are for Gaussians; Poisson is handled in Appendix B.
-5.1 Experiments with detection algorithms that control ARL
-5.1.1 Known pre- and post-change (Setting I)
-We consider Gaussian mean-change from $N(0,1)$ (pre-change) to $N(1,1)$ (post-change) having the true (unknown) changepoint at $T=100,500$.
-We employ the CUSUM change detection method (C.1) with thresholds at $A=1000,10000$ (which controls ARL at level $A$ [23]).
-After detection, we employ Algorithm 2 to construct confidence sets for $T$, with $B=N=100$, $\alpha=0.05$.
-Fig.
-5.1 visualizes the confidence sets and point estimates (2.7) across $5$ runs.
-Table 5.2 has the average size and the coverages across $100$ independent runs.
-Figure 5.1: Setting I (known pre- and post-change Gaussian distributions): The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$ for $T=100$ (top row) and $T=500$ (bottom row).
-Change is detected using the CUSUM detector (C.1) with $A=10000$ (top row) or $A=1000$ (bottom row).
-The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$.
-Results of $5$ independent simulations are shown.
-Table 5.2: Known pre- and post-change (Setting I, finite ARL)
-$T$ A
-Conditional
-coverage
-Unconditional
-coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 10000 0.96 0.96 12.21 2.97 17.63
-100 1000 0.96 0.94 12.34 2.85 13.97
-500 10000 0.95 0.95 12.88 2.59 17.26
-500 1000 0.95 0.90 12.57 2.62 13.22
-For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible.
-However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage.
-Nevertheless, the empirical conditional coverage rates meet the desired level, supporting our theory.
-We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
-5.1.2 Known pre-change and composite post-change (Setting II)
-We perform experiments for Gaussian mean-change scenarios having pre-change distribution as $N(0,1)$, and the true (unknown) post-change data distribution as $N(1,1)$, with the changepoint at $T=100,500.$ For detection, we employ a weighted CUSUM detector (defined in (3.15)) with thresholds at $A=1000$ (which is known to control ARL at level $A$).
-For weighted CUSUM, we employ a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.75\}$ and $\theta=0.9$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
-After detection, we construct confidence sets (as defined in (3.11)) for $T$.
-We report the average size and coverages across $100$ independent runs in Table 5.4.
-Fig.
-5.3 provides visualizations of the confidence sets (3.11) and point estimates (3.5) of the changepoint across $5$ random runs.
-Figure 5.3: Setting II: The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$ with $T=100$ (top row) or $T=500$ (bottom row).
-Change is detected using a weighted CUSUM detector (3.15) with $A=10000$ (top row) or $A=1000$ (bottom row).
-The point estimate (3.5) is shown in a vertical red dashed line and the confidence set (3.11) is shown in red points.
-$F_{0}=N(0,1)$, $\mathcal{P}_{1}=\{N(\mu,1):\mu\geq 0.9\}$, $N=B=100$, $\alpha=0.05,\beta=0.025$ and $L=\infty$.
-Results of $5$ independent simulations are shown.
-Table 5.4: Known pre-change but unknown post-change (Setting II, finite ARL)
-T
-Composite
-post-change
-Conditional
-coverage
-Unconditional
-coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 26.25 3.95 16.87
-100 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 24.11 3.67 16.21
-500 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.89 30.83 3.45 17.63
-500 $\{N(\mu,1):\mu>0.9\}$ 0.97 0.91 28.56 3.48 15.81
-For $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in much lower unconditional coverage.
-However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.2).
-We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
-We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
-5.1.3 Composite pre- and post-change (Setting III)
-Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, with the changepoint at $T=100$ and $T=500$.
-For change detection, we employ a weighted CUSUM-type detector (4.17), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in setting II is considered.
-We set the thresholds at $A=1000,10000$.
-This is a special case of “e-detector” and is known to control ARL at level $A$ [20].
-After detection, we construct confidence sets (4.12) for $T$ (assuming $T\neq 1$), with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$.
-Fig.
-5.5 visualizes the confidence sets (4.12) and point estimates (4.6) of the changepoint across five random runs.
-Figure 5.5: Setting III: The first $T-1$ observations are drawn from $N(0,1)$ and the remaining observations from $N(1,1)$ with $T=100$ (top row) or $T=500$ (bottom row).
-Change is detected using the CUSUM detector defined in (4.17) with $A=1000$.
-The point estimate (4.6) is shown in vertical red dashed line and the confidence set (3.11) is shown in red points.
-$\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.1\}$, $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, $N=B=100$, $\alpha=0.05,\beta=\gamma=0.025$ and $L=\infty$.
-Results of $5$ independent simulations are shown.
-Table 5.6 contains the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs.
-The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table 5.2).
-We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table 5.2), exhibiting both higher coverage and larger size.
-Table 5.6: Composite pre- and post-change (Setting III, finite ARL)
-T
-Composite
-pre-change
-Composite
-post-change
-Conditional
-coverage
-Unconditional
-coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 27.84 4.36 25.81
-100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 25.71 4.19 23.15
-500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 31.45 4.03 24.48
-500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 30.66 4.07 23.07
-5.2 Experiments with detection algorithms that control PFA
-5.2.1 Known pre- and post-change (Setting A)
-We perform experiments for both Normal and Poisson mean-change scenarios having the true changepoint at $T=100,500$, using a likelihood ratio detector defined as
-$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (5.1)
-with thresholds at $A=1000$.
-Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality [24], the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
-After detection, we employ Algorithm 1 to construct confidence sets for $T$, with $B=100$ and $\alpha=0.05$.
-We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.7.
-We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
-Table 5.7: Known pre- and post-change distributions (Setting A, PFA $\leq 0.001$)
-$T$ $F_{0}$ $F_{1}$
-Conditional
-coverage
-Unconditional
-coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 $N(0,1)$ $N(1,1)$ 0.95 0.94 12.60 2.74 110.91
-100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 17.95 4.98 94.33
-500 $N(0,1)$ $N(1,1)$ 0.97 0.97 12.65 2.24 514.12
-500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 16.82 3.18 417.44
-5.2.2 Known pre- and unknown post-change (Setting B)
-We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector defined as
-$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (5.2)
-with thresholds at $A=1000$.
-The weights are chosen just as described in Section 5.1.2.
-Similarly, the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
-After detection, we construct confidence sets for $T$ as defined in (3.12).
-We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table 5.8, with $B=100$ and $\alpha=0.05,\beta=0.025,L=\infty$.
-We also report the average absolute deviation of $\hat{T}$ in (3.5) and delay (given no false detection) in the table.
-Table 5.8: Known pre-change but unknown post-change (Setting B, PFA $\leq 0.001$)
-T
-Composite
-post-change
-Conditional
-coverage
-Unconditional
-coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 $\{N(\mu,1):\mu>0.75\}$ 0.95 0.95 26.03 3.68 118.76
-100 $\{N(\mu,1):\mu>0.9\}$ 0.94 0.94 24.84 3.36 115.64
-500 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 30.95 3.18 530.83
-500 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 30.09 3.23 528.92
-5.2.3 Composite pre- and post-change (Setting C)
-We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) [13, 6, 12] being the representative of the pre-change class) defined as
-$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (5.3)
-with $f^{*}_{\theta}$ being the RIPr of $f_{\theta}$ on the pre-change class $\mathcal{P}_{0}$ and with thresholds at $A=1000$.
-This is a stop time of an ”e-process” [17] and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
-The weights are chosen just as described in Section 5.1.3.
-After detection, we construct confidence sets for $T$ as defined in (4.13) (assuming $T\neq 1$).
-The pre- and post-change data-generating distributions are $N(0,1)$ and $N(0,1)$ respectively, but neither of which is assumed known, and we vary the pre- and post-change classes.
-We use $B=100$ and $\alpha=0.05,\beta=\gamma=0.025,L=\infty$.
-We report the average size and coverage rates across $100$ independent runs in Table 5.9.
-We also report the average absolute deviation of $\hat{T}$ in (4.6) and delay (given no false detection).
-The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
-Table 5.9: Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
-T
-Composite
-pre-change
-Composite
-post-change
-Conditional
-coverage
-Unconditional
-coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.95 0.95 26.72 3.97 125.64
-100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 25.48 3.88 121.28
-500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 32.18 3.82 544.91
-500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 31.45 3.75 536.76
-5.3 Comparison of results with ARL and PFA controlling detection algorithms
-As expected, algorithms that control PFA exhibit a significantly higher average detection delay compared to those controlling ARL.
-This delay arises because stricter false alarm constraints force the detection algorithm to accumulate stronger evidence before signaling a change.
-In terms of inference, we observe that PFA-controlling algorithms often tend to yield slightly lower conditional coverage probabilities with confidence sets of comparable sizes.
-Also, conditional and unconditional coverage rates are identical when PFA-controlling detection algorithms are employed, due to the very low probability of false detections, while the ARL controlling algorithms lead to lower unconditional conditional coverage rates for $T=500$, supporting our theory.
-Moreover, we observe that the average absolute deviations of the point estimates are barely smaller with the PFA controlling algorithms.
-Overall, the major differences between these two types lie in detection delay and unconditional coverage properties.
-We do not observe significant differences in the size and conditional coverage properties of the confidence sets nor in the absolute deviation of the point estimates, across these two types of detection algorithms.
+In Settings I, II, and III, both of our proposed methods — universal (Section 3) and adaptive (Section 4) — are applicable.
+For nonparametric settings (IV and V), the only valid method currently known is the universal method.
+Prior works are valid (asymptotically) only for Setting I.
+As shown in Table 6.2 for known pre- and post-change distributions (Setting I), the adaptive method consistently outperforms the universal method, which tends to be conservative.
+Also, our adaptive method is the only method we know that remains valid under dependence (see experimental results in Supplement G.1).
+However, for the composite settings (II and III), the universal method outperforms the adaptive method.
+Moreover, the universal method is also significantly more efficient — it runs in $O(T)$ time (assuming $M_{t}$ can be computed in $O(1)$ time), while the adaptive method incurs a higher computational cost of $O(T^{2}B)$ (assuming $\mathcal{A}$ runs in $O(T)$ time and $M_{t}$ can be computed in $O(1)$ time).
+See Supplement Sections D, E, and G for further experiments, involving confidence intervals for pre- and post-change parameters, detection procedures with bounded PFA, Poisson rate change problems, settings with dependent data, etc.
 7 Conclusion
-We developed a framework to construct a confidence set for the unknown changepoint $T$ after a sequential detection algorithm declares a change at a stopping time, providing conditional coverage guarantees (conditioned on the event that the stopping occurs after the changepoint) when the pre- and post-change distributions are known.
-We also extended our framework to handle composite pre- and post-change distributions and provide conditional coverage guarantees, demonstrating its good empirical performance.
-We impose no specific conditions on the change detection algorithms; they can function as entirely black-box or heuristic methods; we only need to be able to run the algorithm on simulated data sequences.
-By filling this significant gap in post-detection analysis, our work provides crucial tools for precise and timely changepoint localization in real-time data streams.
+We developed a general framework for constructing confidence sets for the unknown changepoint $T$ after a sequential detector signals a change at a data-dependent stopping time.
+It applies to any general nonparametric post-change class when the pre-change distribution is known, and we further extend it to settings with a composite pre-change class under a suitable assumption.
+For the parametric setting, we proposed an alternative simulation-based method that empirically yields tighter confidence sets when both pre- and post-change distributions are known and can handle dependent data as well.
+However, we note that our primary (universal) method is nonparametric, computationally more efficient, and simpler to implement.
+Crucially, all our methods are agnostic to the detection algorithm used and provide non-asymptotic (conditional) coverage guarantees, whereas prior works are tailored for exponential family distributions with the CUSUM procedure and provide only asymptotic (conditional) coverage.
+By bridging a longstanding gap in the literature, our work offers the first nonparametric, theoretically principled, and broadly applicable method for post-detection inference.
+Supplementary Material
+Outline of Supplementary Material
+Omitted proofs can be found in Appendix A.
+We provide the implementation details of Algorithms 2 and 3 in Appendices B and C respectively, along with two concrete examples in both sections.
+Experimental results for the confidence intervals for pre- and post-change parameters are in Appendix D.
+Experimental results for a different implementation method are in Appendix E.
+Simpler algorithms for PFA-controlling change detectors are presented in Appendix F.
+Some additional experiments with dependent (Markov) pre- and post-change data, with detection procedures with bounded PFA, and with Poisson rate change problems are in Appendix G.
+Some heuristic algorithms are presented in Appendices H and I, pointing at open avenues for further theoretical work.
 Appendix A Omitted proofs
-Here, we provide the formal mathematical details that were omitted in the main paper.
-A.1 Omitted proofs from Section 2
-Proof of Theorem 2.2.
-First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d. under $H_{0,t}$.
-[…]
-$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-\alpha)(B+1)\rceil}{B+1}\geq 1-\alpha.$ (A.1)
-For $L\in\mathbb{N}\cup\{\infty\},t\in\mathbb{N}$, define $C_{t,L}=\operatorname{Quantile}(1-\alpha;M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B}))$. Note that
-$\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\},\tau<T)\leq\mathbb{P}_{T}[\tau<T]=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$
-the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$.
-Also, observe that for any $L\in\mathbb{N}\cup\{\infty\}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$ and hence, $C_{t,L}\geq C_{t,\infty}$, for all $t\in\mathbb{N}$.
-$1-\alpha$ $\leq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})$
-$\leq\mathbb{P}_{T}(M_{T}\leq C_{t,L})$
-$=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})$
-$=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\},\tau\geq T)+\mathbb{P}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\},\tau<T)$
-$\leq\mathbb{P}_{T}(T\in\mathcal{C})+\mathbb{\mathbb{P}}_{\infty}(\tau<T).$
-Therefore, $\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<T)$.
-Now, if $\mathcal{A}$ controls PFA at level $\delta$, i.e., $\mathbb{P}_{\infty}(\tau<\infty)\leq\delta$, we have
-$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<T)\geq 1-\alpha-\mathbb{\mathbb{P}}_{\infty}(\tau<\infty)\geq 1-\alpha-\delta.$
-Proof of Proposition 2.4.
+In this section, we provide the mathematical details that were omitted in the main paper.
+Proof of Proposition 2.1..
 For the sake of contradiction, assume that $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$ for some $\mathcal{C}^{*}\subseteq\{1,\cdots,\tau\}$. Now we observe that
@@
 Proof of Proposition 2.2.
-Note that $\mathcal{C}^{*}=\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\cup\{\tau+1,\tau+2,\cdots\}=\{t\geq 1:M_{t}\leq C_{t,L}\}$, since we defined $M_{t}=-\infty\leq C_{t,L}$, for $t>\tau.$
+First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
+Let $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying $\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\geq 1-\alpha$, for all $t\in\mathbb{N}$.
+Since $\tau$ be a stopping time, $[\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$.
+This implies $\mathbb{P}_{F_{0},\infty}(\tau\geq t)=\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ Therefore,
+$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)$ $=\mathbb{P}_{F_{0},t,F_{1}}(t\notin\mathcal{C})$
+$=1-\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C})$
+$=1-\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t)$
+$\leq 1-(1-\alpha)\mathbb{P}_{F_{0},\infty}(\tau\geq t)$
+$\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t),\leavevmode\nobreak\ \forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
+For the converse part, $\phi^{(t)}$ satisfies $\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$
+So, $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
+$\mathbb{P}_{F_{0},T,F_{1}}(\phi^{(T)}(X_{1},\cdots,X_{\tau})=0)$
+$=\mathbb{P}_{F_{0},T,F_{1}}(\phi^{(T)}(X_{1},\cdots,X_{\tau})=0,\tau\geq T)+\mathbb{P}_{F_{0},T,F_{1}}(\phi^{(T)}(X_{1},\cdots,X_{\tau})=0,\tau<T)$
+$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\})+\mathbb{P}_{F_{0},\infty}(\tau<T)$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})+\mathbb{P}_{F_{0},\infty}(\tau<T)$
+Therefore, $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq T)-\mathbb{P}_{F_{0},\infty}(\tau<T)$, which implies
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C},\tau\geq T)}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
+Lemma A.1 (Ville’s inequality [32]).
+If $\{R_{n}\}_{n\in\mathbb{N}}$ is an e-process for $\mathcal{P}$ with respect to a filtration $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, then
+$\mathbb{P}\left(\exists n\in\mathbb{N}:R_{n}\geq 1/\alpha\right)\leq\alpha,\forall\mathbb{P}\in\mathcal{P}.$ (A.1)
+For a short proof, see [11][Lemma 1].
+Proof of Theorem 3.1.
+Note that by definition of $M_{t}$, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t=\hat{T}\leq\tau<\infty]=0$, $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t>\tau]=0$ and $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}>2/\alpha,\tau=\infty]=0$.
+Therefore, the only cases where $M_{t}>2/\alpha$ might occur are $t<\hat{T}\leq\tau<\infty$ or $\hat{T}<t\leq\tau<\infty$.
+Hence, for any $t\in\mathbb{N}$,
+$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha]$
+$=\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t<\hat{T}\leq\tau<\infty]+\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,\hat{T}<t\leq\tau<\infty]$
+$=\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{\hat{T}-1}\geq 2/\alpha,t<\hat{T}\leq\tau<\infty\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{\hat{T}}\geq 2/\alpha,\hat{T}<t\leq\tau<\infty\right]$
+$\leq\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{(t-1)\wedge\hat{T}}\geq 2/\alpha\right]$
+Now, from Ville’s inequality (Lemma A.1) on the e-processes $\{R^{(t)}_{n}\}_{n\geq t}$,
+$P_{F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
+Therefore,
+$\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha\right]$ $=\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha,t\vee(\hat{T}-1)=n\right]\right)$
+$\leq\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{n}\geq 2/\alpha\right]\right)$
+$=P_{F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
+Similarly, for the e-processes $\{S^{(t)}_{n}\}_{n<t}$, we get $\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{(t-1)\wedge\hat{T}}\geq 2/\alpha\right]\leq\alpha/2.$ Therefore, from the above steps, we finally have
+$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha]\leq\alpha/2+\alpha/2=\alpha.$
+Proof of Theorem 3.2.
+From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other. Hence,
+$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}]=\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t],\forall t\in\mathbb{N}.$ Now,
+$\mathbb{P}_{F_{0},T,F_{1}}(M_{T}<2/(\alpha r_{T}))$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}<2/(\alpha r_{t})\})$
+$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}<2/(\alpha r_{t})\},\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}<2/(\alpha r_{t})\}\mid\tau\geq T)\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T),$
+where we twice used the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$, which holds because $\tau$ is a stopping time and so the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
+So, we obtain
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\mathbb{P}_{F_{0},T,F_{1}}(M_{T}>2/(\alpha r_{T}))}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$ (A.2)
+Proof of Proposition 3.4.
+We only prove the case when $\theta_{0}^{*}=\sup\Theta_{0}\leq\inf\Theta_{1}$ and $\tau$ is the CUSUM type stopping time $\tau_{\text{wcs-ripr}}$ defined in (C.12).
+Proofs for the other cases, when $\theta_{0}^{*}=\inf\Theta_{0}\geq\sup\Theta_{1}$ and/or $\tau$ is the SR-type stopping time (C.13), follow analogously.
+$\mathbb{P}_{\theta,\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}_{\theta,\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta_{1}}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}dW(\theta_{1})<A\right].$
+Since the family $\{F_{\theta}\}$ satisfies the MLR property in the statistic $T(x)$, and since $\frac{f_{\theta_{1}}(x)}{f_{\theta_{0}^{*}}(x)}$ is non-decreasing in $T(x)$, it follows that the random variables $\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}$ are stochastically increasing in $\theta$.
+That is, for $X\sim F_{\theta}$ and $X^{\prime}\sim F_{\theta^{\prime}}$ with $\theta\leq\theta^{\prime}\leq\sup\Theta_{0}$, we have $T(X)\preceq_{\text{st}}T(X^{\prime})$ (see e.g., [16, Chapter 3.4]) and hence,
+$\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}\preceq_{\text{st}}\frac{f_{\theta_{1}}(X^{\prime})}{f_{\theta_{0}^{*}}(X^{\prime})}.$
+By Theorem 1.A.2 of [24], the stochastic ordering implies that there exist a random variable $Z_{\theta_{1}}$ and functions $\psi_{\theta_{1},1}$ and $\psi_{\theta_{1},2}$ such that $\psi_{\theta_{1},1}(z)\leq\psi_{\theta_{1},2}(z)$ for all $z$ and $\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{\theta_{1},1}(Z_{\theta_{1}})$ and $\frac{f_{\theta_{1}}(X^{\prime})}{f_{\theta_{0}^{*}}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{\theta_{1},2}(Z_{\theta_{1}})$.
+Therefore,
+$\mathbb{P}_{\theta,\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\psi_{\theta_{1},1}(Z_{\theta_{1},i})dW(\theta_{1})<A\right]$
+$\geq\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\psi_{\theta_{1},2}(Z_{\theta_{1},i})dW(\theta_{1})<A\right]$
+$=\mathbb{P}_{\theta^{\prime},\infty}[\tau_{\text{wcs-ripr}}\geq t].$
+Proof of Theorem 3.5.
+For all $t\in\mathbb{N},\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}^{*}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}^{*}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}^{*}]=\alpha\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]\geq\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t].$ The last step follows from Assumption 3.3.
+The rest of the proof is the same as in the proof of Theorem 3.2.
+∎
+Proof of Theorem 4.1.
+First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
+Defining $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, and $\hat{\alpha}_{t}=\alpha r_{t}$. Now,
+$\mathbb{P}_{F_{0},T,F_{1}}(M_{T}\leq C_{T,L})$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\})$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau\geq T)+\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\},\tau<T)$
+$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\mid\tau\geq T)\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T),$
+where we twice invoked $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$ as explained earlier. So, we obtain
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{T,L})}{P_{\infty}(\tau\geq T)}.$ (A.3)
+First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d.
+under $H_{0,t}$.
+Hence, $M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-c)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$.
+So, for any fixed $c\in(0,1)$,
+$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c\text{ for all }t.$
+Since $\hat{\alpha}_{t}$ is independent of $M_{t}$, it follows that
+$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t})\geq\frac{\lceil(1-\hat{\alpha}_{t})(B+1)\rceil}{B+1}\geq 1-\hat{\alpha}_{t},\text{ almost surely for all }t.$
 Therefore, we have
-$\mathbb{P}_{T}(T\in\mathcal{C}^{*})=\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t,L}\})=\mathbb{P}_{T}(M_{T}\leq C_{t,L})\geq\mathbb{P}_{T}(M_{T}\leq C_{T,\infty})\geq 1-\alpha,$
-where the last inequality follows from (A.1). ∎
-Proof of Corollary 2.6.
-From Theorem 2.2, we get $\mathbb{P}_{T}(T\in\mathcal{C})\geq\mathbb{P}_{\infty}(\tau\geq T)-\alpha.$
-Also, $\mathbb{P}_{T}(\tau\geq T)=\mathbb{P}_{\infty}(\tau\geq T)$ because of the fact the event $[\tau\geq T]$ does not depend $X_{i}$ for $i\geq T$. Therefore,
-$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)=\frac{\mathbb{P}_{T}(T\in\mathcal{C})}{\mathbb{P}_{T}(\tau\geq T)}\geq 1-\frac{\alpha}{\mathbb{P}_{\infty}(\tau\geq T)}.$
-Therefore, if $\mathcal{A}$ controls PFA at level $\delta$, we have
-$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{\mathbb{P}_{\infty}(\tau\geq T)}\geq 1-\frac{\alpha}{\mathbb{P}_{\infty}(\tau=\infty)}\geq 1-\frac{\alpha}{1-\gamma}.$
-A.2 Omitted proofs from Section 3
+$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}\leq C_{t,\infty})=\mathbb{E}_{t}(\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t}))\geq\mathbb{E}_{t}(1-\hat{\alpha}_{t})=1-\alpha\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t).$
+As noted earlier, $\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t)=\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ and hence, $\mathbb{P}_{F_{0},t,F_{1}}(M_{t}\leq C_{t,\infty})\geq 1-\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ Also, for all $L\in\mathbb{N}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$.
+Therefore, $C_{t,L}\geq C_{t,\infty}$.
+Thus, for all $t$,
+$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ (A.4)
+Therefore, combining (A.3) and (A.4), we obtain $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
 Proof of Theorem 4.4.
@@
 Note that $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}\}),$ for all $t\in\mathbb{N},\alpha\in(0,1)$.
-  1. (a)
-    For $\mathcal{C}_{1}$: For all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
+For all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
 $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))$
@@
 Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha r_{t}})\geq 1-(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta$ from the above inequality.
-  2. (b)
-    For $\mathcal{C}_{2}$: Here, we have $r_{t}=1$.
-The proof for $\mathcal{C}_{2}$ is similar.
-For fixed $\alpha\in(0,1),$ with $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverage $1-\beta$ we show that
-    $P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta.$
-    Then, we have
-    $P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
-    If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, we obtain
-    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha+\beta}{\mathbb{P}_{\theta_{0},\infty}(\tau=\infty)}\geq 1-\frac{\alpha+\beta}{1-\delta}.$
-    And similarly,
-    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)-(\alpha+\beta)\geq 1-\alpha-\beta-\delta.$
-A.3 Omitted proofs from Section 4
-Proof of Theorem 4.2.
+Proof of Theorem 4.5..
 First, observe that $M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}$ are i.i.d. under $\mathbb{P}_{\theta_{0},t,\theta_{1}}$.
@@
 Note that $\theta_{0}\in\mathcal{S}_{t-1}$ and $\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1}$ imply that $C_{t,L}^{\alpha}\geq C_{t,\infty}^{\alpha}\geq\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}\}),$ for any $\alpha\in(0,1).$ By Assumption 4.1, we have that $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$.
-  1. (a)
-    For $\mathcal{C}_{1}$: Also, for all $t$, with $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}$ is independent to $\{X_{n}\}_{n}$, we have
-    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})\mid r_{t}))$
-    $\geq\mathbb{E}(1-\beta r_{t})=1-\beta\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
+Also, for all $t$, with $r_{t}^{*}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, since $r_{t}^{*}$ is independent to $\{X_{n}\}_{n}$,
+$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})\mid r_{t}^{*}))$
+$\geq\mathbb{E}(1-\beta r_{t}^{*})=1-\beta\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
 And similarly,
-    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{o}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t})\mid r_{t}))$
-    $\geq\mathbb{E}(1-\gamma r_{t})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
+$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{o}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})\mid r_{t}^{*}))$
+$\geq\mathbb{E}(1-\gamma r_{t}^{*})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
 Then, we have
-    $\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}})$
-    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}},\theta_{0}\in\mathcal{S}_{t-1},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\notin\mathcal{S}_{t-1}))+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
-    $\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))+(\beta+\gamma)\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$
-    $\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})\mid r_{t}\right)\right)+(\beta+\gamma)\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
-    $\leq\mathbb{E}(\alpha r_{t})+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
+$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}^{*}})$
+$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha r_{t}^{*}},\theta_{0}\in\mathcal{S}_{t-1},\theta_{1}\in\mathcal{S}^{\prime}_{\tau-t+1})+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\notin\mathcal{S}_{t-1}))+\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\notin\mathcal{S}^{\prime}_{\tau-t+1})$
+$\leq\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}>\operatorname{Quantile}(1-\alpha r_{t}^{*};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))+(\beta+\gamma)\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)$
+$\leq\mathbb{E}\left(\mathbb{P}_{\theta_{0},t,\theta_{1}}\left(M_{t}>\operatorname{Quantile}(1-\alpha r_{t}^{*};M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B})\mid r_{t}^{*}\right)\right)+(\beta+\gamma)\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
+$\leq\mathbb{E}(\alpha r_{t}^{*})+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
 $=\alpha\times\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t)+(\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$
@@
 We also have that
-    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
+$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}^{*}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
 the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
-    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\alpha r_{t}})$
-    $=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\})$
-    $=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)$
+$\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\alpha r_{t}^{*}})$
+$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}^{*}}\})$
+$=\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}^{*}}\}\mid\tau\geq T)\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau\geq T)+\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)$
 $\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
 Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L})\geq 1-(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma$ from the above inequality.
-  2. (b)
-    For $\mathcal{C}_{2}$: Proof is similar to part (a).
-Here we have for fixed $r_{t}=1$ with $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ having coverages $1-\beta$ and $1-\gamma$ respectively,
-    $P_{\theta_{0},t,\theta_{1}}(M_{t}>C_{t,L}^{\alpha})\leq\alpha+\beta+\gamma.$
-    Similarly, we have
-    $P_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+1-\mathbb{P}_{\theta_{0},\infty}(\tau\geq T).$
-    If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha})\geq 1-\alpha-\beta$, we obtain
-    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta+\gamma}{\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha+\beta+\gamma}{\mathbb{P}_{\theta_{0},\infty}(\tau=\infty)}\geq 1-\frac{\alpha+\beta+\gamma}{1-\delta}.$
-    Similarly,
-    $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)-(\alpha+\beta+\gamma)\geq 1-\alpha-\beta-\gamma-\delta.$
-Appendix B Another concrete example with Poisson distribution
-B.1 Known pre-change and composite post-change
+Proof of Theorem 5.1.
+First note that $\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CS}(X_{T},\cdots,X_{\tau};1-\eta_{1}r_{T}^{*})\mid r_{T}^{*}\right)\leq\eta_{1}r_{T}^{*}$ and $\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{0}\notin\displaystyle\operatorname{CI}(X_{1},\cdots,X_{T-1};1-\eta_{0}r_{T}^{*})\mid r_{T}^{*}\right)\leq\eta_{0}r_{T}^{*}$, because $r_{T}^{*}$ is independent of $X_{i}$’s.
+Now,
+$\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})\mid\tau\geq T\right)$
+$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*}),T\in\mathcal{C}\mid\tau\geq T\right)+\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(T\notin\mathcal{C}\mid\tau\geq T\right)$
+$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CS}(X_{T},\cdots,X_{\tau};1-\eta_{1}r_{T}^{*})\mid\tau\geq T\right)+\alpha$
+$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CS}(X_{T},\cdots,X_{\tau};1-\eta_{1}r_{T}^{*})\right)/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
+$=\mathbb{E}(\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CS}(X_{T},\cdots,X_{\tau};1-\eta_{1}r_{T}^{*})\mid r_{T}^{*}\right))/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
+$\leq\mathbb{E}(\eta_{1}r_{T}^{*})/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
+$=\alpha+\eta_{1}\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq T)/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$
+$\leq\alpha+\eta_{1},$
+where the last inequality follows from Assumption 3.3. Similarly,
+$\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})\mid\tau\geq T\right)$
+$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}),T\in\mathcal{C}\mid\tau\geq T\right)+\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(T\notin\mathcal{C}\mid\tau\geq T\right)$
+$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CI}(X_{1},\cdots,X_{T-1};1-\eta_{0}r_{T}^{*})\mid\tau\geq T\right)+\alpha$
+$\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CI}(X_{1},\cdots,X_{T-1};1-\eta_{0}r_{T}^{*})\right)/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
+$=\mathbb{E}(\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CI}(X_{1},\cdots,X_{T-1};1-\eta_{0}r_{T}^{*})\mid r_{T}^{*}\right))/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
+$\leq\mathbb{E}(\eta_{0}r_{T})/\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)+\alpha$
+$\leq\alpha+\eta_{0}.$
+Appendix B Computationally and statistically efficient subroutines in Algorithm 2
+Let $L_{t}(\theta_{1};X_{1},\cdots,X_{n})$ be the likelihood of data $X_{1},\cdots,X_{n}$ under $\mathcal{H}_{0,t}$. Then, for $n\geq t$,
+$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (B.1)
+We define $\hat{\theta}_{1,j:n}$ to be a maximum likelihood estimator (MLE) of the post-change parameter $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$.
+The point estimate of $T$ can be naturally extended to:
+$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (B.2)
+The prototypical test statistic that we employ in our simulations is the likelihood-ratio-based test statistic $M_{t}^{(n)}$, where for $t,n\in\mathbb{N},t\leq n$,
+$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (B.3)
+Note that the test statistic $M_{t}$ for testing $\mathcal{H}_{0,t}$ reduces to
+$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (B.4)
+For computational and statistical efficiency, we generate sequences from all $t\in\{1,\cdots,\tau\}$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ simultaneously using a single shared source of randomness.
+To achieve this, we assume that we have a fixed element $\theta^{\prime\prime}\in\Theta$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta$.
+Therefore, for each $j=1,\cdots,B$, we generate only one data sequence $\{X_{n}^{j}\}_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{\theta^{\prime\prime}}$, and then for each $t\in\{1,\cdots,\tau\}$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ (having joint distribution $\mathbb{P}_{\theta_{0},t,\theta}$) by applying the function $G(.,\theta_{0})$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta)$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
+$Y_{n}^{t}(x,\theta):=\begin{cases}G(x,\theta_{0}),&\text{if }n<t\\ G(x,\theta),&\text{ otherwise. }\end{cases}$ (B.5)
+Let us provide a few examples of the function $G$.
+If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}}(x))$ works.
+In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$.
+In the case of discrete distributions, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta)\}_{n}$ across all $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
+Let us now fix $t\in\{1,\cdots,\tau\}$.
+Computing the detection time $\tau_{j,t}^{\theta}$ for each $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set.
+However, it is often straightforward to compute $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$.
+Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient.
+For example, we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$.
+Again, we bound $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)$ by $V_{t,L}^{j}$, where for $L<\infty$,
+$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (B.6)
+And for $L=\infty$,
+$V_{t,\infty}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{otherwise,}\\ \end{cases}$ (B.7)
+which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section B.1). We define the confidence set as
+$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (B.8)
+We observe that $\mathcal{C}_{1}^{\prime}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}$ is the set produced by Algorithm 2, since
+$\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$ $\geq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
+$\geq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}).$
+Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.4.
+Corollary B.1.
+$\mathcal{C}^{\prime}_{1}$ defined in (B.8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.4.
+Next, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with the Poisson distribution in Section G.3.
+B.1 A concrete example: Gaussian mean shift problem
+Let the pre-change distribution $F_{0}$ be $N(\theta_{0},1)$ and the post-change model be $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
+In this setting, we have the following example of a $1-\beta$ confidence sequence [12] for $\theta_{1}$
+$\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (B.9)
+where $\bar{X}_{t:\tau}=\frac{1}{\tau-t+1}\sum_{i=t}^{\tau}X_{i}$ and $s_{n}(\beta)=\sqrt{\log\log(2n)+0.72\log(10.4/\beta)}$.
+We fix $t\in\{1,\cdots,\tau\}$.
+Note that we can consider $\theta^{\prime\prime}=\theta_{0}$ and $G(x,\theta)=x+\theta-\theta_{0}$.
+Therefore, only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(\theta_{0},1)$ for $j=1,\cdots,B$.
+For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
+$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t,\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (B.10)
+where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$.
+We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (B.4), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
+Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$.
+Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.2) and they are finite.
+So, we can have $L=\infty$.
+If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t$, it can be simplified in our setting as below:
+$V_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j}}{t^{\prime}-t+1}+\theta;Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}$
+$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
+$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta_{t^{\prime}}^{i});Y_{1}^{t}(X_{1}^{j},\theta_{t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta_{t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
+where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta)=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta),$ and $\theta_{t^{\prime}}^{i}$ is the unique element that maximizes $L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)$, or equivalently minimizes
+$-2\log L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)=$
+$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{(t-i)(\theta-\theta_{0})}{t^{\prime}-i+1}\right)^{2}+\text{const},\text{ if }i<t,\\ \text{const},\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{ const},\text{ if }i>t,\end{cases}$
+over $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$.
+Here, constant refers to the terms that do not involve $\theta$.
+Note that $\theta_{t^{\prime}}^{i}$ has a simple closed-form expression since the above equation is a quadratic function of $\theta$.
+And hence, $V_{t,\infty}^{j}$ can be computed easily.
+Remark B.2.
+[…]
+$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (B.11)
+[…]
+For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$.
+Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{n}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases.
+Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
+Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
+Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
+One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section G.2.
+The Gaussian mean shift example we provided serves as a concrete and illustrative case to clarify the methodology.
+It is worth noting that $C^{\prime}$, as defined in (B.8), is practically implementable as long as the following relatively weak conditions are met:
+  1. 1.
+    The coupling function $G$ is available (which is always the case when distributions are continuous and in the case of discrete distributions, even if a suitable $G$ is not directly available, our method may still be implementable, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution).
+  2. 2.
+    We can construct a confidence sequence for the parameter of interest (which is always achievable as long as one can construct a confidence interval for the parameter).
+  3. 3.
+    The quantities $t_{j}^{1}$ and $t_{j}^{2}$ are computable (either exactly or numerically).
+  4. 4.
+    If we employ the likelihood-ratio-based test statistic (B.4), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
+For instance, one can easily verify that this approach applies to Gaussian, Laplace, or exponential scale change problems.
+B.2 Another concrete example: Poisson rate change problem
 Here is $F_{0}$ is $\text{Pois}(\theta_{0})$ and the post-change model is $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\in\Theta_{1}\}$.
@@
 Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta_{0})$, for $n<t$ and $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta)$, for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1},n\geq t$.
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (2.8), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
+We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in Section 3, with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
 Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$.
@@
 Remark B.3.
-For weighted CUSUM or SR detectors (defined in (3.15),(3.16)), finding $t_{1}^{j},t_{2}^{j}$ is often straightforward.
+For weighted CUSUM or SR detectors (defined in Remark 3.3.), finding $t_{1}^{j},t_{2}^{j}$ is often straightforward.
 For example, consider $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$.
@@
 Also, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\neq\theta_{0}\}$, if we employ Gamma weights (it is a conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
-B.2 Composite pre- and post-change
+Appendix C Computationally and statistically efficient subroutines in Algorithm 3
+Let $L_{t}^{n}(\theta_{0},\theta_{1})$ be the likelihood of data up to time $n\geq t$ under $H_{0,t}$. Then,
+$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (C.1)
+We define $\hat{\theta}_{0,1:j}$ to be the maximum likelihood estimator (MLE) of the null $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$.
+Similarly, we define $\hat{\theta}_{1,j:n}$ to be the MLE of the alternative $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$.
+The point estimate of $T$ can be naturally extended to the following form:
+$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (C.2)
+As a prototypical example, we consider the likelihood-ratio-based test statistic $M_{t}^{(n)}$, which is defined for $t,n\in\mathbb{N},t\leq n$ as below.
+$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (C.3)
+Note that the test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ reduces to
+$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{0,1:\hat{T}-1},\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau\text{ or }\tau=\infty.\\ \end{cases}$ (C.4)
+As in Appendix B, we assume that we have a fixed element $\theta^{\prime\prime}$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta_{0}\cup\Theta_{1}$.
+Therefore, for each $j=1,\cdots,B$, we generate data sequence $X_{1}^{j},X_{2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta^{\prime\prime}}$ only once.
+Then, for each $t\in\{1,\cdots,\tau\}$ and $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta,\theta^{\prime})\}_{n}$ having joint distribution $\mathbb{P}_{\theta,t,\theta^{\prime}},$ by applying the function $G(.,\theta)$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta^{\prime})$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
+$Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (C.5)
+Note that, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta,\theta^{\prime})\}_{n}$ across all $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
+Let us now fix $t\in\{1,\cdots,\tau\}$.
+Computing the detection time $\tau^{j}_{\theta,\theta^{\prime}}$ for each $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are a finite sets.
+However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$ and $t_{2}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$.
+Also, we can often derive a suitable upper bound on the threshold, which does not require any approximation and is computationally and statistically efficient.
+E.g., we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$.
+Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where for $L<\infty$,
+$U_{t,L}^{j}:=\begin{cases}-\infty,\leavevmode\nobreak\ \leavevmode\nobreak\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (C.6)
+And for $L=\infty$,
+$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (C.7)
+We consider the following confidence set:
+$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (C.8)
+Observe that $\mathcal{C}^{\prime\prime}_{1}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}_{1}$ is the set produced by Algorithm 3, since
+$\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{U_{t,L}^{j}\}_{j=1}^{B})\}$ $\geq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$
+$\geq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B}).$
+Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.5.
+Corollary C.1.
+$\mathcal{C}^{\prime\prime}_{1}$ defined in (C.8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.5.
+In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method.
+We illustrate another example with Poisson distribution in Appendix C of the Supplementary Material.
+C.1 A concrete example: Gaussian mean shift problem
+We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$.
+Then, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$:
+$\mathcal{S}_{t-1}=\left[\bar{X}_{1:t-1}-\frac{q_{\gamma/2}}{\sqrt{t-1}},\bar{X}_{1:t-1}+\frac{q_{\gamma/2}}{\sqrt{{t-1}}}\right]\cap\Theta_{0},$ (C.9)
+and the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (B.9).
+We fix $t\in\{2,\cdots,\tau\}$.
+Note that we can consider $\theta^{\prime\prime}=0$ and $G(x,\theta)=x+\theta$.
+Therefore, only need to generate $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$.
+For $\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}$, define
+$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (C.10)
+We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (C.4), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
+Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark C.2.
+For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
+$\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j}+\theta_{0},\bar{\epsilon}_{t:t^{\prime}}^{j}+\theta_{1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}$
+$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
+$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\bar{Y}_{i:t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i});Y_{1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
+where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime})=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime}),$ and $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ is the one such pair that maximizes $L_{i}\left(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)$, or equivalently minimizes
+$-2\log L_{i}\Big{(}\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\dots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\Big{)}=$
+$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (C.11)
+over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
+Here, constant means that it does not involve $\theta,\theta^{\prime}$.
+Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (C.1) is a quadratic function of $\theta_{1}-\theta_{0}$.
+Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
+Remark C.2.
+[…]
+$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (C.12)
+$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (C.13)
+where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$.
+For example, consider $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\geq b\}$, for some $b\geq a$.
+Observe that for all $\theta\in\Theta_{1},f_{\theta}^{*}=f_{a}$, and $\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}=\exp\{\frac{1}{2}(\theta-a)\times\sum_{i=j}^{n}(2X_{i}-a-\theta)\}$ increses as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases.
+Therefore,
+$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
+$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
+Similarly, one can check that for $\mathcal{P}_{0}=\{N(\theta,1):\theta\geq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\leq b\}$, for $b\leq a$,
+$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
+$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
+The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology.
+Note that $\mathcal{C}^{\prime\prime}$ in (C.8) is computable if similar conditions as listed in Section B.1 are met.
+For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions.
+Another concrete example with the Poisson distribution is in Section G.3.
+C.2 Another concrete example: Poisson rate change problem
 Here we assume the pre and post-change model to be $\mathcal{P}_{i}=\{\text{Pois}(\theta):\theta\in\Theta_{i}\}$, for $i=0,1$ respectively.
@@
 $Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise.
-}\end{cases}$ (B.3)
+}\end{cases}$ (C.14)
 Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta)$, for $n<t$, $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta^{\prime})$, for $n\geq t$ and $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (4.8), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
+We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in earlier, with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
 For Poisson distribution as well, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly for the same detection algorithms, similar to the Gaussian setting discussed in Remark 4.5.
@@
 All the maximum are taken over a finite set; hence, $U_{t}^{j}$ can be computed easily.
-B.3 Experimental results
-B.3.1 Known pre- and post-change (Setting I)
-We consider Poisson the mean-change problem: $\text{Pois}(1)$ (pre-change) vs $\text{Pois}(2)$ (post-change) with the true (unknown) changepoint at $T=100,500$.
-We employ the CUSUM change detection method defined in (C.1) with thresholds at $A=1000,10000$.
-After detection, we employ Algorithm 2 to construct confidence sets for $T$.
-5.1 and B.1 provide visualizations of the confidence sets and point estimates (2.7) of the changepoints across $5$ runs.
-Figure B.1: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois$(1)$ and the rest from Pois$(2)$, with $T=100$ (top row) or $T=500$ (bottom row).
-Change is detected using CUSUM detector (C.1) with $A=10000$ (top row) or $A=1000$ (bottom row).
-The point estimate (2.7) is shown in vertical red dashed line and the confidence set (using Algorithm 2) is shown in red points.
+Appendix D Experiments for confidence intervals for pre/post change parameters
+D.1 Setting II
+Table D.1 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (5.1)) for $\theta_{1}$ across 500500500 independent runs with $\alpha=\beta=\eta=0.05$.
+The data generation and change detection procedure remain the same as in the experiment described in Setting II of Section 6.
+Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of 0.850.850.85, indicating that our method is reliable but conservative.
+Table D.1: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change.
+The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column of the table.
+(Setting II)
+T $\Theta_{1}$
+Conditional coverage
+Marginal coverage
+Conditional Length Conditional CI
+100 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.98 1.37 [0.76, 2.13]
+100 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.96 1.23 [0.90, 2.16]
+500 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.99 0.93 1.48 [0.75, 2.23]
+500 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.91 1.31 [0.90, 2.21]
+D.2 Setting III
+Table D.2 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), which are defined in (5.2) and (5.1)), for $\theta_{0}$ and $\theta_{1}$ across 500500500 independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$.
+he data generation and change detection procedure remain the same as in the experiment described in Setting III of Section 6.
+Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of 0.800.800.80.
+Table D.2: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.850.85(Setting III): unknown pre- and post-change.
+The true (unknown) values of the parameters are $\theta_{0}=0$ and $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns of the table.
+T
+$\Theta_{0}$
+$\Theta_{1}$
+Cond. coverage
+Marg. coverage
+Cond. length Cond. CI
+$\theta_{0}$ $\theta_{1}$ $\theta_{0}$ $\theta_{1}$ $\theta_{0}$ $\theta_{1}$ $\theta_{0}$ $\theta_{1}$
+100 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 1.00 0.99 1.00 0.99 0.55 1.19 [-0.32, 0.23] [0.76, 1.95]
+100 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.99 0.98 0.98 0.61 1.16 [-0.51, 0.10] [0.90, 2.06]
+500 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.99 0.99 0.93 0.93 0.21 1.22 [-0.10, 0.11] [0.75, 1.97]
+500 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.99 0.99 0.94 0.94 0.19 1.24 [-0.09, 0.10] [0.90, 2.14]
+D.3 Comparison with [36]
+Next, we compare our confidence interval for the post-change parameter with that proposed by [39] and discussed in [36].
+We implement Wu’s interval as discussed in Chapter 4.3 of [36] (or Section of [39]), replacing $\theta$ by $\hat{\theta}$ in the bias and variance terms, which is given by $\hat{\theta}-\frac{4}{5\sqrt{d\hat{\theta}(\tau-\hat{\nu})}}\pm z_{\alpha/2}\frac{\sqrt{1-{13}/(16d\hat{\theta})}}{\sqrt{\tau-\hat{\nu}}}$, where $\hat{\theta}=\frac{T_{\tau}}{\tau-\hat{\nu}}$, $\tau=\tau^{\prime}_{\text{cusum}}$, and $z_{\alpha^{\prime}/2}$ is the $\alpha^{\prime}/2$ quantile of the standard normal distribution.
+We use CUSUM [18] (the same change detector as in the last experiment of Section 6), with $d=10$ for both methods.
+For our method, we assume the post-change space to be $\Theta_{1}=[0,\infty)$ with $\alpha=\beta=\eta=0.05$.
+We report the average size and the conditional and unconditional coverage rates of both methods across 500500500 independent runs in Table D.3, with a target conditional coverage of $\alpha^{\prime}=0.85$.
+The results show that the existing method [36], which is based on asymptotic approximations, undercovers $\theta_{1}$ (both conditionally and marginally).
+At the same time, our finite-sample confidence intervals are conservative — achieving higher-than-nominal (conditional) coverage at the expense of larger (conditional) length.
+Table D.3: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage 0.850.850.85) with Wu [36] (Setting II)
+$F_{0}$ $F_{1}$ $T$ Conditional Length Conditional Coverage Marginal Coverage
+Ours Wu (2007) Ours Wu (2007) Ours Wu (2007)
+${N}(-0.25,1)$ ${N}(0.25,1)$ 100 1.46 0.47 0.99 0.77 0.98 0.74
+${N}(-0.3,1)$ ${N}(0.3,1)$ 100 1.43 0.51 0.98 0.76 0.98 0.75
+${N}(-0.25,1)$ ${N}(0.25,1)$ 500 1.48 0.49 0.99 0.78 0.91 0.64
+${N}(-0.3,1)$ ${N}(0.3,1)$ 500 1.49 0.51 0.99 0.79 0.91 0.71
+Appendix E Additional experiments for composite pre/post change
+In this section, we present experimental results for the alternative approach in which each set $\mathcal{S}_{t-1}$ and/or $\mathcal{S}_{\tau-t+1}^{\prime}$ is discretized using 50 equispaced grid points.
+Let, $\{\theta_{1}^{(0)},\cdots,\theta_{I}^{(0)}\}\subseteq\mathcal{S}_{t-1}$ and $\{\theta_{1}^{(1)},\cdots,\theta_{K}^{(1)}\}\subseteq\mathcal{S}_{\tau-t+1}^{\prime}$ be the grid points.
+Here, $I=K=50$.
+Therefore, for $j\in\{1,\cdots,B\}$, we only need to generate $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$. For $i\in[I],k\in[K]$, define
+$Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})=\begin{cases}\theta_{i}^{(0)}+\epsilon_{n}^{j},\text{ if }n<t\\ \theta_{k}^{(1)}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (E.1)
+We run $\mathcal{A}$ on $\{Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{n}$ and denote $\tau_{i,k}^{j}$ as the detection time.
+We compute the test statistics based on $\tau_{i,k}^{j}$ and $Y_{1}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)}),\cdots,Y_{\tau_{i,k}^{j}}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})$.
+Then, we consider the following confidence set:
+$\mathcal{C}=\left\{t\in\{1,\cdots,\tau\}:M_{t}\leq\max_{i\in[I],k\in[K]}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{j=1}^{B})\right\}$ (E.2)
+As a special case, when $\theta_{0}$ is known, then, $\theta_{0}=\theta_{1}^{(0)}=\cdots=\theta_{I}^{(0)}=\mathcal{S}_{t-1}.$
+As shown in Table E.1, the confidence sets (E.2) have lengths and coverages that are comparable to those from the earlier method reported in Table 6.4 and Table 6.6.
+However, this grid-based approach incurs significantly higher computational cost.
+Therefore, for practical purposes, we recommend using the earlier, more computationally efficient method.
+Table E.1: Composite pre/post change
+T
+Pre-change
+Class
+Post-change
+Class
+Conditional
+coverage
+Unconditional
+coverage
+Size Time (s)
+100 $N(0,1)$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 27.05 365.81
+100 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 24.62 360.48
+100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 30.88 17572.26
+100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 30.14 17092.11
+Appendix F Simpler confidence sets for detection algorithms controlling PFA
+In this section, we shall discuss how one can construct simpler confidence sets when the change detection algorithm $\mathcal{A}$ controls the probability of false alarm (PFA) at level $\delta\in(0,1)$.
+Importantly, this method allows us to avoid Assumption 3.3, and our results hold in the fully nonparametric case with arbitrary pre- and post-change classes, $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$, respectively.
+We consider the test statistic $M_{t}$ introduced in (3.4), and define the following confidence set for $T$:
+$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha}\right\},$ (F.1)
+which is a simplified version of Equation 3.8, obtained by directly replacing $\alpha r_{t}^{*}$ with $\alpha$.
+This avoids the need to compute $r_{t}^{*}$ (and hence $\tau_{1}^{*},\cdots,\tau_{N}^{*}$), making the method easier to implement and free of Assumption 3.3.
+We now investigate the coverage properties of the above confidence set.
+Theorem F.1.
+Let $\mathcal{C}$ be the confidence set as defined in (F.1), for some $\alpha\in(0,1)$.
+If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, then for any $F_{0}\in\mathcal{P}_{0}$ and $F_{1}\in\mathcal{P}_{1}$,
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha-\delta\leavevmode\nobreak\ \text{, and }\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{1-\delta},$
+where the latter guarantee holds if $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,$ for all $F_{0}\in\mathcal{P}_{0}.$
+Proof.
+From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha}]\leq\alpha$. Hence,
+$\mathbb{P}_{F_{0},T,F_{1}}(M_{T}<2/\alpha)$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}<2/\alpha\})$
+$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}<2/\alpha\},\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}<2/\alpha\})+\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$
+$\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})+\delta,$
+where the last inequality follows from the fact that the PFA is less than or equal to $\delta$. Thus, we obtain
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha-\delta.$
+Now,
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha}{1-\delta},$
+A similar simplification can be applied to the adaptive confidence sets introduced in Section 4 for the parametric setting.
+By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 4.1, 4.4 and 4.5, we obtain
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}\}_{j=1}^{B})\right\},$ (F.2)
+for known pre- and post-change setting. Similarly, for the known pre-change and composite post-change setting,
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}_{\tau-t+1}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})\right\},$ (F.3)
+And finally,
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (F.4)
+for composite pre- and post-change setting.
+For detection algorithms with PFA $\leq\delta,\leavevmode\nobreak\ \delta\in(0,1)$, these simpler forms retain the coverage guarantees — both conditional and unconditional — similar to what we have shown in the theorem above.
+In the next section, we present the experimental results with the bounded PFA algorithms, along with some other experiments.
+Appendix G Some additional experimental results
+G.1 Experiments with dependent pre- and post-change data
+[…]
+We use the likelihood ratio-based cusum detector:
+$\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(j:n)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(j:n)}\geq A\right\},$ (G.1)
+where $n_{11}(j:n)$ denotes the number of transitions from 1 to 1 and $n_{10}(j:n)$ denotes the number of transitions from 0 to 1 within $X_{j},\cdots,X_{n}$, and $A=1000$.
+After detection, we compute the adaptive confidence sets (4.1) with the likelihood ratio-based test statistic (3.2) and point estimate as in (3.1).
+Note that our adaptive method (4.1) is the only one we know that is valid for dependent data.
+Table G.1 contains the results averaged across 500500500 runs.
+Table G.1: Known pre- and post-change with dependency (Setting I)
+T
+Conditional
+coverage
+Marginal
+coverage
+Size
+$|\hat{T}-T|$
+Delay
+100 0.99 0.98 36.12 4.35 28.99
+500 0.98 0.98 35.83 4.42 27.36
+G.2 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
+G.2.1 Known pre- and post-change (Setting A, PFA 0.001$\leq$ 0.001≤ 0.001)
+We perform experiments for both Gaussian and Poisson mean-change scenarios having the true changepoint at $T=100,500$, using a likelihood ratio detector defined as
+$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (G.2)
+[…]
+Table G.2: Known pre- and post-change distributions (Setting A, PFA 0.001$\leq$ 0.001≤ 0.001)
+$T$ $F_{0}$ $F_{1}$ Conditional Coverage Conditional Size
+$|\hat{T}-T|$
+Delay
+Universal Adaptive Universal Adaptive (Conditional) (Conditional)
+100 $N(0,1)$ $N(1,1)$ 0.95 0.94 15.44 12.60 2.74 110.91
+[…]
+500 $N(0,1)$ $N(1,1)$ 0.97 0.97 15.92 12.65 2.24 514.12
+[…]
+G.2.2 Known pre- and unknown post-change (Setting B, PFA 0.001$\leq$ 0.001≤ 0.001)
+We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector defined as
+$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (G.3)
+with thresholds at $A=1000$.
+The weights are chosen just as described in Table 6.2.
+Similarly, the PFA of this detector is at most $1/A$, i.e., 0.0010.0010.001 here.
+After detection, we construct our universal confidence sets (F.1) for $T$, with $\alpha=0.075$.
+The test statistic and the point estimate are the same as mentioned in Setting II of Section 6.
+We report the average size and the conditional and unconditional coverage rates across 500500500 independent runs in Table G.3.
+We also report the average absolute deviation of $\hat{T}$ in (B.2) and delay (given no false detection) in the table.
+Table G.3: Known pre-change but unknown post-change (Setting B, PFA 0.001$\leq$ 0.001≤ 0.001)
+T
+Composite
+post-change
+Conditional
+coverage
+Unconditional
+coverage
+Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
+100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 23.82 3.68 118.76
+100 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 20.15 3.36 115.64
+500 $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 23.15 3.18 530.83
+500 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 19.27 3.23 528.92
+G.2.3 Composite pre- and post-change (Setting C, PFA 0.001$\leq$ 0.001≤ 0.001)
+We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) [15, 8, 14] being the representative of the pre-change class) defined as
+$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (G.4)
+[…]
+Table G.4: Composite pre- and post-change (Setting C, PFA 0.001$\leq$ 0.001≤ 0.001)
+T
+Composite
+pre-change
+Composite
+post-change
+Conditional
+coverage
+Marginal
+coverage
+Size
+$|\hat{T}-T|$
+Delay
+100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 26.02 3.97 125.64
+100 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 21.96 3.88 121.28
+500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 25.18 3.82 544.91
+500 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 0.98 0.98 21.45 3.75 536.76
+G.2.4 Huber-robust (nonparametric) pre- and post-change (Setting D, PFA 0.001$\leq$ 0.001≤ 0.001)
+We perform experiments for the robust Gaussian mean-change problem under the total variation contamination model, with the true changepoint at $T=100$ and $T=500$.
+The pre- and post-change classes are total variation balls of radius $\epsilon$ around $N(\mu_{0},1)$ and $N(\mu_{1},1)$, respectively.
+$\mathcal{P}_{i}=\{F\in\mathcal{M}:d_{TV}(F,N(\mu_{i},1))\leq\epsilon\};\leavevmode\nobreak\ \leavevmode\nobreak\ i=0,1.$
+In our experiment, $\mu_{0}=0,\mu_{1}=1$ and $\epsilon=0.01,0.001$.
+The true (unknown) pre- and post-change distributions are $(1-\epsilon)\times N(0,1)+\epsilon\times\text{Cauchy}(-1,10)$ and $(1-\epsilon)\times N(1,1)+\epsilon\times\text{Cauchy}(-1,10)$ respectively.
+We use the Huber-robust tests by [23] as change-detectors:
+$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}E_{i}\geq A\right\},$ (G.5)
+[…]
+After detection, we construct our universal confidence sets (F.1) for $T$, with $R_{n}^{(t)}=\prod_{i=t}^{n}E^{\prime}_{i}$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}E_{i}$, for $n<t$ and $\alpha=0.05$, where $E^{\prime}_{i}=\frac{\pi^{\prime}(X_{i})}{\mathbb{E}_{F_{1}}\pi^{\prime}(X_{i})+(c^{\prime\prime}-c^{\prime})\epsilon},$ $\pi^{\prime}(x)=\min\{\max\{\frac{f_{0}(x)}{f_{1}(x)},c^{\prime}\},c^{\prime\prime}\}$ and $E_{i}$ is as defined before.
+We report the average size and coverage rates across 200200200 independent runs in Table G.5.
+We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection).
+The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
+Table G.5: Huber-robust (nonparametric) pre- and post-change (Setting D, PFA 0.001$\leq$ 0.001≤ 0.001)
+T $\epsilon$
+Conditional
+coverage
+Marginal
+coverage
+Size
+$|\hat{T}-T|$
+Delay
+100 0.01 0.97 0.97 16.04 3.23 119.38
+100 0.001 0.97 0.97 14.57 3.18 113.06
+500 0.01 0.96 0.96 16.18 3.87 536.58
+500 0.001 0.97 0.97 15.13 3.25 531.44
+G.2.5 Comparison of experimental results with ARL and PFA controlling detection algorithms
+As expected, algorithms that control PFA exhibit a significantly higher average detection delay compared to those controlling ARL.
+This delay arises because stricter false alarm constraints force the detection algorithm to accumulate stronger evidence before signaling a change.
+In terms of inference, we observe that PFA-controlling algorithms often tend to yield slightly lower conditional coverage probabilities with confidence sets of comparable sizes.
+Also, conditional and unconditional coverage rates are identical when PFA-controlling detection algorithms are employed, due to the very low probability of false detections, while the ARL controlling algorithms lead to lower unconditional conditional coverage rates for $T=500$, supporting our theory.
+Moreover, we observe that the average absolute deviations of the point estimates are barely smaller with the PFA controlling algorithms.
+Overall, the major differences between these two types lie in detection delay and unconditional coverage properties.
+We do not observe significant differences in the size and conditional coverage properties of the confidence sets nor in the absolute deviation of the point estimates, across these two types of detection algorithms.
+G.3 Experimental results with Poisson distribution with detection algorithms that control ARL
+G.3.1 Known pre- and post-change (Setting I)
+[…]
+Figure G.6: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois(1)(1)( 1 ) and the rest from Pois(2)(2)( 2 ), with $T=100,500$.
+The point estimate is shown in a vertical red dashed line, and the confidence set (using Algorithm 2) is shown in red points.
 $B=N=100$, $\alpha=0.05,L=\infty$.
-Results of $5$ independent simulations are shown.
-We report the average size and the conditional and unconditional coverage rates across $100$ independent runs in Table B.2.
-Table B.2: Setting I: Average (of $100$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (2.7) and average detection delay (given a true detection) of the detection algorithm used (C.1) with threshold $A$, while varying $T$ and $A$.
+Results of 555 independent simulations are shown.
+We report the average size and the conditional and unconditional coverage rates across 500500500 independent runs in Table G.7.
+Table G.7: Setting I: Average (of 500500500 independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ and average detection delay (given a true detection) of the detection algorithm used (LABEL:eq:cusum) with threshold $A$, while varying $T$ and $A$.
 $T$ $A$
@@
 Nevertheless, the empirical conditional coverage rates meet the desired level, supporting our theory.
-We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
-B.3.2 Known pre-change and composite post-change (Setting II)
-We perform experiments for Poisson rate-change scenarios having pre-change distribution as $\text{Pois}(1)$, and the true (unknown) post-change data distribution as $\text{Pois}(2)$, with the changepoint at $T=100,200,$ assuming the composite post-change model to be $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$.
-For detection, we employ a weighted CUSUM detector (3.15) with thresholds at $A=1000,10000$.
-For weighted CUSUM, we employ a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, and $\theta=1.9$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
-After detection, we construct confidence sets (as defined in (3.11)) for $T$.
-We report the average size and the conditional and unconditional coverage across $100$ independent runs in (5.4).
+We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table.
+G.3.2 Known pre-change and composite post-change (Setting II)
+[…]
+$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (G.6)
+[…]
+We report the average size and the conditional and unconditional coverage across 500500500 independent runs in (G.8).
 The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table B.2).
-We also report the average absolute deviation of $\hat{T}$ (defined in (2.7)) and delay (given no false detection) in the table.
+It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table G.7).
+We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table.
 We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
-Table B.3: Setting II: Average (of $100$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets (3.11) with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ (3.5) and average detection delay (given no false detection) of the change detection method used (3.15) with threshold $A$, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
+[…]
 T A
@@
 200 1000 0.97 0.93 35.24 5.74 21.27
-B.3.3 Composite pre- and post-change (Setting III)
-Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $200$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively.
-For change detection, we employ a weighted CUSUM-type detector (defined in (4.17)), where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, and for the post-change parameter, the same discrete weight distribution as in Section B.3.2 is considered.
-After detection, we construct confidence sets (as defined in (4.12)) for $T$ (assuming $T\neq 1$).
-We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$ (defined in (2.7)), and delay (given no false detection) across $100$ independent runs in Table B.4.
-Table B.4: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets (4.12) with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ (4.6) and average detection delay (given no false detection) of the change detection method used (4.17) with threshold $A$, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
+G.3.3 Composite pre- and post-change (Setting III)
+Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and 500500500, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively.
+For change detection, we employ a weighted CUSUM-type detector, where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, which is defined as
+$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (G.7)
+where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For the post-change parameter, the same discrete weight distribution as in Section G.3.2 is considered.
+After detection, we construct confidence sets (as defined in Section C.2, assuming $T\neq 1$), with $\alpha=0.05,\beta=\gamma=0.025,B=N=100,L=\infty$.
+We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$, and delay (given no false detection) across 500500500 independent runs in Table G.9.
+[…]
 T A
@@
 However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table B.2).
-We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table B.2), exhibiting both higher coverage and larger size.
-Appendix C Simulations for point estimates
-In this section, we study the empirical performances of our point estimates of the changepoint defined in Equations 2.7, 3.5 and 4.6, which are the sequential analogs of the maximum likelihood estimators in the offline setting.
-Our empirical findings indicate that these estimators perform well in practice, even though we do not have theoretical guarantees on their properties.
-C.1 Known pre- and post-change distributions
-In this experiment, $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$ For detecting the change, we use the usual cumulative sum (CUSUM) and Shiryaev-Roberts (SR) procedures [23] as defined below:
-$\tau_{\text{cusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak$ (C.1)
-$\tau_{\text{sr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (C.2)
-where $A$ is some fixed threshold.
-For our experiments, we set the value of $A$ at $10000$.
-Once the change is detected, we obtain $\hat{T}$, as defined in (2.7), based on observations up to the detection time.
-Fig.
-C.1 shows histograms of the values taken by $\hat{T}$ over $10000$ independent experiments.
-Figure C.1: Histogram of $\hat{T}$, defined in (2.7), for known pre- and post-change over $10000$ independent experiments with $F_{0}=N(0,1),F_{1}=N(1,1)$ and $T=100.$
-C.2 Known pre-change and composite post-change distributions
-In this experiment, $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta>0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
-Here, we use the weighted CUSUM/SR detectors defined in (3.15) and (3.16).
-We considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
-Fig.
-C.2 shows histograms of the values taken by $\hat{T}$ defined in (3.5) over $10000$ independent experiments.
-Figure C.2: Histogram of $\hat{T}$, defined in (3.5), for known pre-change and unknown post-change over $10000$ independent experiments with $F_{0}=N(0,1),\mathcal{P}_{1}=\{N(\theta,1):\theta\neq 0\}$ and $T=100.$ True post-change distribution is $F_{1}=N(1,1)$ (unknown).
-C.3 Composite pre- and post-change distributions
-In this experiment, $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively.
-We use the weighted CUSM/SR-type change detectors defined in (4.17) and (4.18).
-In this experiment, we considered a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ for $\theta=0.6,d=0.2$ with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ Fig.
-C.3 shows histograms of the values taken by $\hat{T}$ defined in (4.6) over $10000$ independent experiments.
-It is crucial to note that the RIPr-based estimator has a higher bias but lower variance.
-Figure C.3: Histogram of $\hat{T}$, defined in (4.6), for unknown pre- and post-change over $10000$ independent experiments with $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq 0.5\},\mathcal{P}_{1}=\{N(\theta,1):\theta>0.5\}$ and $T=100.$ True (unknown) pre- and post-change distributions are $F_{0}=N(0,1)$ and $F_{1}=N(1,1)$ respectively.
-Appendix D A heuristic method for composite pre- or post-change
-This section provides an alternative method for constructing confidence intervals for composite pre/post-change distributions.
-While this approach lacks theoretical guarantees, it performs slightly better in our experiments than our previous approach.
+It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table G.7).
+We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table G.7), exhibiting both higher coverage and larger size.
+Appendix H A heuristic method for composite pre- or post-change
+This section presents an alternative method for constructing confidence intervals for composite pre- and post-change distributions.
 Specifically, we propose a heuristic bootstrap-based approach that estimates the unknown distributions from the observed data and subsequently generates simulations from the estimated distributions.
-D.1 Known pre-change and unknown post-change distributions
+H.1 Known pre-change and unknown post-change distributions
 Here we assume that the pre-change distribution $F_{0}$ is known and the composite post-change model is $\mathcal{P}_{1}$.
@@
 $\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},\text{ for some (unknown) }F_{1}\in\mathcal{P}_{1}.$
-For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined in (2.8)).
+For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$.
 Since the pre-change $F_{0}$ is known, we can generate $N$ many i.i.d.
@@
 To address this, we employ a simple bootstrap-based approach.
-The main idea is to first obtain some point estimate of the changepoint (for example, a likelihood ratio based estimate, as defined in (3.5)).
+The main idea is to first obtain some point estimate of the changepoint (for example, a likelihood ratio based estimate, as defined earlier).
 We denote the estimate as $\hat{T}$.
@@
 Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 4 contains an overview of the method.
-Input: $\alpha$, $N,B$, change detection algorithm $\mathcal{A}$ and data sequence $X_{1},X_{2},\cdots$ until change is detected at $\tau$ using $\mathcal{A}$
-Output: A confidence set $\mathcal{C}$ for changepoint $T$
-1 Compute a point estimator $\hat{T}$ for $T$, (eg., as defined in (3.5))
-2 Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
-3 for $i=1,\cdots,N$ do
-4       Simulate $X_{1}^{i},X_{2}^{i}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first.
-5
-6 end for
+Algorithm 4 CI for known pre-change $F_{0}$ and composite post-change $\mathcal{P}_{1}$
 […]
-8 for $t=1,\cdots,\tau$ do
-9       Compute some test statistics $M_{t}$ (eg., as defined in (3.7)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
-10       for $j=1,\cdots,B$ do
-11             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
-12             Compute the same test statistics $M_{t}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}$
-13
-14       end for
-15      if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
-16            $\mathcal{C}=\mathcal{C}\cup\{t\}$
-17       end if
-18
-19 end for
-Algorithm 5 CI for known pre-change $F_{0}$ and composite post-change $\mathcal{P}_{1}$
-D.2 Composite pre- and post-change distributions
+2:A confidence set $\mathcal{C}$ for changepoint $T$
+3:Compute a point estimator $\hat{T}$ for $T$
+4:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
+5:for $i=1$ to $N$ do
+6:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
+7:end for
+[…]
+9:for $t=1$ to $\tau$ do
+10:  Compute some test statistics $M_{t}$ based on $\tau$ and $X_{1},\cdots,X_{\tau}$
+11:  for $j=1$ to $B$ do
+12:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
+13:   Compute the same test statistics $M_{t}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}$
+14:  end for
+15:  if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
+16:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
+17:  end if
+18:end for
+H.2 Composite pre- and post-change distributions
 We now consider the general case where both pre and post-change distributions follow some composite models $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively.
@@
 $\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0},\text{ for some }F_{0}\in\mathcal{P}_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},\text{ for some }F_{1}\in\mathcal{P}_{1}.$
-For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined in (4.8)).
+For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined earier).
 As in Algorithm 4, we employ a simple bootstrap-based approach, but here we need to estimate both pre- and post-change distributions.
-We first obtain some point estimate $\hat{T}$ of the changepoint $T$ (eg., the likelihhod-ratio-based estimate as defined in (4.6)).
-Then we use the observations from $1$ to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$.
+We first obtain some point estimate $\hat{T}$ of the changepoint $T$.
+Then we use the observations from 111 to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$.
 For some $N\in\mathbb{N}$, we generate $N$ many i.i.d.
@@
 Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 5 contains an overview of the method.
-Input: $\alpha$, $N,B$, change detection algorithm $\mathcal{A}$ and data sequence $X_{1},X_{2},\cdots$ until change is detected at $\tau$ using $\mathcal{A}$
-Output: A confidence set $\mathcal{C}$ for changepoint $T$
-1 Compute some point estimator $\hat{T}$ for $T$ (eg., as defined in (4.6))
-2 Obtain an estimate $\hat{F}_{0}$ of the (unknown) pre-change $F_{0}\in\mathcal{P}_{0}$ based on $X_{1},\cdots,X_{\hat{T}-1}$
-3 Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
-4 for $i=1,\cdots,N$ do
-5       Simulate $X_{1}^{i},X_{2}^{i}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first.
-6
-7 end for
+Algorithm 5 CI for unknown pre- and post-change
 […]
-9 for $t=1,\cdots,\tau$ do
-10       Compute some test statistics $M_{t}$ (eg., as defined in (4.8)) based on $X_{1},\cdots,X_{\tau}$
-11       for $j=1,\cdots,B$ do
-12             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
-13             Compute the same test statistics $M_{t}^{j}$ based on $X_{1}^{j},\cdots,X_{\tau_{j}}^{j}$
-14
-15       end for
-16      if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
-17            $\mathcal{C}=\mathcal{C}\cup\{t\}$
-18       end if
-19
-20 end for
-Algorithm 6 CI for unknown pre- and post-change
-D.3 Experimental Results
-D.3.1 Known pre-change and composite post-change (Setting II)
+2:A confidence set $\mathcal{C}$ for changepoint $T$
+3:Compute some point estimator $\hat{T}$ for $T$
+4:Obtain an estimate $\hat{F}_{0}$ of the (unknown) pre-change $F_{0}\in\mathcal{P}_{0}$ based on $X_{1},\cdots,X_{\hat{T}-1}$
+5:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
+6:for $i=1$ to $N$ do
+7:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
+8:end for
+[…]
+10:for $t=1$ to $\tau$ do
+11:  Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$
+12:  for $j=1$ to $B$ do
+13:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
+14:   Compute the same test statistics $M_{t}^{j}$ based on $X_{1}^{j},\cdots,X_{\tau_{j}}^{j}$
+15:  end for
+16:  if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
+17:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
+18:  end if
+19:end for
+H.3 Experimental Results
+H.3.1 Known pre-change and composite post-change (Setting II)
 We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$.
-For detection, we employ a weighted CUSUM detector (3.15) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
+For detection, we employ a weighted CUSUM detector (B.11) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
 After detection, we employ Algorithm 4 to construct confidence sets.
-Table D.1 summarizes the results, including the average size and the conditional and unconditional coverage across $100$ independent runs.
+Table H.1 summarizes the results, including the average size and the conditional and unconditional coverage across 500500500 independent runs.
 For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible.
@@
 Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
-Table D.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.15) with threshold $A=1000$.
+Table H.1: Setting II: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 4 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (B.11) with threshold $A=1000$.
 T Pre-change
@@
 100 $N(0,1)$$\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.94 0.94 14.76
-100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.95 20.95
+[…]
 500 $N(0,1)$$\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.95 0.91 14.06
-500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.89 20.83
-D.3.2 Composite pre- and post-change (Setting III)
-Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$.
-We use a weighted CUSUM detector with RIPr defined in (4.17) for change detection.
+[…]
+H.3.2 Composite pre- and post-change (Setting III)
+Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and 500500500.
+We use a weighted CUSUM detector with RIPr defined in (C.12) for change detection.
 The thresholds are set at $A=1000$ for both $T=100$ and $T=500$.
@@
 After detection, we employ Algorithm 5 to construct confidence sets.
-We report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.2.
+We report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table H.2.
 The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
-Table D.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 6 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
+Table H.2: Setting III: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (C.12) with threshold $A=1000$.
 T
@@
 100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 19.52
-100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.94 0.94 23.32
+[…]
 500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 18.56
-500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 23.04
-Appendix E Another heuristic method for composite pre- or post-change
+[…]
+Appendix I Another heuristic method for composite pre- or post-change
 In this section, we provide another heuristic method for dealing with composite pre- or post-change distributions.
@@
 Nonetheless, we do not have any theoretical guarantee for coverage for this method.
-For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section D.3.1 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table E.1.
-Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section D.3.2 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table E.1.
-From Table E.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
-Table E.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.15) with threshold $A=1000$.
+For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section H.3.1 and we report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table I.1.
+Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section H.3.2 and we report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table I.1.
+From Table I.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
+Table I.1: Setting II: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (B.11) with threshold $A=1000$.
 T Pre-change
@@
 100 $N(0,1)$$\{N(\mu,1):\mu>0.75\}$ $N(1,1)$ 0.95 0.95 23.16
-100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.94 0.94 25.83
+[…]
 500 $N(0,1)$$\{N(\mu,1):\mu>0.75\}$ $N(1,1)$ 0.95 0.88 23.82
-500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.90 27.17
-Table E.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
+[…]
+Table I.2: Setting III: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (C.12) with threshold $A=1000$.
 T
@@
 100 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 26.03
-100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 29.64
+[…]
 500 $\{N(\mu,1):\mu\leq 0.25\}$$\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 26.52
-500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 30.49
-While we do not have any theoretical guarantees for this method, it is computationally more efficient than the previous approaches with theoretical guarantees, as demonstrated in Table E.3.
-Summary of comparisons of heuristic and non-heuristic methods
-We first compare the execution times of the three types of algorithms and present the results in Table E.3 for the Gaussian mean shift setting.
-Algorithm 3 and Algorithm 4 are collectively referred to as Proposed Method, which comes with theoretical guarantees but requires more computational time compared to the heuristic approaches.
-The first heuristic approach, described in Algorithm 5 and Algorithm 6, is referred to as Heuristic 1.
-Finally, the heuristic introduced in this section (which is Algorithm 2, with $F_{0}$ and $F_{1}$ being the “closest” pair of distributions in the composite pre- and post-change models) is labeled as Heuristic 2.
-Table E.3: Average (of $10$ independent runs) execution times (in seconds) taken by different methods, with $N=B=100$, $\alpha=0.05$, and for the proposed method, $\beta=\gamma=0.05$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(1,1)$ respectively.
-T A
-Pre-change model
-Post-change model
-Proposed Method Heuristic 1 Heuristic 2
-100 1000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 41.25 10.43 9.72
-100 10000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 43.14 11.38 10.97
-100 1000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 66.13 11.49 11.76
-100 10000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 69.47 11.86 11.52
-500 1000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 364.18 143.98 148.08
-500 10000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 380.46 155.89 144.75
-500 1000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 535.82 171.13 167.46
-500 10000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 561.74 177.28 169.41
-From Table E.3, we observe that the computational times are lower for the heuristic algorithms.
-Moreover, comparing the corresponding tables, we observe that the theoretically valid method (Proposed Method) gives higher coverage, also higher size of the confidence set, as compared to both heuristic methods.
-This suggests that while the heuristic approaches offer practical benefits in terms of speed, they may lead to less conservative inference, despite not having theoretical guarantees.
-This points at a natural direction for future work: to establish guarantees for our heuristic algorithms, or demonstrate settings where they empirically do not have coverage.
+[…]

### v4

@@
 This paper addresses a fundamental but largely unexplored challenge in sequential changepoint analysis: conducting inference following a detected change.
-We construct confidence sets for the unknown changepoint using only the data observed up to a data-dependent stopping time at which an arbitrary sequential detection algorithm $\mathcal{A}$ declares a change.
-Our framework is nonparametric, makes no assumption on the composite post-change class.
-We extend it to the composite pre-change class under a suitable assumption.
-We impose no conditions on the observation space or on $\mathcal{A}$.
-We thus present the first general tool for sequential changepoint localization, which is theoretically sound and practically effective.
+We develop a very general framework to construct confidence sets for the unknown changepoint using only the data observed up to a data-dependent stopping time at which an arbitrary sequential detection algorithm declares a change.
+Our framework is nonparametric, making no assumption on the composite post-change class, the observation space, or the sequential detection procedure used, and is nonasymptotically valid.
+We also extend it to handle composite pre-change classes under a suitable assumption, and also derive confidence sets for the change magnitude in parametric settings.
+We provide theoretical guarantees on the width of the confidence intervals.
+Extensive simulations demonstrate that the produced sets have reasonable size, and slightly conservative coverage.
+In summary, we present the first general method for sequential changepoint localization, which is theoretically sound and broadly applicable in practice.
 1 Introduction
@@
 A sequence of $\mathcal{X}$-valued observations ${\bf X}=X_{1},X_{2},\cdots$ arrives sequentially, such that for some unknown time $T\in\mathbb{N}\cup\{\infty\}$ (the changepoint),
-$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1.1)
+$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1.1)
 for some $F_{0}$ in a pre-change class $\mathcal{P}_{0}$ and $F_{1}$ in a post-change class $\mathcal{P}_{1}$, where $\mathcal{P}_{0}$ and $\mathcal{P}_{0}$ denote some non-intersecting classes of probability distributions.
@@
 The fact that we do not require knowledge of internal details of $\mathcal{A}$ is a particularly important feature; it implies that our changepoint localization method can wrap around any existing workflow for changepoint detection — it could be employed both with simple classical statistical algorithms (like CUSUM), and also with complex modern pipelines in the information technology industry that may involve heuristic machine learning algorithms (to deal with complex data types).
+For instance, rapid shifts in public sentiment often occur on online platforms: a product failure may trigger a sudden surge of negative reviews, or a major event may abruptly alter the tone of social-media discussions.
+Detecting when such a shift occurs is vital for understanding emerging issues, tracing public reactions to events, or monitoring product reputation in real time.
+To illustrate the relevance of our methodology in such settings, we conduct an experiment using the Stanford Sentiment Treebank (SST-2) dataset [31], simulating a stream of text reviews whose sentiment abruptly transitions from predominantly positive to predominantly negative.
+Accurately localizing the changepoint is important: it can help investigate why the change may have occurred.
 As an initial goal, one may hope to obtain a set $\mathcal{C}\subseteq\{1,\dots,\tau\}$ such that
@@
 This gap is significant because in most real-world applications, data arrives sequentially, and it can be essential to make timely inferences about the changepoint immediately upon detecting a change.
-[…]
+We are aware of only two directly related works: [5] and [39, Chapter 3].
+They also chose to aim for the guarantee (1.3) like us.
+Both [5, 39] provide asymptotically valid inference after the CUSUM procedure [19] declares a change for known $F_{0}$ and $F_{1}$ in an exponential family; the former only constructs a lower confidence bound for $T$ while the latter also provides a confidence set for $T$.
+Their approaches have notable methodological and theoretical limitations: (i) they are tailored specifically to the CUSUM procedure, (ii) they assume that the pre- and post-change distributions are both known and they belong to an exponential family, (iii) they only provide asymptotic coverage guarantees (valid when both $T\to\infty$ and ARL $\to\infty$).
+These constraints limit practical applicability, and their confidence set exhibits undercoverage in our experiments.
 In sharp contrast, our proposed confidence sets for $T$ are very general: We make no assumptions on the detection procedure $\mathcal{A}$, and our framework acts as a versatile wrapper that can be layered on top of any sequential change detection algorithm.
@@
 Known, Known $\mathcal{A}=$ CUSUM. $F_{0},F_{1}$ Asymptotic No assumption Nonasymptotic
-[…]
+($F_{0},F_{1}$) $\in$ exponential family ($T$, ARL $\to\infty$) (any $\mathcal{A},F_{0},F_{1}$) (Theorem 4.1)
 Known, Unknown No existing work No assumption Nonasymptotic
@@
 These adaptive methods allow the observations to be dependent.
-We establish corresponding coverage guarantees in Theorems 4.1, 4.4 and 4.5.
+We establish corresponding coverage guarantees in Theorems 4.1, 4.3 and 4.4.
   • As a byproduct of our methods, we derive confidence sets for the unknown pre- and post-change parameters and show their coverage guarantees in Theorem 5.1.
@@
 In this section, we highlight some key observations that shape our methodology of post-detection inference for the change-point $T$.
-We begin by clarifying some notational conventions that will be used throughout the paper.
 Notation.
 For $t\in\mathbb{N}$, let $\mathbb{P}_{F_{0},t,F_{1}}$ denote joint distribution of the sequence $\{X_{n}\}_{n}$ when $X_{1},\dots,X_{t-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$ and $X_{t},X_{t+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{1}$.
-333 To understand the key ideas in the paper, it will suffice for the reader to focus on the case when all $X_{i}$’s are all independent, and $F_{0}$ and $F_{1}$ denote the pre- and post-change marginal distributions.
+[^3] To understand the key ideas in the paper, it will suffice for the reader to focus on the case when all $X_{i}$’s are all independent, and $F_{0}$ and $F_{1}$ denote the pre- and post-change marginal distributions.
 However, our framework in Section 4 can actually handle dependent data.
@@
 If, for some $\alpha\in(0,1)$, a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}^{*})\geq 1-\alpha$, $\forall T\in\mathbb{N}$, then $\mathbb{P}_{F_{0},\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL is infinite).
-Said differently, for any algorithm with finite ARL, it is impossible to construct a $1-\alpha$ confidence set that is only a subset of $\{1,\dots,\tau\}$.
+The proposition says that for any algorithm with finite ARL, it is impossible to construct a $1-\alpha$ confidence set that is only a subset of $\{1,\dots,\tau\}$.
 Thus, for post-detection inference, the usual notion of coverage is not an appropriate one.
 Instead, we focus on conditional coverage (1.3), a more reasonable criterion, which was also considered in prior work [5, 39].
-The key idea behind constructing our confidence intervals is simple, but several details are subtle.
-For each candidate $t\in\{1,\cdots,\tau\}$, we test whether $t$ is the true changepoint from some pre-change class $\mathcal{P}_{0}$ to some post-change class $\mathcal{P}_{1}$:
-$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (2.1)
-Let $\phi_{\alpha^{*}}^{(t)}=\mathds{1}(H_{0,t}\text{ is rejected})$ denote some test for the null $H_{0,t}$ such that
-$\mathbb{P}_{F_{0},t,F_{1}}(\phi_{\alpha^{*}}^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha^{*},\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$ (2.2)
-and form the confidence set by inverting these tests:
+The key idea behind our methodology is that for each candidate $t\in\{1,\cdots,\tau\}$, we test whether $t$ is the true changepoint from some pre-change class $\mathcal{P}_{0}$ to some post-change class $\mathcal{P}_{1}$:
+$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (2.1)
+Let $\phi_{\alpha^{*}}^{(t)}=\mathds{1}(H_{0,t}\text{ is rejected})$ denote some test for the null $H_{0,t}$ such that $\mathbb{P}_{F_{0},t,F_{1}}(\phi_{\alpha^{*}}^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha^{*},\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$ and form the confidence set by inverting these tests:
 $\mathcal{C}:=\{t\in\mathbb{N}:t\leq\tau,H_{0,t}\text{ is not rejected}\}$ (2.2)
-[…]
+In classical statistics, inverting level-$\alpha$ tests yields $1-\alpha$ confidence sets.
+However, this relationship does not directly hold for us since we test the hypotheses up to a data-dependent time $\tau$, and we are considering conditional coverage.
 Proposition 2.2.
@@
 Conversely, given any test $\phi^{(t)}$ for $H_{0,t}$ using data up to $\tau$ satisfying (2.3), the inverted confidence set $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
-$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\leavevmode\nobreak\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.5)
-The above result establishes the duality between hypothesis tests and a post-detection confidence set.
-In the next section, we propose a general framework for constructing confidence sets for $T$ by inverting tests based on the above.
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\penalty 10000\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.4)
+The above result establishes a duality between tests and a post-detection confidence set. Next, we construct the latter by inverting tests based on the above.
 3 A general framework for nonparametric post-change
@@
 It simplifies to
-$M_{t}=\begin{cases}\prod_{i=t}^{\hat{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})},&\text{ if }t<\hat{T}\leq\tau<\infty\\ 1,&\text{ if }t=\hat{T}\leq\tau<\infty\\ \prod_{i=\hat{T}}^{t-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})},&\text{ if }\hat{T}<t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (3.2)
+[…]
 One natural way to extend this test statistic for the general nonparametric setting is to replace the likelihood ratio processes by e-processes [20, 8], which have emerged as a central tool in anytime-valid inference.
@@
 Such powered e-processes for $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ (and vice versa) exist if the total variation distance between the two convex hulls of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ is positive (see e.g., Section 3.4 of [22]).
-We require all $X_{i}$’s to be independent for both powered (nontrivial) e-processes to be adapted to their respective filtrations.
+The likelihood ratio process is the simplest example of an e-process.
+There is a substantial literature on constructing e-processes in both parametric (e.g., [36]) and nonparametric settings (e.g., [37]).
+For readers less familiar with these developments, we refer to Chapters 6 and 7 of the recent textbook [22].
+Note that we require all $X_{i}$’s to be independent for both powered (nontrivial) e-processes to be adapted to their respective filtrations.
 The final ingredient we need is $\hat{T}$, which is a preliminary point estimate of $T$.
 [^4] Here, we again note that our confidence set based on a point estimate $\hat{T}$ can be viewed analogously to the conformal prediction set, which is constructed around a preliminary “point prediction”.
-For example, one can construct such an estimator based on e-processes under $\mathcal{P}_{0}$.
-To elaborate, let $\{E^{(t)}_{n}\}_{n\geq t}$ be some forward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$).
+One can easily construct a point estimate.
+For example, let $\{S^{(t)}_{n}\}_{n\geq t}$ be some forward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$).
 Then, one reasonable estimate is
-$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau}E^{(t)}_{\tau}.$ (3.3)
+$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau}S^{(t)}_{\tau}$. (3.3)
 This can be viewed as a nonparametric generalization of the maximum likelihood estimator for known pre- and post-change defined in (3.1).
@@
 Therefore, we use a forward e-process $R^{(t)}_{T-1}$ that accumulates evidence in favor of $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ over this segment.
-Similarly, if $t>T$, we employ a backward e-process $S^{(t)}_{T}$ to accumulate evidence for $\mathcal{P}_{0}$ against $\mathcal{P}_{1}$.
-Since the true change-point $T$ is unknown, we replace it with an estimated change-point $\hat{T}$.
-This naturally leads us to define the following:
-$M_{t}=\begin{cases}R^{(t)}_{\hat{T}-1},&\text{ if }t<\hat{T}\leq\tau<\infty\\ 1,&\text{ if }t=\hat{T}\leq\tau<\infty\\ S^{(t)}_{\hat{T}},&\text{ if }\hat{T}<t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (3.4)
+Similarly, if $t>T$, we use a backward e-process $S^{(t)}_{T}$ to accumulate evidence for $\mathcal{P}_{0}$ against $\mathcal{P}_{1}$, using the segment $X_{T},\dots,X_{t-1}$ .
+Since the true $T$ is unknown, we replace it with an estimated change-point $\hat{T}$.
+Fig.
+3.1 provides an illustration of the key idea.
+This naturally leads us to define:
 […]
+Figure 3.1: Pre-change and post-change parts are shown in black and blue, respectively.
+Outside the time segment $t,\dots,{\hat{T}-1}]$ when $t<\hat{T}$ (or, $\hat{T},\dots,{t-1}]$ when $t>\hat{T}$) both $H_{0,t}$ and $H_{0,\hat{T}}$ agree on the distribution of the observations.
+Although, $M_{t}$ is not an e-variable under $H_{0,t}$ (nor are its constant multiples), thresholding $\frac{M_{t}}{2}$ at $\frac{1}{\alpha}$ yields a level-$\alpha$ test for $H_{0,t}$ — a nontrivial fact shown below.
 Theorem 3.1.
@@
 3.1 Confidence set for $T$ with known pre-change distribution
-Fix some $N\in\mathbb{N}$.
-For each $j=1,\dots,N$, simulate a sequence of i.i.d.
+Let $r_{t}$ be an unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, that is independent of the data. Then, define the confidence set as
+$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}}\right\}.$ (3.5)
+One natural approach to estimate $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$ is via Monte Carlo simulation.
+Fix $N\in\mathbb{N}$ and for each $j=1,\dots,N$, simulate a sequence of i.i.d.
 data points $Z_{1}^{(j)},Z_{2}^{(j)},\dots\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}$ until either the detection algorithm triggers or the sample size reaches $\tau$—whichever occurs first.
 Let $\tau_{j}:=\min\{\tau,\mathcal{A}(\{Z_{n}^{(j)}\}_{n})\}$ denote the stopping time for the $j$-th simulated sequence.
-We then estimate $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$ using
-$r_{t}=\frac{1}{N}\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t).$ (3.5)
-Define the confidence set as
-$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}}\right\}.$ (3.6)
+Then $r_{t}=\frac{1}{N}\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)$ is a natural unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$ ,but it can equal zero with positive probability.
+To avoid this, we fix some $r\geq 2$ and for each $t$, one might define $N_{t}$ to be a random index until $\tau_{j}\geq t$ happens exactly $r$ many times.
+Then $r_{t}=\frac{r-1}{N_{t}-1}$ is an unbiased estimator, which is strictly positive.
+Alternatively, one may use an asymptotically unbiased estimator with $N$, such as, $r_{t}=\frac{1}{N+1}\left(1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)\right)$, which guarantees strict positivity but only provides asymptotic $1-\alpha$ coverage guarantee as $N\to\infty$.
 The next theorem proves the validity of the above confidence set.
@@
 Suppose that the pre-change distribution is known, i.e., $\mathcal{P}_{0}=F_{0}$.
-Then, for any $\alpha\in(0,1),T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfying $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (3.6) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{1}\in\mathcal{P}_{1}.$
-Next, we extend it to composite pre-change under an additional assumption.
-3.2 A partial extension to composite pre-change
+Then, for any $\alpha\in(0,1),T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfying $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (3.5), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{1}\in\mathcal{P}_{1}.$
+Now we analyze the size of the confidence set (3.5) for known pre-change $F_{0}$ and post-change $F_{1}$, with $r_{t}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)}{N+1}$, $M_{t}$ as defined in (3.2) with $\hat{T}$ as defined in (3.1).
+We provide a bound on the expected size of the confidence set as $N\to\infty$, where $N$ is the number of simulations used to calculate $r_{t}$.
+Define, $p_{t}=\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$.
+Note that $r_{t}\to p_{t}$ as $N\to\infty$.
+Define $\rho_{0}(s)=\mathbb{E}_{F_{0}}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}$ and $\rho_{1}(s)=\mathbb{E}_{F_{1}}\left(\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right)^{s}$.
+These two quantities measure closeness between $F_{0}$ and $F_{1}$, and thus the problem hardness.
+Note that for $s=1/2,$ this coincides with Bhattacharyya distance.
+At $s=0,1$, $\rho_{0}$ and $\rho_{1}$ take value $1.$ Since for $s\in(0,1)$, $h(x)=x^{s}$ is a strictly concave function and since $F_{1}\neq F_{0}$, $\frac{f_{1}(X)}{f_{0}(X)}$ is not almost everywhere constant, we get a strict inequality using Jensen’s: $\rho_{0}(s)=\mathbb{E}_{F_{0}}\left[\left(\frac{f_{1}(X)}{f_{0}(X)}\right)^{s}\right]<\left(\mathbb{E}_{F_{0}}\left[\frac{f_{1}(X)}{f_{0}(X)}\right]\right)^{s}=1.$ Similarly, for $s\in(0,1),\rho_{1}(s)<1$.
+We now define $s_{0}=\arg\min_{s\in[0,1]}\rho_{0}(s)$ and $s_{1}=\arg\min_{s\in[0,1]}\rho_{1}(s)$, and note that the preceding facts imply that $0<s_{0}<1$ and $0<s_{1}<1$.
+Our length bound will depend on the expected detection delay, $\Delta_{F_{0},T,F_{1}}:=\mathbb{E}_{F_{0},T,F_{1}}(|\tau-T|\mid\tau\geq T).$ For typical detection algorithms, this quantity often scales like the logarithm of the average run length divided by the KL divergence between $F_{1},F_{0}$.
+Finally, we call the stopping time $\tau$ as sensitive if
+$\penalty 10000\ \mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]\geq\mathbb{P}_{F_{0},\infty}[\tau\leq t],$ (3.6)
+for all $t$ and $T$.
+Equivalently, $\tau$ under a change is stochastically smaller than $\tau$ under no change.
+We can expect all practically used detection algorithms to be sensitive, as they should stop later when a change is absent than when it is present.
+Theorem 3.3.
+In the setting described above, for any change detection method with stopping time $\tau$ satisfying and $p_{T}\neq 0$, the size of $\mathcal{C}$} can be bounded as:
+$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\cdot\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{F_{0},T,F_{1}}.$ (3.7)
+Additionally, if $\tau$ is sensitive in the sense of (3.6), the last term in (3.7) can be improved to $\Psi_{F_{0},T,F_{1}}:=\min\left\{\Delta_{F_{0},T,F_{1}},\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right)\right\}$.
+Remark 3.4.
+Suppose $F_{1}=N(\mu_{1},\sigma^{2})$ vs $F_{0}=N(\mu_{0},\sigma^{2})$ and we are using a CUSUM detector with the likelihood ratio (which satisfies the assumption) with average run length $A$.
+Consider the asymptotic regime when $A\to\infty$, and $T=CA$, for some $C>0$.
+Using the property of the CUSUM detector that $\frac{\tau}{\mathbb{E}_{F_{0},\infty}(\tau)}\stackrel{{\scriptstyle d}}{{\to}}\exp(1)$ as $A\to\infty$, we have $p_{T}=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq CA_{T})\to\exp(-C)$.
+Also, for the Gaussian distribution, it is easy to verify that $\rho_{0}(s)=\rho_{1}(s)=\exp\left(-s(1-s)K\right)$, where $K=\frac{(\mu_{1}-\mu_{0})^{2}}{2\sigma^{2}}$ and $s_{0}=s_{1}=1/2$.
+Substituting, we get
+$\limsup_{T,A\to\infty,T=CA}\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}\left(|\mathcal{C}|\mid\tau\geq T\right)\leq$ $\sqrt{\frac{2}{\alpha}}\left[\frac{\exp(\frac{3C}{2}-\frac{K}{4})}{1-\exp\left(-\frac{K}{4}\right)}+\frac{\exp\left(C-\frac{K}{8}\right)}{1-\exp\left(-\frac{K}{8}\right)}\right]$
+$+1+\frac{\exp(C-\frac{K}{4})}{1-\exp\left(-\frac{K}{4}\right)}.$
+To simplify further, consider the difficult regime of small $K$, where $\left(\exp\left(\frac{K}{4}\right)-1\right)\approx K/4$, and $\left(\exp\left(\frac{K}{8}\right)-1\right)\approx K/8$, and take $C=\frac{1}{3}\log(\frac{2}{\alpha})$, so that the above expression is bounded by, ignoring constants, $K^{-1}[\frac{1}{\alpha}+\frac{1}{\alpha^{5/6}}++\frac{1}{\alpha^{1/3}}]$.
+The fact that the bound grows with $C=T/A$ is empirically verified in Supplementary Section G.1.
+This behavior is not a limitation of our methodology; rather, it reflects a fundamental difficulty of the problem: $p_{t}$ decays as $C$ increases, and hence, by our duality result (Proposition 2.2), the tests corresponding to any confidence set will have smaller levels $\alpha p_{t}$, leading to reduced power and, consequently, wider confidence sets.
+3.2 An extension to composite pre-change
 To extend our method to composite pre-change classes $\mathcal{P}_{0}$, we require the following assumption on the change detection method and $\mathcal{P}_{0}$:
 Assumption 3.5.
-$\exists F_{0}^{*}\in\mathcal{P}_{0}$ such that $\inf_{F_{0}\in\mathcal{P}_{0}}\mathbb{P}_{F_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t)$, $\forall t\in\mathbb{N}$.
-This assumption posits the existence of a “least favorable” distribution $F_{0}^{*}$, in the sense that the detection algorithm is most prone to false alarms under $F_{0}^{*}$.
-We show that it holds for monotone likelihood ratio families with CUSUM/SR type change detectors when the pre- and post-change parameter spaces, $\Theta_{0}$ and $\Theta_{1}$, are ordered, e.g., $\Theta_{0}=(-\infty,a],\Theta_{1}=(b,\infty)$, for some $a<b$.
-[^5] For detection algorithms with bounded PFA, one can avoid Assumption 3.3 and obtain confidence sets with coverage guarantees without any additional assumption (i.e., for arbitrary $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$); see Section F of the Supplementary.
+$\exists P_{0}^{*}\in\mathcal{P}_{0}$ such that $\inf_{P_{0}\in\mathcal{P}_{0}}\mathbb{P}_{P_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{P_{0}^{*},\infty}(\tau\geq t)$, $\forall t\in\mathbb{N}$.
+We justify this assumption by appealing to Huber’s notion of “least favorable distribution pair” [13] in terms of risk, where the risk $R$ for a hypothesis test $\phi$ between $\mathcal{P}_{0}$ vs.
+$\mathcal{P}_{1}$ is defined as
+$R(P_{0},\phi)=C_{0}P_{0}(\phi\text{ accepts }\mathcal{P}_{1}),\penalty 10000\ \penalty 10000\ R(P_{1},\phi)=C_{1}P_{1}(\phi\text{ accepts }\mathcal{P}_{0}).$ (3.8)
+The formal definition of a least favorable distribution pair is given below.
+Definition 3.6.
+A pair $(P_{0}^{*},P_{1}^{*})\in\mathcal{P}_{0}\times\mathcal{P}_{1}$ is called least favorable distribution (LFD) pair in terms of risk $R$ for testing $\mathcal{P}_{0}$ vs.
+$\mathcal{P}_{1}$, if for every likelihood ratio test $\phi$ between $P_{0}^{*}$ and $P_{1}^{*}$, $R(P_{j}^{*},\phi)\geq R(P_{j},\phi),\penalty 10000\ \text{ for all }P_{j}\in\mathcal{P}_{j},\penalty 10000\ j=0,1.$
+The following proposition establishes that the existence of an LFD pair is sufficient for our assumption to hold.
 Proposition 3.7.
-Suppose that $\{F_{\theta}:\theta\in\Theta\}$ is some monotone likelihood ratio family with respect to some sufficient statistic for the parameter $\theta$ and $\sup\Theta_{0}\leq\inf\Theta_{1}$.
-For the CUSUM type stopping time $\tau$ defined in (C.12), and for any $\theta\leq\theta^{\prime}\leq\sup\Theta_{0}$, $\mathbb{P}_{\theta,\infty}(\tau\geq t)\geq\mathbb{P}_{\theta^{\prime},\infty}(\tau\geq t),$ and hence, $\theta_{0}^{*}=\sup\Theta_{0}$ satisfies Assumption 3.3.
-Similarly, if $\inf\Theta_{0}\geq\sup\Theta_{1}$, then $\theta_{0}^{*}=\inf\Theta_{0}$ satisfies Assumption 3.3.
-For each $j=1,\dots,N$, we now simulate a sequence of i.i.d.
-data points $Z_{1}^{(j)},Z_{2}^{(j)},\dots\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}^{*}$ until either the detection algorithm triggers or the sample size reaches $\tau$—whichever occurs first.
-Let $\tau_{j}^{*}:=\min\{\tau,\mathcal{A}(\{Z_{n}^{(j)}\}_{n})\}$ denote the stopping time for $\{Z_{n}^{(j)}\}_{n}$.
-We estimate the probability $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$ using
-$r_{t}^{*}=\frac{1}{N}\sum_{j=1}^{N}\mathds{1}(\tau_{j}^{*}\geq t).$ (3.7)
-Define the confidence set as
+$(P_{0}^{*},P_{1}^{*})$ is a least favourable distribution pair in terms of risk for testing $\mathcal{P}_{0}$ vs.
+$\mathcal{P}_{1}$, then Assumption 3.5 holds with $F^{*}=P^{*}$, and $\tau$ being the stopping time of a CUSUM/SR change detector with the likelihood ratio of $P_{1}^{*}$ to $P_{0}^{*}$.
+LFD pairs are known to exist for various parametric and nonparametric models, e.g., the monotone likelihood ratio (MLR) families, and Huber’s robust models (including $\epsilon$ neighborhood model, total variation model, and more generally with 2-alternating capacities [14]).
+Suppose we have estimators $r_{t}^{*}$ of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$, for each $t$ (which can be constructed using simulations from $F^{*}_{0}$ as described in Section 3.1).
+Now, we analogously define the confidence set as
 $\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}^{*}}\right\}.$ (3.9)
@@
 Suppose that the pre-change class $\mathcal{P}_{0}$ and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfy Assumption 3.5 and $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,\forall F_{0}\in\mathcal{P}_{0}$.
-Then, for any $\alpha\in(0,1),$ $T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, the confidence set $\mathcal{C}$ in (3.8) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
+Then, for any $\alpha\in(0,1),$ $T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, the confidence set $\mathcal{C}$ defined in (3.9), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
+We now analyze the size of the confidence set, generalizing Theorem 3.3 to the case where an LFD pair $(F_{0}^{*},F_{1}^{*})$ exists for $\mathcal{P}_{0}$ vs $\mathcal{P}_{1}$.
+We consider the confidence set in (3.9) with $r_{t}^{*}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}^{*}\geq t)}{N+1}$, $M_{t}$ as defined in (3.4) with $\hat{T}$ as defined in (3.3) using $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})},$ and $S_{n}^{(t)}=\prod_{i=n}^{t-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}$.
+We redefine $p_{t},\rho_{0}$ and $\rho_{1}$ in terms of the LFD pair: $p_{t}=\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t)$, $\rho_{0}(s)=\mathbb{E}_{F_{0}^{*}}\left(\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\right)^{s}$ and $s_{0}=\arg\min_{s\in(0,1)}\rho_{0}(s)$ and $\rho_{1}(s)=\mathbb{E}_{F_{1}}\left(\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\right)^{s}$ and $s_{1}=\arg\min_{s\in(0,1)}\rho_{1}(s)$.
+Finally, we call the stopping time $\tau$ well-behaved if it satisfies
+$\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ $\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t),\forall t\in\mathbb{N},\forall F_{0}\in\mathcal{P}_{0};$ (3.10a)
+$\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},\infty}[\tau\leq t],\forall t,T\in\mathbb{N}\penalty 10000\ \text{ and }$ (3.10b)
+$\mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t],\forall t,T\in\mathbb{N},\text{ and }\forall(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ (3.10c)
+Note that (3.10b) is analogous to (3.6), while (3.10c) requires that the stopping time $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically larger than $\tau$ under change from $F_{0}$ to $F_{1}$, for any $(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ To justify that (3.10c) is sensible, we note that the LFD pair $(F_{0}^{*},F_{1}^{*})$ is the closest pair within $\mathcal{P}_{0}\times\mathcal{P}_{1}$ (according to various metrics or divergences such as Kullback-Leibler), i.e., it renders the change detection problem most difficult.
+Therefore, it is reasonable to assume that $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically largest.
+(3.10a) can be justified similarly because $F_{0}^{*}$ is the “closest” element of the pre-change class to the post-change class.
+Theorem 3.9.
+In the setting described above, for any change detection method with stop time $\tau$ that is well-behaved in the sense of (3.10a) and $p_{T}\neq 0$,
+$\limsup_{N\to\infty}\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0}\\ F_{1}\in\mathcal{P}_{1}\end{subarray}}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\frac{1-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}},$ (3.11)
+where $\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}:=\displaystyle\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}\end{subarray}}\Delta_{F_{0},T,F_{1}}$.
+Additionally, if the stopping time $\tau$ is well-behaved in the sense of (3.10b) and (3.10c) as well, the last term in (3.11) can be improved to $\min\left\{\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right),\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}\right\}.$
 Next, we explore how, in parametric settings, one can obtain adaptive thresholds via simulation.
@@
 In this section, we introduce an alternative approach that, while restricted to the parametric setting, allows observations to be dependent, and often yields tighter confidence sets than the method proposed in the previous section, as shown by our simulations later.
-We begin with the most basic case, where the pre- and post-change distributions are known, and then extend it to more general parametric settings subsequently.
+We begin with the most basic case, where the pre- and post-change distributions are known, and then extend it to more general parametric settings.
 4.1 Known pre- and post-change distributions
@@
 For the $j$-th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at $\tau_{j}^{\prime}=\mathcal{A}(\{X_{n}^{j}\}_{n})$ or until a fixed time $L$, whichever happens first.
-Then, we compute the test statistic $M_{t,L}^{j}$ as described in Section 4.1.1, based on $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$.
+Then, we compute a test statistic $M_{t,L}^{j}$, based on $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$ (details are described in Supplementary Section B).
 Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$:
 $\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (4.1)
-where $r_{t}$ is as defined in (3.7). Algorithm 1 contains an overview of the method.
+where $r_{t}$ is any unbiased estimator of $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; see Algorithm 1.
 Algorithm 1 CI for known pre-change $F_{0}$ and post-change $F_{1}$
@@
 […]
-4:for $i=1$ to $N$ do
+4:for $i=1$ to NN do
 5:  Draw $\{Z_{n}^{i}\}_{n}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$, until $\mathcal{A}$ detects at $\tau_{i}$, or until $\tau$, whichever happens first
@@
 Theorem 4.1.
-For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (4.1) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
+For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (4.1), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
 Remark 4.2.
@@
 The only necessity is a stopping rule that can be applied to generated data sequences.
-4.1.1 Subtleties regarding Subroutine $\mathcal{L}$
-The problem with $L=\infty$ is that the stopping time may never be reached for some sampled sequence.
-One can set $L=\infty$ if $\mathcal{A}$ is known to satisfy $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely).
-This is a very mild assumption because any reasonable change detector should eventually raise an alarm if there is a change; indeed, for example, this holds for all algorithms with finite ARL.
-If $\mathcal{A}$ is not known to satisfy this condition, one can simply pick $L<\infty$.
-If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual.
-If $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it).
-For $L<\infty$,
-$M^{j}_{t,L}:=\begin{cases}-\infty,&\text{if }t>\tau_{j}^{\prime},\\ M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),&\text{if }t\leq\tau_{j}^{\prime}\leq L,\\ \infty,&\text{if }t\leq\tau_{j}^{\prime}\text{ and }\tau_{j}^{\prime}>L.\\ \end{cases}$ (4.2)
-We still need a definition of $M^{j}_{t,\infty}$ in the case that $L=\infty$ and $\tau_{j}=\infty$.
-We define $M^{j}_{t,\infty}:=M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),\text{if }t\leq\tau_{j}^{\prime}<\infty,$ and $M^{j}_{t,\infty}:=-\infty,\text{if }t>\tau_{j}^{\prime}\text{ or }\tau_{j}^{\prime}=\infty.$
-Remark 4.3.
-The value of $L$ does not need to be a fixed constant; it can be set adaptively depending on the observed data sequence, for instance, one can choose $L$ to be $2\tau$.
 4.2 Adaptive confidence set for known pre-change distribution and parametric post-change class
 We now assume that the post-change distribution follows some parametric composite model $\mathcal{P}_{1}=\{F_{\theta}:{\theta}\in\Theta_{1}\}$ and the pre-change distribution, $F_{\theta_{0}}(\theta_{0}\notin\Theta_{1})$ is known.
-[…]
+Since the post-change parameter is unknown, for each $t$, we estimate it based on $X_{t},\cdots,X_{\tau}$.
+However, the sample size used for the estimate might be too small to have a reliable estimate.
+To address this issue, we would use interval estimates instead of point estimates, allowing us to achieve a predetermined control over the error.
+Notably, conventional confidence intervals are invalid here, since $\tau$ is random.
+Therefore, we rely on confidence sequences [4, 12].
+A confidence sequence having coverage 1c1-c is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform guarantee:
 $\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ (4.2)
-We construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage, where $r_{t}$ is as defined in (3.5).
-Despite $r_{t}$ being random, it turns out that marginalizing over $r_{t}$ yields that the CS has coverage $\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; we formalize this more explicitly in the proof of the Theorem 4.4.
+We construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage, where $r_{t}$ is any unbiased estimator of $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$.
 Fix $B\in\mathbb{N}$.
@@
 sequences $\{X^{j}_{n}(\theta_{0},\theta)\}_{n}$ are allowed to be dependent across different values of $t$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$.
-A specific method for generating these sequences is detailed in Section B of the Supplementary.
+A specific method for generating these sequences is detailed in Section C of the Supplementary.
 We draw the data until a change is detected at $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until a fixed time $L$, whichever occurs first.
@@
 […]
-$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}$ (4.4)
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}.$ (4.3)
 Algorithm 2 summarizes the method.
 Note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ can be approximated by discretizing $\mathcal{S}^{\prime}_{\tau-t+1}$ into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum.
-While this approach is conceptually straightforward, it is computationally intensive; related experiments are in Section E of the Supplementary.
+While this approach is conceptually straightforward, it is computationally intensive; related experiments are in Section F of the Supplementary.
 Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient and does not require any approximation.
-Details are provided in Section B of the Supplementary, along with two concrete examples of the Gaussian and Poisson mean change problems.
+Details are provided in Section C of the Supplementary, along with two concrete examples of the Gaussian and Poisson mean change problems.
 Algorithm 2 CI for known pre-change ($F_{\theta_{0}}$) and parametric post-change
@@
 Theorem 4.3.
-Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (4.4) for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$.
+Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (4.3), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{\theta_{0},\infty}[\tau\geq t]$, for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$.
 Then for any given detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$
@@
 We now assume that both pre- and post-change distributions follow some parametric composite models: $\mathcal{P}_{0}=\{F_{\theta}:\theta\in\Theta_{0}\}$ and $\mathcal{P}_{1}=\{F_{\theta}:\theta\in\Theta_{1}\}$, $\Theta_{0}\cap\Theta_{1}=\emptyset$.
-[…]
+Since both pre- and post-change parameters are unknown, for each $t$, we estimate them based on $X_{1},\cdots,X_{t-1}$ and $X_{t},\cdots,X_{\tau}$ respectively.
+Instead of point estimates, we rely on interval estimates to achieve a predetermined control over the error.
+Suppose that we have some confidence interval (CI) procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$ that constructs CI using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{0}$ with coverage 1c1-c and a confidence sequence (CS) procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, that constructs CS using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{1}$ having coverage 1c1-c.
+We construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})$, where $r_{t}^{*}$ is computed under Assumption 3.5.
 For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ having joint distribution as $\mathbb{P}_{\theta,t,\theta^{\prime}}$.
@@
 sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ can be dependent across $t$ and $\theta\in\mathcal{S}_{t-1},\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$.
-A specific method for generating these sequences is detailed in Section C of the Supplementary.
+A specific method for generating these sequences is detailed in Section D of the Supplementary.
 We draw the data until a change is detected at the stopping time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever occurs first.
@@
 Algorithm 3 contains an overview.
-If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 3.3 and provides coverage guarantees, as elaborated in Section F of the Supplementary.
-As before, one can always adopt a grid-based approach, which is conceptually straightforward, but it is computationally intensive; corresponding experimental results are reported in Section E of the Supplementary.
+If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 3.5 and provides coverage guarantees, as elaborated in Section G of the Supplementary.
+As before, one can always adopt a grid-based approach, which is conceptually straightforward, but it is computationally intensive; corresponding experimental results are reported in Section F of the Supplementary.
 Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient and exact, but that can be a bit more conservative.
-We provide the details in Section C of the Supplementary, along with two concrete examples.
+We provide the details in Section D of the Supplementary, along with two concrete examples.
 Algorithm 3 CI for parametric pre-change and post-change
@@
 […]
-4:for $i=1$ to $N$ do
+4:for $i=1$ to NN do
 5:  Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F_{\theta_{0}^{*}}$, until $\mathcal{A}$ detects at $\tau_{i}^{*}$ or until $\tau$, whichever happens first.
@@
 Theorem 4.4.
-Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.5), for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$.
+Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.4), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{\theta_{0}^{*},\infty}[\tau\geq t]$, for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$.
 For any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1}$, and any change detection algorithm $\mathcal{A}$ satisfying Assumption 3.5 and having $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, the following guarantee holds: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
@@
 As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets, $i=0,1$, we have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$.
-So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 4.4 and 4.5 reduces to Theorem 4.1.
+So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 4.3 and 4.4 reduces to Theorem 4.1.
 5 Confidence set for pre- and post-change parameters
@@
 Analogously, for the pre-change parameter $\theta_{0}$, we fix each $t\in\mathcal{C}$ and construct a confidence interval for $\theta_{0}$ based on the data $X_{1},\ldots,X_{t-1}$}, denoted as $\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})$, for some $\eta_{0}\in(0,1)$, chosen to be smaller than the target (conditional) non-coverage level, $\alpha^{\prime\prime}$.
-The final confidence set for $\theta_{0}$ is the following:
-$\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$ (5.2)
-We formalize its coverage guarantee in the next theorem, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods [39, 36].
+The final confidence set for $\theta_{0}$ is $\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$ Next, we formalize its coverage guarantee, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods.
 Theorem 5.1.
@@
 $\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{0}\in\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})\mid\tau\geq T\right)\geq 1-\alpha-\eta_{0},\text{ and }$
-$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\in\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})\mid\tau\geq T\right)\geq 1-\alpha-\eta_{1}.$
+$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\in\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})\mid\tau\geq T\right)\geq 1-\alpha-\eta_{1}.$
 6 Experiments
-We validate our methods across various settings and demonstrate their practicality.
 Known pre- and post-change (Setting I): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$, with $T$ being the only unknown parameter.
@@
 Fig.
-6.1 visualizes the confidence sets (4.1) and point estimates (3.1) across 555 runs.
-Table 6.2 has the (conditional) average size and coverages of the confidence sets, the (conditional) average absolute deviation of $\hat{T}$, and delay of detection, i.e, $\tau-T$, across 500500500 runs.
-Figure 6.1: Setting I: The first $T-1$ observations are drawn from $N(0,1)$ and the rest from $N(1,1)$.
+6.1 visualizes the confidence sets (4.1) and point estimates (3.1) across $5$ runs.
+Table 6.2 has the (conditional) average size and coverages of the confidence sets, the (conditional) average absolute deviation of $\hat{T}$, and delay of detection, i.e, $\tau-T$, across $500$ runs.
+Figure 6.1: Setting I: The first T1T-1 observations are drawn from $N(0,1)$ and the rest from $N(1,1)$.
 The point estimates (3.1) are shown in a vertical red dashed line, and confidence sets (adaptive (4.1)) are shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$.
-Results of 555 independent simulations are shown.
+Results of $5$ independent simulations are shown.
 Table 6.2: Setting I: Known pre- and post-change.
@@
 where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$.
-After detection, we construct both the proposed confidence sets for $T$ (universal (3.6) with $\alpha=0.075$ and adaptive (4.4) [^6] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section B.1 (and Section C.1) of the Supplementary.
-with $\alpha=0.05,\beta=0.025$) with target conditional coverage 0.9250.9250.925.
+After detection, we construct both the proposed confidence sets for $T$ (universal (3.5) with $\alpha=0.075$ and adaptive (4.3) [^5] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section C.1 (and Section D.1) of the Supplementary.
+with $\alpha=0.05,\beta=0.025$) with target conditional coverage $0.925$.
 For implementing (3.5) , we use $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})},$ where $\theta_{1}^{*}$} is the closest element of $\Theta_{0}$ from $\Theta_{1}$, and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w$ being the same weight mentioned above.
-In Table 6.4, we report the (conditional) average size and coverages, absolute deviation of $\hat{T}$ , and detection delay $\tau-T$ across 500500500 runs.
+In Table 6.3, we report the (conditional) average size and coverages, absolute deviation of $\hat{T}$ , and detection delay $\tau-T$ across $500$ runs.
 Fig.
-6.3 visualizes the confidence sets (3.6) and point estimates (B.2) across 555 random runs.
-Figure 6.3: Setting II: First $T-1$ samples are drawn from $N(0,1)$ and the rest from $N(1,1)$.
-The point estimates (B.2) are shown in a vertical red dashed line, and confidence sets (universal (3.6)) are shown in red points.
-$F_{0}=N(0,1)$, $\mathcal{P}_{1}=\{N(\mu,1):\mu\geq 0.9\}$, $N=100$, $\alpha=0.075$.
-Results of 555 independent simulations are shown.
+6.4(a) shows the confidence sets (3.5) and point estimates (S2) across $5$ random runs.
 Table 6.3: Known pre-change but unknown post-change.
@@
 Universal Adaptive Universal Adaptive (Conditional) (Conditional)
-100 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 22.21 25.72 3.95 16.87
-100 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.97 17.85 24.11 3.67 16.21
-500 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 22.89 30.83 3.45 17.63
-500 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.98 18.14 28.56 3.48 15.81
+100 $0.75,\infty)$ 0.98 0.97 22.21 25.72 3.95 16.87
+100 $0.9,\infty)$ 0.97 0.97 17.85 24.11 3.67 16.21
+500 $0.75,\infty)$ 0.98 0.97 22.89 30.83 3.45 17.63
+500 $0.9,\infty)$ 0.97 0.98 18.14 28.56 3.48 15.81
 Composite parametric pre- and post-change (Setting III): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$.
@@
 where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$ and $\hat{\theta}_{0,:j-1}$ is the MLE of $\theta_{0}\in\Theta_{0}$ based on $X_{1},\cdots,X_{j-1}$.
-After detection, we construct both the proposed confidence sets (universal (3.8) with $\alpha=0.1$ and adaptive (4.5) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage 0.90.90.9.
+After detection, we construct both the proposed confidence sets (universal (3.9) with $\alpha=0.1$ and adaptive (4.4) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage $0.9$.
 For implementing (3.9), we use $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})}w^{\prime}(\theta)d\theta,$ with $w^{\prime}(\theta)$ being the discrete distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta-d,\theta_{3}=\theta-2d,\cdots,\theta_{10}=\theta-9d$, $d=0.2$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}w(\theta)d\theta$, with $w(\theta)$ being the same as mentioned in Setting II.
@@
 Fig.
-6.5 visualizes the confidence sets (3.8) and point estimates (C.2) across 5 random runs.
-Table 6.6 reports the results averaged across 500500500 runs.
-Figure 6.5: Setting III: First $T-1$ samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1)$.
-The point estimates (C.2) are shown in a vertical red dashed line.
+6.4(b) visualizes the confidence sets (3.9) and point estimates (S2) across 5 random runs.
+Table 6.5 reports the results averaged across $500$ runs.
+Figure 6.4: First T1T-1 samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1),T=100$.
+The point estimates (S2) are shown in a vertical red dashed line.
 Confidence sets (universal (3.9)) are shown in red points.
-$\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.1\}$, $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.9\}$, $N=100$, $\alpha=0.1$.
-Results of 555 random simulations are shown.
+$N=100$, $\alpha=0.1$.
+Results of $5$ random simulations are shown.
 Table 6.5: Confidence interval for $T$ (Setting III): unknown pre- and post-change.
@@
 Universal Adaptive Universal Adaptive (Cond.) (Cond.)
-100 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 26.91 26.84 4.36 25.81
-100 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.97 18.63 25.71 4.19 23.15
-500 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.97 26.12 31.45 4.03 24.48
-500 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.97 0.98 18.86 30.66 4.07 23.07
-Known pre- and nonparametric post-change (Setting IV): Here, the pre-change samples are from known $F_{0}=\text{U}[0,1]$ (uniform distribution on [0,1][0,1][ 0 , 1 ]) and let the post-change class be $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[X]\leq\mu\right\},$ where $\mathcal{M}$ is the set of all probability measures on $([0,1],\mathcal{B}([0,1]))$, with $\mu=0.25$.
+100 $(\infty,0.25]$ $0.75,\infty)$ 0.98 0.97 26.91 26.84 4.36 25.81
+100 $(\infty,0.1]$ $0.9,\infty)$ 0.97 0.97 18.63 25.71 4.19 23.15
+500 $(\infty,0.25]$ $0.75,\infty)$ 0.98 0.97 26.12 31.45 4.03 24.48
+500 $(\infty,0.1]$ $0.9,\infty)$ 0.97 0.98 18.86 30.66 4.07 23.07
+Known pre- and nonparametric post-change (Setting IV): Here, the pre-change samples are from known $F_{0}=\text{U}[0,1]$ (uniform distribution on $[0,1]$) and let the post-change class be $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[X]\leq\mu\right\},$ where $\mathcal{M}$ is the set of all probability measures on $([0,1],\mathcal{B}([0,1]))$, with $\mu=0.25$.
 For detection, we use an e-detector [28] $\tau_{e}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}\geq A\right\},$ where $\hat{p}_{i-1}^{\prime}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=t}^{i-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},$ for $x\in[0,1]$, with $A=1000$.
@@
 The forward $t$-delay e-process $\{R^{(t)}_{n}\}_{n\geq t}$, we use numeraire e-variable for $\mathcal{P}_{1}$ against $F_{0}$, as in Section 5.2 of [17] to get $R^{(t)}_{n}=\prod_{i=t}^{n}(1+\lambda^{*}(X_{i}-\mu)),$ where $\lambda^{*}\in(0,1/\mu)$ is the solution of the equation $e^{\lambda^{*}}=\frac{1+\lambda^{*}(1-\mu)}{1-\lambda^{*}\mu}.$ For the backward e-processes $\{S^{(t)}_{n}\}_{n<t}$, we adopt the predictable plug-in method [20].
-We divide the interval [0,1][0,1][ 0 , 1 ] into 101010 equal-sized bins $B_{i}=[\frac{i-1}{10},\frac{i}{10}),i=1,\cdots,10$ and perform histogram-based estimation.
+We divide the interval $[0,1]$ into $10$ equal-sized bins $B_{i}=[\frac{i-1}{10},\frac{i}{10}),i=1,\cdots,10$ and perform histogram-based estimation.
 We construct $S^{(t)}_{n}=\prod_{i=n}^{t-1}\frac{\hat{p}_{i-1}(X_{i})}{{p}_{0}(X_{i})},$ where ${p}_{0}(x)=\frac{1}{10},$ and $\hat{p}_{i-1}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=i-1}^{t-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},x\in[0,1]$.
@@
 $\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}.$ (6.5)
-Table 6.7 contains average results across 500500500 runs and Fig. 6.8 visualizes across 555 runs.
-[…]
+Table 6.6 shows average results across $500$ runs and Fig. 6.7(a) visualizes across $5$ runs.
+Table 6.6: Setting IV: known pre-change ($\text{U}[0,1]$) but nonparametric post-change.
 T
@@
 100 $4(1-x)^{3}$ 0.98 0.97 24.14 4.06 20.65
-[…]
+100 $0.8\text{U}[0,0.2]+0.2\text{U}[0.2,1]$ 0.98 0.97 20.92 2.86 14.08
 500 $4(1-x)^{3}$ 0.98 0.94 24.63 4.11 19.57
-[…]
-Figure 6.8: Setting IV: First $T-1$ samples are drawn from $U(0,1)$ and the remaining observations from the density $f(x)=4(1-x)^{3},x\in[0,1]$.
-The point estimate (6.5) is shown in a vertical red dashed line, and the confidence set (3.8) is shown in red points.
-Results of 555 random simulations are shown.
-[…]
+500 $0.8\text{U}[0,0.2]+0.2\text{U}[0.2,1]$ 0.99 0.93 21.27 2.94 15.78
+Composite parametric pre- and nonparametric post-change (Setting V): Here, we consider the true (unknown) pre- and post-change distributions to be $\text{N}(1,1)$ and $\text{U}[-1.2,0.8]$ respectively.
+Let the pre-change class be $\mathcal{P}_{0}=\left\{\text{N}(\theta,1):\theta\geq 0.5\right\},$ and the post-change class be the set of all SubGaussians with a nonpositive mean: $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[e^{\lambda X}]\leq e^{-\lambda^{2}/2},\forall\lambda\in[0,\infty)\right\},$ $\mathcal{M}$ is the set of all probability measures on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$.
+We use an e-detector [28],
 $\tau_{e}^{\prime}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}\geq A\right\},$
-with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}-0.5.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 3.3 with $F_{0}^{*}=\text{N}(0.5,1)$.
+with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}-0.5.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 3.5 with $P_{0}^{*}=\text{N}(0.5,1)$.
 After detection, we construct the confidence sets (3.9).
@@
 $\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}^{\prime}}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}.$ (6.6)
-Figure 6.9: Setting V.
-First $T-1$ samples are drawn from $N(1,1)$ and the remaining samples are from $U(-1.2,0.8)$.
-Vertical red dashed lines show the point estimates (6.6).
-The confidence set (3.8) is marked in red points.
-Results of 555 independent simulations are shown.
+Figure 6.7: Vertical red dashed lines show the point estimates (6.6). The confidence set (3.9) is marked in red points. Results of $5$ independent simulations are shown.
 Table 6.8: Setting V: Parametric pre-change, nonparametric post-change.
@@
 500 0.98 0.94 28.03 3.15 25.12
-Comparison with [36]: We now compare our method with Ch.
-3.5 of [36] on the Gaussian mean-change, having the (unknown) changepoint at $T=100$ or 500500500.
-We recall that, unlike our general nonasymptotic guarantees for any $\mathcal{A}$, Wu’s method only has asymptotic approximate guarantees in specific parametric settings, and only for CUSUM procedure [18], $\tau^{\prime}_{\text{cusum}}:=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}T_{j}\geq d\right\},$ where $T_{n}=\max\{0,T_{n-1}+X_{n}\}$ and $d$ is the threshold.
-We used $d=8.59$ and 7.567.567.56 for ${N}(-0.25,1)$ vs.
-${N}(0.25,1)$ and ${N}(-0.3,1)$ vs.
-${N}(0.3,1)$ respectively.
-We implement Wu’s confidence set of the form $V_{c}\cup[L_{s},\hat{\nu})$, where $\hat{\nu}=\max\{k<\tau:T_{k}=0\}$, $V_{c}=\{k\geq\hat{\nu}:T_{k}\leq c\}$, and $L_{k}$ denote the $k$-th last zero point of $T_{n}$ counted backward starting from $L_{0}:=\hat{\nu}$, with their proposed approximations (for the case when $\theta_{1}=-\theta_{0}$) $s=\log(\alpha)\times(1/(\sqrt{2}\theta_{1})+0.088)$ and $c=-\log(1-\sqrt{1-\alpha})/(2\theta_{1})-0.583$ with $\alpha=0.05$.
-Table 6.11 reports the average size and coverage rates (for target conditional coverage 0.950.950.95) across 500500500 runs.
+Huber-robust (nonparametric) pre- and post-change (Setting VI):
+We perform experiments for the robust Gaussian mean-change problem under the $\epsilon$ contamination model, with the true changepoint at $T=100,500$.
+The pre- and post-change classes are $\epsilon$ neighbourhoods around $N(\mu_{0},1)$ and $N(\mu_{1},1)$, respectively: $\mathcal{P}_{i}=\{(1-\epsilon)N(\mu_{i},1)+\epsilon F:F\in\mathcal{M}\};\penalty 10000\ \penalty 10000\ i=0,1.$ In our experiment, $\mu_{0}=0,\mu_{1}=1$ and $\epsilon=0.01,0.001$.
+The true (unknown) pre- and post-change distributions are $(1-\epsilon)\times N(0,1)+\epsilon\times\text{Cauchy}(-1,10)$ and $(1-\epsilon)\times N(1,1)+\epsilon\times\text{Cauchy}(-1,10)$ respectively.
+We use Huber’s LFD (which are also log-optimal e-values [24]) to construct a CUSUM-style change-detector as below:
+$\tau=\inf\left\{t\in\mathbb{N}:\max_{1\leq j\leq t}\prod_{i=j}^{t}\pi(X_{i})\geq A\right\},$ (6.7)
+where $\pi(x)=\min\{\max\{\frac{f_{1}(x)}{f_{0}(x)},c^{\prime}\},c^{\prime\prime}\}$ for constants $c^{\prime\prime}>c^{\prime}\geq 0$ as mentioned in [13, 24].
+We set the threshold at $A=1000$.
+$\hat{T}$ is as defined in (3.3) with the same e-process, $\prod_{i=1}^{t}\pi(X_{i})$, as mentioned above.
+After detection, we construct our universal confidence sets (3.9) for $T$, with $R_{n}^{(t)}=1/\prod_{i=t}^{n}\pi(X_{i})$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}\pi(X_{i})$, for $n<t$ and $\alpha=0.05$.
+We report the average size and coverage rates across $200$ independent runs in Table 6.9.
+We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection).
+Table 6.9: Setting VI: Huber-robust (nonparametric) pre- and post-change
+T $\epsilon$
+Conditional
+coverage
+Marginal
+coverage
+Size
+$|\hat{T}-T|$
+Delay
+100 0.01 0.98 0.98 16.74 3.23 18.38
+100 0.001 0.96 0.96 15.57 3.18 13.96
+500 0.01 0.97 0.93 16.18 3.87 16.58
+500 0.001 0.97 0.92 15.13 3.25 14.44
+[…]
 Table 6.10: Comparison of the confidence sets for $T$ with [39] (Setting I)
@@
 Ours Wu (2007) Ours Wu (2007) Ours Wu (2007)
-${N}(-0.25,1)$ ${N}(0.25,1)$ 100 46.39 47.83 0.96 0.86 0.88 0.79
-${N}(-0.3,1)$ ${N}(0.3,1)$ 100 33.74 34.14 0.96 0.83 0.82 0.76
-${N}(-0.25,1)$ ${N}(0.25,1)$ 500 46.58 47.21 0.96 0.86 0.57 0.53
-${N}(-0.3,1)$ ${N}(0.3,1)$ 500 35.19 34.81 0.97 0.82 0.56 0.49
+N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 100 46.39 47.83 0.96 0.86 0.88 0.79
+N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 100 33.74 34.14 0.96 0.83 0.82 0.76
+N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 500 46.58 47.21 0.96 0.86 0.57 0.53
+N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 500 35.19 34.81 0.97 0.82 0.56 0.49
 The results show that [39] (which is only asymptotically valid and relies on approximations) exhibits noticeable undercoverage, while the sizes for both methods are comparable.
 However, our method (4.1) achieves the target conditional coverage.
+Sentiment change in language data: We consider the Stanford Sentiment Treebank (SST-2) dataset of movie reviews labeled with a binary sentiment [31].
+We want to localize the change from $60\%$ positive reviews to $40\%$ positive reviews.
+In our experimental setup, the true changepoint is at 500; the pre-change data are i.i.d.
+draws of $75\%$ positive reviews, and the post-change data are i.i.d.
+draws of $25\%$ negative reviews (these are all unknown to our algorithm).
+We construct a change detection method and a point estimate $\hat{T}$, using a pre-trained DistilBERT-base model fine-tuned on the uncased SST-2 [25], which is a binary sentiment classifier that maps each review to either 0 (negative) or $1$ (positive), thereby producing a sequence of Bernoulli observations.
+After obtaining the binary sequence and $\hat{T}$, we construct the test statistic (3.4) using the likelihood ratio process for $\text{Bern}(0.4)$ against $\text{Bern}(0.6)$ (forward) and for $\text{Bern}(0.6)$ against $\text{Bern}(0.4)$ (backward).
+Note that given the change detector (and as a byproduct, the sequence of Bernoulli) and $\hat{T}$, it is simple to compute (3.4) in $O(T)$ time per $t$.
+So, the confidence set (3.9) takes total $O(T^{2})$ time.
+Fig.
+6.11 provides a visualization.
+Figure 6.11: Vertical red dashed lines show the point estimate.
+The confidence set (3.9) is marked in red points.
+The true changepoint $T=500$ is marked with the black line.
+Texts are plotted from time $t=470$ upto detection time, $\tau=522$.
 Remark 6.1.
 In Settings I, II, and III, both of our proposed methods — universal (Section 3) and adaptive (Section 4) — are applicable.
-For nonparametric settings (IV and V), the only valid method currently known is the universal method.
+For nonparametric settings (IV, V, and VI), the only valid method currently known is the universal method.
 Prior works are valid (asymptotically) only for Setting I.
 As shown in Table 6.2 for known pre- and post-change distributions (Setting I), the adaptive method consistently outperforms the universal method, which tends to be conservative.
-Also, our adaptive method is the only method we know that remains valid under dependence (see experimental results in Supplement G.1).
+Also, our adaptive method is the only method we know that remains valid under dependence (see experimental results in Supplement H.1).
 However, for the composite settings (II and III), the universal method outperforms the adaptive method.
-Moreover, the universal method is also significantly more efficient — it runs in $O(T)$ time (assuming $M_{t}$ can be computed in $O(1)$ time), while the adaptive method incurs a higher computational cost of $O(T^{2}B)$ (assuming $\mathcal{A}$ runs in $O(T)$ time and $M_{t}$ can be computed in $O(1)$ time).
-See Supplement Sections D, E, and G for further experiments, involving confidence intervals for pre- and post-change parameters, detection procedures with bounded PFA, Poisson rate change problems, settings with dependent data, etc.
+Moreover, the universal method is also significantly more efficient — it runs in $O(T^{2})$ time (assuming $M_{t}$ can be computed in $O(T)$ time), while the adaptive method incurs a higher computational cost of $O(T^{3}B)$ (assuming $\mathcal{A}$ runs in $O(T)$ time and $M_{t}$ can be computed in $O(T)$ time).
+See Supplement Sections E, F, and H for further experiments, involving confidence intervals for pre- and post-change parameters, detection procedures with bounded PFA, settings with dependent data, etc.
 7 Conclusion
@@
 Omitted proofs can be found in Appendix A.
-We provide the implementation details of Algorithms 2 and 3 in Appendices B and C respectively, along with two concrete examples in both sections.
-Experimental results for the confidence intervals for pre- and post-change parameters are in Appendix D.
-Experimental results for a different implementation method are in Appendix E.
-Simpler algorithms for PFA-controlling change detectors are presented in Appendix F.
-Some additional experiments with dependent (Markov) pre- and post-change data, with detection procedures with bounded PFA, and with Poisson rate change problems are in Appendix G.
-Some heuristic algorithms are presented in Appendices H and I, pointing at open avenues for further theoretical work.
+We provide the implementation details of Algorithm 1 in Appendix B and of Algorithms 2 and 3 in Appendices C and D respectively, along with two concrete examples in both sections.
+Experimental results for the confidence intervals for pre- and post-change parameters are in Appendix E.
+Experimental results for a different implementation method are in Appendix F.
+Simpler algorithms for PFA-controlling change detectors are presented in Appendix G.
+Some additional experiments with dependent (Markov) pre- and post-change data, with detection procedures with bounded PFA, and with Poisson rate change problems are in Appendix H.
+Some heuristic algorithms are presented in Appendices I and J, pointing at open avenues for further theoretical work.
 Appendix A Omitted proofs
@@
 Proof of Proposition 2.2.
-First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
+First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
 Let $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying $\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\geq 1-\alpha$, for all $t\in\mathbb{N}$.
-Since $\tau$ be a stopping time, $[\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$.
+Since $\tau$ be a stopping time, $\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$.
 This implies $\mathbb{P}_{F_{0},\infty}(\tau\geq t)=\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ Therefore,
@@
 $\leq 1-(1-\alpha)\mathbb{P}_{F_{0},\infty}(\tau\geq t)$
-$\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t),\leavevmode\nobreak\ \forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
+$\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t),\penalty 10000\ \forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
 For the converse part, $\phi^{(t)}$ satisfies $\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$
@@
 Therefore, $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq T)-\mathbb{P}_{F_{0},\infty}(\tau<T)$, which implies
-$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C},\tau\geq T)}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C},\tau\geq T)}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
 Lemma A.1 (Ville’s inequality [35]).
 If $\{R_{n}\}_{n\in\mathbb{N}}$ is an e-process for $\mathcal{P}$ with respect to a filtration $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, then
-$\mathbb{P}\left(\exists n\in\mathbb{N}:R_{n}\geq 1/\alpha\right)\leq\alpha,\forall\mathbb{P}\in\mathcal{P}.$ (A.1)
+$\mathbb{P}\left(\exists n\in\mathbb{N}:R_{n}\geq 1/\alpha\right)\leq\alpha,\forall\mathbb{P}\in\mathcal{P}.$ (S1)
 For a short proof, see [11][Lemma 1].
@@
 $\leq\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{n}\geq 2/\alpha\right]\right)$
-$=P_{F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
+$=\mathbb{P}_{F_{0},t,F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
 Similarly, for the e-processes $\{S^{(t)}_{n}\}_{n<t}$, we get $\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{(t-1)\wedge\hat{T}}\geq 2/\alpha\right]\leq\alpha/2.$ Therefore, from the above steps, we finally have
@@
 Proof of Theorem 3.2.
-From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other. Hence,
-$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}]=\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t],\forall t\in\mathbb{N}.$ Now,
+Observe that
 $\mathbb{P}_{F_{0},T,F_{1}}(M_{T}<2/(\alpha r_{T}))$
@@
 $=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T),$
-where we twice used the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$, which holds because $\tau$ is a stopping time and so the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
-So, we obtain
-$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\mathbb{P}_{F_{0},T,F_{1}}(M_{T}>2/(\alpha r_{T}))}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$ (A.2)
+where we twice used the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$, which holds because $\tau$ is a stopping time and so the event $\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
+From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other.
+Now, since $r_{t}$ is an unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}]=\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\forall t\in\mathbb{N}.$ So, we obtain
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1$- $\frac{\mathbb{P}_{F_{0},T,F_{1}}(M_{T}>2/(\alpha r_{T}))}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
+Proof of Theorem 3.3.
+We first decompose the length of the confidence set into contributions from the pre- and post-change regions, and then control each contribution separately.
+Decomposition of the confidence set.
+Since $M_{t}=\frac{L_{\hat{T}}^{\tau}}{L_{t}^{\tau}}$ and $\hat{T}=\arg\max_{1\leq i\leq\tau}L_{i}^{\tau}$ and hence, whenever $\tau\geq T>t,$
+$M_{t}\geq\frac{L_{{T}}^{\tau}}{L_{t}^{\tau}}=\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}.$
+Similarly, whenever $T<t\leq\tau,$ $M_{t}=\frac{L_{\hat{T}}^{\tau}}{L_{t}^{\tau}}\geq\frac{L_{{T}}^{\tau}}{L_{t}^{\tau}}=\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$. Therefore,
+$|{\mathcal{C}}|\times\mathds{1}(\tau\geq T)$
+$\leq\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+\sum_{t=T+1}^{{\tau}}\mathds{1}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t})\right)$
+$=\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+\sum_{t=T+1}^{{\infty}}\mathds{1}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),t\leq\tau\right).$
+We also have
+$|{\mathcal{C}}|\leq\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+|\tau-T|.$ (S2)
+So, taking the conditional expectation on both sides and combining both bounds, we get
+$\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq 1+\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$
+$+\min\left\{\sum_{t=T+1}^{{\infty}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),t\leq\tau\mid\tau\geq T\right),\Delta_{F_{0},T,F_{1}}\right\}.$ (S3)
+Controlling the Pre-Change Tail.
+Fix some $s\in(0,1)$. Then, using Markov’s equality,
+$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$ $\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{T})\right)\quad\text{[Since }r_{t}\leq r_{T}]$
+$=\mathbb{E}\left(\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}\geq(\alpha r_{T})^{s}/2^{s}\mid r_{T}\right)\right)$
+$\leq\mathbb{E}\left(\left(\frac{2}{\alpha r_{T}}\right)^{s}\mathbb{E}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}\right)\right)$
+$=\mathbb{E}\left(\left(\frac{2}{\alpha r_{T}}\right)^{s}\left(\mathbb{E}_{F_{0}}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}\right)^{T-t}\right)$
+$=\mathbb{E}\left(\frac{2}{\alpha r_{T}}\right)^{s}\rho_{0}(s)^{T-t}.$
+This holds for all $s\in(0,1)$ and hence for $s=s_{0}.$ As noted earlier, $\rho_{0}(s_{0})<1$. So, we have
+$\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$ $\leq\frac{1}{p_{T}}\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$
+$\leq\mathbb{E}\left(\frac{2}{\alpha r_{T}}\right)^{s}\frac{1}{p_{T}}\sum_{t=1}^{{T-1}}\rho_{0}(s)^{T-t}$
+$\to\frac{2^{s}}{\alpha^{s}p_{{T}}^{s+1}}\times\frac{\rho_{0}(s)-\rho_{0}(s)^{T-1}}{1-\rho_{0}(s)},$
+as $N\to\infty.$ To justify the convergence, we note that $\left(\frac{2}{\alpha r_{T}}\right)^{s}$ is uniformly integrable as a sequence in $N\in\mathbb{N}$ (follows from Lemma A.2), and by the strong law of large numbers, $\left(\frac{2}{\alpha r_{T}}\right)^{s}\to\left(\frac{2}{\alpha p_{T}}\right)^{s}$ almost surely.
+Since the above holds for any $s\in(0,1)$, it also holds for $s=s_{0}.$
+Deriving the first bound.
+So, taking the conditional expectation on both sides of (S2), we get
+$\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq 1+\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)+\Delta_{F_{0},T,F_{1}}.$
+Therefore, it follows from the above calculation that
+$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\times\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{F_{0},T,F_{1}}.$ (S4)
+Controlling the “Post-Change” Tail when $\tau$ is sensitive.
+Similarly, for $t\geq T+1$, we have
+$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),t\leq\tau\right)$
+$\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{\tau})\right)$
+$\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{\tau}),r_{\tau}>\rho_{1}(s)^{t-T}\right)+\mathbb{P}_{F_{0},T,F_{1}}[r_{\tau}\leq\rho_{1}(s)^{t-T}]$
+$\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\left(\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right)^{s}\geq(\alpha\rho_{1}(s)^{t-T})^{s}/2^{s}\right)+\mathbb{P}_{F_{0},T,F_{1}}[r_{\tau}\leq\rho_{1}(s)^{t-T}]$
+$\leq\left(\frac{2}{\alpha\rho_{1}(s)^{t-T}}\right)^{s}\mathbb{E}_{F_{1}}\left(\prod_{i=T}^{{t}-1}\left(\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right)^{s}\right)+\mathbb{P}_{F_{0},T,F_{1}}[r_{\tau}\leq\rho_{1}(s)^{t-T}]$
+$\leq\left(\frac{2}{\alpha\rho_{1}(s)^{t-T}}\right)^{s}\rho_{1}(s)^{t-T}+\rho_{1}(s)^{t-T}.$
+The last step follows from the assumption that $\tau$ is sensible (recall that if $\tau$ is sensitive, then $\tau_{j}$ is stochastically larger than $\tau$) and from Lemma A.3.
+Also, as we noted before, $\rho_{1}(s_{1})<1.$ Therefore,
+$\sum_{t=T+1}^{{\infty}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),\tau\geq t\mid\tau\geq T\right)$
+$\leq\frac{1}{p_{T}}\sum_{t=T+1}^{{\infty}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),\tau\geq t\right)$
+$\leq\frac{1}{p_{T}}\sum_{t=T+1}^{{\infty}}\left(\left(\frac{2}{\alpha}\right)^{s}\rho_{1}(s)^{s(t-T)}+\rho_{1}(s)^{t-T}\right)$
+$\leq\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s}\times\frac{\rho_{1}(s)^{s}}{1-\rho_{1}(s)^{s}}+\frac{\rho_{1}(s)}{1-\rho_{1}(s)}\right).$
+Since the above holds for any $s\in(0,1)$, it also holds for $s=s_{1}.$
+Deriving the final bound.
+Finally, it follows from (A) and the above calculations that
+$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\times\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1$
+$+\min\left\{\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right),\Delta_{F_{0},T,F_{1}}\right\}.$ (S5)
 Proof of Proposition 3.7.
-We only prove the case when $\theta_{0}^{*}=\sup\Theta_{0}\leq\inf\Theta_{1}$ and $\tau$ is the CUSUM type stopping time $\tau_{\text{wcs-ripr}}$ defined in (C.12).
-Proofs for the other cases, when $\theta_{0}^{*}=\inf\Theta_{0}\geq\sup\Theta_{1}$ and/or $\tau$ is the SR-type stopping time (C.13), follow analogously.
-$\mathbb{P}_{\theta,\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}_{\theta,\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta_{1}}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}dW(\theta_{1})<A\right].$
-Since the family $\{F_{\theta}\}$ satisfies the MLR property in the statistic $T(x)$, and since $\frac{f_{\theta_{1}}(x)}{f_{\theta_{0}^{*}}(x)}$ is non-decreasing in $T(x)$, it follows that the random variables $\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}$ are stochastically increasing in $\theta$.
-That is, for $X\sim F_{\theta}$ and $X^{\prime}\sim F_{\theta^{\prime}}$ with $\theta\leq\theta^{\prime}\leq\sup\Theta_{0}$, we have $T(X)\preceq_{\text{st}}T(X^{\prime})$ (see e.g., [16, Chapter 3.4]) and hence,
-$\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}\preceq_{\text{st}}\frac{f_{\theta_{1}}(X^{\prime})}{f_{\theta_{0}^{*}}(X^{\prime})}.$
-By Theorem 1.A.2 of [24], the stochastic ordering implies that there exist a random variable $Z_{\theta_{1}}$ and functions $\psi_{\theta_{1},1}$ and $\psi_{\theta_{1},2}$ such that $\psi_{\theta_{1},1}(z)\leq\psi_{\theta_{1},2}(z)$ for all $z$ and $\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{\theta_{1},1}(Z_{\theta_{1}})$ and $\frac{f_{\theta_{1}}(X^{\prime})}{f_{\theta_{0}^{*}}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{\theta_{1},2}(Z_{\theta_{1}})$.
+We only prove the case when $\tau$ is the CUSUM type stopping time $\tau_{\text{wcs-ripr}}$ defined as
+$\tau_{\text{cs}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\prod_{i=j}^{n}\frac{p_{1}^{*}(X_{i})}{p_{0}^{*}(X_{i})}\geq A\right\}.$
+Proofs for the other case, when $\tau$ is the SR-type stopping time, follow analogously.
+$\mathbb{P}_{P,\infty}[\tau_{\text{cs}}\geq t]$ $=\mathbb{P}_{P,\infty}\left[\displaystyle\max_{1\leq k\leq t-1}\max_{1\leq j\leq k}\prod_{i=j}^{k}\frac{p_{1}^{*}(X_{i})}{p_{0}^{*}(X_{i})}<A\right].$
+Since $(P_{0}^{*},P_{1}^{*})$ is an LFD pair for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, for any $P_{0}\in\mathcal{P}_{0},$ we have
+${R(P_{0},\phi)\leq R(P_{0}^{*},\phi)}$
+[…]
+${\implies{P_{0}}\!\left(\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}>\eta\right)\leq{P_{0}^{*}}\!\left(\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}>\eta\right),\text{ for all }\eta}$
+which is same as
+$\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\preceq_{\text{st}}\frac{p_{1}^{*}(X^{\prime})}{p_{0}^{*}(X^{\prime})},$
+where $X\sim P$ and $X^{\prime}\sim P^{*}$.
+By Theorem 1.A.2 of [26], the stochastic ordering implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\frac{p_{1}^{*}(X^{\prime})}{p_{0}^{*}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$.
 Therefore,
-$\mathbb{P}_{\theta,\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\psi_{\theta_{1},1}(Z_{\theta_{1},i})dW(\theta_{1})<A\right]$
-$\geq\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\psi_{\theta_{1},2}(Z_{\theta_{1},i})dW(\theta_{1})<A\right]$
-$=\mathbb{P}_{\theta^{\prime},\infty}[\tau_{\text{wcs-ripr}}\geq t].$
+$\mathbb{P}_{P,\infty}[\tau_{\text{cs}}\geq t]$ $=\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t-1}\max_{1\leq j\leq k}\prod_{i=j}^{k}\psi_{1}(Z_{i})dW(\theta_{1})<A\right]$
+$\geq\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t-1}\max_{1\leq j\leq k}\prod_{i=j}^{k}\psi_{2}(Z_{i})dW(\theta_{1})<A\right]$
+$=\mathbb{P}_{P^{*},\infty}[\tau_{\text{cs}}\geq t].$
 Proof of Theorem 3.8.
-For all $t\in\mathbb{N},\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}^{*}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}^{*}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}^{*}]=\alpha\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]\geq\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t].$ The last step follows from Assumption 3.3.
-The rest of the proof is the same as in the proof of Theorem 3.2.
-∎
+Since $r_{t}$ is an unbiased estimator of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$ and is independent of data, we have $\forall t\in\mathbb{N}$,
+$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/(\alpha r_{t}^{*})]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/(\alpha r_{t}^{*})\mid r_{t}^{*}]]\leq\mathbb{E}[\alpha r_{t}^{*}]=\alpha\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]\leq\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t],$ $\forall F_{0}\in\mathcal{P}_{0}$.
+The last steps follow from Assumption 3.5. The rest of the proof is the same as in the proof of Theorem 3.2. ∎
 Proof of Theorem 3.9.
-First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
+It follows from the definition of the least favorable distribution that
+$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right).$
+Therefore, for any $t\leq T-1$, using the above inequality and the assumption (3.10a) we get
+$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$
+$\leq\frac{1}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right)$
+$\leq\frac{1}{\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)}\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right).$
+Similarly, for any $t\geq T+1$,
+$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha r_{t}),\tau\geq t\mid\tau\geq T\right)$
+$\leq\frac{1}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha r_{\tau})\right)$
+$\leq\frac{1}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\left[\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)+\mathbb{P}_{F_{0},T,F_{1}}(r_{\tau}\leq\eta)\right]$
+$\leq\frac{1}{\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)}\left[\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)+\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}(r_{\tau}\leq\eta)\right].$
+In the last step, $\mathbb{P}_{F_{0},\infty}(\tau\geq T)\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)$ follows from (3.10a), $\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)$ follows from the fact that $(F_{0}^{*},F_{1}^{*})$ is an LFD pair, and $\mathbb{P}_{F_{0},T,F_{1}}(r_{\tau}\leq\eta)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}(r_{\tau}\leq\eta)$ follows from (3.10c).
+The remaining calculations are the same as those done in the proof of the previous theorem, with $f_{i}$ replaced by $f_{i}^{*}.$
+Proof of Theorem 4.1.
+First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
 Defining $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, and $\hat{\alpha}_{t}=\alpha r_{t}$. Now,
@@
 where we twice invoked $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$ as explained earlier. So, we obtain
-$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{T,L})}{P_{\infty}(\tau\geq T)}.$ (A.3)
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{T,L})}{P_{\infty}(\tau\geq T)}.$ (S6)
 First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d.
@@
 Thus, for all $t$,
-$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ (A.4)
-Therefore, combining (A.3) and (A.4), we obtain $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
+$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ (S7)
+Therefore, combining (S6) and (S7), we obtain $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
 Proof of Theorem 4.3.
@@
 Hence, for any fixed $c\in(0,1)$,
-$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (A.5)
+$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (S8)
 Define, $C_{t,L}^{\alpha}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\alpha;M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$.
@@
 $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
-the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
+the last equality holds because the event $\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
 $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha r_{t}})$
@@
 Hence, for any fixed $c\in(0,1)$,
-$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (A.6)
+$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (S9)
 Define, $C_{t,L}^{\alpha}=\operatorname{Quantile}(1-\alpha;M_{t},\{\max_{\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$.
@@
 And similarly,
-$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{o}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})\mid r_{t}^{*}))$
+$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})\mid r_{t}^{*}))$
 $\geq\mathbb{E}(1-\gamma r_{t}^{*})=1-\gamma\mathbb{P}_{\theta_{0}^{*},\infty}(\tau\geq t).$
@@
 $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\{t\geq 1:M_{t}\leq C_{t,L}^{\alpha r_{t}^{*}}\},\tau<T)\leq\mathbb{P}_{\theta_{0},T,\theta_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau<T),$
-the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
+the last equality holds because the event $\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
 $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L}^{\alpha r_{t}^{*}})$
@@
 $\leq\alpha+\eta_{0}.$
-Appendix B Computationally and statistically efficient subroutines in Algorithm 2
+A.1 Auxilary Lemmas
+Lemma A.2.
+Define, $r_{T}=\frac{1+\sum_{i=1}^{N}\mathds{1}(\tau\geq T)}{N+1}$. Then, for fixed $s\in(0,1)$, $\{r_{T}^{-s}\}_{N\geq 1}$ is uniformly integrable.
+Proof.
+[…]
+$\mathbb{E}[r_{T}^{-1}]=(N+1)\mathbb{E}\left[\frac{1}{K_{N}+1}\right]=\frac{1-(1-p_{T})^{N+1}}{p_{T}}\leq\frac{1}{p_{T}}.$
+This bound $\frac{1}{p_{T}}$ is a constant that does not depend on $N$. Hence, $r_{t}^{-s}$ is uniformly integrable. ∎
+Lemma A.3.
+Suppose, $\tau_{1},\cdots,\tau_{N}$ are i.i.d.
+random variables.
+Define, $r_{t}=\frac{1+\sum_{i=1}^{N}\mathds{1}(\tau_{i}\geq t)}{N+1}$, for $t\in\mathbb{N}.$ If each $\tau_{j}$ is stochastically larger than a random variable $\tau$, then $r_{\tau}$ is a p-value.
+Proof.
+Suppose $\tau_{N+1}$ is independent and identically distributed (i.i.d.) with $\tau_{1},\cdots,\tau_{N}$.
+By Theorem 1.A.2 of [26], $\tau_{N+1}$ is stochastically larger than $\tau$ implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\tau\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\tau_{N+1}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$.
+By direct calculation,
+$\mathbb{P}[r_{\tau}\leq(1+k)/(N+1)]=\mathbb{P}[\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq\tau)\leq k]$ $=\mathbb{P}[\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq\psi_{1}(Z))\leq k]$
+$\leq\mathbb{P}[\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq\psi_{2}(Z))\leq k]$
+$=\mathbb{P}[\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq\tau_{N+1})\leq k]\leq\frac{1+k}{N+1},$
+for all $k=0,1,\cdots,N.$ Therefore, $r_{\tau}$ is a p-value. ∎
+Appendix B Subtleties regarding Subroutine $\mathcal{L}$ in Algorithm 1
+The problem with $L=\infty$ is that the stopping time may never be reached for some sampled sequence.
+One can set $L=\infty$ if $\mathcal{A}$ is known to satisfy $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely).
+This is a very mild assumption because any reasonable change detector should eventually raise an alarm if there is a change; indeed, for example, this holds for all algorithms with finite ARL.
+If $\mathcal{A}$ is not known to satisfy this condition, one can simply pick $L<\infty$ (The value of $L$ does not need to be a fixed constant; it can be set adaptively depending on the observed data sequence, for instance, one can choose $L$ to be $2\tau$.) If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual.
+If $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it).
+For $L<\infty$,
+[…]
+We still need a definition of $M^{j}_{t,\infty}$ in the case that $L=\infty$ and $\tau_{j}=\infty$.
+We define $M^{j}_{t,\infty}:=M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),\text{if }t\leq\tau_{j}^{\prime}<\infty,$ and $M^{j}_{t,\infty}:=-\infty,\text{if }t>\tau_{j}^{\prime}\text{ or }\tau_{j}^{\prime}=\infty.$
+Appendix C Computationally and statistically efficient subroutines in Algorithm 2
 Let $L_{t}(\theta_{1};X_{1},\cdots,X_{n})$ be the likelihood of data $X_{1},\cdots,X_{n}$ under $\mathcal{H}_{0,t}$. Then, for $n\geq t$,
-$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (B.1)
+$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (S1)
 We define $\hat{\theta}_{1,j:n}$ to be a maximum likelihood estimator (MLE) of the post-change parameter $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$.
 The point estimate of $T$ can be naturally extended to:
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (B.2)
+$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (S2)
 The prototypical test statistic that we employ in our simulations is the likelihood-ratio-based test statistic $M_{t}^{(n)}$, where for $t,n\in\mathbb{N},t\leq n$,
-$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (B.3)
+$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (S3)
 Note that the test statistic $M_{t}$ for testing $\mathcal{H}_{0,t}$ reduces to
-$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (B.4)
+[…]
 For computational and statistical efficiency, we generate sequences from all $t\in\{1,\cdots,\tau\}$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ simultaneously using a single shared source of randomness.
@@
 Therefore, for each $j=1,\cdots,B$, we generate only one data sequence $\{X_{n}^{j}\}_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{\theta^{\prime\prime}}$, and then for each $t\in\{1,\cdots,\tau\}$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ (having joint distribution $\mathbb{P}_{\theta_{0},t,\theta}$) by applying the function $G(.,\theta_{0})$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta)$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
-$Y_{n}^{t}(x,\theta):=\begin{cases}G(x,\theta_{0}),&\text{if }n<t\\ G(x,\theta),&\text{ otherwise. }\end{cases}$ (B.5)
+$Y_{n}^{t}(x,\theta):=\begin{cases}G(x,\theta_{0}),&\text{if }n<t\\ G(x,\theta),&\text{ otherwise. }\end{cases}$ (S5)
 Let us provide a few examples of the function $G$.
-If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}}(x))$ works.
+If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists, and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}_{t}}(x))$ works.
 In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$.
@@
 Again, we bound $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)$ by $V_{t,L}^{j}$, where for $L<\infty$,
-$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (B.6)
+$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (S6)
 And for $L=\infty$,
-$V_{t,\infty}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{otherwise,}\\ \end{cases}$ (B.7)
-which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section B.1). We define the confidence set as
-$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (B.8)
+[…]
+which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section C.1). We define the confidence set as
+$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (S8)
 We observe that $\mathcal{C}_{1}^{\prime}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}$ is the set produced by Algorithm 2, since
@@
 Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.3.
-Corollary B.1.
-$\mathcal{C}^{\prime}_{1}$ defined in (B.8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.4.
-Next, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with the Poisson distribution in Section G.3.
-B.1 A concrete example: Gaussian mean shift problem
+Corollary C.1.
+$\mathcal{C}^{\prime}_{1}$ defined in (S8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.3.
+Next, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with the Poisson distribution in Section H.4.
+C.1 A concrete example: Gaussian mean shift problem
 Let the pre-change distribution $F_{0}$ be $N(\theta_{0},1)$ and the post-change model be $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
 In this setting, we have the following example of a $1-\beta$ confidence sequence [12] for $\theta_{1}$
-$\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (B.9)
+$\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (S9)
 where $\bar{X}_{t:\tau}=\frac{1}{\tau-t+1}\sum_{i=t}^{\tau}X_{i}$ and $s_{n}(\beta)=\sqrt{\log\log(2n)+0.72\log(10.4/\beta)}$.
@@
 For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
-$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t,\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (B.10)
+$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t,\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (S10)
 where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$.
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (B.4), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
+We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (S4), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
 Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$.
-Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.2) and they are finite.
+Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark C.2) and they are finite.
 So, we can have $L=\infty$.
@@
 And hence, $V_{t,\infty}^{j}$ can be computed easily.
-Remark B.2.
-[…]
-$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (B.11)
-[…]
+Remark C.2.
+For weighted CUSUM or SR detectors, finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. For some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$, these are defined as
+$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (S11)
+ $\text{ and }\penalty 10000\ \tau_{\text{wsr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}.$ (S12)
 For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$.
@@
 Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
-One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section G.2.
+One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section H.3.
 The Gaussian mean shift example we provided serves as a concrete and illustrative case to clarify the methodology.
-It is worth noting that $C^{\prime}$, as defined in (B.8), is practically implementable as long as the following relatively weak conditions are met:
+It is worth noting that $C^{\prime}$, as defined in (S8), is practically implementable as long as the following relatively weak conditions are met:
   1. 1.
@@
   4. 4.
-    If we employ the likelihood-ratio-based test statistic (B.4), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
+    If we employ the likelihood-ratio-based test statistic (S4), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
 For instance, one can easily verify that this approach applies to Gaussian, Laplace, or exponential scale change problems.
-B.2 Another concrete example: Poisson rate change problem
+C.2 Another concrete example: Poisson rate change problem
 Here is $F_{0}$ is $\text{Pois}(\theta_{0})$ and the post-change model is $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\in\Theta_{1}\}$.
 In this setting, we have the following example of a $1-\beta$ confidence sequence [2] for $\theta_{1}$
-$\mathcal{S}^{\prime}_{n}=\{\theta:n\theta-\sum_{i=t}^{t+n-1}X_{i}\log\theta\leq\log\left(\frac{1}{\beta}\right)+\log I(c,c\theta)-\log I(n+c,\sum_{i=t}^{t+n-1}X_{i}+c\theta)\}\cap\Theta_{1},$ (B.13)
-[…]
+$\mathcal{S}^{\prime}_{n}=\{\theta:n\theta-\sum_{i=t}^{t+n-1}X_{i}\log\theta\leq\log\left(\frac{1}{\beta}\right)+\log I(c,c\theta)-\log I(n+c,\sum_{i=t}^{t+n-1}X_{i}+c\theta)\}\cap\Theta_{1},$ (S13)
+where $c$ is any constant and $I(a,b)=\int_{-\infty}^{\infty}e^{-ae^{x}+bx}dx$.
 We fix $t\in\{1,\cdots,\tau\}$.
@@
 $Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta_{0}/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ otherwise.
-}\end{cases}$ (B.14)
+}\end{cases}$ (S14)
 Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta_{0})$, for $n<t$ and $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta)$, for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1},n\geq t$.
@@
 Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$.
-Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.3) and they are finite.
+Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark C.3) and they are finite.
 So, we can choose $L=\infty$.
@@
 All the maximum are taken over a finite set; hence, $V_{t}^{j}$ can be computed easily.
-Remark B.3.
+Remark C.3.
 For weighted CUSUM or SR detectors (defined in Remark 3.3.), finding $t_{1}^{j},t_{2}^{j}$ is often straightforward.
@@
 Also, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\neq\theta_{0}\}$, if we employ Gamma weights (it is a conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
-Appendix C Computationally and statistically efficient subroutines in Algorithm 3
+Appendix D Computationally and statistically efficient subroutines in Algorithm 3
 Let $L_{t}^{n}(\theta_{0},\theta_{1})$ be the likelihood of data up to time $n\geq t$ under $H_{0,t}$. Then,
-$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (C.1)
+$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (S1)
 We define $\hat{\theta}_{0,1:j}$ to be the maximum likelihood estimator (MLE) of the null $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$.
@@
 The point estimate of $T$ can be naturally extended to the following form:
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (C.2)
+$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (S2)
 As a prototypical example, we consider the likelihood-ratio-based test statistic $M_{t}^{(n)}$, which is defined for $t,n\in\mathbb{N},t\leq n$ as below.
-$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (C.3)
+$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (S3)
 Note that the test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ reduces to
-$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{0,1:\hat{T}-1},\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau\text{ or }\tau=\infty.\\ \end{cases}$ (C.4)
-As in Appendix B, we assume that we have a fixed element $\theta^{\prime\prime}$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta_{0}\cup\Theta_{1}$.
+[…]
+As in Appendix C, we assume that we have a fixed element $\theta^{\prime\prime}$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta_{0}\cup\Theta_{1}$.
 Therefore, for each $j=1,\cdots,B$, we generate data sequence $X_{1}^{j},X_{2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta^{\prime\prime}}$ only once.
 Then, for each $t\in\{1,\cdots,\tau\}$ and $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta,\theta^{\prime})\}_{n}$ having joint distribution $\mathbb{P}_{\theta,t,\theta^{\prime}},$ by applying the function $G(.,\theta)$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta^{\prime})$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
-$Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (C.5)
+$Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (S5)
 Note that, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta,\theta^{\prime})\}_{n}$ across all $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
@@
 Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where for $L<\infty$,
-$U_{t,L}^{j}:=\begin{cases}-\infty,\leavevmode\nobreak\ \leavevmode\nobreak\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (C.6)
+$U_{t,L}^{j}:=\begin{cases}-\infty,\penalty 10000\ \penalty 10000\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\penalty 10000\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (S6)
 And for $L=\infty$,
-$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (C.7)
+$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \penalty 10000\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (S7)
 We consider the following confidence set:
-$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (C.8)
+$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (S8)
 Observe that $\mathcal{C}^{\prime\prime}_{1}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}_{1}$ is the set produced by Algorithm 3, since
@@
 Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.4.
-Corollary C.1.
-$\mathcal{C}^{\prime\prime}_{1}$ defined in (C.8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.5.
+Corollary D.1.
+$\mathcal{C}^{\prime\prime}_{1}$ defined in (S8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.4.
 In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method.
 We illustrate another example with Poisson distribution in Appendix C of the Supplementary Material.
-C.1 A concrete example: Gaussian mean shift problem
+D.1 A concrete example: Gaussian mean shift problem
 We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$.
 Then, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$:
-$\mathcal{S}_{t-1}=\left[\bar{X}_{1:t-1}-\frac{q_{\gamma/2}}{\sqrt{t-1}},\bar{X}_{1:t-1}+\frac{q_{\gamma/2}}{\sqrt{{t-1}}}\right]\cap\Theta_{0},$ (C.9)
-and the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (B.9).
+$\mathcal{S}_{t-1}=\left[\bar{X}_{1:t-1}-\frac{q_{\gamma/2}}{\sqrt{t-1}},\bar{X}_{1:t-1}+\frac{q_{\gamma/2}}{\sqrt{{t-1}}}\right]\cap\Theta_{0},$ (S9)
+and the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (S9).
 We fix $t\in\{2,\cdots,\tau\}$.
@@
 For $\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}$, define
-$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (C.10)
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (C.4), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
-Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark C.2.
+$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S10)
+We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (S4), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
+Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark D.2.
 For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
@@
 $-2\log L_{i}\Big(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\dots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\Big)=$
-$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (C.11)
+$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (S11)
 over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
 Here, constant means that it does not involve $\theta,\theta^{\prime}$.
-Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (C.1) is a quadratic function of $\theta_{1}-\theta_{0}$.
+Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (D.1) is a quadratic function of $\theta_{1}-\theta_{0}$.
 Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
-Remark C.2.
-[…]
-$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (C.12)
-$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (C.13)
+Remark D.2.
+For weighted CUSUM or SR detectors, finding $t_{1}^{j},t_{2}^{j}$ is often straightforward.
+So, for some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$, the weighted CUSUM or SR detectors (with the Reverse Information Projection (RIPr) [17, 8, 16] being the representative of the pre-change class) are defined as
+$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S12)
+$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S13)
 where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$.
@@
 The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology.
-Note that $\mathcal{C}^{\prime\prime}$ in (C.8) is computable if similar conditions as listed in Section B.1 are met.
+Note that $\mathcal{C}^{\prime\prime}$ in (S8) is computable if similar conditions as listed in Section C.1 are met.
 For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions.
-Another concrete example with the Poisson distribution is in Section G.3.
-C.2 Another concrete example: Poisson rate change problem
+Another concrete example with the Poisson distribution is in Section H.4.
+D.2 Another concrete example: Poisson rate change problem
 Here we assume the pre and post-change model to be $\mathcal{P}_{i}=\{\text{Pois}(\theta):\theta\in\Theta_{i}\}$, for $i=0,1$ respectively.
-In this setting, we have the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (B.13) and the same method can be employed for constructing $1-\gamma$ confidence interval for $\theta_{0}$ as well.
+In this setting, we have the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (S13) and the same method can be employed for constructing $1-\gamma$ confidence interval for $\theta_{0}$ as well.
 We fix $t\in\{1,\cdots,\tau\}$.
@@
 $Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise.
-}\end{cases}$ (C.14)
+}\end{cases}$ (S14)
 Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta)$, for $n<t$, $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta^{\prime})$, for $n\geq t$ and $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
@@
 All the maximum are taken over a finite set; hence, $U_{t}^{j}$ can be computed easily.
-Appendix D Experiments for confidence intervals for pre/post change parameters
-D.1 Setting II
-Table D.1 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (5.1)) for $\theta_{1}$ across 500500500 independent runs with $\alpha=\beta=\eta=0.05$.
+Appendix E Experiments for confidence intervals for pre/post change parameters
+E.1 Setting II
+Table E.1 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (5.1)) for $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta=0.05$.
 The data generation and change detection procedure remain the same as in the experiment described in Setting II of Section 6.
-Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of 0.850.850.85, indicating that our method is reliable but conservative.
-Table D.1: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change.
+Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.85$, indicating that our method is reliable but conservative.
+Table E.1: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change.
 The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column of the table.
@@
 Conditional Length Conditional CI
-100 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.98 1.37 [0.76, 2.13]
-100 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.96 1.23 [0.90, 2.16]
-500 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.99 0.93 1.48 [0.75, 2.23]
-500 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.91 1.31 [0.90, 2.21]
-D.2 Setting III
-Table D.2 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), which are defined in (5.2) and (5.1)), for $\theta_{0}$ and $\theta_{1}$ across 500500500 independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$.
+100 $0.75,\infty)$ 0.98 0.98 1.37 [0.76, 2.13]
+100 $0.9,\infty)$ 0.98 0.96 1.23 [0.90, 2.16]
+500 $0.75,\infty)$ 0.99 0.93 1.48 [0.75, 2.23]
+500 $0.9,\infty)$ 0.98 0.91 1.31 [0.90, 2.21]
+E.2 Setting III
+Table E.2 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), and (5.1)), for $\theta_{0}$ and $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$.
 he data generation and change detection procedure remain the same as in the experiment described in Setting III of Section 6.
-Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of 0.800.800.80.
-Table D.2: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.850.85(Setting III): unknown pre- and post-change.
+Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.80$.
+Table E.2: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.85(Setting III): unknown pre- and post-change.
 The true (unknown) values of the parameters are $\theta_{0}=0$ and $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns of the table.
-T
-$\Theta_{0}$
-$\Theta_{1}$
+T $\Theta_{0}$ $\Theta_{1}$
 Cond. coverage
@@
 $\theta_{0}$ $\theta_{1}$ $\theta_{0}$ $\theta_{1}$ $\theta_{0}$ $\theta_{1}$ $\theta_{0}$ $\theta_{1}$
-100 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 1.00 0.99 1.00 0.99 0.55 1.19 [-0.32, 0.23] [0.76, 1.95]
-100 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.99 0.98 0.98 0.61 1.16 [-0.51, 0.10] [0.90, 2.06]
-500 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.99 0.99 0.93 0.93 0.21 1.22 [-0.10, 0.11] [0.75, 1.97]
-500 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.99 0.99 0.94 0.94 0.19 1.24 [-0.09, 0.10] [0.90, 2.14]
-D.3 Comparison with [36]
+100 $(\infty,0.25]$ $0.75,\infty)$ 1.00 0.99 1.00 0.99 0.55 1.19 [-0.32, 0.23] [0.76, 1.95]
+100 $(\infty,0.1]$ $0.9,\infty)$ 0.98 0.99 0.98 0.98 0.61 1.16 [-0.51, 0.10] [0.90, 2.06]
+500 $(\infty,0.25]$ $0.75,\infty)$ 0.99 0.99 0.93 0.93 0.21 1.22 [-0.10, 0.11] [0.75, 1.97]
+500 $(\infty,0.1]$ $\infty) 0.99 0$.99 0.94 0.94 0.19 1.24 [-0.09, 0.10] [0.90, 2.14]
+E.3 Comparison with [39]
 Next, we compare our confidence interval for the post-change parameter with that proposed by [42] and discussed in [39].
@@
 For our method, we assume the post-change space to be $\Theta_{1}=[0,\infty)$ with $\alpha=\beta=\eta=0.05$.
-We report the average size and the conditional and unconditional coverage rates of both methods across 500500500 independent runs in Table D.3, with a target conditional coverage of $\alpha^{\prime}=0.85$.
+We report the average size and the conditional and unconditional coverage rates of both methods across $500$ independent runs in Table E.3, with a target conditional coverage of $\alpha^{\prime}=0.85$.
 The results show that the existing method [39], which is based on asymptotic approximations, undercovers $\theta_{1}$ (both conditionally and marginally).
 At the same time, our finite-sample confidence intervals are conservative — achieving higher-than-nominal (conditional) coverage at the expense of larger (conditional) length.
-Table D.3: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage 0.850.850.85) with Wu [36] (Setting II)
+Table E.3: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage $0.85$) with Wu [39] (Setting II)
 $F_{0}$ $F_{1}$ $T$ Conditional Length Conditional Coverage Marginal Coverage
 Ours Wu (2007) Ours Wu (2007) Ours Wu (2007)
-${N}(-0.25,1)$ ${N}(0.25,1)$ 100 1.46 0.47 0.99 0.77 0.98 0.74
-${N}(-0.3,1)$ ${N}(0.3,1)$ 100 1.43 0.51 0.98 0.76 0.98 0.75
-${N}(-0.25,1)$ ${N}(0.25,1)$ 500 1.48 0.49 0.99 0.78 0.91 0.64
-${N}(-0.3,1)$ ${N}(0.3,1)$ 500 1.49 0.51 0.99 0.79 0.91 0.71
-Appendix E Additional experiments for composite pre/post change
+N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 100 1.46 0.47 0.99 0.77 0.98 0.74
+N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 100 1.43 0.51 0.98 0.76 0.98 0.75
+N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 500 1.48 0.49 0.99 0.78 0.91 0.64
+N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 500 1.49 0.51 0.99 0.79 0.91 0.71
+Appendix F Additional experiments for composite pre/post change
 In this section, we present experimental results for the alternative approach in which each set $\mathcal{S}_{t-1}$ and/or $\mathcal{S}_{\tau-t+1}^{\prime}$ is discretized using 50 equispaced grid points.
@@
 Therefore, for $j\in\{1,\cdots,B\}$, we only need to generate $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$. For $i\in[I],k\in[K]$, define
-$Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})=\begin{cases}\theta_{i}^{(0)}+\epsilon_{n}^{j},\text{ if }n<t\\ \theta_{k}^{(1)}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (E.1)
+$Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})=\begin{cases}\theta_{i}^{(0)}+\epsilon_{n}^{j},\text{ if }n<t\\ \theta_{k}^{(1)}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S1)
 We run $\mathcal{A}$ on $\{Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{n}$ and denote $\tau_{i,k}^{j}$ as the detection time.
@@
 Then, we consider the following confidence set:
-$\mathcal{C}=\left\{t\in\{1,\cdots,\tau\}:M_{t}\leq\max_{i\in[I],k\in[K]}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{j=1}^{B})\right\}$ (E.2)
+$\mathcal{C}=\left\{t\in\{1,\cdots,\tau\}:M_{t}\leq\max_{i\in[I],k\in[K]}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{j=1}^{B})\right\}$ (S2)
 As a special case, when $\theta_{0}$ is known, then, $\theta_{0}=\theta_{1}^{(0)}=\cdots=\theta_{I}^{(0)}=\mathcal{S}_{t-1}.$
-As shown in Table E.1, the confidence sets (E.2) have lengths and coverages that are comparable to those from the earlier method reported in Table 6.4 and Table 6.6.
+As shown in Table F.1, the confidence sets (S2) have lengths and coverages that are comparable to those from the earlier method reported in Table 6.3 and Table 6.5.
 However, this grid-based approach incurs significantly higher computational cost.
 Therefore, for practical purposes, we recommend using the earlier, more computationally efficient method.
-Table E.1: Composite pre/post change
+Table F.1: Composite pre/post change
 T
@@
 100 $\{N(\mu,1):\mu\leq 0.1\}$ $\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 30.14 17092.11
-Appendix F Simpler confidence sets for detection algorithms controlling PFA
+Appendix G Simpler confidence sets for detection algorithms controlling PFA
 In this section, we shall discuss how one can construct simpler confidence sets when the change detection algorithm $\mathcal{A}$ controls the probability of false alarm (PFA) at level $\delta\in(0,1)$.
@@
 We consider the test statistic $M_{t}$ introduced in (3.4), and define the following confidence set for $T$:
-$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha}\right\},$ (F.1)
+$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha}\right\},$ (S1)
 which is a simplified version of Equation 3.9, obtained by directly replacing $\alpha r_{t}^{*}$ with $\alpha$.
@@
 We now investigate the coverage properties of the above confidence set.
-Theorem F.1.
-Let $\mathcal{C}$ be the confidence set as defined in (F.1), for some $\alpha\in(0,1)$.
+Theorem G.1.
+Let $\mathcal{C}$ be the confidence set as defined in (S1), for some $\alpha\in(0,1)$.
 If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, then for any $F_{0}\in\mathcal{P}_{0}$ and $F_{1}\in\mathcal{P}_{1}$,
-$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha-\delta\leavevmode\nobreak\ \text{, and }\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{1-\delta},$
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha-\delta\penalty 10000\ \text{, and }\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{1-\delta},$
 where the latter guarantee holds if $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,$ for all $F_{0}\in\mathcal{P}_{0}.$
@@
 A similar simplification can be applied to the adaptive confidence sets introduced in Section 4 for the parametric setting.
-By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 4.1, 4.4 and 4.5, we obtain
-$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}\}_{j=1}^{B})\right\},$ (F.2)
+By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 4.1, 4.3 and 4.4, we obtain
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}\}_{j=1}^{B})\right\},$ (S2)
 for known pre- and post-change setting. Similarly, for the known pre-change and composite post-change setting,
-$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}_{\tau-t+1}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})\right\},$ (F.3)
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}_{\tau-t+1}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})\right\},$ (S3)
 And finally,
-$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (F.4)
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (S4)
 for composite pre- and post-change setting.
-For detection algorithms with PFA $\leq\delta,\leavevmode\nobreak\ \delta\in(0,1)$, these simpler forms retain the coverage guarantees — both conditional and unconditional — similar to what we have shown in the theorem above.
+For detection algorithms with PFA $\leq\delta,\penalty 10000\ \delta\in(0,1)$, these simpler forms retain the coverage guarantees — both conditional and unconditional — similar to what we have shown in the theorem above.
 In the next section, we present the experimental results with the bounded PFA algorithms, along with some other experiments.
-Appendix G Some additional experimental results
-G.1 Experiments with dependent pre- and post-change data
+Appendix H Some additional experimental results
+H.1 Experiments varying $T$ and ARL
+We have seen in deriving the asymptotic length bound that it increases with the ratio $T/A_{T}$.
+Now, we experimentally validate the fact by conducting the same experiment as in Setting I (known pre- and post-change) in Section 6 with varying $T$ thresholds $A$ of the CUSUM detector (larger threshold means large ARL).
+Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 3.5 are reported in Table H.1.
+The results show a clear pattern: as the ratio $T/A$ increases, the resulting confidence sets become wider, reflecting growing conservativeness.
+Table H.1: Setting I: Known pre- and post-change.
+Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 3.5 with $F_{0}=N(0,1),F_{1}=N(1,1),N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (3.1) and average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
+$T$ A
+Conditional
+coverage
+Marginal
+coverage
+Size
+$|\hat{T}-T|$
+Delay
+50 100 0.98 0.91 17.28 2.58 9.33
+1000 100 0.99 0.21 20.88 2.45 9.14
+2000 100 1 0.07 27.02 2.60 8.35
+500 1000 0.98 0.97 15.77 2.62 13.22
+10000 1000 1 0.27 24.29 2.36 13.25
+20000 1000 1 0.07 30.16 2.29 13.22
+H.2 Experiments with dependent pre- and post-change data
 […]
 We use the likelihood ratio-based cusum detector:
-$\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(j:n)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(j:n)}\geq A\right\},$ (G.1)
+$\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(j:n)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(j:n)}\geq A\right\},$ (S1)
 where $n_{11}(j:n)$ denotes the number of transitions from 1 to 1 and $n_{10}(j:n)$ denotes the number of transitions from 0 to 1 within $X_{j},\cdots,X_{n}$, and $A=1000$.
@@
 Note that our adaptive method (4.1) is the only one we know that is valid for dependent data.
-Table G.1 contains the results averaged across 500500500 runs.
-Table G.1: Known pre- and post-change with dependency (Setting I)
+Table H.2 contains the results averaged across $500$ runs.
+Table H.2: Known pre- and post-change with dependency (Setting I)
 T
@@
 500 0.98 0.98 35.83 4.42 27.36
-G.2 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
-G.2.1 Known pre- and post-change (Setting A, PFA 0.001$\leq$ 0.001≤ 0.001)
+H.3 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
+H.3.1 Known pre- and post-change (Setting A, PFA $\leq 0.001$)
 We perform experiments for both Gaussian and Poisson mean-change scenarios having the true changepoint at $T=100,500$, using a likelihood ratio detector defined as
-$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (G.2)
-[…]
-Table G.2: Known pre- and post-change distributions (Setting A, PFA 0.001$\leq$ 0.001≤ 0.001)
+$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (S2)
+with thresholds at $A=1000$.
+Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality [35], the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
+After detection, we construct both the proposed confidence sets (universal (S1) and adaptive (S2)), with $B=100$, $\alpha=0.05$.
+We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.3.
+We also report the average absolute deviation of $\hat{T}$ (defined in (3.1)) and delay (given no false detection) in the table.
+Table H.3: Known pre- and post-change distributions (Setting A, PFA $\leq 0.001$)
 $T$ $F_{0}$ $F_{1}$ Conditional Coverage Conditional Size
@@
 100 $N(0,1)$ $N(1,1)$ 0.95 0.94 15.44 12.60 2.74 110.91
-[…]
+100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 19.36 17.95 4.98 94.33
 500 $N(0,1)$ $N(1,1)$ 0.97 0.97 15.92 12.65 2.24 514.12
-[…]
-G.2.2 Known pre- and unknown post-change (Setting B, PFA 0.001$\leq$ 0.001≤ 0.001)
+500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 18.89 16.82 3.18 417.44
+H.3.2 Known pre- and unknown post-change (Setting B, PFA $\leq 0.001$)
 We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector defined as
-$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (G.3)
+$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (S3)
 with thresholds at $A=1000$.
 The weights are chosen just as described in Table 6.2.
-Similarly, the PFA of this detector is at most $1/A$, i.e., 0.0010.0010.001 here.
-After detection, we construct our universal confidence sets (F.1) for $T$, with $\alpha=0.075$.
+Similarly, the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
+After detection, we construct our universal confidence sets (S1) for $T$, with $\alpha=0.075$.
 The test statistic and the point estimate are the same as mentioned in Setting II of Section 6.
-We report the average size and the conditional and unconditional coverage rates across 500500500 independent runs in Table G.3.
-We also report the average absolute deviation of $\hat{T}$ in (B.2) and delay (given no false detection) in the table.
-Table G.3: Known pre-change but unknown post-change (Setting B, PFA 0.001$\leq$ 0.001≤ 0.001)
+We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.4.
+We also report the average absolute deviation of $\hat{T}$ in (S2) and delay (given no false detection) in the table.
+Table H.4: Known pre-change but unknown post-change (Setting B, PFA $\leq 0.001$)
 T
@@
 500 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 19.27 3.23 528.92
-G.2.3 Composite pre- and post-change (Setting C, PFA 0.001$\leq$ 0.001≤ 0.001)
+H.3.3 Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
 We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) [17, 8, 16] being the representative of the pre-change class) defined as
-$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (G.4)
-[…]
-Table G.4: Composite pre- and post-change (Setting C, PFA 0.001$\leq$ 0.001≤ 0.001)
+$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (S4)
+with $f^{*}_{\theta}$ being the RIPr of $f_{\theta}$ on the pre-change class $\mathcal{P}_{0}$ and with thresholds at $A=1000$.
+This is a stopping time of an “e-process” [20] and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
+The weights are chosen just as described in Table 6.3.
+After detection, we construct our universal confidence sets (S1) for $T$, with $\alpha=0.1$.
+The test statistic and the point estimate are the same as mentioned in Setting III of Section 6.
+We report the average size and coverage rates across $500$ independent runs in Table H.5.
+We also report the average absolute deviation of $\hat{T}$ in (S2) and delay (given no false detection).
+The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
+Table H.5: Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
 T
@@
 500 $\{N(\mu,1):\mu\leq 0.1\}$ $\{N(\mu,1):\mu>0.9\}$ 0.98 0.98 21.45 3.75 536.76
-G.2.4 Huber-robust (nonparametric) pre- and post-change (Setting D, PFA 0.001$\leq$ 0.001≤ 0.001)
-We perform experiments for the robust Gaussian mean-change problem under the total variation contamination model, with the true changepoint at $T=100$ and $T=500$.
-The pre- and post-change classes are total variation balls of radius $\epsilon$ around $N(\mu_{0},1)$ and $N(\mu_{1},1)$, respectively.
-$\mathcal{P}_{i}=\{F\in\mathcal{M}:d_{TV}(F,N(\mu_{i},1))\leq\epsilon\};\leavevmode\nobreak\ \leavevmode\nobreak\ i=0,1.$
-In our experiment, $\mu_{0}=0,\mu_{1}=1$ and $\epsilon=0.01,0.001$.
-The true (unknown) pre- and post-change distributions are $(1-\epsilon)\times N(0,1)+\epsilon\times\text{Cauchy}(-1,10)$ and $(1-\epsilon)\times N(1,1)+\epsilon\times\text{Cauchy}(-1,10)$ respectively.
-We use the Huber-robust tests by [23] as change-detectors:
-$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}E_{i}\geq A\right\},$ (G.5)
-[…]
-After detection, we construct our universal confidence sets (F.1) for $T$, with $R_{n}^{(t)}=\prod_{i=t}^{n}E^{\prime}_{i}$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}E_{i}$, for $n<t$ and $\alpha=0.05$, where $E^{\prime}_{i}=\frac{\pi^{\prime}(X_{i})}{\mathbb{E}_{F_{1}}\pi^{\prime}(X_{i})+(c^{\prime\prime}-c^{\prime})\epsilon},$ $\pi^{\prime}(x)=\min\{\max\{\frac{f_{0}(x)}{f_{1}(x)},c^{\prime}\},c^{\prime\prime}\}$ and $E_{i}$ is as defined before.
-We report the average size and coverage rates across 200200200 independent runs in Table G.5.
-We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection).
-The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
-Table G.5: Huber-robust (nonparametric) pre- and post-change (Setting D, PFA 0.001$\leq$ 0.001≤ 0.001)
-T $\epsilon$
-Conditional
-coverage
-Marginal
-coverage
-Size
-$|\hat{T}-T|$
-Delay
-100 0.01 0.97 0.97 16.04 3.23 119.38
-100 0.001 0.97 0.97 14.57 3.18 113.06
-500 0.01 0.96 0.96 16.18 3.87 536.58
-500 0.001 0.97 0.97 15.13 3.25 531.44
-G.2.5 Comparison of experimental results with ARL and PFA controlling detection algorithms
+H.3.4 Comparison of experimental results with ARL and PFA controlling detection algorithms
 As expected, algorithms that control PFA exhibit a significantly higher average detection delay compared to those controlling ARL.
@@
 We do not observe significant differences in the size and conditional coverage properties of the confidence sets nor in the absolute deviation of the point estimates, across these two types of detection algorithms.
-G.3 Experimental results with Poisson distribution with detection algorithms that control ARL
-G.3.1 Known pre- and post-change (Setting I)
-[…]
-Figure G.6: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois(1)(1)( 1 ) and the rest from Pois(2)(2)( 2 ), with $T=100,500$.
+H.4 Experimental results with Poisson distribution with detection algorithms that control ARL
+H.4.1 Known pre- and post-change (Setting I)
+We consider Poisson the mean-change problem: $\text{Pois}(1)$ (pre-change) vs $\text{Pois}(2)$ (post-change) with the true (unknown) changepoint at $T=100,500$.
+We employ the likelihood-based CUSUM change detection method with thresholds at $A=1000,10000$.
+After detection, we construct confidence sets (adaptive, Equation 4.1) for $T$.
+Fig.
+H.6 provides visualizations of the confidence sets and point estimates of the changepoints across $5$ runs.
+Figure H.6: Setting I (known pre- and post-change Poisson distributions): The first T1T-1 observations are drawn from Pois(1)(1) and the rest from Pois(2)(2), with $T=100,500$.
 The point estimate is shown in a vertical red dashed line, and the confidence set (using Algorithm 2) is shown in red points.
 $B=N=100$, $\alpha=0.05,L=\infty$.
-Results of 555 independent simulations are shown.
-We report the average size and the conditional and unconditional coverage rates across 500500500 independent runs in Table G.7.
-Table G.7: Setting I: Average (of 500500500 independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ and average detection delay (given a true detection) of the detection algorithm used (LABEL:eq:cusum) with threshold $A$, while varying $T$ and $A$.
-$T$ $A$
+Results of $5$ independent simulations are shown.
+We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.7.
+Table H.7: Setting I: Average (of $500$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ and average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
+$T$ AA
 Conditional
@@
 We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table.
-G.3.2 Known pre-change and composite post-change (Setting II)
-[…]
-$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (G.6)
-[…]
-We report the average size and the conditional and unconditional coverage across 500500500 independent runs in (G.8).
+H.4.2 Known pre-change and composite post-change (Setting II)
+We perform experiments for Poisson rate-change scenarios having pre-change distribution as $\text{Pois}(1)$, and the true (unknown) post-change data distribution as $\text{Pois}(2)$, with the changepoint at $T=100,200,$ assuming the composite post-change model to be $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$.
+For detection, we employ a weighted CUSUM detector defined as
+$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (S5)
+for some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$.
+We use thresholds $A=1000,10000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, and $\theta=1.9$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
+After detection, we construct confidence sets (as defined in Section C.2), with $\alpha=0.05,\beta=0.025,B=N=100,L=\infty$.
+We report the average size and the conditional and unconditional coverage across $500$ independent runs in (H.8).
 The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table G.7).
+It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table H.7).
 We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table.
 We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
-[…]
+Table H.8: Setting II: Average (of $500$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets Section 3.2.
+with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ and average detection delay (given no false detection) of the change detection method used (S11) with threshold AA, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
 T A
@@
 200 1000 0.97 0.93 35.24 5.74 21.27
-G.3.3 Composite pre- and post-change (Setting III)
-Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and 500500500, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively.
+H.4.3 Composite pre- and post-change (Setting III)
+Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively.
 For change detection, we employ a weighted CUSUM-type detector, where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, which is defined as
-$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (G.7)
-where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For the post-change parameter, the same discrete weight distribution as in Section G.3.2 is considered.
-After detection, we construct confidence sets (as defined in Section C.2, assuming $T\neq 1$), with $\alpha=0.05,\beta=\gamma=0.025,B=N=100,L=\infty$.
-We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$, and delay (given no false detection) across 500500500 independent runs in Table G.9.
-[…]
+$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S6)
+where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For the post-change parameter, the same discrete weight distribution as in Section H.4.2 is considered.
+After detection, we construct confidence sets (as defined in Section D.2, assuming $T\neq 1$), with $\alpha=0.05,\beta=\gamma=0.025,B=N=100,L=\infty$.
+We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$, and delay (given no false detection) across $500$ independent runs in Table H.9.
+Table H.9: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ and average detection delay (given no false detection) of the change detection method with threshold AA, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
 T A
@@
 However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table G.7).
-We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table G.7), exhibiting both higher coverage and larger size.
-Appendix H A heuristic method for composite pre- or post-change
+It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table H.7).
+We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table H.7), exhibiting both higher coverage and larger size.
+Appendix I A heuristic method for composite pre- or post-change
 This section presents an alternative method for constructing confidence intervals for composite pre- and post-change distributions.
 Specifically, we propose a heuristic bootstrap-based approach that estimates the unknown distributions from the observed data and subsequently generates simulations from the estimated distributions.
-H.1 Known pre-change and unknown post-change distributions
+I.1 Known pre-change and unknown post-change distributions
 Here we assume that the pre-change distribution $F_{0}$ is known and the composite post-change model is $\mathcal{P}_{1}$.
@@
 4:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
-5:for $i=1$ to $N$ do
+5:for $i=1$ to NN do
 6:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
@@
 11:  for $j=1$ to $B$ do
-12:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
+12:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\penalty 10000\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
 13:   Compute the same test statistics $M_{t}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}$
@@
 18:end for
-H.2 Composite pre- and post-change distributions
+I.2 Composite pre- and post-change distributions
 We now consider the general case where both pre and post-change distributions follow some composite models $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively.
 Following the approach from the previous sections, we define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
-$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0},\text{ for some }F_{0}\in\mathcal{P}_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},\text{ for some }F_{1}\in\mathcal{P}_{1}.$
+[…]
 For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined earier).
@@
 We first obtain some point estimate $\hat{T}$ of the changepoint $T$.
-Then we use the observations from 111 to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$.
+Then we use the observations from $1$ to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$.
 For some $N\in\mathbb{N}$, we generate $N$ many i.i.d.
@@
 5:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
-6:for $i=1$ to $N$ do
+6:for $i=1$ to NN do
 7:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
@@
 12:  for $j=1$ to $B$ do
-13:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
+13:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\penalty 10000\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
 14:   Compute the same test statistics $M_{t}^{j}$ based on $X_{1}^{j},\cdots,X_{\tau_{j}}^{j}$
@@
 19:end for
-H.3 Experimental Results
-H.3.1 Known pre-change and composite post-change (Setting II)
+I.3 Experimental Results
+I.3.1 Known pre-change and composite post-change (Setting II)
 We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$.
-For detection, we employ a weighted CUSUM detector (B.11) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
+For detection, we employ a weighted CUSUM detector (S11) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
 After detection, we employ Algorithm 4 to construct confidence sets.
-Table H.1 summarizes the results, including the average size and the conditional and unconditional coverage across 500500500 independent runs.
+Table I.1 summarizes the results, including the average size and the conditional and unconditional coverage across $500$ independent runs.
 For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible.
@@
 Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
-Table H.1: Setting II: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 4 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (B.11) with threshold $A=1000$.
+Table I.1: Setting II: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 4 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (S11) with threshold $A=1000$.
 T Pre-change
@@
 100 $N(0,1)$ $\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.94 0.94 14.76
-[…]
+100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.95 20.95
 500 $N(0,1)$ $\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.95 0.91 14.06
-[…]
-H.3.2 Composite pre- and post-change (Setting III)
-Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and 500500500.
-We use a weighted CUSUM detector with RIPr defined in (C.12) for change detection.
+500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.89 20.83
+I.3.2 Composite pre- and post-change (Setting III)
+Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$.
+We use a weighted CUSUM detector with RIPr defined in (S12) for change detection.
 The thresholds are set at $A=1000$ for both $T=100$ and $T=500$.
@@
 After detection, we employ Algorithm 5 to construct confidence sets.
-We report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table H.2.
+We report the average size and the conditional and unconditional coverage across $500$ independent runs in Table I.2.
 The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
-Table H.2: Setting III: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (C.12) with threshold $A=1000$.
+Table I.2: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (S12) with threshold $A=1000$.
 T
@@
 100 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 19.52
-[…]
+100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.94 0.94 23.32
 500 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 18.56
-[…]
-Appendix I Another heuristic method for composite pre- or post-change
+500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 23.04
+Appendix J Another heuristic method for composite pre- or post-change
 In this section, we provide another heuristic method for dealing with composite pre- or post-change distributions.
@@
 Nonetheless, we do not have any theoretical guarantee for coverage for this method.
-For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section H.3.1 and we report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table I.1.
-Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section H.3.2 and we report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table I.1.
-From Table I.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
-Table I.1: Setting II: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (B.11) with threshold $A=1000$.
+For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section I.3.1 and we report the average size and the conditional and unconditional coverage across $500$ independent runs in Table J.1.
+Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section I.3.2 and we report the average size and the conditional and unconditional coverage across $500$ independent runs in Table J.1.
+From Table J.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
+Table J.1: Setting II: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (S11) with threshold $A=1000$.
 T Pre-change
@@
 100 $N(0,1)$ $\{N(\mu,1):\mu>0.75\}$ $N(1,1)$ 0.95 0.95 23.16
-[…]
+100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.94 0.94 25.83
 500 $N(0,1)$ $\{N(\mu,1):\mu>0.75\}$ $N(1,1)$ 0.95 0.88 23.82
-[…]
-Table I.2: Setting III: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (C.12) with threshold $A=1000$.
+500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.90 27.17
+Table J.2: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (S12) with threshold $A=1000$.
 T
@@
 100 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 26.03
-[…]
+100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 29.64
 500 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 26.52
-[…]
+500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 30.49

### v6

@@
 A sequence of $\mathcal{X}$-valued observations ${\bf X}=X_{1},X_{2},\cdots$ arrives sequentially, such that for some unknown time $T\in\mathbb{N}\cup\{\infty\}$ (the changepoint),
-$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1.1)
+$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1)
 for some $F_{0}$ in a pre-change class $\mathcal{P}_{0}$ and $F_{1}$ in a post-change class $\mathcal{P}_{1}$, where $\mathcal{P}_{0}$ and $\mathcal{P}_{0}$ denote some non-intersecting classes of probability distributions.
@@
 Detecting when such a shift occurs is vital for understanding emerging issues, tracing public reactions to events, or monitoring product reputation in real time.
-To illustrate the relevance of our methodology in such settings, we conduct an experiment using the Stanford Sentiment Treebank (SST-2) dataset [31], simulating a stream of text reviews whose sentiment abruptly transitions from predominantly positive to predominantly negative.
+To illustrate the relevance of our methodology in such settings, we conduct an experiment using the Stanford Sentiment Treebank (SST-2) dataset [socher2013recursive], simulating a stream of text reviews whose sentiment abruptly transitions from predominantly positive to predominantly negative.
 Accurately localizing the changepoint is important: it can help investigate why the change may have occurred.
 As an initial goal, one may hope to obtain a set $\mathcal{C}\subseteq\{1,\dots,\tau\}$ such that
-$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha,$ (1.2)
-where $\mathbb{P}_{F_{0},T,F_{1}}$ is the data distribution in (1.1), i.e. having the true changepoint at $T$. But this goal is harder to achieve than may initally appear; we elaborate below.
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha,$ (2)
+where $\mathbb{P}_{F_{0},T,F_{1}}$ is the data distribution in (1), i.e. having the true changepoint at $T$. But this goal is harder to achieve than may initally appear; we elaborate below.
 There are two classes of algorithms $\mathcal{A}$ in the literature: (i) algorithms with probability of false alarm (PFA) bounded by $\delta<1$; we call these “bounded PFA” for short, and (ii) algorithms with PFA equal to one, but have a bounded average run length to false alarm (ARL2FA, or just ARL for short); we call these “bounded ARL”.
@@
 In this paper, we remain agnostic to the choice of $\mathcal{A}$.
-For algorithms with a bounded ARL, we prove that no set $\mathcal{C}$ can satisfy guarantee (1.2) (for algorithms with bounded PFA, we show how to achieve (1.2) in the Supplementary).
+For algorithms with a bounded ARL, we prove that no set $\mathcal{C}$ can satisfy guarantee (2) (for algorithms with bounded PFA, we show how to achieve (2) in the Supplementary Section F).
 This impossibility result (which we prove formally later) intuitively holds because it is meaningless to ask for coverage if the algorithm stopped before the changepoint; indeed, on the event $\{\tau<T\}$, we have no information to rule out the possibility that $T=\infty$, and thus the only way to cover $T$ is to output the set of all natural numbers $\mathbb{N}$.
 Thus, an arguably more sensible goal is to have
-$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,$ (1.3)
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,$ (3)
 meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$.
-This definition has been considered by, for example, [5] and [42, 39], who achieved it (in a certain asymptotic sense) in much more restricted settings than the current investigation.
-The aforesaid impossibility result no longer applies to the new goal (1.3).
+This definition has been considered by, for example, [ding2003lower] and [WU20063625, wu2007inference], who achieved it (in a certain asymptotic sense) in much more restricted settings than the current investigation.
+The aforesaid impossibility result no longer applies to the new goal (3).
 It is important to note that we do not prescribe which $\mathcal{A}$ to use in which situation, but simply enable post-hoc inference on $T$ for any user-chosen $\mathcal{A}$.
-One can view our techniques as providing a wrapper to convert $\mathcal{A}$ from producing only a stopping time $\tau$ to also producing a confidence set for $T$  [^2] In some sense, our goal can be viewed in an analogous way to conformal inference, which wraps around an arbitrary prediction algorithm to convert a “point prediction” into a prediction set, by running the same algorithm on many hypothetical datasets (for full conformal prediction).
-However, our methods are not distribution-free, and rely on smartly using simulations to accomplish the goal..
-As we survey below, there is a rich literature on designing change detection algorithms $\mathcal{A}$, but a substantial paucity of inference methodology for localizing the change after detection; our focus is on the latter goal, assuming the user has chosen a suitable algorithm $\mathcal{A}$ for detection.
+One can view our techniques as providing a wrapper to convert $\mathcal{A}$ from producing only a stopping time $\tau$ to also producing a confidence set for $T$ As we survey below, there is a rich literature on designing change detection algorithms $\mathcal{A}$, but a substantial paucity of inference methodology for localizing the change after detection; our focus is on the latter goal, assuming the user has chosen a suitable algorithm $\mathcal{A}$ for detection.
 While inference for $T$ is intrinsically interesting and important, a side benefit of confidence sets for $T$ is that one can use these to derive post-hoc confidence sets for unknown parameters of $F_{0}$ and $F_{1}$.
 To clarify, suppose $\mathcal{P}_{i}=\{F_{\theta}\}_{\theta\in\Theta_{i}};i=0,1$.
-Our method that delivers (1.3) also delivers confidence sets $\mathcal{C}^{\prime}\subseteq\Theta_{0}\times\Theta_{1}$ such that $\mathbb{P}_{\theta_{0},T,\theta_{1}}((\theta_{0},\theta_{1})\in\mathcal{C}^{\prime}\mid\tau\geq T)\geq 1-\alpha,\forall\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},$ where $\mathbb{P}_{\theta_{0},T,\theta_{1}}$ is the distribution of the sequence $X_{1},\dots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{0}}$, $X_{T},X_{T+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{1}}$.
+Our method that delivers (3) also delivers confidence sets $\mathcal{C}^{\prime}\subseteq\Theta_{0}\times\Theta_{1}$ such that $\mathbb{P}_{\theta_{0},T,\theta_{1}}((\theta_{0},\theta_{1})\in\mathcal{C}^{\prime}\mid\tau\geq T)\geq 1-\alpha,\forall\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},$ where $\mathbb{P}_{\theta_{0},T,\theta_{1}}$ is the distribution of the sequence $X_{1},\dots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{0}}$, $X_{T},X_{T+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{1}}$.
 We stress that our confidence sets do not rely on asymptotics and are valid in finite samples.
@@
 This gap is significant because in most real-world applications, data arrives sequentially, and it can be essential to make timely inferences about the changepoint immediately upon detecting a change.
-We are aware of only two directly related works: [5] and [39, Chapter 3].
-They also chose to aim for the guarantee (1.3) like us.
-Both [5, 39] provide asymptotically valid inference after the CUSUM procedure [19] declares a change for known $F_{0}$ and $F_{1}$ in an exponential family; the former only constructs a lower confidence bound for $T$ while the latter also provides a confidence set for $T$.
+We are aware of only two directly related works: [ding2003lower] and [wu2007inference, Chapter 3].
+They also chose to aim for the guarantee (3) like us.
+Both [ding2003lower, wu2007inference] provide asymptotically valid inference after the CUSUM procedure [page1954continuous] declares a change for known $F_{0}$ and $F_{1}$ in an exponential family; the former only constructs a lower confidence bound for $T$ while the latter also provides a confidence set for $T$.
 Their approaches have notable methodological and theoretical limitations: (i) they are tailored specifically to the CUSUM procedure, (ii) they assume that the pre- and post-change distributions are both known and they belong to an exponential family, (iii) they only provide asymptotic coverage guarantees (valid when both $T\to\infty$ and ARL $\to\infty$).
@@
 Thus, our framework is the first nonparametric, theoretically sound, and widely applicable practical tool for sequential change localization tasks.
-We summarize the comparison in Table 1.1.
-Table 1.1: Comparison of existing methods and our proposed approach
-Pre-, Post- Existing Methods [5, 39] Our Proposed Method
+We summarize the comparison in Table 1.
+Table 1: Comparison of existing methods and our proposed approach
+Pre-, Post- Existing Methods [ding2003lower, wu2007inference] Our Proposed Method
 change distributions Assumptions Guarantee Assumptions Guarantee
 Known, Known $\mathcal{A}=$ CUSUM. $F_{0},F_{1}$ Asymptotic No assumption Nonasymptotic
-($F_{0},F_{1}$) $\in$ exponential family ($T$, ARL $\to\infty$) (any $\mathcal{A},F_{0},F_{1}$) (Theorem 4.1)
+($F_{0},F_{1}$) $\in$ exponential family ($T$, ARL $\to\infty$) (any $\mathcal{A},F_{0},F_{1}$) (Theorem 6)
 Known, Unknown No existing work No assumption Nonasymptotic
-($F_{0},\mathcal{P}_{1}$) (any $\mathcal{A},F_{0},\mathcal{P}_{1}$) (Theorem 3.2)
-Unknown, Unknown No existing work Assumption 3.5 Nonasymptotic
-($\mathcal{P}_{0},\mathcal{P}_{1}$) on $\mathcal{A},\mathcal{P}_{0}$ (any $\mathcal{P}_{1}$) (Theorem 3.8)
+($F_{0},\mathcal{P}_{1}$) (any $\mathcal{A},F_{0},\mathcal{P}_{1}$) (Theorem 2)
+Unknown, Unknown No existing work Assumption 1 Nonasymptotic
+($\mathcal{P}_{0},\mathcal{P}_{1}$) on $\mathcal{A},\mathcal{P}_{0}$ (any $\mathcal{P}_{1}$) (Theorem 4)
 Literature on sequential change detection.
@@
 This methodology has proved crucial in many applications, such as quality control, cybersecurity, medical diagnostics, and so on.
-Classical methods such as Shewhart control charts [27], the CUSUM procedure [19], and the Shiryaev-Roberts approach [29, 23] assume known pre- and post-change distributions.
-To address scenarios with unknown post-change distributions, generalized likelihood ratio (GLR) procedures [18, 30] were developed.
-Comprehensive reviews of these methods are provided by [33].
-More recently, kernel-based techniques [9], e-detectors [28], etc., have emerged to handle nonparametric settings.
+Classical methods such as Shewhart control charts [shewhart1925application], the CUSUM procedure [page1954continuous], and the Shiryaev-Roberts approach [shiryaev1963optimum, roberts1966comparison] assume known pre- and post-change distributions.
+To address scenarios with unknown post-change distributions, generalized likelihood ratio (GLR) procedures [lorden1971procedures, siegmund1995using] were developed.
+Comprehensive reviews of these methods are provided by [tartakovsky2014sequential].
+More recently, kernel-based techniques [harchaoui2008kernel], e-detectors [shin2022detectors], etc., have emerged to handle nonparametric settings.
+All these methods focus on the quickest change detection, with average run length (ARL) control.
+In contrast, another strand of the literature emphasizes controlling the false alarm rate instead of ARL, often at the cost of a longer detection delay [chu1996monitoring, kirch2022sequential, aue2024state].
+These approaches typically rely on an initial historical period of data to calibrate the detection thresholds.
 Offline vs.
@@
 As far as we are aware, most existing approaches to changepoint localization focus on offline settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times.
-Classical works, e.g., [10] study the asymptotic behavior of the MLE for the changepoint and construct likelihood-ratio-based confidence sets.
-[3] proposed some nonparametric estimators.
-[38] developed likelihood ratio-based confidence sets for exponential family distributions, while [6] used bootstrap-based tests for each candidate time to construct nonparametric confidence sets with asymptotic coverage.
-Recently, [34] studies optimal rates for both changepoint detection and localization.
-[15] constructs confidence sets by evaluating a local test statistic.
+Classical works, e.g., [hinkley1970inferencce] study the asymptotic behavior of the MLE for the changepoint and construct likelihood-ratio-based confidence sets.
+[darkhovskh1976a] proposed some nonparametric estimators.
+[worsley1986confidence] developed likelihood ratio-based confidence sets for exponential family distributions, while [dumbgen1991asymptotic] used bootstrap-based tests for each candidate time to construct nonparametric confidence sets with asymptotic coverage.
+Recently, [verzelen2023optimal] studies optimal rates for both changepoint detection and localization.
+[jang2024fast] constructs confidence sets by evaluating a local test statistic.
 To re-emphasize, all these methods are for the offline setting.
-In the sequential setting, [32, 7, 40, 1] develop point estimators for $T$; but very few works construct confidence sets [5, 39], whose drawbacks were discussed earlier.
+In the sequential setting, [Srivastava01011999, Gombay10012003, wu2004bias, Brodsky30042010] develop point estimators for $T$; but very few works construct confidence sets [ding2003lower, wu2007inference], whose drawbacks were discussed earlier.
 Inference on $\theta_{0},\theta_{1}$ after stopping.
 As with the problem of post-detection inference on the changepoint $T$, there is limited literature on valid estimation or inference for the pre- and post-change parameters $\theta_{0}$ and $\theta_{1}$ post-stopping.
-[42] constructs asymptotic bias-corrected point estimates and confidence intervals for the post-change mean after a mean change is detected by a CUSUM procedure.
-[41] generalizes these results for the case when there is a possible change in the variance as well.
-These methods are also discussed in [39].
+[WU20063625] constructs asymptotic bias-corrected point estimates and confidence intervals for the post-change mean after a mean change is detected by a CUSUM procedure.
+[wu2005inference] generalizes these results for the case when there is a possible change in the variance as well.
+These methods are also discussed in [wu2007inference].
 However, these approaches share several limitations with the existing post-detection changepoint inference methods, as mentioned earlier: they rely on asymptotics, assume specific parametric forms, and apply only to the CUSUM procedure.
@@
   • We propose a general framework for constructing a confidence set for the changepoint based on inverting tests obtained using a universal threshold, assuming all observations to be independent.
-We show that it provides $1-\alpha$ conditional coverage for known pre-change and arbitrary (nonparametric) post-change classes under no additional assumption in Theorem 3.2.
-We also extend it for composite pre-change under a suitable assumption in Theorem 3.8.
+We show that it provides $1-\alpha$ conditional coverage for known pre-change and arbitrary (nonparametric) post-change classes under no additional assumption in Theorem 2.
+We also extend it for composite pre-change under a suitable assumption in Theorem 4.
   • We propose an alternative method of constructing the tests using adaptive thresholds (which is often tighter than the universal threshold), assuming pre- and post-change distributions to be known, in Algorithm 1.
@@
 These adaptive methods allow the observations to be dependent.
-We establish corresponding coverage guarantees in Theorems 4.1, 4.3 and 4.4.
-  • As a byproduct of our methods, we derive confidence sets for the unknown pre- and post-change parameters and show their coverage guarantees in Theorem 5.1.
+We establish corresponding coverage guarantees in Theorems 6, 7 and 8.
+  • As a byproduct of our methods, we derive confidence sets for the unknown pre- and post-change parameters and show their coverage guarantees in Theorem 9.
 In short, we present the first nonparametric, theoretically grounded, and widely applicable post-detection confidence sets that are agnostic to the detection method.
@@
 In Section 5, we derive confidence sets for the changepoint, as well as the pre- and post-change parameters.
-Section 6 presents an extensive simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach.
+Section 6 presents an extensive simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach  [^2] Codes are available at https://github.com/Aytijhya/Seq-changepoint-localization.
 This article is concluded in Section 7.
@@
 In this section, we highlight some key observations that shape our methodology of post-detection inference for the change-point $T$.
+We begin by clarifying some notational conventions that will be used throughout the paper.
 Notation.
@@
 Under dependence, $F_{0}$ and $F_{1}$ denotes the joint distributions of $X_{1},\dots,X_{t-1}$ and $X_{t},X_{t+1},\dots$ respectively, and is in fact ambivalent to whether $X_{t}$ is drawn independently of $X_{1},\dots,X_{t-1}$ or conditional on it, so we do not emphasize the distinction in our notation.
-As a corner cases, let $\mathbb{P}_{F_{0},\infty}$ be the joint distribution when $X_{1},X_{2},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{0}$.
+As a corner case, let $\mathbb{P}_{F_{0},\infty}$ be the joint distribution when $X_{1},X_{2},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{0}$.
 In parametric settings, we write $\mathbb{P}_{\theta_{0},t,\theta_{1}}$ and $\mathbb{P}_{\theta_{0},\infty}$ as shorthand for $\mathbb{P}_{F_{\theta_{0}},t,F_{\theta_{1}}}$ and $\mathbb{P}_{F_{\theta_{0}},\infty}$ respectively.
@@
 The next proposition formalizes this fact.
-Proposition 2.1.
+Proposition 1.
 If, for some $\alpha\in(0,1)$, a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}^{*})\geq 1-\alpha$, $\forall T\in\mathbb{N}$, then $\mathbb{P}_{F_{0},\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL is infinite).
@@
 Thus, for post-detection inference, the usual notion of coverage is not an appropriate one.
-Instead, we focus on conditional coverage (1.3), a more reasonable criterion, which was also considered in prior work [5, 39].
+Instead, we focus on conditional coverage (3), a more reasonable criterion, which was also considered in prior work [ding2003lower, wu2007inference].
 The key idea behind our methodology is that for each candidate $t\in\{1,\cdots,\tau\}$, we test whether $t$ is the true changepoint from some pre-change class $\mathcal{P}_{0}$ to some post-change class $\mathcal{P}_{1}$:
-$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (2.1)
+$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (4)
 Let $\phi_{\alpha^{*}}^{(t)}=\mathds{1}(H_{0,t}\text{ is rejected})$ denote some test for the null $H_{0,t}$ such that $\mathbb{P}_{F_{0},t,F_{1}}(\phi_{\alpha^{*}}^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha^{*},\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$ and form the confidence set by inverting these tests:
-$\mathcal{C}:=\{t\in\mathbb{N}:t\leq\tau,H_{0,t}\text{ is not rejected}\}$ (2.2)
+$\mathcal{C}:=\{t\in\mathbb{N}:t\leq\tau,H_{0,t}\text{ is not rejected}\}.$ (5)
 In classical statistics, inverting level-$\alpha$ tests yields $1-\alpha$ confidence sets.
 However, this relationship does not directly hold for us since we test the hypotheses up to a data-dependent time $\tau$, and we are considering conditional coverage.
-Proposition 2.2.
-Let $\tau$ be a stopping time and $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying (2.4).
-Then, the test defined by $\phi^{(t)}(X_{1},\cdots,X_{\tau})=\mathds{1}(t\notin\mathcal{C})$ satisfies
-$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall t\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.3)
-Conversely, given any test $\phi^{(t)}$ for $H_{0,t}$ using data up to $\tau$ satisfying (2.3), the inverted confidence set $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
-$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\penalty 10000\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.4)
+Proposition 2.
+Let $\tau$ be a stopping time and $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying (7).
+Then, the test defined by $\phi^{(t)}(X_{1},\cdots,X_{\tau})=\mathds{1}(t\notin\mathcal{C},t\leq\tau)$ satisfies
+$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall t\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (6)
+Conversely, given any test $\phi^{(t)}$ for $H_{0,t}$ using data up to $\tau$ satisfying (6), the inverted confidence set $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
+$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\penalty 10000\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (7)
 The above result establishes a duality between tests and a post-detection confidence set. Next, we construct the latter by inverting tests based on the above.
@@
 Let $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ denote the pre- and post-change classes respectively, $\mathcal{P}_{0}\cap\mathcal{P}_{1}=\emptyset$.
-Our goal is to construct post-detection confidence sets for the changepoint $T$ by inverting tests for $H_{0,t}$, as formalized in Proposition 2.2.
-To motivate the general construction, we begin with the simplest case when both pre- and post-change distributions $F_{0}$ and $F_{1}$ (with densities $f_{0}$ and $f_{1}$) are known.
-Then, the likelihood of the data $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}$, $X_{T},\cdots,X_{n}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}$ is given by $L_{T}^{n}:=\prod_{i=1}^{T-1}f_{0}(X_{i})\prod_{i=T}^{n}f_{1}(X_{i})$, and the maximum likelihood estimator of $T$ [10] can be written as $\hat{T}_{n}=\operatorname*{arg\,max}_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ In our setting, having data is observed up to a stopping time $\tau$, the analogous estimator is obtained by replacing $n$ with $\tau$ [40, 39]:
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ (3.1)
-Now, a prototypical likelihood-ratio-based test statistic is the following: $M_{t}=\displaystyle\max_{1\leq i\leq\tau}{L_{i}^{\tau}}/{L_{t}^{\tau}},\text{ if }t\leq\tau<\infty$ and $M_{t}=-\infty\text{ if }t>\tau\text{ or }\tau=\infty$.
-It simplifies to
+Our goal is to construct post-detection confidence sets for the changepoint $T$ by inverting tests for $H_{0,t}$, as formalized in Proposition 2.
 […]
-One natural way to extend this test statistic for the general nonparametric setting is to replace the likelihood ratio processes by e-processes [20, 8], which have emerged as a central tool in anytime-valid inference.
-While e-processes have been used for sequential testing [36, 21], estimation [37], and change detection [28], we employ them for the first time for changepoint localization.
-For $t\in\mathbb{N}$, we define two different types of e-processes:
+$M_{t}=\begin{cases}\displaystyle\max\left\{\max_{1\leq j\leq t-1}\prod_{i=j}^{t-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})},\max_{t\leq j\leq\tau}\prod_{i=t}^{j}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right\},&\text{ if }t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (8)
+One natural way to extend this test statistic for the general nonparametric setting is to replace the likelihood ratio by e-processes [ramdas2022game, grunwald2024safe], which have emerged as a central tool in anytime-valid inference.
+While e-processes have been used for sequential testing [wasserman2020universal, ramdas2022testing, saha2024testing], estimation [waudby2023estimating], and change detection [shin2022detectors], we employ them for the first time for changepoint localization.
+For $t\in\mathbb{N}$, we use two different types of e-processes:
   • Forward $t$-delay e-process under $\mathcal{P}_{1}$ (against $\mathcal{P}_{0}$).
@@
 This is a stochastic process $\{S^{(t)}_{n}\}_{n<t}$ adapted to the backward filtration $\{\mathcal{G}^{(t)}_{n}\}_{n=t-1,t-2,\cdots}$, where $\mathcal{G}^{(t)}_{n}:=\sigma(X_{n},X_{n+1},\cdots,X_{t-1}),$ such that for all stopping times $\kappa$ with respect to $\{\mathcal{G}^{(t)}_{n}\}_{n<t}$, $\sup_{F_{0}\in\mathcal{P}_{0}}\mathbb{E}_{F_{0}}[S^{(t)}_{\kappa}]\leq 1.$ Ideally, this process should have power against $\mathcal{P}_{1}$, i.e., $S^{(t)}_{n}$ should grow to infinity (as $n\to\infty$) under $\mathcal{P}_{0}$.
-Such powered e-processes for $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ (and vice versa) exist if the total variation distance between the two convex hulls of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ is positive (see e.g., Section 3.4 of [22]).
+Such powered e-processes for $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ (and vice versa) exist if the total variation distance between the two convex hulls of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ is positive (see e.g., Section 3.4 of [ramdas2024hypothesis]).
 The likelihood ratio process is the simplest example of an e-process.
-There is a substantial literature on constructing e-processes in both parametric (e.g., [36]) and nonparametric settings (e.g., [37]).
-For readers less familiar with these developments, we refer to Chapters 6 and 7 of the recent textbook [22].
-Note that we require all $X_{i}$’s to be independent for both powered (nontrivial) e-processes to be adapted to their respective filtrations.
-The final ingredient we need is $\hat{T}$, which is a preliminary point estimate of $T$.
-[^4] Here, we again note that our confidence set based on a point estimate $\hat{T}$ can be viewed analogously to the conformal prediction set, which is constructed around a preliminary “point prediction”.
-One can easily construct a point estimate.
-For example, let $\{S^{(t)}_{n}\}_{n\geq t}$ be some forward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$).
-Then, one reasonable estimate is
-$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau}S^{(t)}_{\tau}$. (3.3)
-This can be viewed as a nonparametric generalization of the maximum likelihood estimator for known pre- and post-change defined in (3.1).
+There is a substantial literature on constructing e-processes in both parametric (e.g., [wasserman2020universal]) and nonparametric settings (e.g., [waudby2023estimating]).
+For readers less familiar with these developments, we refer to Chapters 6 and 7 of the recent textbook [ramdas2024hypothesis].
 With the above ingredients in place, we construct a test statistic for $H_{0,t}$ by comparing it against a competing hypothesis $H_{0,T}$.
@@
 Similarly, if $t>T$, we use a backward e-process $S^{(t)}_{T}$ to accumulate evidence for $\mathcal{P}_{0}$ against $\mathcal{P}_{1}$, using the segment $X_{T},\dots,X_{t-1}$ .
-Since the true $T$ is unknown, we replace it with an estimated change-point $\hat{T}$.
-Fig.
-3.1 provides an illustration of the key idea.
-This naturally leads us to define:
-[…]
-Figure 3.1: Pre-change and post-change parts are shown in black and blue, respectively.
-Outside the time segment $t,\dots,{\hat{T}-1}]$ when $t<\hat{T}$ (or, $\hat{T},\dots,{t-1}]$ when $t>\hat{T}$) both $H_{0,t}$ and $H_{0,\hat{T}}$ agree on the distribution of the observations.
+Since the true $T$ is unknown, we consider the maximum over all valid choices of $T$ for both the e-processes, i.e., $\max_{t\leq n\leq\tau}R^{(t)}_{n}$ and $\max_{1\leq n\leq t-1}S^{(t)}_{n}$.
+We define the final test statistic as the maximum of these two quantities:
+$M_{t}=\begin{cases}\displaystyle\max\left\{\max_{t\leq n\leq\tau}R^{(t)}_{n},\max_{1\leq n\leq t-1}S^{(t)}_{n}\right\},&\text{ if }t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (9)
 Although, $M_{t}$ is not an e-variable under $H_{0,t}$ (nor are its constant multiples), thresholding $\frac{M_{t}}{2}$ at $\frac{1}{\alpha}$ yields a level-$\alpha$ test for $H_{0,t}$ — a nontrivial fact shown below.
-Theorem 3.1.
+Theorem 1.
 $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha]\leq\alpha,$ for any $t\in\mathbb{N},F_{0}\in\mathcal{P}_{0},$ and $F_{1}\in\mathcal{P}_{1}.$
 While this bound may be conservative, it is remarkably universal: it holds without any assumptions on the distribution classes $\mathcal{P}_{0}$, $\mathcal{P}_{1}$, or on the detection method $\mathcal{A}$.
-For constructing a valid confidence set, Proposition 2.2 implies that we require tests of level $\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t)$, whose value is of course not known to us.
+For constructing a valid confidence set, Proposition 2 implies that we require tests of level $\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t)$, whose value is of course not known to us.
 To circumvent this issue, we first assume that the pre-change distribution $F_{0}$ is known.
@@
 Let $r_{t}$ be an unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, that is independent of the data. Then, define the confidence set as
-$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}}\right\}.$ (3.5)
+$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}}\right\}.$ (10)
 One natural approach to estimate $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$ is via Monte Carlo simulation.
@@
 The next theorem proves the validity of the above confidence set.
-Theorem 3.2.
+Theorem 2.
 Suppose that the pre-change distribution is known, i.e., $\mathcal{P}_{0}=F_{0}$.
-Then, for any $\alpha\in(0,1),T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfying $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (3.5), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{1}\in\mathcal{P}_{1}.$
-Now we analyze the size of the confidence set (3.5) for known pre-change $F_{0}$ and post-change $F_{1}$, with $r_{t}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)}{N+1}$, $M_{t}$ as defined in (3.2) with $\hat{T}$ as defined in (3.1).
+Then, for any $\alpha\in(0,1),T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfying $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (10), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{1}\in\mathcal{P}_{1}.$
+Now we analyze the size of the confidence set (10) for known pre-change $F_{0}$ and post-change $F_{1}$, with $r_{t}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)}{N+1}$, $M_{t}$ as defined in (8).
 We provide a bound on the expected size of the confidence set as $N\to\infty$, where $N$ is the number of simulations used to calculate $r_{t}$.
@@
 Finally, we call the stopping time $\tau$ as sensitive if
-$\penalty 10000\ \mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]\geq\mathbb{P}_{F_{0},\infty}[\tau\leq t],$ (3.6)
+$\penalty 10000\ \mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]\geq\mathbb{P}_{F_{0},\infty}[\tau\leq t],$ (11)
 for all $t$ and $T$.
@@
 We can expect all practically used detection algorithms to be sensitive, as they should stop later when a change is absent than when it is present.
-Theorem 3.3.
-In the setting described above, for any change detection method with stopping time $\tau$ satisfying and $p_{T}\neq 0$, the size of $\mathcal{C}$} can be bounded as:
-$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\cdot\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{F_{0},T,F_{1}}.$ (3.7)
-Additionally, if $\tau$ is sensitive in the sense of (3.6), the last term in (3.7) can be improved to $\Psi_{F_{0},T,F_{1}}:=\min\left\{\Delta_{F_{0},T,F_{1}},\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right)\right\}$.
-Remark 3.4.
+Theorem 3.
+In the setting described above, for any change detection method with stopping time $\tau$ satisfying $p_{T}\neq 0$, the size of $\mathcal{C}$} can be bounded as:
+$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\cdot\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{F_{0},T,F_{1}}.$ (12)
+Additionally, if $\tau$ is sensitive in the sense of (11), the last term in (12) can be improved to $\Psi_{F_{0},T,F_{1}}:=\min\left\{\Delta_{F_{0},T,F_{1}},\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right)\right\}$.
+Remark 1.
 Suppose $F_{1}=N(\mu_{1},\sigma^{2})$ vs $F_{0}=N(\mu_{0},\sigma^{2})$ and we are using a CUSUM detector with the likelihood ratio (which satisfies the assumption) with average run length $A$.
-Consider the asymptotic regime when $A\to\infty$, and $T=CA$, for some $C>0$.
+Consider the asymptotic regime when $A\to\infty$, and $T=CA$, for some constant $C>0$.
 Using the property of the CUSUM detector that $\frac{\tau}{\mathbb{E}_{F_{0},\infty}(\tau)}\stackrel{{\scriptstyle d}}{{\to}}\exp(1)$ as $A\to\infty$, we have $p_{T}=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq CA_{T})\to\exp(-C)$.
@@
 The fact that the bound grows with $C=T/A$ is empirically verified in Supplementary Section G.1.
-This behavior is not a limitation of our methodology; rather, it reflects a fundamental difficulty of the problem: $p_{t}$ decays as $C$ increases, and hence, by our duality result (Proposition 2.2), the tests corresponding to any confidence set will have smaller levels $\alpha p_{t}$, leading to reduced power and, consequently, wider confidence sets.
+This behavior is not a limitation of our methodology; rather, it reflects a fundamental difficulty of the problem: $p_{t}$ decays as $C$ increases, and hence, by our duality result (Proposition 2), the tests corresponding to any confidence set will have smaller levels $\alpha p_{t}$, leading to reduced power and, consequently, wider confidence sets.
 3.2 An extension to composite pre-change
 To extend our method to composite pre-change classes $\mathcal{P}_{0}$, we require the following assumption on the change detection method and $\mathcal{P}_{0}$:
-Assumption 3.5.
+Assumption 1.
 $\exists P_{0}^{*}\in\mathcal{P}_{0}$ such that $\inf_{P_{0}\in\mathcal{P}_{0}}\mathbb{P}_{P_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{P_{0}^{*},\infty}(\tau\geq t)$, $\forall t\in\mathbb{N}$.
-We justify this assumption by appealing to Huber’s notion of “least favorable distribution pair” [13] in terms of risk, where the risk $R$ for a hypothesis test $\phi$ between $\mathcal{P}_{0}$ vs.
+We justify this assumption by appealing to Huber’s notion of “least favorable distribution pair” [huber1965robust] in terms of risk, where the risk $R$ for a hypothesis test $\phi$ between $\mathcal{P}_{0}$ vs.
 $\mathcal{P}_{1}$ is defined as
-$R(P_{0},\phi)=C_{0}P_{0}(\phi\text{ accepts }\mathcal{P}_{1}),\penalty 10000\ \penalty 10000\ R(P_{1},\phi)=C_{1}P_{1}(\phi\text{ accepts }\mathcal{P}_{0}).$ (3.8)
+$R(P_{0},\phi)=C_{0}P_{0}(\phi\text{ accepts }\mathcal{P}_{1}),\penalty 10000\ \penalty 10000\ R(P_{1},\phi)=C_{1}P_{1}(\phi\text{ accepts }\mathcal{P}_{0}).$ (13)
 The formal definition of a least favorable distribution pair is given below.
-Definition 3.6.
+Definition 1.
 A pair $(P_{0}^{*},P_{1}^{*})\in\mathcal{P}_{0}\times\mathcal{P}_{1}$ is called least favorable distribution (LFD) pair in terms of risk $R$ for testing $\mathcal{P}_{0}$ vs.
@@
 The following proposition establishes that the existence of an LFD pair is sufficient for our assumption to hold.
-Proposition 3.7.
+Proposition 3.
 $(P_{0}^{*},P_{1}^{*})$ is a least favourable distribution pair in terms of risk for testing $\mathcal{P}_{0}$ vs.
-$\mathcal{P}_{1}$, then Assumption 3.5 holds with $F^{*}=P^{*}$, and $\tau$ being the stopping time of a CUSUM/SR change detector with the likelihood ratio of $P_{1}^{*}$ to $P_{0}^{*}$.
-LFD pairs are known to exist for various parametric and nonparametric models, e.g., the monotone likelihood ratio (MLR) families, and Huber’s robust models (including $\epsilon$ neighborhood model, total variation model, and more generally with 2-alternating capacities [14]).
+$\mathcal{P}_{1}$, then Assumption 1 holds with $F^{*}=P^{*}$, and $\tau$ being the stopping time of a CUSUM/SR change detector with the likelihood ratio of $P_{1}^{*}$ to $P_{0}^{*}$.
+LFD pairs are known to exist for various parametric and nonparametric models, e.g., the monotone likelihood ratio (MLR) families, and Huber’s robust models (including $\epsilon$ neighborhood model, total variation model, and more generally with 2-alternating capacities [huber1973minimax]).
 Suppose we have estimators $r_{t}^{*}$ of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$, for each $t$ (which can be constructed using simulations from $F^{*}_{0}$ as described in Section 3.1).
 Now, we analogously define the confidence set as
-$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}^{*}}\right\}.$ (3.9)
+$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}^{*}}\right\}.$ (14)
 The next theorem proves the validity of the above confidence set.
-Theorem 3.8.
-Suppose that the pre-change class $\mathcal{P}_{0}$ and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfy Assumption 3.5 and $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,\forall F_{0}\in\mathcal{P}_{0}$.
-Then, for any $\alpha\in(0,1),$ $T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, the confidence set $\mathcal{C}$ defined in (3.9), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
-We now analyze the size of the confidence set, generalizing Theorem 3.3 to the case where an LFD pair $(F_{0}^{*},F_{1}^{*})$ exists for $\mathcal{P}_{0}$ vs $\mathcal{P}_{1}$.
-We consider the confidence set in (3.9) with $r_{t}^{*}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}^{*}\geq t)}{N+1}$, $M_{t}$ as defined in (3.4) with $\hat{T}$ as defined in (3.3) using $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})},$ and $S_{n}^{(t)}=\prod_{i=n}^{t-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}$.
+Theorem 4.
+Suppose that the pre-change class $\mathcal{P}_{0}$ and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfy Assumption 1 and $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,\forall F_{0}\in\mathcal{P}_{0}$.
+Then, for any $\alpha\in(0,1),$ $T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, the confidence set $\mathcal{C}$ defined in (14), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
+We now analyze the size of the confidence set, generalizing Theorem 3 to the case where an LFD pair $(F_{0}^{*},F_{1}^{*})$ exists for $\mathcal{P}_{0}$ vs $\mathcal{P}_{1}$.
+We consider the confidence set in (14) with $r_{t}^{*}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}^{*}\geq t)}{N+1}$, $M_{t}$ as defined in (9) using $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})},$ and $S_{n}^{(t)}=\prod_{i=n}^{t-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}$.
 We redefine $p_{t},\rho_{0}$ and $\rho_{1}$ in terms of the LFD pair: $p_{t}=\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t)$, $\rho_{0}(s)=\mathbb{E}_{F_{0}^{*}}\left(\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\right)^{s}$ and $s_{0}=\arg\min_{s\in(0,1)}\rho_{0}(s)$ and $\rho_{1}(s)=\mathbb{E}_{F_{1}}\left(\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\right)^{s}$ and $s_{1}=\arg\min_{s\in(0,1)}\rho_{1}(s)$.
 Finally, we call the stopping time $\tau$ well-behaved if it satisfies
-$\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ $\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t),\forall t\in\mathbb{N},\forall F_{0}\in\mathcal{P}_{0};$ (3.10a)
-$\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},\infty}[\tau\leq t],\forall t,T\in\mathbb{N}\penalty 10000\ \text{ and }$ (3.10b)
-$\mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t],\forall t,T\in\mathbb{N},\text{ and }\forall(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ (3.10c)
-Note that (3.10b) is analogous to (3.6), while (3.10c) requires that the stopping time $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically larger than $\tau$ under change from $F_{0}$ to $F_{1}$, for any $(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ To justify that (3.10c) is sensible, we note that the LFD pair $(F_{0}^{*},F_{1}^{*})$ is the closest pair within $\mathcal{P}_{0}\times\mathcal{P}_{1}$ (according to various metrics or divergences such as Kullback-Leibler), i.e., it renders the change detection problem most difficult.
+$\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ $\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t),\forall t\in\mathbb{N},\forall F_{0}\in\mathcal{P}_{0};$ (15a)
+$\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},\infty}[\tau\leq t],\forall t,T\in\mathbb{N}\penalty 10000\ \text{ and }$ (15b)
+$\mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t],\forall t,T\in\mathbb{N},\text{ and }\forall(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ (15c)
+Note that (15b) is analogous to (11), while (15c) requires that the stopping time $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically larger than $\tau$ under change from $F_{0}$ to $F_{1}$, for any $(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ To justify that (15c) is sensible, we note that the LFD pair $(F_{0}^{*},F_{1}^{*})$ is the closest pair within $\mathcal{P}_{0}\times\mathcal{P}_{1}$ (according to various metrics or divergences such as Kullback-Leibler), i.e., it renders the change detection problem most difficult.
 Therefore, it is reasonable to assume that $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically largest.
-(3.10a) can be justified similarly because $F_{0}^{*}$ is the “closest” element of the pre-change class to the post-change class.
-Theorem 3.9.
-In the setting described above, for any change detection method with stop time $\tau$ that is well-behaved in the sense of (3.10a) and $p_{T}\neq 0$,
-$\limsup_{N\to\infty}\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0}\\ F_{1}\in\mathcal{P}_{1}\end{subarray}}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\frac{1-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}},$ (3.11)
+(15a) can be justified similarly because $F_{0}^{*}$ is the “closest” element of the pre-change class to the post-change class.
+Theorem 5.
+In the setting described above, for any change detection method with stop time $\tau$ that is well-behaved in the sense of (15a) and $p_{T}\neq 0$,
+$\limsup_{N\to\infty}\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0}\\ F_{1}\in\mathcal{P}_{1}\end{subarray}}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\frac{1-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}},$ (16)
 where $\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}:=\displaystyle\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}\end{subarray}}\Delta_{F_{0},T,F_{1}}$.
-Additionally, if the stopping time $\tau$ is well-behaved in the sense of (3.10b) and (3.10c) as well, the last term in (3.11) can be improved to $\min\left\{\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right),\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}\right\}.$
+Additionally, if the stopping time $\tau$ is well-behaved in the sense of (15b) and (15c) as well, the last term in (16) can be improved to
+$\min\left\{\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right),\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}\right\}.$
 Next, we explore how, in parametric settings, one can obtain adaptive thresholds via simulation.
@@
 We develop a simulation-based scheme to approximate the exact $\alpha$-th quantile of the distribution of the test statistic $M_{t}$.
-Here, $M_{t}$ need not be of the form (3.4), it can be any test statistic based on $X_{1},\cdots,X_{\tau}$.
+Here, $M_{t}$ need not be of the form (9), it can be any test statistic based on $X_{1},\cdots,X_{\tau}$.
 Assuming known pre- and post-change distributions, we draw $B$ many independent streams of data under $H_{0,t}$.
@@
 Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$:
-$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (4.1)
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (17)
 where $r_{t}$ is any unbiased estimator of $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; see Algorithm 1.
@@
 7:for $t=1$ to $\tau$ do
-8:  Compute test statistic $M_{t}$ (e.g. (3.2)) based on $\tau$ and $X_{1},\dots,X_{\tau}$
+8:  Compute test statistic $M_{t}$ (e.g. (8)) based on $\tau$ and $X_{1},\dots,X_{\tau}$
 9:  for $j=1$ to $B$ do
@@
 16:end for
-Theorem 4.1.
-For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (4.1), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
-Remark 4.2.
+Theorem 6.
+For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (17), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
+Remark 2.
 Notably, this method allows arbitrary (known) dependence among observations and is valid for any choice of $B,N,L$, test statistic $M_{t}$, and detection algorithm $\mathcal{A}$, which can function as an entirely black-box or heuristic method.
@@
 Notably, conventional confidence intervals are invalid here, since $\tau$ is random.
-Therefore, we rely on confidence sequences [4, 12].
+Therefore, we rely on confidence sequences [darling1967confidence, howard2021time].
 A confidence sequence having coverage 1c1-c is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform guarantee:
-$\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ (4.2)
+$\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ (18)
 We construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage, where $r_{t}$ is any unbiased estimator of $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$.
@@
 […]
-$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}.$ (4.3)
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}.$ (19)
 Algorithm 2 summarizes the method.
 Note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ can be approximated by discretizing $\mathcal{S}^{\prime}_{\tau-t+1}$ into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum.
-While this approach is conceptually straightforward, it is computationally intensive; related experiments are in Section F of the Supplementary.
+While this approach is conceptually straightforward, it is computationally intensive.
 Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient and does not require any approximation.
-Details are provided in Section C of the Supplementary, along with two concrete examples of the Gaussian and Poisson mean change problems.
+Details are provided in Section C of the Supplementary, along with a concrete example of the Gaussian mean change problems.
 Algorithm 2 CI for known pre-change ($F_{\theta_{0}}$) and parametric post-change
@@
 22:end for
-Theorem 4.3.
-Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (4.3), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{\theta_{0},\infty}[\tau\geq t]$, for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$.
+Theorem 7.
+Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (19), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{\theta_{0},\infty}[\tau\geq t]$, for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$.
 Then for any given detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$
@@
 Suppose that we have some confidence interval (CI) procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$ that constructs CI using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{0}$ with coverage 1c1-c and a confidence sequence (CS) procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, that constructs CS using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{1}$ having coverage 1c1-c.
-We construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})$, where $r_{t}^{*}$ is computed under Assumption 3.5.
+We construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})$, where $r_{t}^{*}$ is computed under Assumption 1.
 For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ having joint distribution as $\mathbb{P}_{\theta,t,\theta^{\prime}}$.
@@
 Then, we compute the same test statistics truncated at $L$, based on $\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$, which is denoted as $M^{j}_{t,L}(\theta,\theta^{\prime})$.
-Similar to (4.3), the adaptive confidence set is defined as
-$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha}r_{t}^{*};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (4.4)
+Similar to (19), the adaptive confidence set is defined as
+$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha}r_{t}^{*};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (20)
 where $\mathcal{\mathbf{S}}^{(t)}=\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
 Algorithm 3 contains an overview.
-If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 3.5 and provides coverage guarantees, as elaborated in Section G of the Supplementary.
+If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 1 and provides coverage guarantees, as elaborated in Section F of the Supplementary.
 As before, one can always adopt a grid-based approach, which is conceptually straightforward, but it is computationally intensive; corresponding experimental results are reported in Section F of the Supplementary.
 Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient and exact, but that can be a bit more conservative.
-We provide the details in Section D of the Supplementary, along with two concrete examples.
+We provide the details in Section D of the Supplementary, along with a concrete example.
 Algorithm 3 CI for parametric pre-change and post-change
@@
 23:end for
-Theorem 4.4.
-Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.4), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{\theta_{0}^{*},\infty}[\tau\geq t]$, for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$.
-For any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1}$, and any change detection algorithm $\mathcal{A}$ satisfying Assumption 3.5 and having $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, the following guarantee holds: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
-Remark 4.5.
+Theorem 8.
+Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (20), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{\theta_{0}^{*},\infty}[\tau\geq t]$, for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$.
+For any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1}$, and any change detection algorithm $\mathcal{A}$ satisfying Assumption 1 and having $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, the following guarantee holds: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
+Remark 3.
 As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets, $i=0,1$, we have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$.
-So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 4.3 and 4.4 reduces to Theorem 4.1.
+So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 7 and 8 reduces to Theorem 6.
 5 Confidence set for pre- and post-change parameters
@@
 But since both starting and ending times are random here, this strategy does not yield valid coverage guarantees.
-Instead, we fix each $t\in\mathcal{C}$ and construct a confidence sequence based on $X_{t},\cdots,X_{\tau}$ for $\theta_{1}$, denoted as $\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})$ (recall the definition in (4.2)), for some $\eta_{1}\in(0,1)$, whose value is less than the target non-coverage, $\alpha^{\prime}$.
+Instead, we fix each $t\in\mathcal{C}$ and construct a confidence sequence based on $X_{t},\cdots,X_{\tau}$ for $\theta_{1}$, denoted as $\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})$ (recall the definition in (18)), for some $\eta_{1}\in(0,1)$, whose value is less than the target non-coverage, $\alpha^{\prime}$.
 Finally, the union of these CS across $t\in\mathcal{C}$, yields the post-detection confidence interval for $\theta_{1}$:
-$\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*}).$ (5.1)
+$\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*}).$ (21)
 There is a natural tension between how quickly $\mathcal{A}$ detects a change from $\theta_{0}$ and the tightness of inference on $\theta_{1}$ — if $\mathcal{A}$ stops very quickly, we will have very few data points with which to estimate $\theta_{1}$, leading to wider intervals.
@@
 Analogously, for the pre-change parameter $\theta_{0}$, we fix each $t\in\mathcal{C}$ and construct a confidence interval for $\theta_{0}$ based on the data $X_{1},\ldots,X_{t-1}$}, denoted as $\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})$, for some $\eta_{0}\in(0,1)$, chosen to be smaller than the target (conditional) non-coverage level, $\alpha^{\prime\prime}$.
-The final confidence set for $\theta_{0}$ is $\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$ Next, we formalize its coverage guarantee, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods.
-Theorem 5.1.
-Let $\mathcal{C}$ be a confidence set, e.g., (3.9) or (4.4), such that for any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},T\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ satisfying Assumption 3.5 and $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, we have: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha$.
+The final confidence set for $\theta_{0}$ is $\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$
+Next, we formalize its coverage guarantee, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods.
+Theorem 9.
+Let $\mathcal{C}$ be a confidence set, e.g., (14) or (20), such that for any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},T\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ satisfying Assumption 1 and $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, we have: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha$.
 Then, for any $T\in\mathbb{N}$,
@@
 6 Experiments
+We validate our methods across various simulation and real-data settings, demonstrating their practicality.
 Known pre- and post-change (Setting I): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$, with $T$ being the only unknown parameter.
 We employ the CUSUM detector, $\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ with $A=1000$.
-After detection, we construct both the proposed confidence sets (universal (3.5) and adaptive (4.1)), with $B=N=100$, $\alpha=0.05$.
+After detection, we construct both the proposed confidence sets (universal (10) and adaptive (17)), with $B=100,N=50$, $\alpha=0.1$.
 Fig.
-6.1 visualizes the confidence sets (4.1) and point estimates (3.1) across $5$ runs.
-Table 6.2 has the (conditional) average size and coverages of the confidence sets, the (conditional) average absolute deviation of $\hat{T}$, and delay of detection, i.e, $\tau-T$, across $500$ runs.
-Figure 6.1: Setting I: The first T1T-1 observations are drawn from $N(0,1)$ and the rest from $N(1,1)$.
-The point estimates (3.1) are shown in a vertical red dashed line, and confidence sets (adaptive (4.1)) are shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$.
-Results of $5$ independent simulations are shown.
-Table 6.2: Setting I: Known pre- and post-change.
+1 visualizes the confidence sets (17) across $5$ runs.
+Table 2 has the (conditional on the event $\tau\geq T$) average size and coverages of the confidence sets and delay of detection, i.e, $\tau-T$, across $500$ runs.
+Figure 1: Setting I.
+The first T1T-1 observations are drawn from $N(0,1)$ and the rest from $N(1,1)$.
+The confidence sets (adaptive (17)) are shown in red points, with $B=N=100$, $\alpha=0.1,L=\infty$.
+Results of 55 independent simulations are shown.
+Table 2: Setting I: Known pre- and post-change.
 $T$ A Conditional Coverage Conditional Size
-$|\hat{T}-T|$
 Delay
-Universal Adaptive Universal Adaptive (Conditional) (Conditional)
-100 1000 0.98 0.95 15.63 12.34 2.85 13.97
-500 1000 0.98 0.95 15.77 12.57 2.62 13.22
+Universal Adaptive Universal Adaptive (Conditional)
+100 1000 0.948 0.906 12.412 9.694 12.803
+500 1000 0.942 0.908 12.176 9.278 13.227
 Known pre- and parametric post-change (Setting II): We perform experiments with $F_{0}=N(0,1)$ (known pre-change), post-change class $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$, ; we vary $\Theta_{1}$ and $T$.
 The post-change samples are drawn from $N(1,1)$ (unknown).
-For detection, we use a weighted CUSUM detector:
-$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (6.1)
-with $A=1000$, a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized mixture over $\mathcal{P}_{1}$.
-We use
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})},$ (6.2)
-where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$.
-After detection, we construct both the proposed confidence sets for $T$ (universal (3.5) with $\alpha=0.075$ and adaptive (4.3) [^5] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section C.1 (and Section D.1) of the Supplementary.
-with $\alpha=0.05,\beta=0.025$) with target conditional coverage $0.925$.
-For implementing (3.5) , we use $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})},$ where $\theta_{1}^{*}$} is the closest element of $\Theta_{0}$ from $\Theta_{1}$, and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w$ being the same weight mentioned above.
-In Table 6.3, we report the (conditional) average size and coverages, absolute deviation of $\hat{T}$ , and detection delay $\tau-T$ across $500$ runs.
+For detection, we use the CUSUM detector with $A=1000$, a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized mixture over $\mathcal{P}_{1}$.
+After detection, we construct both the proposed confidence sets for $T$ (universal (10) with $\alpha=0.075$ and adaptive (19) [^4] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section C.1 (and Section D.1) of the Supplementary.
+with $\alpha=0.1,\beta=0.025$) with target conditional coverage $0.925$.
+For implementing (10) , we use $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})},$ where $\theta_{1}^{*}$} is the closest element of $\Theta_{0}$ from $\Theta_{1}$, and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w$ being the same weight mentioned above.
+In Table 3, we report the results averaged across $500$ runs.
 Fig.
-6.4(a) shows the confidence sets (3.5) and point estimates (S2) across $5$ random runs.
-Table 6.3: Known pre-change but unknown post-change.
+2(a) shows the confidence sets (10) across $5$ random runs.
+Table 3: Known pre-change but unknown post-change.
 The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column.
@@
 $T$ $\Theta_{1}$ Conditional Coverage Conditional Size
-$|\hat{T}-T|$
 Delay
-Universal Adaptive Universal Adaptive (Conditional) (Conditional)
-100 $0.75,\infty)$ 0.98 0.97 22.21 25.72 3.95 16.87
-100 $0.9,\infty)$ 0.97 0.97 17.85 24.11 3.67 16.21
-500 $0.75,\infty)$ 0.98 0.97 22.89 30.83 3.45 17.63
-500 $0.9,\infty)$ 0.97 0.98 18.14 28.56 3.48 15.81
+Universal Adaptive Universal Adaptive (Conditional)
+100 $0.75,\infty)$ 0.972 0.966 16.870 14.824 13.878
+100 $0.9,\infty)$ 0.976 0.972 12.860 11.552 12.212
+500 $0.75,\infty)$ 0.969 0.968 16.168 14.266 13.638
+500 $0.9,\infty)$ 0.972 0.971 14.734 12.602 12.814
 Composite parametric pre- and post-change (Setting III): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$.
@@
 For detection, we employ a weighted CUSUM-type detector
-$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta^{*}_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (6.3)
+$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta^{*}_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (22)
 with $A=1000$, $\theta^{*}_{0}$} is the “closest” element of $\Theta_{0}$ from $\Theta_{1}$, and for the post-change parameter, the same discrete weight distribution as in setting II is considered.
 We define $\hat{\theta}_{0,1:j}$ to be the MLE of $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$.
-We use
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})},$ (6.4)
-where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$ and $\hat{\theta}_{0,:j-1}$ is the MLE of $\theta_{0}\in\Theta_{0}$ based on $X_{1},\cdots,X_{j-1}$.
-After detection, we construct both the proposed confidence sets (universal (3.9) with $\alpha=0.1$ and adaptive (4.4) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage $0.9$.
-For implementing (3.9), we use $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})}w^{\prime}(\theta)d\theta,$ with $w^{\prime}(\theta)$ being the discrete distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta-d,\theta_{3}=\theta-2d,\cdots,\theta_{10}=\theta-9d$, $d=0.2$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}w(\theta)d\theta$, with $w(\theta)$ being the same as mentioned in Setting II.
+After detection, we construct both the proposed confidence sets (universal (14) with $\alpha=0.1$ and adaptive (20) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage $0.9$.
+For implementing (14), we use $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})}w^{\prime}(\theta)d\theta,$ with $w^{\prime}(\theta)$ being the discrete distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta-d,\theta_{3}=\theta-2d,\cdots,\theta_{10}=\theta-9d$, $d=0.2$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}w(\theta)d\theta$, with $w(\theta)$ being the same as mentioned in Setting II.
 $\theta_{i}^{*}$ is the closest element of $\Theta_{i}$ from the other space.
 Fig.
-6.4(b) visualizes the confidence sets (3.9) and point estimates (S2) across 5 random runs.
-Table 6.5 reports the results averaged across $500$ runs.
-Figure 6.4: First T1T-1 samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1),T=100$.
-The point estimates (S2) are shown in a vertical red dashed line.
-Confidence sets (universal (3.9)) are shown in red points.
+2(b) visualizes the confidence sets (14) across 5 random runs.
+Table 4 reports the results averaged across $500$ runs.
+Figure 2: First T1T-1 samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1),T=100$.
+Confidence sets (universal (14)) are shown in red points.
 $N=100$, $\alpha=0.1$.
-Results of $5$ random simulations are shown.
-Table 6.5: Confidence interval for $T$ (Setting III): unknown pre- and post-change.
+Results of 55 random simulations are shown.
+Table 4: Confidence interval for $T$ (Setting III): unknown pre- and post-change.
 The true (unknown) values of the parameters are $\theta_{0}=0$, $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns.
 $T$ $\Theta_{0}$ $\Theta_{1}$ Conditional Coverage Conditional Size
-$|\hat{T}-T|$
 Delay
-Universal Adaptive Universal Adaptive (Cond.) (Cond.)
-100 $(\infty,0.25]$ $0.75,\infty)$ 0.98 0.97 26.91 26.84 4.36 25.81
-100 $(\infty,0.1]$ $0.9,\infty)$ 0.97 0.97 18.63 25.71 4.19 23.15
-500 $(\infty,0.25]$ $0.75,\infty)$ 0.98 0.97 26.12 31.45 4.03 24.48
-500 $(\infty,0.1]$ $0.9,\infty)$ 0.97 0.98 18.86 30.66 4.07 23.07
-Known pre- and nonparametric post-change (Setting IV): Here, the pre-change samples are from known $F_{0}=\text{U}[0,1]$ (uniform distribution on $[0,1]$) and let the post-change class be $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[X]\leq\mu\right\},$ where $\mathcal{M}$ is the set of all probability measures on $([0,1],\mathcal{B}([0,1]))$, with $\mu=0.25$.
-For detection, we use an e-detector [28] $\tau_{e}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}\geq A\right\},$ where $\hat{p}_{i-1}^{\prime}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=t}^{i-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},$ for $x\in[0,1]$, with $A=1000$.
-Post-detection, we construct the confidence sets (3.5).
-The forward $t$-delay e-process $\{R^{(t)}_{n}\}_{n\geq t}$, we use numeraire e-variable for $\mathcal{P}_{1}$ against $F_{0}$, as in Section 5.2 of [17] to get $R^{(t)}_{n}=\prod_{i=t}^{n}(1+\lambda^{*}(X_{i}-\mu)),$ where $\lambda^{*}\in(0,1/\mu)$ is the solution of the equation $e^{\lambda^{*}}=\frac{1+\lambda^{*}(1-\mu)}{1-\lambda^{*}\mu}.$ For the backward e-processes $\{S^{(t)}_{n}\}_{n<t}$, we adopt the predictable plug-in method [20].
-We divide the interval $[0,1]$ into $10$ equal-sized bins $B_{i}=[\frac{i-1}{10},\frac{i}{10}),i=1,\cdots,10$ and perform histogram-based estimation.
-We construct $S^{(t)}_{n}=\prod_{i=n}^{t-1}\frac{\hat{p}_{i-1}(X_{i})}{{p}_{0}(X_{i})},$ where ${p}_{0}(x)=\frac{1}{10},$ and $\hat{p}_{i-1}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=i-1}^{t-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},x\in[0,1]$.
-We use the point estimator
-$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}.$ (6.5)
-Table 6.6 shows average results across $500$ runs and Fig. 6.7(a) visualizes across $5$ runs.
-Table 6.6: Setting IV: known pre-change ($\text{U}[0,1]$) but nonparametric post-change.
-T
-True (unknown) post-
-change density, $f_{1}(x)$
-Coverage
-Coverage
-Size
-$|\hat{T}-T|$
-Delay
-100 $4(1-x)^{3}$ 0.98 0.97 24.14 4.06 20.65
-100 $0.8\text{U}[0,0.2]+0.2\text{U}[0.2,1]$ 0.98 0.97 20.92 2.86 14.08
-500 $4(1-x)^{3}$ 0.98 0.94 24.63 4.11 19.57
-500 $0.8\text{U}[0,0.2]+0.2\text{U}[0.2,1]$ 0.99 0.93 21.27 2.94 15.78
-Composite parametric pre- and nonparametric post-change (Setting V): Here, we consider the true (unknown) pre- and post-change distributions to be $\text{N}(1,1)$ and $\text{U}[-1.2,0.8]$ respectively.
-Let the pre-change class be $\mathcal{P}_{0}=\left\{\text{N}(\theta,1):\theta\geq 0.5\right\},$ and the post-change class be the set of all SubGaussians with a nonpositive mean: $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[e^{\lambda X}]\leq e^{-\lambda^{2}/2},\forall\lambda\in[0,\infty)\right\},$ $\mathcal{M}$ is the set of all probability measures on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$.
-We use an e-detector [28],
-$\tau_{e}^{\prime}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}\geq A\right\},$
-with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}-0.5.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 3.5 with $P_{0}^{*}=\text{N}(0.5,1)$.
-After detection, we construct the confidence sets (3.9).
-The forward $t$-delay e-process we use is $\{R^{(t)}_{n}\}_{n\geq t}=\prod_{i=t}^{n}e^{\lambda_{i}X_{i}-\lambda_{i}^{2}/2}$, where $\lambda_{i}=\max\{0.5,\frac{1}{i-t}\sum_{j=t}^{i-1}X_{i}\}$ and the backward e-processes is constructed similarly: $\{S^{(t)}_{n}\}_{n<t}=\prod_{i=n}^{t}e^{\mu_{i}X_{i}-\mu_{i}^{2}/2}$, where $\mu_{i}=\min\{0,\frac{1}{t-i}\sum_{j=i+1}^{t}X_{j}\}-0.5$ (see, e.g., Ch.
-6.7.5 of [22]).
-We use the following point estimator:
-$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}^{\prime}}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}.$ (6.6)
-Figure 6.7: Vertical red dashed lines show the point estimates (6.6). The confidence set (3.9) is marked in red points. Results of $5$ independent simulations are shown.
-Table 6.8: Setting V: Parametric pre-change, nonparametric post-change.
+Universal Adaptive Universal Adaptive (Conditional)
+100 $(\infty,0.25]$ $0.75,\infty)$ 0.984 0.982 23.683 23.827 22.810
+100 $(\infty,0.1]$ $0.9,\infty)$ 0.968 0.978 15.552 15.294 15.282
+500 $(\infty,0.25]$ $0.75,\infty)$ 0.982 0.976 23.825 24.482 24.060
+500 $(\infty,0.1]$ $0.9,\infty)$ 0.970 0.942 16.860 16.956 15.576
+Composite parametric pre- and nonparametric post-change (Setting IV): Here, we consider the true (unknown) pre- and post-change distributions to be $\text{N}(1,1)$ and $\text{U}[-1.2,0.8]$ respectively.
+Let the pre-change class be $\mathcal{P}_{0}=\left\{\text{N}(\theta,1):\theta\geq\mu_{1}\right\},\mu_{1}=0.75$ and the post-change class be the set of all SubGaussians with a nonpositive mean: $\mathcal{P}_{1}=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[e^{\lambda X}]\leq e^{-\frac{\lambda^{2}}{2}},\forall\lambda\in[0,\infty)\right\}$, $\mathcal{M}$ is the set of all probability measures on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$.
+We use an e-detector [shin2022detectors],
+$\tau_{e}^{\prime}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}e^{(\nu_{i}-\mu_{1})X_{i}-\nu_{i}^{2}/2+\mu_{1}^{2}/2}\geq A\right\},$
+with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 1 with $P_{0}^{*}=\text{N}(\mu_{1},1)$.
+After detection, we construct the confidence sets (14).
+The forward $t$-delay e-process we use is $\{R^{(t)}_{n}\}_{n\geq t}=\prod_{i=t}^{n}e^{\lambda_{i}X_{i}-\lambda_{i}^{2}/2}$, where $\lambda_{i}=\max\{\mu_{1},\frac{1}{i-t}\sum_{j=t}^{i-1}X_{i}\}$ and the backward e-processes is constructed similarly: $\{S^{(t)}_{n}\}_{n<t}=\prod_{i=n}^{t}e^{(\eta_{i}-\mu_{1})X_{i}-\eta_{i}^{2}/2+\mu_{1}^{2}/2}$, where $\eta_{i}=\min\{0,\frac{1}{t-i}\sum_{j=i+1}^{t}X_{j}\}$ (see, e.g., Ch.
+6.7.5 of [ramdas2024hypothesis]).
+Table 5 contains the results across $500$ runs and Fig.
+3 visualizes across $5$ runs.
+Figure 3: The confidence set (14) is marked in red points. Results of 55 independent simulations are shown.
+Table 5: Setting IV: Parametric pre-change, nonparametric post-change.
 T
@@
 Size
-$|\hat{T}-T|$
 Delay
-100 0.98 0.97 27.65 3.08 27.34
-500 0.98 0.94 28.03 3.15 25.12
-Huber-robust (nonparametric) pre- and post-change (Setting VI):
+100 0.984 0.982 23.859 9.959
+500 0.992 0.930 25.381 9.606
+Huber-robust (nonparametric) pre- and post-change (Setting V):
 We perform experiments for the robust Gaussian mean-change problem under the $\epsilon$ contamination model, with the true changepoint at $T=100,500$.
@@
 The true (unknown) pre- and post-change distributions are $(1-\epsilon)\times N(0,1)+\epsilon\times\text{Cauchy}(-1,10)$ and $(1-\epsilon)\times N(1,1)+\epsilon\times\text{Cauchy}(-1,10)$ respectively.
-We use Huber’s LFD (which are also log-optimal e-values [24]) to construct a CUSUM-style change-detector as below:
-$\tau=\inf\left\{t\in\mathbb{N}:\max_{1\leq j\leq t}\prod_{i=j}^{t}\pi(X_{i})\geq A\right\},$ (6.7)
-where $\pi(x)=\min\{\max\{\frac{f_{1}(x)}{f_{0}(x)},c^{\prime}\},c^{\prime\prime}\}$ for constants $c^{\prime\prime}>c^{\prime}\geq 0$ as mentioned in [13, 24].
+We use Huber’s LFD (which are also log-optimal e-values [saha2024huber]) to construct a CUSUM-style change-detector as below:
+$\tau=\inf\left\{t\in\mathbb{N}:\max_{1\leq j\leq t}\prod_{i=j}^{t}\pi(X_{i})\geq A\right\},$ (23)
+where $\pi(x)=\min\{\max\{\frac{f_{1}(x)}{f_{0}(x)},c^{\prime}\},c^{\prime\prime}\}$ for constants $c^{\prime\prime}>c^{\prime}\geq 0$ as mentioned in [huber1965robust, saha2024huber].
 We set the threshold at $A=1000$.
-$\hat{T}$ is as defined in (3.3) with the same e-process, $\prod_{i=1}^{t}\pi(X_{i})$, as mentioned above.
-After detection, we construct our universal confidence sets (3.9) for $T$, with $R_{n}^{(t)}=1/\prod_{i=t}^{n}\pi(X_{i})$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}\pi(X_{i})$, for $n<t$ and $\alpha=0.05$.
-We report the average size and coverage rates across $200$ independent runs in Table 6.9.
-We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection).
-Table 6.9: Setting VI: Huber-robust (nonparametric) pre- and post-change
+After detection, we construct our universal confidence sets (14) for $T$, with $R_{n}^{(t)}=1/\prod_{i=t}^{n}\pi(X_{i})$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}\pi(X_{i})$, for $n<t$ and $\alpha=0.1$.
+We report the results averaged across $500$ independent runs in Table 6, and Fig.
+3 visualizes the results across $5$ runs.
+Table 6: Setting V: Huber-robust (nonparametric) pre- and post-change
 T $\epsilon$
-Conditional
-coverage
-Marginal
-coverage
-Size
-$|\hat{T}-T|$
-Delay
-100 0.01 0.98 0.98 16.74 3.23 18.38
-100 0.001 0.96 0.96 15.57 3.18 13.96
-500 0.01 0.97 0.93 16.18 3.87 16.58
-500 0.001 0.97 0.92 15.13 3.25 14.44
+Conditional coverage
+Conditional size
+Conditional delay
+100 0.01 0.939 13.298 14.140
+500 0.01 0.948 13.634 13.943
+100 0.001 0.944 12.274 12.638
+500 0.001 0.958 11.927 12.496
 […]
-Table 6.10: Comparison of the confidence sets for $T$ with [39] (Setting I)
-$F_{0}$ $F_{1}$ $T$ Conditional Size Conditional Coverage Marginal Coverage
+Table 7: Comparison of the confidence sets for $T$ with [wu2007inference] (Setting I)
+$F_{0}$ $F_{1}$ $T$ Conditional Coverage Conditional Size Marginal Coverage
 Ours Wu (2007) Ours Wu (2007) Ours Wu (2007)
-N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 100 46.39 47.83 0.96 0.86 0.88 0.79
-N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 100 33.74 34.14 0.96 0.83 0.82 0.76
-N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 500 46.58 47.21 0.96 0.86 0.57 0.53
-N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 500 35.19 34.81 0.97 0.82 0.56 0.49
-The results show that [39] (which is only asymptotically valid and relies on approximations) exhibits noticeable undercoverage, while the sizes for both methods are comparable.
-However, our method (4.1) achieves the target conditional coverage.
-Sentiment change in language data: We consider the Stanford Sentiment Treebank (SST-2) dataset of movie reviews labeled with a binary sentiment [31].
+N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 100 0.916 0.838 34.595 39.706 0.864 0.792
+N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 100 0.896 0.774 23.827 26.678 0.828 0.736
+N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 500 0.902 0.841 34.382 38.195 0.546 0.578
+N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 500 0.906 0.796 23.188 27.343 0.512 0.442
+The results show that [wu2007inference] (which is only asymptotically valid and relies on approximations) exhibits noticeable undercoverage, while the sizes for both methods are comparable.
+However, our method (17) achieves the target conditional coverage.
+Sentiment change in language data: We consider the Stanford Sentiment Treebank (SST-2) dataset of movie reviews labeled with a binary sentiment [socher2013recursive].
 We want to localize the change from $60\%$ positive reviews to $40\%$ positive reviews.
@@
 draws of $25\%$ negative reviews (these are all unknown to our algorithm).
-We construct a change detection method and a point estimate $\hat{T}$, using a pre-trained DistilBERT-base model fine-tuned on the uncased SST-2 [25], which is a binary sentiment classifier that maps each review to either 0 (negative) or $1$ (positive), thereby producing a sequence of Bernoulli observations.
-After obtaining the binary sequence and $\hat{T}$, we construct the test statistic (3.4) using the likelihood ratio process for $\text{Bern}(0.4)$ against $\text{Bern}(0.6)$ (forward) and for $\text{Bern}(0.6)$ against $\text{Bern}(0.4)$ (backward).
-Note that given the change detector (and as a byproduct, the sequence of Bernoulli) and $\hat{T}$, it is simple to compute (3.4) in $O(T)$ time per $t$.
-So, the confidence set (3.9) takes total $O(T^{2})$ time.
+We construct a change detection method using a pre-trained DistilBERT-base model fine-tuned on the uncased SST-2 [sanh2019distilbert], which is a binary sentiment classifier that maps each review to either 0 (negative) or $1$ (positive), thereby producing a sequence of Bernoulli observations.
+After obtaining the binary sequence, we construct the test statistic (9) using the likelihood ratio process for $\text{Bern}(0.4)$ against $\text{Bern}(0.6)$ (forward) and for $\text{Bern}(0.6)$ against $\text{Bern}(0.4)$ (backward).
+Note that given the change detector (and as a byproduct, the sequence of Bernoulli), it is simple to compute (9) in $O(T)$ time per $t$.
+So, the confidence set (14) takes total $O(T^{2})$ time.
 Fig.
-6.11 provides a visualization.
-Figure 6.11: Vertical red dashed lines show the point estimate.
-The confidence set (3.9) is marked in red points.
-The true changepoint $T=500$ is marked with the black line.
-Texts are plotted from time $t=470$ upto detection time, $\tau=522$.
-Remark 6.1.
+4 provides a visualization.
+Figure 4: The confidence set (14) is marked in red points. The true changepoint $T=500$ is marked with the black line. Texts are plotted from time $t=470$ upto detection time, $\tau=534$.
+Remark 4.
 In Settings I, II, and III, both of our proposed methods — universal (Section 3) and adaptive (Section 4) — are applicable.
-For nonparametric settings (IV, V, and VI), the only valid method currently known is the universal method.
+For nonparametric settings (IV and V), the only valid method currently known is the universal method.
 Prior works are valid (asymptotically) only for Setting I.
-As shown in Table 6.2 for known pre- and post-change distributions (Setting I), the adaptive method consistently outperforms the universal method, which tends to be conservative.
-Also, our adaptive method is the only method we know that remains valid under dependence (see experimental results in Supplement H.1).
-However, for the composite settings (II and III), the universal method outperforms the adaptive method.
+As shown in Table 2 for known pre-distributions (Settings I and II), the adaptive method consistently outperforms the universal method, which tends to be conservative.
+Also, our adaptive method is the only method we know that remains valid under dependence (see experimental results in Supplementary Section G.2).
+However, for setting III, the universal and adaptive methods perform similarly.
 Moreover, the universal method is also significantly more efficient — it runs in $O(T^{2})$ time (assuming $M_{t}$ can be computed in $O(T)$ time), while the adaptive method incurs a higher computational cost of $O(T^{3}B)$ (assuming $\mathcal{A}$ runs in $O(T)$ time and $M_{t}$ can be computed in $O(T)$ time).
-See Supplement Sections E, F, and H for further experiments, involving confidence intervals for pre- and post-change parameters, detection procedures with bounded PFA, settings with dependent data, etc.
+See Supplement Sections E, F, and G for further experiments, involving confidence intervals for pre- and post-change parameters, settings with dependent data, settings with detection algorithms controlling PFA, etc.
 7 Conclusion
@@
 By bridging a longstanding gap in the literature, our work offers the first nonparametric, theoretically principled, and broadly applicable method for post-detection inference.
+A particularly fruitful direction for future work is to extend this framework to the distribution-free setting.
+Acknowledgement
+Part of this work was carried out while AS was a student at the Indian Statistical Institute.
+The authors thank the Editor, the Associate Editor, and the reviewers for their careful reading of an earlier version of the article and for their helpful comments.
+The authors also thank Swapnaneel Bhattacharyya for identifying and helping eliminate an inconsistency in an earlier version of the article.
+AR acknowledges support from NSF DMS-1916320.
 Supplementary Material
-Outline of Supplementary Material
+Outline of the Supplementary Material
 Omitted proofs can be found in Appendix A.
@@
 Experimental results for the confidence intervals for pre- and post-change parameters are in Appendix E.
-Experimental results for a different implementation method are in Appendix F.
-Simpler algorithms for PFA-controlling change detectors are presented in Appendix G.
-Some additional experiments with dependent (Markov) pre- and post-change data, with detection procedures with bounded PFA, and with Poisson rate change problems are in Appendix H.
-Some heuristic algorithms are presented in Appendices I and J, pointing at open avenues for further theoretical work.
+Simpler algorithms for PFA-controlling change detectors are presented in Appendix F, along with their implementations in simulation experiments.
+Additional experiments with varying $T$ vs.
+ARL and scenarios with dependent (Markov) pre- and post-change data are presented in Appendix G.
 Appendix A Omitted proofs
 In this section, we provide the mathematical details that were omitted in the main paper.
-Proof of Proposition 2.1..
-For the sake of contradiction, assume that $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$ for some $\mathcal{C}^{*}\subseteq\{1,\cdots,\tau\}$. Now we observe that
+Proof of Proposition 1..
+Assume that $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$ for some $\mathcal{C}^{*}\subseteq\{1,\cdots,\tau\}$. Now we observe that
 $\mathbb{P}_{\infty}(\tau\geq T)=\mathbb{P}_{T}(\tau\geq T)\geq\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha.$
-Threfore, $\mathbb{P}_{\infty}(\tau=\infty)=1-\mathbb{P}_{\infty}(\cup_{n=1}^{\infty}[\tau<n])=1-\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau<n)=\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau\geq n)\geq 1-\alpha.$
-Proof of Proposition 2.2.
+Therefore, $\mathbb{P}_{\infty}(\tau=\infty)=1-\mathbb{P}_{\infty}(\cup_{n=1}^{\infty}[\tau<n])=1-\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau<n)=\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau\geq n)\geq 1-\alpha.$
+Proof of Proposition 2..
 First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
 Let $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying $\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\geq 1-\alpha$, for all $t\in\mathbb{N}$.
-Since $\tau$ be a stopping time, $\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$.
-This implies $\mathbb{P}_{F_{0},\infty}(\tau\geq t)=\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ Therefore,
-$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)$ $=\mathbb{P}_{F_{0},t,F_{1}}(t\notin\mathcal{C})$
-$=1-\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C})$
-$=1-\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t)$
-$\leq 1-(1-\alpha)\mathbb{P}_{F_{0},\infty}(\tau\geq t)$
+Since $\tau$ is a stopping time, $\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$.
+This implies $\mathbb{P}_{F_{0},\infty}(\tau\geq t)=\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t).$ Therefore,
+$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)$ $=\mathbb{P}_{F_{0},t,F_{1}}(t\notin\mathcal{C},t\leq\tau)$
+$=\mathbb{P}_{F_{0},t,F_{1}}(t\notin\mathcal{C}\mid\tau\geq t)\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t)$
 $\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t),\penalty 10000\ \forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
-For the converse part, $\phi^{(t)}$ satisfies $\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$
+For the converse part, $\phi^{(t)}$ satisfies $\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$.
 So, $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
@@
 $\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\})+\mathbb{P}_{F_{0},\infty}(\tau<T)$
-$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})+\mathbb{P}_{F_{0},\infty}(\tau<T)$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})+\mathbb{P}_{F_{0},\infty}(\tau<T).$
 Therefore, $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq T)-\mathbb{P}_{F_{0},\infty}(\tau<T)$, which implies
 $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C},\tau\geq T)}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
-Lemma A.1 (Ville’s inequality [35]).
+This completes the proof. ∎
+Lemma 1 (Ville’s inequality [ville1939etude]).
 If $\{R_{n}\}_{n\in\mathbb{N}}$ is an e-process for $\mathcal{P}$ with respect to a filtration $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, then
 $\mathbb{P}\left(\exists n\in\mathbb{N}:R_{n}\geq 1/\alpha\right)\leq\alpha,\forall\mathbb{P}\in\mathcal{P}.$ (S1)
-For a short proof, see [11][Lemma 1].
-Proof of Theorem 3.1.
+For a short proof, see [howard2020time][Lemma 1].
+Proof of Theorem 1..
 Note that by definition of $M_{t}$, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t=\hat{T}\leq\tau<\infty]=0$, $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t>\tau]=0$ and $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}>2/\alpha,\tau=\infty]=0$.
@@
 $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha]$
-$=\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t<\hat{T}\leq\tau<\infty]+\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,\hat{T}<t\leq\tau<\infty]$
-$=\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{\hat{T}-1}\geq 2/\alpha,t<\hat{T}\leq\tau<\infty\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{\hat{T}}\geq 2/\alpha,\hat{T}<t\leq\tau<\infty\right]$
-$\leq\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{(t-1)\wedge\hat{T}}\geq 2/\alpha\right]$
-Now, from Ville’s inequality (Lemma A.1) on the e-processes $\{R^{(t)}_{n}\}_{n\geq t}$,
+$=\mathbb{P}_{F_{0},t,F_{1}}\left[\max\{\max_{t\leq n\leq\tau}R^{(t)}_{n},\max_{1\leq n\leq t-1}S^{(t)}_{n}\}\geq 2/\alpha,t\leq\tau<\infty\right]$
+$\leq\mathbb{P}_{F_{0},t,F_{1}}\left[\max_{t\leq n\leq\tau}R^{(t)}_{n}\geq 2/\alpha,t\leq\tau<\infty\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[\max_{1\leq n\leq t-1}S^{(t)}_{n}\geq 2/\alpha,t\leq\tau<\infty\right].$
+Now, from Ville’s inequality (Lemma 1) on the e-processes $\{R^{(t)}_{n}\}_{n\geq t}$,
 $P_{F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
 Therefore,
-$\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha\right]$ $=\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha,t\vee(\hat{T}-1)=n\right]\right)$
-$\leq\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{n}\geq 2/\alpha\right]\right)$
+$\mathbb{P}_{F_{0},t,F_{1}}\left[\max_{t\leq n\leq\tau}R^{(t)}_{n}\geq 2/\alpha,t\leq\tau<\infty\right]$ $\leq\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{n}\geq 2/\alpha\right]\right)$
 $=\mathbb{P}_{F_{0},t,F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
@@
 $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha]\leq\alpha/2+\alpha/2=\alpha.$
-Proof of Theorem 3.2.
+This completes the proof. ∎
+Proof of Theorem 2..
 Observe that
@@
 $\leq\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:M_{t}<2/(\alpha r_{t})\},\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
-$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}<2/(\alpha r_{t})\}\mid\tau\geq T)\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
+$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
 $=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T),$
 where we twice used the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$, which holds because $\tau$ is a stopping time and so the event $\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
-From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other.
+From Theorem 1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other.
 Now, since $r_{t}$ is an unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}]=\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\forall t\in\mathbb{N}.$ So, we obtain
 $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1$- $\frac{\mathbb{P}_{F_{0},T,F_{1}}(M_{T}>2/(\alpha r_{T}))}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
-Proof of Theorem 3.3.
+This completes the proof. ∎
+Proof of Theorem 3..
 We first decompose the length of the confidence set into contributions from the pre- and post-change regions, and then control each contribution separately.
 Decomposition of the confidence set.
-Since $M_{t}=\frac{L_{\hat{T}}^{\tau}}{L_{t}^{\tau}}$ and $\hat{T}=\arg\max_{1\leq i\leq\tau}L_{i}^{\tau}$ and hence, whenever $\tau\geq T>t,$
-$M_{t}\geq\frac{L_{{T}}^{\tau}}{L_{t}^{\tau}}=\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}.$
-Similarly, whenever $T<t\leq\tau,$ $M_{t}=\frac{L_{\hat{T}}^{\tau}}{L_{t}^{\tau}}\geq\frac{L_{{T}}^{\tau}}{L_{t}^{\tau}}=\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$. Therefore,
+Whenever $\tau\geq T>t,$
+$M_{t}\geq\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}.$
+Similarly, whenever $T<t\leq\tau,$
+$M_{t}\geq\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$
+Therefore, the following holds almost surely:
 $|{\mathcal{C}}|\times\mathds{1}(\tau\geq T)$
@@
 $=\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+\sum_{t=T+1}^{{\infty}}\mathds{1}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),t\leq\tau\right).$
-We also have
-$|{\mathcal{C}}|\leq\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+|\tau-T|.$ (S2)
+Also, the following holds almost surely:
+$|{\mathcal{C}}|\times\mathds{1}(\tau\geq T)\leq\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+|\tau-T|.$ (S2)
 So, taking the conditional expectation on both sides and combining both bounds, we get
@@
 Fix some $s\in(0,1)$. Then, using Markov’s equality,
-$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$ $\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{T})\right)\quad\text{[Since }r_{t}\leq r_{T}]$
+$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$
+$\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{T})\right)\quad\text{[Since }r_{t}\leq r_{T}]$
 $=\mathbb{E}\left(\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}\geq(\alpha r_{T})^{s}/2^{s}\mid r_{T}\right)\right)$
@@
 This holds for all $s\in(0,1)$ and hence for $s=s_{0}.$ As noted earlier, $\rho_{0}(s_{0})<1$. So, we have
-$\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$ $\leq\frac{1}{p_{T}}\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$
+$\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$
+$\leq\frac{1}{p_{T}}\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$
 $\leq\mathbb{E}\left(\frac{2}{\alpha r_{T}}\right)^{s}\frac{1}{p_{T}}\sum_{t=1}^{{T-1}}\rho_{0}(s)^{T-t}$
 $\to\frac{2^{s}}{\alpha^{s}p_{{T}}^{s+1}}\times\frac{\rho_{0}(s)-\rho_{0}(s)^{T-1}}{1-\rho_{0}(s)},$
-as $N\to\infty.$ To justify the convergence, we note that $\left(\frac{2}{\alpha r_{T}}\right)^{s}$ is uniformly integrable as a sequence in $N\in\mathbb{N}$ (follows from Lemma A.2), and by the strong law of large numbers, $\left(\frac{2}{\alpha r_{T}}\right)^{s}\to\left(\frac{2}{\alpha p_{T}}\right)^{s}$ almost surely.
+as $N\to\infty.$ To justify the convergence, we note that $\left(\frac{2}{\alpha r_{T}}\right)^{s}$ is uniformly integrable as a sequence in $N\in\mathbb{N}$ (follows from Lemma 2), and by the strong law of large numbers, $\left(\frac{2}{\alpha r_{T}}\right)^{s}\to\left(\frac{2}{\alpha p_{T}}\right)^{s}$ almost surely.
 Since the above holds for any $s\in(0,1)$, it also holds for $s=s_{0}.$
@@
 $\leq\left(\frac{2}{\alpha\rho_{1}(s)^{t-T}}\right)^{s}\rho_{1}(s)^{t-T}+\rho_{1}(s)^{t-T}.$
-The last step follows from the assumption that $\tau$ is sensible (recall that if $\tau$ is sensitive, then $\tau_{j}$ is stochastically larger than $\tau$) and from Lemma A.3.
+The last step follows from the assumption that $\tau$ is sensible (recall that if $\tau$ is sensitive, then $\tau_{j}$ is stochastically larger than $\tau$) and from Lemma 3.
 Also, as we noted before, $\rho_{1}(s_{1})<1.$ Therefore,
@@
 $+\min\left\{\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right),\Delta_{F_{0},T,F_{1}}\right\}.$ (S5)
-Proof of Proposition 3.7.
+This completes the proof. ∎
+Proof of Proposition 3..
 We only prove the case when $\tau$ is the CUSUM type stopping time $\tau_{\text{wcs-ripr}}$ defined as
@@
 where $X\sim P$ and $X^{\prime}\sim P^{*}$.
-By Theorem 1.A.2 of [26], the stochastic ordering implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\frac{p_{1}^{*}(X^{\prime})}{p_{0}^{*}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$.
+By Theorem 1.A.2 of [shaked2007stochastic], the stochastic ordering implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\frac{p_{1}^{*}(X^{\prime})}{p_{0}^{*}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$.
 Therefore,
@@
 $=\mathbb{P}_{P^{*},\infty}[\tau_{\text{cs}}\geq t].$
-Proof of Theorem 3.8.
+This completes the proof. ∎
+Proof of Theorem 4..
 Since $r_{t}$ is an unbiased estimator of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$ and is independent of data, we have $\forall t\in\mathbb{N}$,
 $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/(\alpha r_{t}^{*})]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/(\alpha r_{t}^{*})\mid r_{t}^{*}]]\leq\mathbb{E}[\alpha r_{t}^{*}]=\alpha\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]\leq\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t],$ $\forall F_{0}\in\mathcal{P}_{0}$.
-The last steps follow from Assumption 3.5. The rest of the proof is the same as in the proof of Theorem 3.2. ∎
-Proof of Theorem 3.9.
+The last steps follow from Assumption 1. The rest of the proof is the same as in the proof of Theorem 2. ∎
+Proof of Theorem 5..
 It follows from the definition of the least favorable distribution that
 $\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right).$
-Therefore, for any $t\leq T-1$, using the above inequality and the assumption (3.10a) we get
+Therefore, for any $t\leq T-1$, using the above inequality and the assumption (15a) we get
 $\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$
@@
 $\leq\frac{1}{\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)}\left[\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)+\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}(r_{\tau}\leq\eta)\right].$
-In the last step, $\mathbb{P}_{F_{0},\infty}(\tau\geq T)\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)$ follows from (3.10a), $\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)$ follows from the fact that $(F_{0}^{*},F_{1}^{*})$ is an LFD pair, and $\mathbb{P}_{F_{0},T,F_{1}}(r_{\tau}\leq\eta)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}(r_{\tau}\leq\eta)$ follows from (3.10c).
+In the last step above, we used the inequality $\mathbb{P}_{F_{0},\infty}(\tau\geq T)\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)$, which follows from (15a), and $\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)$ $\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)$ follows from the fact that $(F_{0}^{*},F_{1}^{*})$ is an LFD pair, and $\mathbb{P}_{F_{0},T,F_{1}}(r_{\tau}\leq\eta)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}(r_{\tau}\leq\eta)$ follows from (15c).
 The remaining calculations are the same as those done in the proof of the previous theorem, with $f_{i}$ replaced by $f_{i}^{*}.$
-Proof of Theorem 4.1.
+Proof of Theorem 6..
 First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
@@
 under $H_{0,t}$.
-Hence, $M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-c)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$.
+Hence, $M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-c)(B+1)\rceil$ smallest of the collection $\{M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}\}$.
 So, for any fixed $c\in(0,1)$,
@@
 Thus, for all $t$,
-$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ (S7)
+$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ (S7)
 Therefore, combining (S6) and (S7), we obtain $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
-Proof of Theorem 4.3.
+Proof of Theorem 7..
 First, observe that $M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}$ are i.i.d.
@@
 Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha r_{t}})\geq 1-(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta$ from the above inequality.
-Proof of Theorem 4.4..
+∎
+Proof of Theorem 8..
 First, observe that $M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}$ are i.i.d. under $\mathbb{P}_{\theta_{0},t,\theta_{1}}$.
@@
 Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L})\geq 1-(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma$ from the above inequality.
-Proof of Theorem 5.1.
+∎
+Proof of Theorem 9..
 First note that $\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\operatorname{CS}(X_{T},\cdots,X_{\tau};1-\eta_{1}r_{T}^{*})\mid r_{T}^{*}\right)\leq\eta_{1}r_{T}^{*}$ and $\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{0}\notin\displaystyle\operatorname{CI}(X_{1},\cdots,X_{T-1};1-\eta_{0}r_{T}^{*})\mid r_{T}^{*}\right)\leq\eta_{0}r_{T}^{*}$, because $r_{T}^{*}$ is independent of $X_{i}$’s.
@@
 $\leq\alpha+\eta_{1},$
-where the last inequality follows from Assumption 3.5. Similarly,
+where the last inequality follows from Assumption 1. Similarly,
 $\mathbb{P}_{\theta_{0},T,\theta_{1}}\left(\theta_{1}\notin\displaystyle\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})\mid\tau\geq T\right)$
@@
 $\leq\alpha+\eta_{0}.$
+This completes the proof. ∎
 A.1 Auxilary Lemmas
-Lemma A.2.
+Lemma 2.
 Define, $r_{T}=\frac{1+\sum_{i=1}^{N}\mathds{1}(\tau\geq T)}{N+1}$. Then, for fixed $s\in(0,1)$, $\{r_{T}^{-s}\}_{N\geq 1}$ is uniformly integrable.
@@
 This bound $\frac{1}{p_{T}}$ is a constant that does not depend on $N$. Hence, $r_{t}^{-s}$ is uniformly integrable. ∎
-Lemma A.3.
+Lemma 3.
 Suppose, $\tau_{1},\cdots,\tau_{N}$ are i.i.d.
@@
 Suppose $\tau_{N+1}$ is independent and identically distributed (i.i.d.) with $\tau_{1},\cdots,\tau_{N}$.
-By Theorem 1.A.2 of [26], $\tau_{N+1}$ is stochastically larger than $\tau$ implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\tau\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\tau_{N+1}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$.
+By Theorem 1.A.2 of [shaked2007stochastic], $\tau_{N+1}$ is stochastically larger than $\tau$ implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\tau\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\tau_{N+1}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$.
 By direct calculation,
@@
 Appendix C Computationally and statistically efficient subroutines in Algorithm 2
-Let $L_{t}(\theta_{1};X_{1},\cdots,X_{n})$ be the likelihood of data $X_{1},\cdots,X_{n}$ under $\mathcal{H}_{0,t}$. Then, for $n\geq t$,
-$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (S1)
-We define $\hat{\theta}_{1,j:n}$ to be a maximum likelihood estimator (MLE) of the post-change parameter $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$.
-The point estimate of $T$ can be naturally extended to:
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (S2)
-The prototypical test statistic that we employ in our simulations is the likelihood-ratio-based test statistic $M_{t}^{(n)}$, where for $t,n\in\mathbb{N},t\leq n$,
-$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (S3)
 Note that the test statistic $M_{t}$ for testing $\mathcal{H}_{0,t}$ reduces to
@@
 In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$.
-In the case of discrete distributions, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta)\}_{n}$ across all $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
-Let us now fix $t\in\{1,\cdots,\tau\}$.
+Let us now fix $t\in\{1,\cdots,\tau\}$ and consider the test statistic $M_{t}$ of the form in (9), with some $R^{(t)}_{n}=R^{(t)}_{n}(\{X_{n}\}_{n})$ and $S^{(t)}_{n}=S^{(t)}_{n}(\{X_{n}\}_{n})$.
 Computing the detection time $\tau_{j,t}^{\theta}$ for each $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set.
-However, it is often straightforward to compute $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$.
+However, it is often straightforward to compute $t_{1}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$.
 Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient.
-For example, we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$.
-Again, we bound $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)$ by $V_{t,L}^{j}$, where for $L<\infty$,
-$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (S6)
-And for $L=\infty$,
-[…]
-which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section C.1). We define the confidence set as
-$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (S8)
-We observe that $\mathcal{C}_{1}^{\prime}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}$ is the set produced by Algorithm 2, since
-$\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$ $\geq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
+We consider the following upper bound:
+$V_{t,L}^{j}:=\begin{cases}\displaystyle\max\left\{\max_{t\leq n\leq t_{1}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}R^{(t)}_{n}\left(\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}\right),\max_{1\leq n\leq t-1}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}S^{(t)}_{n}\left(\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}\right)\right\},\\ \hskip 341.43306pt\text{if }t\leq t_{1}^{j}\leq L;\\ -\infty,\text{ if }t>t_{1}^{j};\\ \infty,\text{ if }t\leq t_{1}^{j}\text{ and }t_{1}^{j}>L.\end{cases}$ (S13)
+We consider the following confidence set:
+$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B}\}\text{ and}$ (S14)
+We observe that $\mathcal{C}_{1}^{\prime}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}_{1}$ is the set produced by Algorithm 2, since
+$\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B}\}$ $\geq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
 $\geq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}).$
-Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.3.
-Corollary C.1.
-$\mathcal{C}^{\prime}_{1}$ defined in (S8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.3.
-Next, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with the Poisson distribution in Section H.4.
+Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 7.
+Corollary 1.
+$\mathcal{C}^{\prime}_{1}$ defined in (S14) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 7.
+Next, we present a concrete example of the Gaussian mean shift problem to clarify our method.
 C.1 A concrete example: Gaussian mean shift problem
-Let the pre-change distribution $F_{0}$ be $N(\theta_{0},1)$ and the post-change model be $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
-In this setting, we have the following example of a $1-\beta$ confidence sequence [12] for $\theta_{1}$
+Let the pre-change distribution $F_{0}$ be $N(\theta_{0},1)$ and the post-change model be $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$, with $\Theta_{1}=\{\theta:\theta\geq\theta_{1}\}$ such that $\theta_{1}>\theta_{0}$.
+In this setting, we have the following example of a $1-\beta$ confidence sequence [howard2021time] for $\theta_{1}$
 $\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (S15)
@@
 where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$.
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (S4), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
+We consider the test statistic $M_{t}$, as defined in (9), with $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}}(X_{i})},$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w\geq 0$ being some weight function such that $\int_{\theta\in\Theta_{1}}w(\theta)d\theta=1$.
 Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$.
-Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark C.2) and they are finite.
+Suppose, we can find $t_{1}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark 5) and they are finite.
 So, we can have $L=\infty$.
-If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t$, it can be simplified in our setting as below:
-$V_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j}}{t^{\prime}-t+1}+\theta;Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}$
-$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
-$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta_{t^{\prime}}^{i});Y_{1}^{t}(X_{1}^{j},\theta_{t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta_{t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
-where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta)=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta),$ and $\theta_{t^{\prime}}^{i}$ is the unique element that maximizes $L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)$, or equivalently minimizes
-$-2\log L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)=$
-$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{(t-i)(\theta-\theta_{0})}{t^{\prime}-i+1}\right)^{2}+\text{const},\text{ if }i<t,\\ \text{const},\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{ const},\text{ if }i>t,\end{cases}$
-over $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$.
-Here, constant refers to the terms that do not involve $\theta$.
-Note that $\theta_{t^{\prime}}^{i}$ has a simple closed-form expression since the above equation is a quadratic function of $\theta$.
-And hence, $V_{t,\infty}^{j}$ can be computed easily.
-Remark C.2.
-For weighted CUSUM or SR detectors, finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. For some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$, these are defined as
-$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (S11)
- $\text{ and }\penalty 10000\ \tau_{\text{wsr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}.$ (S12)
-For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$.
+Note that, here we have
+$\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}R^{(t)}_{n}\left(\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}\right)=\exp\left((\theta_{0}-\theta_{1})\sum_{i=t}^{n}(\epsilon_{n}^{j}+\theta^{\prime}_{t,\min})+\frac{n-t+1}{2}(\theta_{1}^{2}-\theta_{0}^{2})\right),$
+and
+$\sup_{\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}}S^{(t)}_{n}\left(\{Y_{n}^{t}(X_{n}^{j},\theta^{\prime})\}_{n}\right)=\displaystyle\underset{\theta\in\Theta_{1}}{\int}\exp\left((\theta-\theta_{0})\sum_{i=n}^{t-1}(\epsilon_{n}^{j}+\theta^{\prime}_{t,\min})+\frac{t-n}{2}(\theta_{0}^{2}-\theta^{2})\right)w(\theta)d\theta,$
+where $\theta^{\prime}_{t,\min}=\inf\mathcal{S}^{\prime}_{\tau-t+1}$. And hence, $V_{t,\infty}^{j}$ in (S13) can be computed easily.
+Remark 5.
+For weighted CUSUM or SR detectors, finding $t_{1}^{j}$ is often straightforward. For some weight $w(\theta)$ such that $\int_{\theta\in\Theta_{1}}w(\theta)=1$, these are defined as
+$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)\geq A\right\}$ (S17)
+ $\text{ and }\penalty 10000\ \tau_{\text{wsr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)\geq A\right\}.$ (S18)
+For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta_{1}\}$, for some $\theta_{1}\geq\theta_{0}$.
 Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{n}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases.
-Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
-Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
-Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
-One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section H.3.
+Therefore, $t_{1}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
+Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta_{1}\}$, for some $\theta_{1}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
+One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors.
 The Gaussian mean shift example we provided serves as a concrete and illustrative case to clarify the methodology.
@@
   1. 1.
-    The coupling function $G$ is available (which is always the case when distributions are continuous and in the case of discrete distributions, even if a suitable $G$ is not directly available, our method may still be implementable, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution).
+    The coupling function $G$ is available (which is always the case when distributions are continuous).
   2. 2.
@@
   3. 3.
-    The quantities $t_{j}^{1}$ and $t_{j}^{2}$ are computable (either exactly or numerically).
-  4. 4.
-    If we employ the likelihood-ratio-based test statistic (S4), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
+    The quantities $t_{j}^{1}$ are computable (either exactly or numerically).
 For instance, one can easily verify that this approach applies to Gaussian, Laplace, or exponential scale change problems.
-C.2 Another concrete example: Poisson rate change problem
-Here is $F_{0}$ is $\text{Pois}(\theta_{0})$ and the post-change model is $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\in\Theta_{1}\}$.
-In this setting, we have the following example of a $1-\beta$ confidence sequence [2] for $\theta_{1}$
-$\mathcal{S}^{\prime}_{n}=\{\theta:n\theta-\sum_{i=t}^{t+n-1}X_{i}\log\theta\leq\log\left(\frac{1}{\beta}\right)+\log I(c,c\theta)-\log I(n+c,\sum_{i=t}^{t+n-1}X_{i}+c\theta)\}\cap\Theta_{1},$ (S13)
-where $c$ is any constant and $I(a,b)=\int_{-\infty}^{\infty}e^{-ae^{x}+bx}dx$.
-We fix $t\in\{1,\cdots,\tau\}$.
-Let $\lambda_{t}=\max\{\theta_{0},\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$.
-We only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Pois}(\lambda_{t})$ and $U_{n,1}^{j},\cdots,U_{n,X_{n}^{j}}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Uniform}(0,1)$, independently for each $n\in\mathbb{N},j=1,\cdots,B$.
-For $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$,
-$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta_{0}/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ otherwise.
-}\end{cases}$ (S14)
-Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta_{0})$, for $n<t$ and $Y_{n}^{t}(X_{n}^{j},\theta)\sim\text{Pois}(\theta)$, for $\theta\in\mathcal{S}^{\prime}_{\tau-t+1},n\geq t$.
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in Section 3, with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
-Let $\tau^{\theta}_{j}$ be the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$.
-Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark C.3) and they are finite.
-So, we can choose $L=\infty$.
-If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,V_{t,\infty}^{j}$ can be simplified as below:
-$V_{t,\infty}^{j}$ $=\max_{\begin{subarray}{c}i,t^{\prime}\in\mathbb{N}:1\leq i\leq t^{\prime},\\ \max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}\end{subarray}}\max_{\theta\in{S}^{j}_{t,t^{\prime}}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-t+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}.$
-Here ${S}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},t\leq n\leq t^{\prime},\inf\mathcal{S}^{\prime}_{\tau-t+1}\leq\lambda_{t}U_{n,i}^{j}\leq\sup\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\inf\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$ is the set consisting of one candidate element from is a finite set with at most $2+\sum_{n=t}^{t^{\prime}}X_{n}^{j}$ many elements.
-All the maximum are taken over a finite set; hence, $V_{t}^{j}$ can be computed easily.
-Remark C.3.
-For weighted CUSUM or SR detectors (defined in Remark 3.3.), finding $t_{1}^{j},t_{2}^{j}$ is often straightforward.
-For example, consider $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$.
-Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{(\tau-j+1)\times(\theta-\theta_{0})\}(\theta/\theta_{0})^{\sum_{i=j}^{\tau}X_{i}}$ increases as $X_{i}$’s increases and hence both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases.
-Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
-Similarly, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$.
-Also, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\neq\theta_{0}\}$, if we employ Gamma weights (it is a conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
 Appendix D Computationally and statistically efficient subroutines in Algorithm 3
-Let $L_{t}^{n}(\theta_{0},\theta_{1})$ be the likelihood of data up to time $n\geq t$ under $H_{0,t}$. Then,
-$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (S1)
-We define $\hat{\theta}_{0,1:j}$ to be the maximum likelihood estimator (MLE) of the null $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$.
-Similarly, we define $\hat{\theta}_{1,j:n}$ to be the MLE of the alternative $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$.
-The point estimate of $T$ can be naturally extended to the following form:
-$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (S2)
-As a prototypical example, we consider the likelihood-ratio-based test statistic $M_{t}^{(n)}$, which is defined for $t,n\in\mathbb{N},t\leq n$ as below.
-$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (S3)
-Note that the test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ reduces to
-[…]
 As in Appendix C, we assume that we have a fixed element $\theta^{\prime\prime}$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta_{0}\cup\Theta_{1}$.
@@
 $Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (S19)
-Note that, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta,\theta^{\prime})\}_{n}$ across all $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
-Let us now fix $t\in\{1,\cdots,\tau\}$.
+Let us now fix $t\in\{1,\cdots,\tau\}$ and consider the test statistic $M_{t}$ of the form in (9), with some $R^{(t)}_{n}=R^{(t)}_{n}(\{X_{n}\}_{n})$ and $S^{(t)}_{n}=S^{(t)}_{n}(\{X_{n}\}_{n})$..
 Computing the detection time $\tau^{j}_{\theta,\theta^{\prime}}$ for each $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are a finite sets.
-However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$ and $t_{2}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$.
+However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$.
 Also, we can often derive a suitable upper bound on the threshold, which does not require any approximation and is computationally and statistically efficient.
 E.g., we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$.
-Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where for $L<\infty$,
-$U_{t,L}^{j}:=\begin{cases}-\infty,\penalty 10000\ \penalty 10000\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\penalty 10000\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (S6)
-And for $L=\infty$,
-$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \penalty 10000\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (S7)
+Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where
+[…]
 We consider the following confidence set:
@@
 $\geq\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B}).$
-Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.4.
-Corollary D.1.
-$\mathcal{C}^{\prime\prime}_{1}$ defined in (S8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.4.
+Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 8.
+Corollary 2.
+$\mathcal{C}^{\prime\prime}_{1}$ defined in (S21) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 8.
 In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method.
-We illustrate another example with Poisson distribution in Appendix C of the Supplementary Material.
 D.1 A concrete example: Gaussian mean shift problem
-We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$.
+We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$, with $\Theta_{0}=\{\theta:\theta\leq\theta_{0}\}$ and $\Theta_{1}=\{\theta:\theta\geq\theta_{1}\}$, for some $\theta_{1}>\theta_{0}$.
 Then, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$:
@@
 For $\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}$, define
-$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S10)
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (S4), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
-Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark D.2.
-For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
-$\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j}+\theta_{0},\bar{\epsilon}_{t:t^{\prime}}^{j}+\theta_{1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}$
-$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
-$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\displaystyle\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{1:i-1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\bar{Y}_{i:t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i});Y_{1}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{j}(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j},\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
-where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime})=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime}),$ and $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ is the one such pair that maximizes $L_{i}\left(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)$, or equivalently minimizes
-$-2\log L_{i}\Big(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\dots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\Big)=$
-$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (S11)
-over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
-Here, constant means that it does not involve $\theta,\theta^{\prime}$.
-Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (D.1) is a quadratic function of $\theta_{1}-\theta_{0}$.
-Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
-Remark D.2.
-For weighted CUSUM or SR detectors, finding $t_{1}^{j},t_{2}^{j}$ is often straightforward.
-So, for some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$, the weighted CUSUM or SR detectors (with the Reverse Information Projection (RIPr) [17, 8, 16] being the representative of the pre-change class) are defined as
-$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S12)
-$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S13)
+$Y_{n}^{t,j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S23)
+We consider the test statistic $M_{t}$, as defined in (9), with $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}}(X_{i})}w^{\prime}(\theta)d\theta,$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w,w^{\prime}\geq 0$ being some weight functions such that $\int_{\theta\in\Theta_{1}}w(\theta)d\theta=1$ and $\int_{\theta\in\Theta_{0}}w^{\prime}(\theta)d\theta=1$.
+Note that $t_{1}^{j}$ is easy to compute for some CUSUM/SR type detection algorithms (see Remark 6) and they are finite.
+So, we can have $L=\infty$.
+Note that, here the term $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}R^{(t)}_{n}\left(\{Y_{n}^{t,j}(\theta,\theta^{\prime})\}_{n}\right)$ simplifies to
+$\underset{\theta\in\Theta_{0}}{\int}\exp\left((\theta-\theta_{1})\sum_{i=t}^{n}(\epsilon_{n}^{j}+\theta^{\prime}_{t,\min})+\frac{n-t+1}{2}(\theta_{1}^{2}-\theta^{2})\right)w^{\prime}(\theta)d\theta,$
+and the term $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}S^{(t)}_{n}\left(\{Y_{n}^{t,j}(\theta,\theta^{\prime})\}_{n}\right)$ simplifies to
+$\underset{\theta\in\Theta_{1}}{\int}\exp\left((\theta-\theta_{0})\sum_{i=n}^{t-1}(\epsilon_{n}^{j}+\theta^{\prime}_{t,\min})+\frac{t-n}{2}(\theta_{0}^{2}-\theta^{2})\right)w(\theta)d\theta,$
+where $\theta^{\prime}_{t,\min}=\inf\mathcal{S}^{\prime}_{\tau-t+1}$. And hence, $U_{t,\infty}^{j}$ in (S20) can be computed easily.
+Remark 6.
+For weighted CUSUM or SR detectors, finding $t_{1}^{j}$ is often straightforward.
+So, for some weight $w(\theta)$ such that $\int_{\theta\in\Theta_{1}}w(\theta)=1$, the weighted CUSUM or SR detectors (with the Reverse Information Projection (RIPr) [larsson2024numeraire, grunwald2024safe, lardy2023universal] being the representative of the pre-change class) are defined as
+$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}w(\theta)\geq A\right\},$ (S24)
+$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}w(\theta)\geq A\right\},$ (S25)
 where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$.
@@
 Therefore,
-$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
-$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
+$t_{1}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
 Similarly, one can check that for $\mathcal{P}_{0}=\{N(\theta,1):\theta\geq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\leq b\}$, for $b\leq a$,
-$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
-$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
+$t_{1}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
 The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology.
@@
 For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions.
-Another concrete example with the Poisson distribution is in Section H.4.
-D.2 Another concrete example: Poisson rate change problem
-Here we assume the pre and post-change model to be $\mathcal{P}_{i}=\{\text{Pois}(\theta):\theta\in\Theta_{i}\}$, for $i=0,1$ respectively.
-In this setting, we have the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (S13) and the same method can be employed for constructing $1-\gamma$ confidence interval for $\theta_{0}$ as well.
-We fix $t\in\{1,\cdots,\tau\}$.
-Let $\lambda_{t}=\max\{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$.
-We only need to generate $\{X_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Pois}(\lambda_{t})$ and $U_{n,1}^{j},\cdots,U_{n,X_{n}^{j}}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Uniform}(0,1)$, independently for each $n\in\mathbb{N},j=1,\cdots,B$.
-For $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$,
-$Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise.
-}\end{cases}$ (S14)
-Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta)$, for $n<t$, $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta^{\prime})$, for $n\geq t$ and $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
-We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in earlier, with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$.
-For Poisson distribution as well, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly for the same detection algorithms, similar to the Gaussian setting discussed in Remark 4.5.
-If $t_{2}^{j}<t$, we have $U_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified as below:
-$U_{t,\infty}^{j}$ $=\max_{\begin{subarray}{c}i,t^{\prime}\in\mathbb{N}:1\leq i\leq t^{\prime},\\ \max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}\end{subarray}}\max_{\begin{subarray}{c}\theta\in{A}^{i,j}_{t,t^{\prime}}\\ \theta^{\prime}\in{B}^{i,j}_{t,t^{\prime}}\end{subarray}}\frac{L_{i}\left(\frac{\sum_{k=1}^{i-1}Y_{k}^{j}(\theta,\theta^{\prime})}{i-1},\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime})}{t^{\prime}-i+1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)}{L_{t}\left(\frac{\sum_{k=1}^{t-1}Y_{k}^{j}(\theta,\theta^{\prime})}{t-1},\frac{\sum_{k=t}^{t^{\prime}}Y_{k}^{j}(\theta,\theta^{\prime})}{t^{\prime}-t+1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)}.$
-Here ${A}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},n\leq t-1,\inf\mathcal{S}_{t-1}\leq\lambda_{t}U_{n,k}^{j}\leq\sup\mathcal{S}_{t-1}\}\cup\{\inf\mathcal{S}_{t-1}\}\cup\{\sup\mathcal{S}_{t-1}\}$ and ${B}^{j}_{t,t^{\prime}}=\{\lambda_{t}U_{n,k}^{j}:k,n\in\mathbb{N},k\leq X_{n}^{j},t\leq n\leq t^{\prime},\inf\mathcal{S}^{\prime}_{\tau-t+1}\leq\lambda_{t}U_{n,k}^{j}\leq\sup\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\inf\mathcal{S}^{\prime}_{\tau-t+1}\}\cup\{\sup\mathcal{S}^{\prime}_{\tau-t+1}\}$ are finite sets with at most $(2+\sum_{n=1}^{t^{\prime}}X_{n}^{j})$ many elements.
-All the maximum are taken over a finite set; hence, $U_{t}^{j}$ can be computed easily.
 Appendix E Experiments for confidence intervals for pre/post change parameters
-E.1 Setting II
-Table E.1 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (5.1)) for $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta=0.05$.
+E.1 Known pre-change but unknown post-change
+Table 8 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (21)) for $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta=0.05$.
 The data generation and change detection procedure remain the same as in the experiment described in Setting II of Section 6.
 Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.85$, indicating that our method is reliable but conservative.
-Table E.1: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change.
+Table 8: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change.
 The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column of the table.
@@
 500 $0.9,\infty)$ 0.98 0.91 1.31 [0.90, 2.21]
-E.2 Setting III
-Table E.2 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), and (5.1)), for $\theta_{0}$ and $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$.
+E.2 Unknown pre- and post-change
+Table 9 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), and (21)), for $\theta_{0}$ and $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$.
 he data generation and change detection procedure remain the same as in the experiment described in Setting III of Section 6.
 Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.80$.
-Table E.2: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.85(Setting III): unknown pre- and post-change.
+Table 9: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.85(Setting III): unknown pre- and post-change.
 The true (unknown) values of the parameters are $\theta_{0}=0$ and $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns of the table.
@@
 500 $(\infty,0.1]$ $\infty) 0.99 0$.99 0.94 0.94 0.19 1.24 [-0.09, 0.10] [0.90, 2.14]
-E.3 Comparison with [39]
-Next, we compare our confidence interval for the post-change parameter with that proposed by [42] and discussed in [39].
-We implement Wu’s interval as discussed in Chapter 4.3 of [39] (or Section of [42]), replacing $\theta$ by $\hat{\theta}$ in the bias and variance terms, which is given by $\hat{\theta}-\frac{4}{5\sqrt{d\hat{\theta}(\tau-\hat{\nu})}}\pm z_{\alpha/2}\frac{\sqrt{1-{13}/(16d\hat{\theta})}}{\sqrt{\tau-\hat{\nu}}}$, where $\hat{\theta}=\frac{T_{\tau}}{\tau-\hat{\nu}}$, $\tau=\tau^{\prime}_{\text{cusum}}$, and $z_{\alpha^{\prime}/2}$ is the $\alpha^{\prime}/2$ quantile of the standard normal distribution.
-We use CUSUM [19] (the same change detector as in the last experiment of Section 6), with $d=10$ for both methods.
+E.3 Comparison with Wu [WU20063625]
+Next, we compare our confidence interval for the post-change parameter with that proposed by [WU20063625] and discussed in [wu2007inference].
+We implement Wu’s interval as discussed in Chapter 4.3 of [wu2007inference] (or Section 3 of [WU20063625]), replacing $\theta$ by $\hat{\theta}$ in the bias and variance terms, which is given by $\hat{\theta}-\frac{4}{5\sqrt{d\hat{\theta}(\tau-\hat{\nu})}}\pm z_{\alpha/2}\frac{\sqrt{1-{13}/(16d\hat{\theta})}}{\sqrt{\tau-\hat{\nu}}}$, where $\hat{\theta}=\frac{T_{\tau}}{\tau-\hat{\nu}}$, $\tau=\tau^{\prime}_{\text{CUSUM}}$, and $z_{\alpha^{\prime}/2}$ is the $\alpha^{\prime}/2$ quantile of the standard normal distribution.
+We use CUSUM [page1954continuous] (the same change detector as in the second last experiment of Section 6), with $d=10$ for both methods.
 For our method, we assume the post-change space to be $\Theta_{1}=[0,\infty)$ with $\alpha=\beta=\eta=0.05$.
-We report the average size and the conditional and unconditional coverage rates of both methods across $500$ independent runs in Table E.3, with a target conditional coverage of $\alpha^{\prime}=0.85$.
-The results show that the existing method [39], which is based on asymptotic approximations, undercovers $\theta_{1}$ (both conditionally and marginally).
+We report the average size and the conditional and unconditional coverage rates of both methods across $500$ independent runs in Table 10, with a target conditional coverage of $\alpha^{\prime}=0.85$.
+The results show that the existing method [WU20063625], which is based on asymptotic approximations, undercovers $\theta_{1}$ (both conditionally and marginally).
 At the same time, our finite-sample confidence intervals are conservative — achieving higher-than-nominal (conditional) coverage at the expense of larger (conditional) length.
-Table E.3: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage $0.85$) with Wu [39] (Setting II)
+Table 10: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage $0.85$) with Wu [WU20063625] (Setting II)
 $F_{0}$ $F_{1}$ $T$ Conditional Length Conditional Coverage Marginal Coverage
@@
 N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 500 1.49 0.51 0.99 0.79 0.91 0.71
-Appendix F Additional experiments for composite pre/post change
-In this section, we present experimental results for the alternative approach in which each set $\mathcal{S}_{t-1}$ and/or $\mathcal{S}_{\tau-t+1}^{\prime}$ is discretized using 50 equispaced grid points.
-Let, $\{\theta_{1}^{(0)},\cdots,\theta_{I}^{(0)}\}\subseteq\mathcal{S}_{t-1}$ and $\{\theta_{1}^{(1)},\cdots,\theta_{K}^{(1)}\}\subseteq\mathcal{S}_{\tau-t+1}^{\prime}$ be the grid points.
-Here, $I=K=50$.
-Therefore, for $j\in\{1,\cdots,B\}$, we only need to generate $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$. For $i\in[I],k\in[K]$, define
-$Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})=\begin{cases}\theta_{i}^{(0)}+\epsilon_{n}^{j},\text{ if }n<t\\ \theta_{k}^{(1)}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S1)
-We run $\mathcal{A}$ on $\{Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{n}$ and denote $\tau_{i,k}^{j}$ as the detection time.
-We compute the test statistics based on $\tau_{i,k}^{j}$ and $Y_{1}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)}),\cdots,Y_{\tau_{i,k}^{j}}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})$.
-Then, we consider the following confidence set:
-$\mathcal{C}=\left\{t\in\{1,\cdots,\tau\}:M_{t}\leq\max_{i\in[I],k\in[K]}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{j=1}^{B})\right\}$ (S2)
-As a special case, when $\theta_{0}$ is known, then, $\theta_{0}=\theta_{1}^{(0)}=\cdots=\theta_{I}^{(0)}=\mathcal{S}_{t-1}.$
-As shown in Table F.1, the confidence sets (S2) have lengths and coverages that are comparable to those from the earlier method reported in Table 6.3 and Table 6.5.
-However, this grid-based approach incurs significantly higher computational cost.
-Therefore, for practical purposes, we recommend using the earlier, more computationally efficient method.
-Table F.1: Composite pre/post change
-T
-Pre-change
-Class
-Post-change
-Class
-Conditional
-coverage
-Unconditional
-coverage
-Size Time (s)
-100 $N(0,1)$ $\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 27.05 365.81
-100 $N(0,1)$ $\{N(\mu,1):\mu>0.9\}$ 0.95 0.95 24.62 360.48
-100 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ 0.96 0.96 30.88 17572.26
-100 $\{N(\mu,1):\mu\leq 0.1\}$ $\{N(\mu,1):\mu>0.9\}$ 0.96 0.95 30.14 17092.11
-Appendix G Simpler confidence sets for detection algorithms controlling PFA
+Appendix F Simpler confidence sets for detection algorithms controlling PFA
 In this section, we shall discuss how one can construct simpler confidence sets when the change detection algorithm $\mathcal{A}$ controls the probability of false alarm (PFA) at level $\delta\in(0,1)$.
-Importantly, this method allows us to avoid Assumption 3.5, and our results hold in the fully nonparametric case with arbitrary pre- and post-change classes, $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$, respectively.
-We consider the test statistic $M_{t}$ introduced in (3.4), and define the following confidence set for $T$:
+Importantly, this method allows us to avoid Assumption 1, and our results hold in the fully nonparametric case with arbitrary pre- and post-change classes, $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$, respectively.
+We consider the test statistic $M_{t}$ introduced in (9), and define the following confidence set for $T$:
 $\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha}\right\},$ (S26)
-which is a simplified version of Equation 3.9, obtained by directly replacing $\alpha r_{t}^{*}$ with $\alpha$.
-This avoids the need to compute $r_{t}^{*}$ (and hence $\tau_{1}^{*},\cdots,\tau_{N}^{*}$), making the method easier to implement and free of Assumption 3.5.
+which is a simplified version of Equation 14, obtained by directly replacing $\alpha r_{t}^{*}$ with $\alpha$.
+This avoids the need to compute $r_{t}^{*}$ (and hence $\tau_{1}^{*},\cdots,\tau_{N}^{*}$), making the method easier to implement and free of Assumption 1.
 We now investigate the coverage properties of the above confidence set.
-Theorem G.1.
+Theorem 10.
 Let $\mathcal{C}$ be the confidence set as defined in (S26), for some $\alpha\in(0,1)$.
@@
 Proof.
-From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha}]\leq\alpha$. Hence,
+From Theorem 1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha}]\leq\alpha$. Hence,
 $\mathbb{P}_{F_{0},T,F_{1}}(M_{T}<2/\alpha)$
@@
 $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha-\delta.$
-Now,
+And
 $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\frac{\alpha}{1-\delta},$
+completing the proof. ∎
 A similar simplification can be applied to the adaptive confidence sets introduced in Section 4 for the parametric setting.
-By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 4.1, 4.3 and 4.4, we obtain
+By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 17, 19 and 20, we obtain
 $\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}\}_{j=1}^{B})\right\},$ (S27)
@@
 For detection algorithms with PFA $\leq\delta,\penalty 10000\ \delta\in(0,1)$, these simpler forms retain the coverage guarantees — both conditional and unconditional — similar to what we have shown in the theorem above.
-In the next section, we present the experimental results with the bounded PFA algorithms, along with some other experiments.
-Appendix H Some additional experimental results
-H.1 Experiments varying $T$ and ARL
-We have seen in deriving the asymptotic length bound that it increases with the ratio $T/A_{T}$.
-Now, we experimentally validate the fact by conducting the same experiment as in Setting I (known pre- and post-change) in Section 6 with varying $T$ thresholds $A$ of the CUSUM detector (larger threshold means large ARL).
-Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 3.5 are reported in Table H.1.
-The results show a clear pattern: as the ratio $T/A$ increases, the resulting confidence sets become wider, reflecting growing conservativeness.
-Table H.1: Setting I: Known pre- and post-change.
-Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 3.5 with $F_{0}=N(0,1),F_{1}=N(1,1),N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (3.1) and average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
-$T$ A
-Conditional
-coverage
-Marginal
-coverage
-Size
-$|\hat{T}-T|$
-Delay
-50 100 0.98 0.91 17.28 2.58 9.33
-1000 100 0.99 0.21 20.88 2.45 9.14
-2000 100 1 0.07 27.02 2.60 8.35
-500 1000 0.98 0.97 15.77 2.62 13.22
-10000 1000 1 0.27 24.29 2.36 13.25
-20000 1000 1 0.07 30.16 2.29 13.22
-H.2 Experiments with dependent pre- and post-change data
-[…]
-We use the likelihood ratio-based cusum detector:
-$\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(j:n)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(j:n)}\geq A\right\},$ (S1)
-where $n_{11}(j:n)$ denotes the number of transitions from 1 to 1 and $n_{10}(j:n)$ denotes the number of transitions from 0 to 1 within $X_{j},\cdots,X_{n}$, and $A=1000$.
-After detection, we compute the adaptive confidence sets (4.1) with the likelihood ratio-based test statistic (3.2) and point estimate as in (3.1).
-Note that our adaptive method (4.1) is the only one we know that is valid for dependent data.
-Table H.2 contains the results averaged across $500$ runs.
-Table H.2: Known pre- and post-change with dependency (Setting I)
-T
-Conditional
-coverage
-Marginal
-coverage
-Size
-$|\hat{T}-T|$
-Delay
-100 0.99 0.98 36.12 4.35 28.99
-500 0.98 0.98 35.83 4.42 27.36
-H.3 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
-H.3.1 Known pre- and post-change (Setting A, PFA $\leq 0.001$)
+In the next subsection, we present the experimental results with the bounded PFA algorithms, along with some other experiments.
+F.1 Experiments with detection algorithms that control PFA
+F.1.1 Known pre- and post-change (Setting A, PFA $\leq 0.001$)
 We perform experiments for both Gaussian and Poisson mean-change scenarios having the true changepoint at $T=100,500$, using a likelihood ratio detector defined as
@@
 with thresholds at $A=1000$.
-Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality [35], the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
-After detection, we construct both the proposed confidence sets (universal (S1) and adaptive (S2)), with $B=100$, $\alpha=0.05$.
-We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.3.
-We also report the average absolute deviation of $\hat{T}$ (defined in (3.1)) and delay (given no false detection) in the table.
-Table H.3: Known pre- and post-change distributions (Setting A, PFA $\leq 0.001$)
+Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality [ville1939etude], the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
+After detection, we construct both the proposed confidence sets (universal (S26) and adaptive (S27)), with $B=100$, $\alpha=0.1$.
+We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table 11.
+We also report the average delay (given no false detection) in the table.
+Table 11: Known pre- and post-change distributions (Setting A, PFA $\leq 0.001$)
 $T$ $F_{0}$ $F_{1}$ Conditional Coverage Conditional Size
-$|\hat{T}-T|$
 Delay
-Universal Adaptive Universal Adaptive (Conditional) (Conditional)
-100 $N(0,1)$ $N(1,1)$ 0.95 0.94 15.44 12.60 2.74 110.91
-100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 19.36 17.95 4.98 94.33
-500 $N(0,1)$ $N(1,1)$ 0.97 0.97 15.92 12.65 2.24 514.12
-500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 18.89 16.82 3.18 417.44
-H.3.2 Known pre- and unknown post-change (Setting B, PFA $\leq 0.001$)
+Universal Adaptive Universal Adaptive (Conditional)
+100 $N(0,1)$ $N(1,1)$ 0.95 0.94 12.65 9.86 110.91
+100 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 14.22 11.85 94.33
+500 $N(0,1)$ $N(1,1)$ 0.97 0.97 12.83 9.76 514.12
+500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 14.79 12.13 417.44
+F.1.2 Known pre- and unknown post-change (Setting B, PFA $\leq 0.001$)
 We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector defined as
@@
 with thresholds at $A=1000$.
-The weights are chosen just as described in Table 6.2.
+The weights are chosen just as described in Table 2.
 Similarly, the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
-After detection, we construct our universal confidence sets (S1) for $T$, with $\alpha=0.075$.
-The test statistic and the point estimate are the same as mentioned in Setting II of Section 6.
-We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.4.
-We also report the average absolute deviation of $\hat{T}$ in (S2) and delay (given no false detection) in the table.
-Table H.4: Known pre-change but unknown post-change (Setting B, PFA $\leq 0.001$)
+After detection, we construct our universal confidence sets (S26) for $T$, with $\alpha=0.1$.
+The test statistic is the same as mentioned in Setting II of Section 6.
+We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table 12.
+We also report the average delay (given no false detection) in the table.
+Table 12: Known pre-change but unknown post-change (Setting B, PFA $\leq 0.001$)
 T
@@
 coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 23.82 3.68 118.76
-100 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 20.15 3.36 115.64
-500 $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 23.15 3.18 530.83
-500 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 19.27 3.23 528.92
-H.3.3 Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
-We perform experiments for Gaussian mean-change scenarios having the true changepoint at $T=100,500$, using a mixture likelihood ratio detector (with the Reverse Information Projection (RIPr) [17, 8, 16] being the representative of the pre-change class) defined as
-$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (S4)
-with $f^{*}_{\theta}$ being the RIPr of $f_{\theta}$ on the pre-change class $\mathcal{P}_{0}$ and with thresholds at $A=1000$.
-This is a stopping time of an “e-process” [20] and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here.
-The weights are chosen just as described in Table 6.3.
-After detection, we construct our universal confidence sets (S1) for $T$, with $\alpha=0.1$.
-The test statistic and the point estimate are the same as mentioned in Setting III of Section 6.
-We report the average size and coverage rates across $500$ independent runs in Table H.5.
-We also report the average absolute deviation of $\hat{T}$ in (S2) and delay (given no false detection).
-The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
-Table H.5: Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
-T
-Composite
-pre-change
-Composite
-post-change
+Size
+Delay
+100 $\{N(\mu,1):\mu>0.75\}$ 0.96 0.95 16.38 118.76
+100 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 13.11 115.64
+500 $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 16.92 530.83
+500 $\{N(\mu,1):\mu>0.9\}$ 0.96 0.96 14.85 528.92
+F.1.3 Composite pre- and post-change (Setting C)
+We generate pre-change data from $N(0,1)$ and post-change data from $N(0,1)$, having a true changepoint at $T=100$ or $500$ — these are all unknown to the algorithm.
+We assume access to a historic/training data $\{Y_{1},\cdots,Y_{m}\}$, $m=200$.
+We detect change using the following change detector, as in Sections 2 and 3 of [aue2024state]:
+$\tau=\min\left\{k\in\mathbb{N}:\frac{1}{\hat{\sigma}_{m}\sqrt{m}}w(m,k)\left|\sum_{j=1}^{k}(X_{j}-\bar{Y}_{m})\right|>A\right\},$ (S32)
+where $\bar{Y}_{m}=\frac{1}{m}\sum_{i=1}^{m}Y_{i}$ is the sample mean of the training data, $\hat{\sigma}_{m}=\sqrt{\frac{1}{m-1}\sum_{i=1}^{m}(Y_{i}-\bar{Y}_{m})^{2}}$ is the estimated standard deviation from the training data, $w(m,k)=\left(1+\frac{k}{m}\right)^{-1}$ and $A$ is the threshold, which we vary in our experiment.
+After detection, we construct our universal confidence sets (S26) for $T$, with $\alpha=0.1$, $\mathcal{P}_{0}=\{N(\mu,1):\mu\leq 0.25\}$ and $\mathcal{P}_{1}=\{N(\mu,1):\mu>0.75\}$.
+The test statistic is the same as mentioned in Setting III of Section 6.
+We report the average size and coverage rates across $500$ independent runs in Table 13.
+We also report the average delay (given no false detection).
+Table 13: Composite pre- and post-change (Setting C)
+T A
 Conditional
@@
 Size
-$|\hat{T}-T|$
 Delay
-100 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 26.02 3.97 125.64
-100 $\{N(\mu,1):\mu\leq 0.1\}$ $\{N(\mu,1):\mu>0.9\}$ 0.97 0.97 21.96 3.88 121.28
-500 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ 0.97 0.97 25.18 3.82 544.91
-500 $\{N(\mu,1):\mu\leq 0.1\}$ $\{N(\mu,1):\mu>0.9\}$ 0.98 0.98 21.45 3.75 536.76
-H.3.4 Comparison of experimental results with ARL and PFA controlling detection algorithms
+100 2 0.97 0.97 23.54 50.64
+100 3 0.97 0.97 21.82 82.95
+500 2 0.97 0.97 21.37 122.84
+500 3 0.98 0.98 23.22 193.59
+F.2 Comparison of experimental results with ARL and PFA controlling detection algorithms
 As expected, algorithms that control PFA exhibit a significantly higher average detection delay compared to those controlling ARL.
@@
 In terms of inference, we observe that PFA-controlling algorithms often tend to yield slightly lower conditional coverage probabilities with confidence sets of comparable sizes.
-Also, conditional and unconditional coverage rates are identical when PFA-controlling detection algorithms are employed, due to the very low probability of false detections, while the ARL controlling algorithms lead to lower unconditional conditional coverage rates for $T=500$, supporting our theory.
-Moreover, we observe that the average absolute deviations of the point estimates are barely smaller with the PFA controlling algorithms.
+Also, conditional and unconditional coverage rates are identical when PFA-controlling detection algorithms are employed, due to the very low probability of false detections, while the ARL controlling algorithms lead to lower unconditional conditional coverage rates for larger $T$, ie.e., $T=500$, supporting our theory.
 Overall, the major differences between these two types lie in detection delay and unconditional coverage properties.
-We do not observe significant differences in the size and conditional coverage properties of the confidence sets nor in the absolute deviation of the point estimates, across these two types of detection algorithms.
-H.4 Experimental results with Poisson distribution with detection algorithms that control ARL
-H.4.1 Known pre- and post-change (Setting I)
-We consider Poisson the mean-change problem: $\text{Pois}(1)$ (pre-change) vs $\text{Pois}(2)$ (post-change) with the true (unknown) changepoint at $T=100,500$.
-We employ the likelihood-based CUSUM change detection method with thresholds at $A=1000,10000$.
-After detection, we construct confidence sets (adaptive, Equation 4.1) for $T$.
-Fig.
-H.6 provides visualizations of the confidence sets and point estimates of the changepoints across $5$ runs.
-Figure H.6: Setting I (known pre- and post-change Poisson distributions): The first T1T-1 observations are drawn from Pois(1)(1) and the rest from Pois(2)(2), with $T=100,500$.
-The point estimate is shown in a vertical red dashed line, and the confidence set (using Algorithm 2) is shown in red points.
-$B=N=100$, $\alpha=0.05,L=\infty$.
-Results of $5$ independent simulations are shown.
-We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.7.
-Table H.7: Setting I: Average (of $500$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ and average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
-$T$ AA
-Conditional
-coverage
-Unconditional
-coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 10000 0.94 0.94 18.04 4.02 22.15
-100 1000 0.94 0.94 17.45 3.98 15.72
-500 10000 0.95 0.95 18.53 4.30 22.10
-500 1000 0.96 0.89 18.72 3.96 16.14
-For $A=10000$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible.
-However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage.
-Nevertheless, the empirical conditional coverage rates meet the desired level, supporting our theory.
-We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table.
-H.4.2 Known pre-change and composite post-change (Setting II)
-We perform experiments for Poisson rate-change scenarios having pre-change distribution as $\text{Pois}(1)$, and the true (unknown) post-change data distribution as $\text{Pois}(2)$, with the changepoint at $T=100,200,$ assuming the composite post-change model to be $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$.
-For detection, we employ a weighted CUSUM detector defined as
-$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (S5)
-for some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$.
-We use thresholds $A=1000,10000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, and $\theta=1.9$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
-After detection, we construct confidence sets (as defined in Section C.2), with $\alpha=0.05,\beta=0.025,B=N=100,L=\infty$.
-We report the average size and the conditional and unconditional coverage across $500$ independent runs in (H.8).
-The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table H.7).
-We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table.
-We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
-Table H.8: Setting II: Average (of $500$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets Section 3.2.
-with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ and average detection delay (given no false detection) of the change detection method used (S11) with threshold AA, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
+We do not observe significant differences in the size and conditional coverage properties of the confidence sets, nor in the absolute deviation of the point estimates, across these two types of detection algorithms.
+Appendix G Some additional experimental results
+G.1 Experiments varying $T$ and ARL
+We have seen in the derivation of the asymptotic length bound that it increases with the ratio $T/A_{T}$.
+Now, we experimentally validate the fact by conducting the same experiment as in Setting I (known pre- and post-change) in Section 6 with varying $T$ thresholds $A$ of the CUSUM detector (larger threshold means large ARL).
+Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 10 are reported in Table 14.
+The results show a clear pattern: as the ratio $T/A$ increases, the resulting confidence sets become wider, reflecting growing conservativeness.
+Table 14: Setting I: Known pre- and post-change.
+Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 10 with $F_{0}=N(0,1),F_{1}=N(1,1),N=100$, $\alpha=0.1$, the average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
 $T$ A
@@
 coverage
-Unconditional
-coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 10000 0.97 0.97 33.13 6.11 27.68
-100 1000 0.97 0.96 33.85 5.59 19.14
-200 10000 0.98 0.98 36.08 5.98 26.81
-200 1000 0.97 0.93 35.24 5.74 21.27
-H.4.3 Composite pre- and post-change (Setting III)
-Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively.
-For change detection, we employ a weighted CUSUM-type detector, where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, which is defined as
-$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S6)
-where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For the post-change parameter, the same discrete weight distribution as in Section H.4.2 is considered.
-After detection, we construct confidence sets (as defined in Section D.2, assuming $T\neq 1$), with $\alpha=0.05,\beta=\gamma=0.025,B=N=100,L=\infty$.
-We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$, and delay (given no false detection) across $500$ independent runs in Table H.9.
-Table H.9: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ and average detection delay (given no false detection) of the change detection method with threshold AA, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
-T A
-Conditional
-coverage
-Unconditional
-coverage
-Size $\mathbb{E}_{T}(|\hat{T}-T||\tau\geq T)$ $\mathbb{E}_{T}(\tau-T|\tau\geq T)$
-100 10000 0.97 0.97 35.04 6.34 31.56
-100 1000 0.97 0.97 35.19 6.27 26.33
-200 10000 0.98 0.98 37.54 5.92 30.45
-200 1000 0.98 0.97 37.86 6.18 25.82
-However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments).
-It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table H.7).
-We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table H.7), exhibiting both higher coverage and larger size.
-Appendix I A heuristic method for composite pre- or post-change
-This section presents an alternative method for constructing confidence intervals for composite pre- and post-change distributions.
-Specifically, we propose a heuristic bootstrap-based approach that estimates the unknown distributions from the observed data and subsequently generates simulations from the estimated distributions.
-I.1 Known pre-change and unknown post-change distributions
-Here we assume that the pre-change distribution $F_{0}$ is known and the composite post-change model is $\mathcal{P}_{1}$.
-Building on the approach from previous sections, we define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
-$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},\text{ for some (unknown) }F_{1}\in\mathcal{P}_{1}.$
-For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$.
-Since the pre-change $F_{0}$ is known, we can generate $N$ many i.i.d.
-sequences under $F_{0}$ until a change is detected, recording the detection times as $\tau_{1},\cdots,\tau_{N}$, for some $N\in\mathbb{N}$.
-Now, the challenge in conducting a level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq t)$ test for each $t$ lies in the composite nature of $\mathcal{H}_{0,t}$, which prevents us from straightforwardly simulating data to determine a rejection threshold as we did in Algorithm 1.
-To address this, we employ a simple bootstrap-based approach.
-The main idea is to first obtain some point estimate of the changepoint (for example, a likelihood ratio based estimate, as defined earlier).
-We denote the estimate as $\hat{T}$.
-Then we use the observations from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$.
-Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with known pre-change distribution $F_{0}$ and estimated post-change distribution $\hat{F}_{1}$ having changepoint at $t$.
-For each sampled stream $j$, we compute the same test statistic $M_{t}^{j}$ under the null hypothesis, resulting in statistics $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if
-$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
-Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 4 contains an overview of the method.
-Algorithm 4 CI for known pre-change $F_{0}$ and composite post-change $\mathcal{P}_{1}$
-[…]
-2:A confidence set $\mathcal{C}$ for changepoint $T$
-3:Compute a point estimator $\hat{T}$ for $T$
-4:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
-5:for $i=1$ to NN do
-6:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
-7:end for
-[…]
-9:for $t=1$ to $\tau$ do
-10:  Compute some test statistics $M_{t}$ based on $\tau$ and $X_{1},\cdots,X_{\tau}$
-11:  for $j=1$ to $B$ do
-12:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\penalty 10000\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
-13:   Compute the same test statistics $M_{t}^{j}$ based on $\tau_{j}^{\prime}$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}$
-14:  end for
-15:  if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
-16:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
-17:  end if
-18:end for
-I.2 Composite pre- and post-change distributions
-We now consider the general case where both pre and post-change distributions follow some composite models $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively.
-Following the approach from the previous sections, we define the null hypothesis that a changepoint occurs at time $t\in\{1,\cdots,\tau\}$ as:
-[…]
-For testing the null hypothesis, we define some test statistic $M_{t}$, based on $X_{1},\cdots,X_{\tau}$ (eg., the likelihood-ratio-based test statistic as defined earier).
-As in Algorithm 4, we employ a simple bootstrap-based approach, but here we need to estimate both pre- and post-change distributions.
-We first obtain some point estimate $\hat{T}$ of the changepoint $T$.
-Then we use the observations from $1$ to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$.
-For some $N\in\mathbb{N}$, we generate $N$ many i.i.d.
-sequences under $\hat{F}_{0}$ until a change is detected and denote the detection times as $\tau_{1},\cdots,\tau_{N}$.
-Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with the estimated pre- and post-change distributions $\hat{F}_{0}$ and $\hat{F}_{1}$ respectively, having changepoint at $t$.
-For each sampled stream $j$, we calculate the test statistic $M_{t}^{j}$ to obtain $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$.
-We reject $\mathcal{H}_{0,t}$ if we have
-$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
-Our confidence set is the collection of all $t$ before $\tau$, for which we fail to reject $\mathcal{H}_{0,t}$. Algorithm 5 contains an overview of the method.
-Algorithm 5 CI for unknown pre- and post-change
-[…]
-2:A confidence set $\mathcal{C}$ for changepoint $T$
-3:Compute some point estimator $\hat{T}$ for $T$
-4:Obtain an estimate $\hat{F}_{0}$ of the (unknown) pre-change $F_{0}\in\mathcal{P}_{0}$ based on $X_{1},\cdots,X_{\hat{T}-1}$
-5:Obtain an estimate $\hat{F}_{1}$ of the (unknown) post-change $F_{1}\in\mathcal{P}_{1}$ based on $X_{\hat{T}},\cdots,X_{\tau}$
-6:for $i=1$ to NN do
-7:  Simulate $X_{1}^{i},X_{2}^{i},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0}$, until change is detected at $\tau_{i}$ using $\mathcal{A}$, or until $\tau$, whichever happens first
-8:end for
-[…]
-10:for $t=1$ to $\tau$ do
-11:  Compute some test statistics $M_{t}$ based on $X_{1},\cdots,X_{\tau}$
-12:  for $j=1$ to $B$ do
-13:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\penalty 10000\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
-14:   Compute the same test statistics $M_{t}^{j}$ based on $X_{1}^{j},\cdots,X_{\tau_{j}}^{j}$
-15:  end for
-16:  if $M_{t}\leq\operatorname{Quantile}(1-\alpha;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
-17:   $\mathcal{C}=\mathcal{C}\cup\{t\}$
-18:  end if
-19:end for
-I.3 Experimental Results
-I.3.1 Known pre-change and composite post-change (Setting II)
-We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$.
-For detection, we employ a weighted CUSUM detector (S11) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
-After detection, we employ Algorithm 4 to construct confidence sets.
-Table I.1 summarizes the results, including the average size and the conditional and unconditional coverage across $500$ independent runs.
-For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible.
-However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage.
-Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
-Table I.1: Setting II: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 4 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (S11) with threshold $A=1000$.
-T Pre-change
-Composite
-post-change
-True (unknown)
-post-change
-Conditional
-coverage
-Unconditional
+Marginal
 coverage
 Size
-100 $N(0,1)$ $\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.94 0.94 14.76
-100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.95 20.95
-500 $N(0,1)$ $\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.95 0.91 14.06
-500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.89 20.83
-I.3.2 Composite pre- and post-change (Setting III)
-Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$.
-We use a weighted CUSUM detector with RIPr defined in (S12) for change detection.
-The thresholds are set at $A=1000$ for both $T=100$ and $T=500$.
-We considered the same discrete weight distribution as above, taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=0.8,d=0.2$ and for Poisson distribution $\theta=1.8,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class.
-After detection, we employ Algorithm 5 to construct confidence sets.
-We report the average size and the conditional and unconditional coverage across $500$ independent runs in Table I.2.
-The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
-Table I.2: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (S12) with threshold $A=1000$.
+Conditional
+Delay
+50 100 0.978 0.91 14.382 9.33
+1000 100 0.989 0.21 18.283 9.14
+2000 100 0.992 0.07 19.627 8.35
+500 1000 0.925 0.97 12.618 13.22
+10000 1000 0.947 0.27 18.254 13.25
+20000 1000 1 0.07 22.113 13.22
+G.2 Experiments with dependent pre- and post-change data
+[…]
+We use the likelihood ratio-based CUSUM detector:
+$\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(j:n)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(j:n)}\geq A\right\},$ (S33)
+where $n_{11}(j:n)$ denotes the number of transitions from 1 to 1 and $n_{10}(j:n)$ denotes the number of transitions from 0 to 1 within $X_{j},\cdots,X_{n}$, and $A=1000$.
+After detection, we compute the adaptive confidence sets (17) with the test statistic in (9) with $R^{(t)}_{n}=\frac{1}{2}\left(\frac{p_{0}}{p_{1}}\right)^{n_{10}(t:n)}\left(\frac{1-p_{0}}{1-p_{1}}\right)^{n_{11}(t:n)}$, $S^{(t)}_{n}=\frac{1}{2}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(n:t-1)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(n:t-1)}$, $B=N=100$ and $\alpha=0.1$.
+Note that $\{S^{(t)}_{n}\}_{n<t}$ is not a valid e-process in the backward filtration.
+In contrast to our universal method, the adaptive procedure in Algorithm 1 (or Equation 17) accommodates arbitrary test statistics and does not rely on e-process structure; consequently, it remains applicable in settings with dependent data.
+Table 15 contains the results averaged across $100$ random runs.
+Table 15: Known pre- and post-change with dependency
 T
-Composite
-post-change
-Composite
-post-change
-True
-pre-change
-True
-post-change
-Conditional
-coverage
-Unconditional
-coverage
+Conditional coverage
 Size
-100 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 19.52
-100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.94 0.94 23.32
-500 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 18.56
-500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 23.04
-Appendix J Another heuristic method for composite pre- or post-change
-In this section, we provide another heuristic method for dealing with composite pre- or post-change distributions.
-We use our algorithm for known pre- and post-change (Algorithm 2) with $F_{0}$ and $F_{1}$ being the “closest” pair of distributions in the composite pre and post-change models.
-For instance, when the pre-change model is given by $\{F_{\theta}:\theta\leq a\}$ and the post-change model is $\{F_{\theta}:\theta\geq b\}$, we set $F_{0}=F_{a}$ and $F_{1}=F_{b}$ in Algorithm 2.
-The intuition behind this choice is that the least separated pair of distributions might form a “worst-case” scenario.
-Nonetheless, we do not have any theoretical guarantee for coverage for this method.
-For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section I.3.1 and we report the average size and the conditional and unconditional coverage across $500$ independent runs in Table J.1.
-Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section I.3.2 and we report the average size and the conditional and unconditional coverage across $500$ independent runs in Table J.1.
-From Table J.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
-Table J.1: Setting II: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (S11) with threshold $A=1000$.
-T Pre-change
-Composite
-post-change
-True (unknown)
-post-change
-Conditional
-coverage
-Unconditional
-coverage
-Size
-100 $N(0,1)$ $\{N(\mu,1):\mu>0.75\}$ $N(1,1)$ 0.95 0.95 23.16
-100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.94 0.94 25.83
-500 $N(0,1)$ $\{N(\mu,1):\mu>0.75\}$ $N(1,1)$ 0.95 0.88 23.82
-500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.90 27.17
-Table J.2: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (S12) with threshold $A=1000$.
-T
-Composite
-post-change
-Composite
-post-change
-True
-pre-change
-True
-post-change
-Conditional
-coverage
-Unconditional
-coverage
-Size
-100 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 26.03
-100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 29.64
-500 $\{N(\mu,1):\mu\leq 0.25\}$ $\{N(\mu,1):\mu>0.75\}$ $N(0,1)$ $N(1,1)$ 0.95 0.95 26.52
-500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 30.49
+Conditional delay
+100 0.918 23.245 31.929
+500 0.907 23.698 34.625
+Remark 7.
+We emphasize again that the universal approach described in Section 3 cannot handle dependent data, but the adaptive method in Section 4 can.
+To elaborate, our test statistic in Section 3 requires constructing non-trivial e-processes in both forward (usual) and backward (nonstandard) filtrations; in settings with general dependence, constructing the latter is generally infeasible.